### v1

# A flexible and general semi-supervised approach to multiple hypothesis testing

Abstract
Standard multiple testing procedures are designed to report a list of discoveries, or suspected false null hypotheses, given the hypotheses’ p-values or test scores. Recently there has been a growing interest in enhancing such procedures by combining additional information with the primary p-value or score. Specifically, such so-called “side information” can be leveraged to improve the separation between true and false nulls along additional “dimensions” thereby increasing the overall sensitivity. In line with this idea, we develop RESET (REScoring via Estimating and Training) which uses a unique data-splitting protocol that subsequently allows any semi-supervised learning approach to factor in the available side-information while maintaining finite-sample error rate control. Our practical implementation, RESET Ensemble, selects from an ensemble of classification algorithms so that it is compatible to a range of multiple testing scenarios without the need for the user to select the appropriate one. We apply RESET to both p-value and competition based multiple testing problems and show that RESET is (1) power-wise competitive, (2) fast compared to most tools and (3) is able to uniquely achieve finite sample FDR or FDP control, depending on the user’s preference.
* Corresponding authors: J.Freestone@maths.usyd.edu.au, uri.keich@sydney.edu.au
1 Introduction
Scientists are often interested in testing many null hypotheses, $\{H_{i}:i=1,\dots,m\}$. As an example, a scientist may be interested in assessing which genes exhibit a change in gene expression levels in response to a particular drug treatment. In this case, the $i$th null hypothesis $H_{i}$ states that “the $i$th gene exhibits no change in gene expression level”. This type of large-scale hypothesis testing is usually achieved by assigning a p-value, $p_{i}$, to each null hypothesis. The collection of p-values then undergoes a filtering process that rejects a subset of these null hypotheses that have sufficiently small p-values subject to a type-1 error rate control to minimize the number of misreported true null hypotheses.
The most common choice of error rate in this setup is the false discovery rate (FDR), which is defined as the expectation of the false discovery proportion (FDP) — the fraction of true null hypotheses in the reported list of rejections, or discoveries [2]. In this case, the filtering procedure ensures that the FDR is $\leq\alpha$ where $\alpha\in\left(0,1\right)$ is a prespecified threshold. Alternatively, we might be interested in controlling the FDP rather than its expectation in the following sense. Given a confidence parameter $\gamma$ and a threshold $\alpha$, the FDP is said to be controlled if $\mathbb{P}(FDP>\alpha)\leq\gamma$. Evidently, such FDP control reduces the risk that the realized FDP exceeds $\alpha$ which is not guaranteed by FDR control [39, 27, 19, 37, 16].
Determining a p-value for each null hypothesis can be challenging because (a) it requires knowledge of the underlying null distribution, and (b) it can be computationally intensive. A recently developed competition framework offers an alternative approach to FDP/FDR control that does not require computation of canonical p-values. Instead, all that these procedures require is a single draw, $\tilde{Z}_{i}$, from each null distribution. The null drawn $\tilde{Z}_{i}$ is then compared with $Z_{i}$, the originally observed score (test statistic) for $H_{i}$, thus conceptually providing a crude 1-bit p-value.
Those p-values are too crude to be used in typical p-value based filters. Instead, assuming that larger values of $Z_{i}$ provide more evidence against $H_{i}$, we compete each pair of scores $(Z_{i},\tilde{Z}_{i})$ by recording a label $L_{i}\in\{\pm 1\}$ indicating which was the largest of the two scores, as well as a winning score, $W_{i}=f(Z_{i},\tilde{Z}_{i})$. The function $f$ needs to be symmetric in its arguments and should be large if $Z_{i}$ is large, e.g., $W_{i}=Z_{i}\vee\tilde{Z}_{i}$, or if $Z_{i}$ is larger than its counterpart $\tilde{Z}_{i}$, e.g., $W_{i}=\lvert Z_{i}-\tilde{Z}_{i}\rvert$ [^1] In general, the functions can be hypothesis-dependent, i.e., $f_{i}$ instead of $f$.. For each true null hypothesis, $L_{i}$ is equally likely to be ±$\pm 1$ independent of all the scores $W=\{W_{i}\}$ and all other labels $L_{-i}$ because both scores were sampled from the same null distribution and $f$ is symmetric. Hence, the number of negative labels, those with $L_{i}=-1$, can be used to estimate the number of true nulls with positive labels among the top scoring hypotheses. Using this fact, the scores and labels $\{(W_{i},L_{i}):i=1,\dots,m\}$ are passed through a filter to reject a subset of the top scoring hypotheses, with either the FDR or FDP controlled depending on the user’s preference.
More recently, research has been focused on improving these filtering methods by leveraging side-information $\mathbf{x}_{i}\in\mathcal{X}$ that complements each p-value $p_{i}$ or score $W_{i}$ to better discriminate between the true and false null hypotheses. These methods utilize a variety of strategies but ultimately operate with the same goal in mind: the side-information is used to weight or rescore the hypotheses according to how confident we believe each hypothesis is a false null. Indeed, there are several such tools that augment p-values with side-information [25, 34, 33, 36, 50, 35]. When it comes to using side-information in the competition framework, we are aware of only one method with proper error-rate control along with a precise implementation, namely the recently published Adaptive Knockoffs [43] (AdaPT [33] offers a meta-procedure, but without a precise implementation).
In this paper, we offer a novel approach for multiple testing with side-information. The idea is to randomly split a suspected subsets of the true nulls into two sets: one set is used for training a semi-supervised learning model to discriminate the true and false null hypotheses, while the second is reserved for estimating the number of false discoveries. Accordingly, we refer to our new approach as RESET—REScoring via Estimating and Training. Our method flexibly incorporates any semi-supervised approach to help distinguish between true and false null hypotheses. Moreover, it works the same in the competition and p-value context and allows the user to choose between the choice of FDR or FDP control. We evaluate RESET in a range of simulations and real data applications to verify that it is a competitive alternative to existing methods in terms of statistical power and speed.
2 Background: competition-based multiple testing
Competition-based multiple testing was first established in the proteomics community where the goal is to infer which proteins, i.e. long chains of amino acids, are present in a biological sample [17]. Direct protein identification is challenging and so a bottom-up approach is taken where the proteins are broken into small chains of amino acids, called peptides, and the objective is to identify these peptides instead. Then, a process called liquid chromatography tandem mass spectrometry (LC-MS/MS) is used to isolate each sample peptide and subsequently generate a fragmentation spectrum. Each such spectrum can be thought of as a fingerprint of the peptide that generated the spectrum, allowing us to hypothesize which peptides are present in the sample.
This is done by searching each spectrum against a database of candidate peptides called targets to find its optimal peptide-spectrum match (PSM), along with a score indicating the strength of the match. Unfortunately, due to the incomplete nature of the database and the noise associated with generating the spectra, the PSM can be incorrect. Hence, we wish to employ a testing procedure for each null hypothesis $H_{i}$ that “the $i$th target does not belong to the sample”. In order to decide which null hypotheses to reject, we pair each target peptide with a decoy peptide that is generated by randomly shuffling or reversing the target peptide’s sequence. Accordingly, a second search is performed for each spectrum against the database of decoys to obtain a second PSM and associated score, and only the highest scoring PSM out of the two is retained.
As outlined in the introduction, we accompany $H_{i}$ with a score $Z_{i}$, defined as the maximum scoring PSM associated to the $i$th target peptide, and the higher the score, the more likely $H_{i}$ is false, i.e., the peptide is in the sample. Similarly we define $\tilde{Z}_{i}$ as the maximum scoring PSM associated to the $i$th target’s corresponding decoy peptide. For each true null hypothesis, we expect $Z_{i}$ and $\tilde{Z}_{i}$ to be drawn from the same (hypothesis-specific) null distribution, independently of everything else. Finally, we obtain a winning score $W_{i}=Z_{i}\vee\tilde{Z}_{i}$ and a label $L_{i}$ indicating whether the winning score was from the $i$th target or its decoy [41].
The statistics community independently formalized this idea in the context of variable selection where the goal is to identify relevant variables from $\{X_{i}:i=1,\dots,p\}$ that are associated with a response $y$. The procedure that achieves this while controlling the FDR is called the knockoff filter, which begins by pairing each variable $X_{i}$ with an artificial variable $\tilde{X}_{i}$, called a knockoff. In this paper, we consider the model-X version of the knockoff filter [7] which makes the following assumptions [^2] However, our results are equally applicable to the fixed-X model as well [1].. First, the joint distribution of the variables, $X$, is assumed to be known. Second, each observation $(X_{i,1},\dots,X_{i,p},y_{i})$ is sampled i.i.d from the joint distribution $F_{XY}$. Because $X$ is random and its distribution known, the knockoffs $\tilde{X}$ are also constructed randomly based on $X$, requiring that they satisfy the following conditions:
  1. 1.
[…]
  2. 2.
    We construct $\tilde{X}$ independently of $y$ by only looking at $X$, i.e., $\tilde{X}\perp\!\!\!\perp y\mid X$.
Once the knockoffs have been constructed, each variable is given a user-defined score $W_{i}:=W_{i}([X,\tilde{X}],y)\in\mathbb{R}$ and a label $L_{i}([X,\tilde{X}],y)\in\{\pm 1\}$. The idea is that the label $L_{i}$ indicates whether the variable $X_{i}$ or the knockoff $\tilde{X}_{i}$ has greater evidence for being relevant to the model, and the score $W_{i}$ indicates by how much. Commonly, the score and label are combined by setting $sign(W_{i})=L_{i}$, however as it will become clear in our later presentation of RESET, it is instructive to separate them. In this paper, we consider the Lasso Coefficent Difference (LCD) score, which is the difference in the magnitude of the coefficients of the Lasso regression [48]. Specifically, cross-validation is used to to determine the $\lambda$ penalty when regresson on the augmented design matrix $[X,\tilde{X}]$. We denote the absolute values of the Lasso($\lambda$)-fitted coefficients as $Z_{i}$ for the variables and $\tilde{Z}_{i}$ for the knockoffs. The scores and labels are then defined as:
$W_{i}:=\big{|}Z_{i}-\tilde{Z_{i}}\big{|},\quad L_{i}:=sign(Z_{i}-\tilde{Z_{i}}).$ (1)
Hypotheses with zero scores are typically ignored. In either context, peptide inference or variable selection, we expect that the labels of the true null hypotheses are independently and equally likely to be ±$\pm 1$, independently of everything else. This is formalized in the next assumption.
Assumption 1.
Let $N$ be the indices of the true null hypotheses. The labels $\{L_{i}:i\in N\}$ are i.i.d ±$\pm 1$ uniform random variables independent of all the scores $W$ and the labels of the false null hypotheses.
For the knockoff filter, Assumption 1 follows from the knockoff construction while in the peptide detection problem, Assumption 1 is based on our previous discussion that the scores of a null target and its decoy, $Z_{i}$ and $\tilde{Z}_{i}$, are independent draws from the same null distribution, and morever, this happens independently of all other pairs of scores. This assumption is commonly applied in the literature, with several analyses validating this claim empirically [38, 21].
Given the above assumption, the number of hypotheses with $L_{i}=-1$ (i.e. decoy or knockoff wins) that score above a score cutoff of $\tau$ can be used to estimate the number of true null hypotheses that score above $\tau$ with $L_{i}=1$. Using this strategy, we can report a list of discoveries while satisfying a type-1 error rate control in one of two ways. The first and by far the more common way is to apply Selective SeqStep+ (SSS+) [1] which controls the FDR at a prespecified threshold $\alpha$ by reporting the following list of discoveries $R_{\alpha}$:
$R_{\alpha}:=\{i\in[p]:W_{i}\geq\tau,L_{i}=1\},\quad\text{where }\tau:=\min\{t\in\mathbb{R}:\frac{\#\{W_{i}\geq t,L_{i}=-1\}+1}{\#\{W_{i}\geq t,L_{i}=1\}\vee 1}\leq\alpha\}.$ (2)
SSS+ is made precise in Algorithm S1 and works in a more general setting when the probabilities of the positive and negative labels of the true null hypotheses are no longer uniform, i.e., $\mathbb{P}(L_{i}=1)=c$ and $\mathbb{P}(L_{i}=-1)=1-c$. In the proteomics community, the application of Equation (2) in the peptide detection problem is referred to as target-decoy competition or TDC for short. As such, we will refer to the method as SSS+ in the variable selection setting, and TDC in the peptide detection setting.
Alternatively, we can control the FDP by using, for example, FDP-stepdown (FDP-SD, Algorithm S2[39]. FDP-SD works by first reranking the labels $L$ according to the scores $W$ in descending order. Next, it compares the number of decoy (knockoff) wins in the first $i$ hypotheses, denoted as $D_{i}:=\#\{L_{j}=-1:j\leq i\}$, with a precomputed bound $\delta_{i}$ which is determined by $\alpha$ and $\gamma$. Each comparison is made in a ‘stepdown’ fashion, i.e, if the first $i$ hypotheses satisfy $D_{1}\leq\delta_{1},\dots,D_{i}\leq\delta_{i}$, then it moves onto the next comparison at $i+1$. For small $i$, it is impossible for $D_{i}\leq\delta_{i}$ and so FDP-SD begins this analysis at $i_{0}$, the smallest possible index for which $D_{i}\leq\delta_{i}$. Under Assumption 1, reporting the following list of discoveries $R_{\alpha,\gamma}$ controls the FDP.
$R_{\alpha,\gamma}:=\{i\in[p]:i\leq k_{FDP},L_{i}=1\},\quad k_{FDP}:=\max\{i:\Pi_{j=i_{0}}^{i}1_{D_{j}\leq\delta_{j}}=1\text{ or }i=0\}.$ (3)
2.1 Extension to side information
In proteomics, so-called post-processors such as PeptideProphet [28] and Percolator [26] have popularized the use of side-information to improve the number of detected peptides. These tools use machine learning procedures, leveraging features associated with the PSMs, such as the charge state associated with the spectrum, and the matched peptide’s length, to better distinguish correct PSMs, and hence to deliver more discoveries. These tools have no theoretical guarantees and having recently pointed out scenarious where such post-processors can fail to control the FDR [18], we will not analyze these tools in this paper.
The only ‘intrinsic’ competition-based multiple testing procedure that uses arbitrary side information in a data-driven manner, that we are aware of, is the recent Adaptive Knockoffs (AdaKO) by Ren and Candès [43]. AdaKO is a specialization of the p-value based procedure AdaPT [33] that we discuss in the next section. By ‘intrinsic’, we mean that while other p-value based approaches may be modified to be used in the competition setting, they are significantly less powerful and do not really qualify in this sense, as demonstrated in [43]. Adaptive Knockoffs works by iteratively pruning $\mathcal{S}$, the candidate set of hypotheses, selecting each time the hypothesis it deems most likely to be a true null. To do that, it fits a classification algorithm using a strict set of information that partially masks the data: the scores $W$, the side information $\mathbf{x}$, the labels $L_{i}$ of the hypotheses that were pruned up to that point ($i\in[p]\setminus\mathcal{S}$), as well as #$\#\{L_{i}=1:i\in S\}$ and #$\#\{L_{i}=-1:i\in S\}$. In their paper, Ren and Candès investigate this prunning strategy using different types of “filters”: a logistic regression (LR), a generative additive model (GAM) which is the default filter in their R package, a random forest (RF) and a two-group probabilistic model which is fitted using the expectation-maximization algorithm (EM) [13]. The pruning process terminates when the estimated FDR is less than or equal to $\alpha$:
$\widehat{FDR}_{\mathcal{S}}:=\frac{\#\{i\in\mathcal{S},L_{i}=-1\}+1}{\#\{i\in\mathcal{S},L_{i}=1\}\vee 1}\leq\alpha.$ (4)
The remaining positively labelled hypotheses in $\mathcal{S}$ are then reported. It was shown that given the following natural extension of Assumption 1, Adaptive Knockoffs controls the FDR.
Assumption 2.
Let $N$ be the indices of the true null hypotheses. The labels $\{L_{i}:i\in N\}$ are i.i.d ±$\pm 1$ uniform random variables independent of all the scores $W$, the side information $\mathbf{x}$, and the labels of the false null hypotheses.
3 Background: p-value based multiple testing
Multiple testing with FDR control was popularized by Benjamini and Hochberg when they developed the first testing procedure of this kind, which we refer to as the “BH procedure” [2]. Since then, many p-value based methods that control the FDR have been developed to improve statistical power.
One such strategy for improving on BH is to estimate the fraction of true null hypotheses, denoted as $\pi_{0}$. Indeed, the BH procedure controls the FDR conservatively at level $\pi_{0}\cdot\alpha$ instead of $\alpha$. Hence, in cases when $\pi_{0}\ll 1$, the BH procedure loses a considerable amount of power. Methods like those by Storey et al. [45] and Benjamini et al. [3] estimate $\pi_{0}$, and essentially ‘plug-in’ the estimate $\hat{\pi}_{0}$, by applying the BH procedure at a threshold of $\alpha/\hat{\pi}_{0}$ instead of $\alpha$.
Instead of the FDR, an alternative to BH that controls the FDP, is Guo and Romano’s stepdown approach [20], which we refer to as GR-SD. In fact, this stepdown approach inspired the development of FDP-SD in the competition setting. GR-SD works by first sorting the p-values in ascending order, i.e. $p_{1}\leq p_{2}\leq...\leq p_{m}$. Then it constructs a sequence of thresholds, $\delta_{i}$ for each p-value $p_{i}$ and searches for the largest $k$ such that $p_{1}\leq\delta_{1},\dots,p_{k}\leq\delta_{k}$ mutually holds, and then reports the $k$ smallest p-values. We provide the pseudocode for this method in Algorithm S3.
3.1 Extension to side information
Other methods utilize side-information in one of two ways: (a) they provide an improved ordering of the p-values based on how confident we believe each hypothesis is a false null, or (b) by assigning weights to the p-values, placing higher weights on hypotheses that we believe to be false nulls. However, the shortcoming of these side-information approaches is that the majority of them are not data-driven, or ‘adaptive’, that is, the ordering or weighting is fixed before the actual testing procedure takes place.
Method Error control Speed
AdaPT (2018) Finite FDR Slow
AdaPTg (2021) Finite FDR Slow
AdaPT-GMMg (2021) Finite FDR Slow
ZAP-asymp (2022) Asymptotic FDR Fast
AdaFDR (2019) Asymptotic FDP Fast
Table 1: A table summarizing some popular and recent p-value based multiple testing procedures that use side information in terms of its type-1 error control and the computational speed.
More recently, several methods have been developed that take an adaptive approach instead. Such methods include Adaptive P-value Thresholding (AdaPT) and its derivatives [33, 8], Z-value Adaptive Procedures (ZAP) [35] and AdaFDR [50]. Table 1 provides a qualitative summary of each method in terms of its type-1 error control, its error control guarantees, and its computational speed. On balance, none of these mentioned tools enjoy both finite type-1 error control while also being fast, and as we will see, some of these methods have variable power when considering a range of simulated and real datasets. We next provide some background on each of those methods.
AdaFDR, splits the pairs $(p_{i},\mathbf{x}_{i})_{i=1}^{m}$ into two folds. For each fold, it fits a mixture model to optimally rescore the hypotheses, which is then applied to the other fold where a list of discoveries is obtained. The two list of discoveries are joined together and then reported. AdaFDR’s mixture model is a composition of Gaussian components to model local ‘bumps’ of genuine signals (false nulls) within the data plus a generalized linear model to capture a monotonic relationship between the side information and the signals.
AdaPT iteratively prunes a candidate set of hypotheses $\mathcal{S}$ similar to Adaptive Knockoffs. At each iteration, AdaPT fits a two group mixture model on a strict set of information: the p-values outside $\mathcal{S}$, the masked p-values in $\mathcal{S}$ given by $\{p_{i}\wedge(1-p_{i}):i\in\mathcal{S}\}$, the side information $\mathbf{x}$, as well as #$\#\{p_{i}>1/2:i\in\mathcal{S}\}$ and #$\#\{p_{i}<1/2:i\in\mathcal{S}\}$. It then removes the hypothesis from $\mathcal{S}$ that is deemed most likely to be a true null based on the fitted model. To fit the mixture model, the EM algorithm is used. The M-step of the EM algorithm results in a regression problem which allows the user to flexibly choose whichever regression method they like. They considered a generative additive model (GAM), a generalized linear model (GLM) and a generalized linear model with $L_{1}$ regularization (GLMnet). If we identify each p-value $p_{i}>1/2$ with a label $L_{i}=-1$ and each p-value $p_{i}<1/2$ with a label $L_{i}=1$, then AdaPT terminates this pruning procedure when the estimated FDR, like in Equation (4), is $\leq\alpha$. The rationale is that for true null hypotheses, the p-values are usually uniformly distributed, so #$\#\{L_{i}=-1:i\in\mathcal{S}\}$ estimates the number of true null hypotheses with positive labels, #$\#\{L_{i}=1:i\in\mathcal{S}\}$. It then reports the remaining positively labelled hypotheses in $\mathcal{S}$ [^3] AdaPT and AdaPTg do not use labels explicitly – we introduce those here for notational convenience.
AdaPTg generalizes AdaPT by constructing asymmetric regions to define the positive and negative labels. For example, the p-values $p_{i}\in\left(0.3,0.9\right)$ can be identified with a label $L_{i}=-1$ and the p-values $p_{i}<0.3$ can be identified with a label $L_{i}=1$. One advantage of defining these asymmetric regions, is to avoid overestimating the number of true null hypotheses when there is a concentration of p-values towards 1. Subsequently, the estimated FDR needs to be adjusted to account for the fact that we expect twice as many true null p-values with negative labels than positive labels. Moreover, the masked p-values are also no longer $p_{i}\wedge(1-p_{i})$ and also need adjustment. In essence, the rest of the procedure remains the same. Lastly, AdaPT-GMMg replaces the two group mixture model in AdaPT/AdaPTg for a Gaussian mixture model (GMM). Similar to AdaPT/AdaPTg, it also allows the user decide from a collection of regression methods to fit the model at the M-step of the EM algorithm.
Lastly, ZAP-asymp directly fits a beta mixture model to the data. ZAP-asymp’s direct application of the model means that it is only able to offer asymptotic FDR control. Importantly, ZAP-asymp can operate directly on the test statistic instead of just the corresponding p-values to rerank the hypotheses. This is particularly advantageous for two-sided tests because the test statistic is mapped to its p-value in a non-bijective manner, losing information regarding the sign of the original test statistic along the way. Note that AdaPT-GMMg can also operate on test statistics, but in this paper, we focus on the p-value information.
The list of methods we examine here is incomplete. Others include Independent Hypothesis Weighting [25, 24] which splits the p-values and side information into several folds and learns a weight for each hypothesis within that fold using out-of-fold information. This was shown to control the FDR asymptotically, and with some later adjustments to the method, finite sample FDR control was achieved. Another is Structure Adaptive Benjamini Hochberg Algorithm (SABHA) [36], which performs a reweighting of the p-values (similar to IHW) with finite sample FDR control. In addition, ZAP-finite is the analogue of ZAP-asymp which offers finite-sample FDR control. It uses the same principles of the AdaPT methods where each hypothesis is pruned one at a time. However, in previous simulations and real data experiments, IHW and SABHA appeared to be consistently outperformed by some of the methods in Table 1. Because we consider the same simulations and real data experiments, we did not include those methods in our comparisons. In addition, ZAP-finite is presented as a less powerful alternative to ZAP-asymp with the computational costs of the AdaPT methods. Given our comparisons to ZAP-asymp and the AdaPT methods, we decided to exclude ZAP-finite from our comparisons.
4 RESET (REScoring via Estimating and Training)
RESET is designed to control the FDR or the FDP in either the competition or p-value based setting. We next describe RESET’s main steps, with further details given in Algorithm S4. In Section S.3, we prove that RESET controls the FDR/FDP in the finite-sample setting. For simplicity, we use the target-decoy terminology, i.e., we refer to a hypothesis with $L=1$ as a ‘target win’ or a target for short, and $L=-1$ as a ‘decoy win’ or a decoy for short.
4.1 RESET outline
  1. 1.
    Intrinsically, RESET’s input consists of the labels, winning scores and side information $(L,W,\mathbf{x})$. Hence, in the p-value setting, each $p_{i}$ is first converted to a pair $(L_{i},W_{i})$ in the following way:
    $L_{i}=\begin{cases}+1&p_{i}\in[0,a)\\ -1&p_{i}\in(b_{1},b_{2}]\\ \end{cases},\quad W_{i}=\begin{cases}\lvert\Phi^{-1}(p_{i})\rvert&p_{i}\in[0,a)\\ \lvert\Phi^{-1}((b_{2}-p_{i})\cdot\frac{a}{b_{2}-b_{1}})\rvert&p_{i}\in(b_{1},b_{2}]\\ \end{cases},$ (5)
    where $0<a\leq 1/2\leq b_{1}<b_{2}\leq 1$ determine the cutoff regions for defining a positive and negative label, and $\Phi$ is the standard normal CDF. Hypotheses with p-values outside of $[0,a)\cup(b_{1},b_{2}]$ are thrown out and the default is $a=b_{1}=1/2$ and $b_{2}=1$.
  2. 2.
    RESET independently and randomly assigns each winning decoy as a training decoy with probability $s$ (we used $s=1/2$ throughout). The complementary set of decoy wins defines the estimating decoy set, and the set of pseudo targets is the set of all estimating decoys and target wins. The pseudo labels, $\tilde{L}_{i}$, denotes whether the $i$th hypothesis is a pseudo target ($\tilde{L}_{i}=1$) or a training decoy ($\tilde{L}_{i}=-1$).
  3. 3.
    Next, RESET applies a user-selected semi-supervised machine learning model which uses the pseudo labels, winning scores and side information $(\tilde{L},W,\mathbf{x})$. Note that this is a semi-supervised task as the positive pseudo labels will typically contain a mixture of true and false nulls, while the negative pseudo labels will indicate true null hypotheses.
    The output of this step is a rescoring of the training decoys and pseudo targets, denoted as $\tilde{W}$. Ideally, $\tilde{W}$ scores many of the false nulls among the pseudo targets higher than they were scored originally. We provide a specific framework that we developed for this crucial step in Section 4.3.
  4. 4.
    With the training complete, the training decoys are thrown out, and the original labels $L$ of the remaining pseudo targets are revealed.
  5. 5.
    RESET then determines its list of discoveries, by applying to the pseudo-targets using the original labels $L$ and the new scores $\tilde{W}$ either
    • SSS+ (Algorithm S1) at the desired level $\alpha$ for FDR control, or
    • FDP-SD (Algorithm S2) at the desired level $\alpha$ and confidence parameter $\gamma$ for FDP control.
    The two algorithms, SSS+ and FDP-SD, require an additional parameter $c$. This parameter is used to define the expected ratio of the number of null targets to decoys, $\frac{c}{1-c}$. Typically, this is set to $c=1/2$ for an expected target-decoy ratio of 1, however, RESET uses approximately half of the decoys for training purposes. Since those decoys are subsequently thrown out, this parameter needs to be adjusted. Indeed, assuming that the labels $L$ of the true nulls are i.i.d uniform ±$\pm 1$ RVs, and $s=1/2$, we set the parameter $c=2/3$, so that the expected target-decoy ratio is 2. More generally, if we know $\mathbb{P}(L_{i}=1)\leq c_{0}$, then we set $c=\frac{c_{0}}{1-s\cdot(1-c_{0})}$. For example in the p-value setting, $c_{0}=\frac{a}{a+b_{2}-b_{1}}$. This choice is made rigorous in Lemma S1.
We note that the above application of $\Phi^{-1}$ in Equation (5) is not strictly necessary for type-1 error rate control. We apply $\Phi^{-1}$ to the p-value $p_{i}$ or the mirrored p-value $(b_{2}-p_{i})\cdot\frac{a}{b_{2}-b_{1}}$ because we found that doing so makes the semi-supervised learning task of Step 3 to be generally less variable and more powerful (data not shown). The idea of mirrored p-value was established by Lei and Fithian [33] and later generalized by [8] as discussed in the previous section.
Crucially in Step 4, the training decoys are thrown out. This is important since during the training phase of Step 3, the semi-supervised machine learning model will have presumably learned to discriminate between the training decoys and the pseudo targets. Hence, the training decoys are no longer suitable for estimating the true nulls among those pseudo targets, because the rescoring will tend to rank them lower than those true nulls.
4.2 RESET controls the FDR or FDP
To establish RESET’s control of the FDR or FDP in both the competition and p-value settings, we generalize Assumption 2 when the distribution of the true null labels $\{L_{i}:i\in N\}$ is non-uniform. This is given below:
Assumption 3.
Let $N$ be the indices of the true null hypotheses. The labels $\{L_{i}:i\in N\}$ are i.i.d ±$\pm 1$ random variables with $\mathbb{P}(L_{i}=1)\leq c_{0}$ independently of all the scores $W$, the side information $\mathbf{x}$, and the labels of the false null hypotheses.
In the p-value setting, Assumption 3 is satisfied under some mild conditions regarding the distribution of the true null p-values. The value of $c_{0}$ is determined by the cutoff regions $[0,a)$ and $(b_{1},b_{2}]$ which are used to assign the positive or negative labels (Equation (5)). This is given by the following Lemma originally established by Chao and Fithian [8].
Lemma 1 (Chao and Fithian).
Assume that the true null p-values are mutually independent and independent of the false null p-values, all the side information, and assume the true null p-values have a non-decreasing density. Then Assumption 3 is satisfied with $c_{0}=\frac{a}{a+b_{2}-b_{1}}$.
The above Lemma is phrased differently by Chao and Fithian and is presented as Lemma A.1 in their paper, but it is materially the same. Most of the time, we will consider symmetric cutoff regions $[0,1/2)$ and $(1/2,1]$. For completeness, we provide a proof of this Lemma in Supplementary Section S.2 using our notation.
In practice, Assumption 3 (or the conditions of Lemma 1) needs to be verified. We provide reasonable justification of this assumption in our real data applications in Sections 5.2 and 6.2. We are now ready to state our two main theorems.
Theorem 1.
Under Assumption 3 RESET controls the FDR at the user-specified threshold of $\alpha$.
Theorem 2.
Under Assumption 3 RESET controls the FDP at the user-specified threshold of $\alpha$ with confidence $1-\gamma$.
We defer the main proof of RESET’s FDR and FDP control to Supplementary Section S.3. The proof requires an intermediate Lemma (Lemma S1) which states that the labels $L$ of the true null pseudo targets in Step 4 remaim i.i.d. ±$\pm 1$ RVs independently of everything else, even after sampling the training decoys in Step 2 and the application of the machine learning model in Step 3.
4.3 Semi-supervised approach
In this section we describe our specific framework for RESET’s Step 3, which as mentioned, can use any semi-supervised machine learning model. It consists of two man steps: (1) we apply a range of classification algorithms to estimate a high quality positive set from the data, containing presumably many false nulls, and (2) this positive set is then subsequently used for a second application of the classification algorithms to rescore the hypotheses.
  1. i.
    Define the negative set as the set of all training decoys and the positive set as a subset of the top scoring pseudo targets (which we elaborate in Step III at the end of this section). We define the features that are subsequently used by a collection of classification algorithms as the combined score and side information $(W,\mathbf{x})$.
  2. ii.
    For each considered classification algorithm, we randomly split all the hypotheses into $K$ folds (we used $K=3$ throughout). For each $k\in\{1,2,\dots,K\}$, we train the classification algorithm using the positive and negative set on the features in the $K\setminus\{k\}$ folds and apply the trained model to all the hypotheses in the $k$th test fold. We consider a random forest (RF), a generative additive model (GAM) and a collection of two-layer neural networks (NN) with decay parameter $\lambda\in\{0,0.1,1\}$ and a number of hidden layer nodes $h\in\{2,5,10\}$.
  3. iii.
    To reduce variability, we repeat Step ii $r$ times (we used $r=10$).
  4. iv.
    To evaluate each classification algorithm, we apply Selective SeqStep (SSS) (Algorithm S1 which is the same as SSS+ of Equation (2) without the “+1”) at an FDR threshold $\alpha$ to each of the $rK$ test folds, using the scores obtained from the classification algorithm and the pseudo labels. If $\mathbb{P}(L_{i}=1)\leq c_{0}$ then we set $c=1-s\cdot(1-c_{0})$. This is to ensure that the ratio $\frac{c}{1-c}$ correctly accounts for the number of null pseudo targets to training decoys [^4] This is not a requirement for error control..
  5. v.
    We record the total number of ‘discoveries’ produced by SSS over the $rK$ test folds from the previous step and select the classification algorithm that maximizes this value.
  6. vi.
    Using the chosen classification algorithm, we assign new scores $\tilde{W}$ to the hypotheses. $\tilde{W}_{i}$ is defined as the average decision value for the $i$th hypothesis, taken over the $K$ test folds that the $i$th hypothesis appears in, and the $r$ repetitions executed in Step iii.
  7. vii.
    We reapply Selective SeqStep (SSS) at $\alpha$ using the pseudo labels and the updated scores, $(\tilde{L},\tilde{W})$ with $c=1-s\cdot(1-c_{0})$.
  8. viii.
    The positive set is redefined as those pseudo targets that are ‘discovered’ in Step vii. Steps ii-vi are then repeated with the new positive set. To ensure the positive set is not too small, we increase $\alpha$ by increments of 0.01 until the positive set reaches min_positive or until all the pseudo targets are in the positive set. We used min_positive = 50.
  9. ix.
    The final scores $\tilde{W}$ are then reported.
We apply the following additional heuristics to our implementation that occur prior to Step 1 and leave some minor details in Section S.4.
  1. I.
    RESET throws out the hypotheses with undefined labels, i.e., those hypotheses with $W_{i}=0$. This effect is negligible when using p-values. However, in variable selection, the LCD score in the knockoff filter forces many of the winning scores to be zero ($W_{i}=0$), leaving typically a few hypotheses to be considered by RESET and consequently losing some information. We try to recover some of this information in the following way. For each hypothesis with $W_{i}\neq 0$, we identify the k-nearest neighbors in terms of the side information (zero-scoring hypotheses included). Then we define an extra side information variable as the number of zero-scoring hypotheses among these identified neighbors (we used the 20-nearest neighbors).
  2. II.
    To reduce the effect of ‘noisy’ side information that is unable to distinguish between true and false nulls, we implement an initial side information selection procedure. That is, we use a generative additive model to fit a smoothing spline using $W\cdot\tilde{L}$ as the response against each user-provided side information variable one at a time. We use tryCatch in case the smoothing spline fails, in which case we fit the response directly to the side information variable, that is, we use a linear fit. Again, we try to make use of the zero-scoring hypotheses where we can by setting $W\cdot\tilde{L}=0$ in these cases. We keep only the side information variables with sufficiently small p-values in the fitted model ($p<0.01$).
  3. III.
    Not all the pseudo targets need to be included in the initial positive set in Step (i). After all, we expect many of the pseudo targets to be true nulls. Hence, to improve the performance of RESET, we consider the following strategy. For each feature in $(W,\mathbf{x})$, reorder the hypotheses according to that feature, and apply SSS to the pseudo labels $\tilde{L}$ at the FDR level $\alpha$. Then select the feature that maximizes the number of pseudo discoveries. Then using the chosen feature, we define the positive set as those pseudo targets discovered at $\alpha_{0}$ (we used $\alpha_{0}=50\%$ throughout), again using SSS with the pseudo labels $\tilde{L}$. Usually the score is selected, but in some cases the side information is highly informative and is selected instead.
    Users interested in a small FDR level, like $\alpha=1\%$, could probably benefit from reducing $\alpha_{0}$ to further improve the performance, and in particular the speed of RESET (though as we mentioned, we kept $\alpha_{0}=50\%$ across the board). To ensure the positive set is not too small, we increase $\alpha_{0}$ by increments of 0.01 until the positive set reaches min_positive or until all the pseudo targets are in the positive set.
We refer to the general RESET wrapper in combination with our semi-supervised approach as RESET Ensemble. We refer to RESET RF/GAM/NN when the ensemble of machine learning methods in Step 2 is reduced to a random forest, generative additive model, and a collection of neural networks as described, respectively.
4.4 Code availability
Our code is available at https://github.com/freejstone/stat_RESET_paper_code. We intend on developing a package for RESET in R.
5 Experiments: competition setting
5.1 Numerical experiments
In this section, we conduct a wide range of simulations in the variable selection problem based on those introducted by Ren and Candès [43]. We compare RESET Ensemble to Adaptive Knockoffs (AdaKO) and the generic Knockoff Filter (KO) (see Section 2). Further implementation details for AdaKO are given in Supplementary Secton S.5. We provide a brief description of each simulation setup, referring to [43] for further details. Each setup is designed to satisfy Assumption 3, which is required for our theoretical guarantees. The labels and scores are obtained according to Equation (1). We focus on FDR control in this section, and study FDP control in Section 5.2.
5.1.1 Simulation 1: Linear model with one-dimensional side information
Our first simulation is based on Simulation 1 by Ren and Candès [43]. They used $n=1000$ independent observations from the joint data distribution of $(X,Y)$, where $X$ is drawn from a hidden markov model (HMM) with $p=900$ variables and $Y|X\sim\mathcal{N}(X\beta,1)$ is a linear model. The $\beta$’s are selected in the following way. Let $k$ denote the number of nonzero $\beta$’s. Then $k$ indices from $\{1,2,\dots,p\}$ are sampled without replacement with indices $i\in\{1,\dots,2k\}$ having probability proportional to $1/i^{2}$ and the indices $i\in\{2k+1,\dots,p\}$ having zero probability. The $k$ chosen indices are the nonzero $\beta$’s and are assigned a value of ±$\pm 3.5/\sqrt{n}$ where the signs are chosen i.i.d uniformly. We consider $k\in\{50,150,300\}$ while the original simulation only considers $k=150$. Each hypothesis is supported by the side information $\mathbf{x}_{i}=i$, which should help in detecting the nonzero $\beta$’s as smaller values of $i$ are more likely to correspond to a nonzero $\beta_{i}$. We used 100 independent runs of the above, where each run consisted of drawing those $n$ observations. The original simulation in [43] sampled the $\beta$’s once and fixed them across the 100 runs. Here we redrew the $\beta$’s for each run so that we may average our results over a range of models, as we noticed significant variability between the draws.
5.1.2 Simulation 2: Linear model with two-dimensional side information
In a second simulation, we considered a larger linear model where $n=2000$ observations are taken from the joint distribution $(X,Y)$ as described in Section 5.1.1, with $p=1800$. Accordingly, we also adjusted the number of false nulls to $k\in\{150,300,450\}$. Each hypothesis is equipped with the following two-dimensional side information. For the $i$th hypothesis, we draw a pair of observations $\mathbf{x}_{i}$ from a bivariate normal $\mathcal{N}(\boldsymbol{\mu},I)$ with $\boldsymbol{\mu}=(0,0)$ for a true null hypothesis and $\boldsymbol{\mu}=\pm(2,2)$ for a false null hypothesis, where the signs are chosen i.i.d uniformly. We used 20 independent runs of this larger simulation instead of 100 where in each run we draw those $n$ observations, $\beta$ coefficients and side information variables.
5.1.3 Simulation 3: Logistic model with two-dimensional side information
Next, we consider Simulation 2 by Ren and Candès [43]. They generate $n=1000$ observations with $p=1600$ variables sampled from the joint data distribution of $(X,Y)$ where $X$ is distributed according to a multivariate normal distribution and $Y|X\sim\text{Bernoulli}\left(e^{\beta\cdot X}/\left(1+e^{\beta\cdot X}\right)\right)$ is a logistic model (see [43] for further details). Each hypothesis’ side information is a unique pair of values $\mathbf{x}_{i}=\left(\mathbf{x}_{i1},\mathbf{x}_{i2}\right)$ from the lattice $[-20,19]\times[-20,19]\subseteq\mathbb{Z}^{2}$. The nonzero $\beta$’s are chosen according to the position of $\left(\mathbf{x}_{i1},\mathbf{x}_{i2}\right)$ in the lattice. An image of the nonzero $\beta$’s is given in Figure 1. Each nonzero $\beta$ is set to ±$\pm 25/\sqrt{n}$ where the signs are chosen i.i.d uniformly. We used 98 runs of the above setup (since in 2 runs, AdaKO EM failed with an error), where each run consisted of drawing those $n$ observations. The original simulation in [43] sampled the sign of the $\beta$’s once, fixing them across the 98 runs. Instead, we redraw the sign of the $\beta$’s in each of the 98 runs since we found the results significantly varied with those draws.
Figure 1: Location of nonzero $\beta$’s according to the pair of values $(x,y)$ in the lattice. The same image is given in [43].
Figure 2: Scatterplot of each hypothesis according to the LCD score denoted as $W$ (x-axis) and side information variable (y-axis). Each hypothesis is colored according to whether it is a true (red) or false null hypothesis (blue). In (A), the side information variable is the index of the hypothesis in Simulation 1 (Section 5.1.1). In (B), we use one of the three side information variables as described in Simulation 4 (Section 5.1.4).
5.1.4 Simulation 4: Large $p$ and $n$ and 3-d side information
Finally, in the last simulation, $n=10K$ observations with $p=10K$ variables and $k=0.15p$ relevant variables are drawn from the joint data distribution $(X,Y)$ exactly as described in Simulation 1. Each variable is equipped with the following three-dimensional side information. The first $k$ variables are randomly assigned a drawn from $i\in\{1,\dots,2k\}$ taken without replacement and having probability proportional to $1/i^{2}$. The remaining $p-k$ variables are uniquely assigned to one of the remaining values from $\{1,2,\dots,p\}$ uniformly at random. We expect that most of the nonzero $\beta$’s to have small values of $i$, but the effect of the random assignment makes the resulting side information variable less informative than in Simulation 1 as depicted in Figure 2. This process is repeated three times, thus associating three side information variables with each hypothesis, which taken together, should be reasonably informative. We conducted 5 independent runs where in each run we drew the $n$ observations, $\beta$ coefficients and side information.
5.1.5 Results
We first compare RESET Ensemble with Adaptive Knockoff’s default method in R, AdaKO GAM, and the generic Knockoff Filter (KO). In Figure 3, we plot the power of each simulation using these methods as a function of the selected FDR thresholds of $\alpha\in\{0.05,0.1,0.2,0.3\}$. The power is calculated as the proportion of relevant variables correctly discovered, averaged over the number of runs used in each simulation. Similarly, in Supplementary Figure S2, we plot the estimated FDR as the proportion of irrelevant variables in the discovery list, averaged over the runs.
Figure 3: RESET Ensemble vs. AdaKO GAM vs. KO. Each panel plots the power for each method at FDR thresholds ranging from 5% to 30%. The first row corresponds to Simulation 1 with three values of $k\in\{50,150,300\}$, the second row corresponds to Simulation 2 with three values of $k\in\{150,300,450\}$, and the last row corresponds to Simulation 3 and 4. For readability, the points are jittered in the horizontal direction. A description of AdaKO GAM and KO can be found in Section 2.
Unsurprisingly, RESET Ensemble uniformly dominates KO, demonstrating it successfully leverages the side information. At the same time, AdaKO GAM mostly improves on KO, with some exceptions.
In Simulation 1, we find that at higher FDR thresholds, RESET Ensemble is the preferred approach in terms of power. On the other hand, AdaKO GAM is more powerful than RESET at lower FDR thresholds, particularly in the case of $k=50$ and $\alpha=5\%$. This is not too surprising given the following two reasons. First, RESET trains its classification algorithms on a relatively small positive set of false nulls. Second, RESET incurs a greater cost in terms of power at lower FDR thresholds if the classification algorithm overits and incorrectly scores a knockoff higher than it should. Indeed, at an FDR threshold of $\alpha=5\%$, each knockoff may cost up to $402/\alpha=$ 40 discoveries.
We find in Simulations 2 and 3 that RESET Ensemble is consistently more powerful than AdaKO GAM; in fact, AdaKO GAM appears to essentially have the same power as KO. The result in Simulation 3 is consistent with the findings of Ren and Candès in their analagous simulation. AdaKO GAM does not implement any smooth terms when the number of side information variables is two or more, and is therefore unable to leverage the side information if the true and false null hypotheses are not “easily” separable. Finally in Simulation 4, we see that RESET Ensemble is generally the most powerful, with essentially the same power as AdaKO GAM at $\alpha=5\%$. We hypothesize that the greater number of variables in this example, particularly relevant variables, appears to overcome the relatively poor performance of RESET Ensemble in Simulation 1 at low FDR thresholds.
Figure 4: Relative power and runtimes for each method. In (A), each boxplot describes the relative powers for each method, computed across the combination of simulations and FDR thresholds in Figure 3. (B) shows the log10 runtimes for each method applied at the 5% FDR level for 5 runs of Simulation 4. For readability, the points are jittered in the horizontal direction.
We compared RESET Ensemble with AdaKO GAM because those are the respective default methods. We complement this comparision with Figure 4A where we summarize the performance of all RESET and Adaptive Knockoff methods. Specifically, for each panel and FDR threshold in Figure 3, we determined the maximum power across all methods and computed the relative power of each method to this maximum. The resulting relative powers are reported as boxplots, giving an indication of the performance of each method across the simulations as a whole. A plot of the actual powers for each of the methods is given in Supplementary Figure S3. While in Supplementary Figure S3 we often find a version of AdaKO that is more powerful, or at least comparable to RESET Ensemble, this varies between the versions. On the other hand, RESET Ensemble’s performance is consistent across these simulations, and overall typically achieves the highest relative power when the results are aggregated. Interestingly, RESET NN seems to be be marginally more powerful than RESET Ensemble. While this may be the case, we still recommend RESET Ensemble as the default approach to alleviate the burden of selecting any singular classification algorithm.
One advantage of RESET is its computational speed. Indeed, our implementation of RESET, as described in Section 4.3, has many steps that can be parallelized: the $r$ evaluations of each classification algorithm on the $K$ folds can all be done in parallel using multiple cores. In contrast, Adaptive Knockoffs must iteratively apply a classifiation algorithm for every nonzero scoring hypothesis in the candidate set $\mathcal{S}$ until the estimated FDR is $\leq\alpha$, or all hypotheses are revealed. Figure 4B demonstrates this advantage and shows the (log10) runtimes of RESET using 10-cores and Adaptive Knockoffs at the 5% FDR level across 5 runs of Simulation 4. We find that each version of RESET is significantly faster than each version of Adaptive Knockoffs. RESET Ensemble is around $10^{3}$ to $10^{4}$ times faster than AdaKO EM, $10^{2}$ to $10^{3}$ times faster than AdaKO RF, and about 3-4 times faster than AdaKO LR and AdaKO GAM. In scenarios where AdaKO EM and AdaKO RF are infeasible, but AdaKO GAM and AdaKO LR are unable to leverage the side information as in Simulations 3 and 4, RESET is arguably more suitable. We investigate such a scenario in the following application.
5.2 Application to peptide detection
In this section we compare RESET to Adaptive Knockoffs in the peptide detection context using mass spectrometry data and show that (a) the Adaptive Knockoff methods are often computationally infeasible, (b) feasibility aside, RESET is more powerful than Adaptive Knockoff’s default filter, AdaKO GAM, when controlling the FDR and (c) RESET is able to typically discover more peptides than the vanilla FDP-SD when controlling the FDP. We provide justification of Assumption 3 which is required for RESET’s theoretical guarantees in Section S.15.
5.2.1 HEK293 data
We show that the application of Adaptive Knockoffs is often too slow to be used effectively in high throughput scenarios such as the peptide detection problem. We demonstrate this, using HEK293 data [9], a popular dataset used in benchmarking the development of new software [30, 49, 32] that we downloaded from the Proteomics Idenitification Database (PRIDE) [40]. After data preparation and searching the combined 24 HEK293 spectrum files, we determined the scores $Z_{i}$ for each target peptide and $\tilde{Z}_{i}$ for each decoy peptide using the XCorr score function (see Supplementary Section S.12 for details). Next, we competed each pair by recording the score $W_{i}=Z_{i}\vee\tilde{Z}_{i}$ and label $L_{i}=1$ if $Z_{i}>\tilde{Z}_{i}$ and $L_{i}=-1$ if $Z_{i}<\tilde{Z}_{i}$ (ties were randomly broken). In addition to the primary score associated with each peptide-spectrum match (PSM) is a set of features, such as the charge state or peptide length. Each winning peptide inherits these features, or side information $\mathbf{x}_{i}$, from its maximally associated PSM, allowing us to define the desired triples $(L,W,\mathbf{x})$. A complete description of the side information used is given in Table S4 of the Supplement.
Method Time
RESET Ensemble 27.5 minutes
AdaKO GAM 1.06 days
AdaKO EM 33.71 days
AdaKO LR 3.35 days
AdaKO RF 836.03 days
Table 2: The average runtimes for each method at 1% FDR on the HEK293 dataset. The average runtimes reported for each Adaptive Knockoffs method is estimated (as described in Section S.13), while we computed the actual average runtime for RESET Ensemble. The average is taken over 5 randomly constructed decoy databases. We used 20 cores for RESET Ensemble.
Finally, using the method described in Supplementary Section S.13, we estimated the time it would take for each method of Adaptive Knockoffs to report discoveries at the 1% FDR level (which is the typical FDR threshold in the peptide detection context). Table 2 shows the average estimated times taken over 5 distinct target-decoy databases, with decoys varying each time by randomly shuffling the target peptides. It is clear that running any implementation of AdaKO is generally too expensive, with the worst being AdaKO RF at an estimate of 836 days, and the default method, AdaKO GAM, taking an estimate of 1.06 days. In contrast, we were able to successfully run RESET Ensemble with FDR control in 27.5 minutes.
5.2.2 PRIDE data
Due to the aforementioned runtime problem, we had to use small datasets when evaluating the power of Adaptive Knockoffs in the peptide detection context. Specifically, we focused on thirteen spectrum files (each with $<$ 33K PSMs) from the PRIDE-20 dataset that we recently used in [18]. In the following results, we only considered searching the spectrum files once against a single target-decoy database since unlike RESET we can only feasibly apply a single application of Adaptive Knockoffs. Data preparation, search settings and details on how each triple $(L,W,\mathbf{x})$ is obtained are given in Supplementary Section S.14 along with the list of side information used in Supplementary Table S5.
Figure 5: Comparison of RESET Ensemble and AdaKO GAM. For each of the 13 PRIDE datasets, we computed the average number of discoveries using 10 applications of RESET Ensemble, varying RESET’s internal seed. The figure shows a boxplot of those average RESET Ensemble discoveries divided by the number of discoveries using AdaKO GAM (left boxplot) and TDC (right boxplot) using the 13 PRIDE-20 datasets (there are total of 13 points for each boxplot).
Figure 5 shows the average number of discoveries reported using 10 applications of RESET Ensemble, each applied at 1% FDR to the same dataset, divided by the number of discoveries reported by AdaKO GAM (left boxplot) and TDC (right boxplot). In each of the 10 applications of RESET, we varied the internal seed which randomly splits the decoys into their training and estimating sets. Clearly, RESET Ensemble is a more powerful alternative than AdaKO GAM and TDC. Interestingly, the relative ratios in the number of discoveries between AdaKO GAM and TDC suggests that AdaKO GAM is reporting less discoveries than the generic TDC.
In a second comparision, we looked at the performance of RESET Ensemble versus AdaKO RF because in this setup it was the most powerful implementation of Adaptive Knockoffs. In Figure 6A, the left boxplot displays the average number of discoveries by RESET Ensemble, divided by the number of discoveries by AdaKO RF. In six of the thirteen PRIDE spectrum files, we found RESET Ensemble to be the superior method, while in the remaining seven spectrum files, AdaKO RF produced more discoveries. In one data set, AdaKO RF discovered 1217 peptides while RESET Ensemble only discovered 1104 peptides, about a 10% increase over RESET Ensemble.
We also assessed the computational times of AdaKO RF in Figure 6B, using both the actual and estimated runtimes as outlined in Section S.13. Reassuringly, our estimated times were less than the actual times taken by each method to complete thus validating our estimated times in the previous section. Although we find that, when analyzing these smaller spectrum sets, AdaKO RF is the most powerful method of Adaptive Knockoffs, and is competitive to RESET Ensemble power-wise, it is again clear that AdaKO RF is generally too slow with some runtimes exceeding multiple days even for these small datasets. In fact, on the dataset where AdaKO RF outperforms RESET Ensemble by about 10%, it takes 3.07 days for AdaKO RF to discover its peptides while RESET Ensemble takes about half a minute to complete. Figure 6C shows that in the worst case, RESET takes roughly 1 minute. The issue of runtimes is especially problematic considering the fact that multiple spectrum files are often analyzed jointly, as in our HEK293 analysis in the previous section.
Figure 6: Comparison of RESET Ensemble and AdaKO RF. For each of the 13 PRIDE datasets, we computed the average number of discoveries using 10 applications of RESET Ensemble, varying RESET’s internal seed. The left panel shows a boxplot of those average RESET Ensemble discoveries divided by the number of discoveries using AdaKO RF (left boxplot) and TDC (right boxplot) using the 13 PRIDE-20 datasets (there are total of 13 points for each boxplot). In the centre panel, we compare the estimated and actual runtimes for AdaKO RF. Lastly, in the right panel, we compare the actual runtimes of RESET Ensemble and AdaKO RF (in log10 scale).
5.2.3 FDP control
Lastly, we evaluated RESET Ensemble’s FDP control using the HEK293 and PRIDE-20 spectrum files from the previous sections. As in Section 5.2.1, the joint HEK293 spectrum files were searched against 5 distinct target-decoy databases, where each target-decoy database was generated by concatenating the target database to a decoy database consisting of shuffled target peptides. Similarly, each PRIDE-20 spectrum file was searched against 10 distinct target-decoy databases. We used the coinflip version of FDP-SD [39], which provides a uniform increase in the number of discoveries (at the cost of some randomness in the discovery list, see Supplementary Algorithm S2). Our results using RESET Ensemble and FDP-SD are given in Table 3 and Figure 7.
Confidence $1-\gamma$
Method 0.5 0.8 0.9
RESET Ensemble 82987 82729 82558
FDP-SD 75977 75464 75254
Table 3: The average number of discoveries at 1% FDP threshold at varying confidence parameters using the HEK293 data. The average number of discoveries reported for RESET Ensemble and FDP-SD on the combined HEK293 spectrum files at different confidence values $0.91-\gamma=0.5,0.8,$ 0.9. The averages are taken over 5 applications of each method, once for each of the 5 combined target-decoy databases.
Analyzing the HEK293 results first, RESET Ensemble is clearly more powerful with an approximately 9-10% increase in the average number of discoveries over FDP-SD at all confidence levels. Analyzing the PRIDE-20 datasets, while RESET Ensemble appears to be generally more powerful, in a few datasets, it fails to make any discoveries at high confidence levels. Upon a closer look, the numbers of discoveries made by FDP-SD in these datasets are relatively small, with the largest average number of discoveries among these datasets being 417. When the number of discoveries is expected to be small and the confidence parameter is high, RESET may yield less power, even if the side information is reasonably informative. We dwell on this fact in the Discussion. Regardless, more often than not RESET Ensemble produces more discoveries and if multiple spectrum files are analyzed jointly, as we do with the HEK293 spectrum files, then the risk of having too few discoveries is reduced.
Figure 7: Comparison of RESET Ensemble and FDP-SD. The left panel shows boxplots of the average number of discoveries using RESET divided by the average number of discoveries using FDP-SD for each of the PRIDE-20 datasets at a 1% FDP threshold and varying confidence parameters. The averages are taken over 10 applications, one for each of the 10 target-decoy databases used for each PRIDE-20 dataset. The right panel is a ‘zoomed’ in version of the left panel.
The results above consider a canonical ‘narrow’ search. Specfically, during the searching phase, each database peptide and spectrum are only allowed to match if the mass of the database peptide and the mass of the sample peptide generating the spectrum, the ‘spectrum mass’, differ by a small amount. Recently, ‘open’ searching, where each spectrum and peptide can match even if their masses differ by hundreds of Daltons, has become a popular method for detecting peptides [30]. Such a search allows for the detection of peptides whose mass have been altered by post-translational modifications (PTMs). We provide further detail in Supplementary Section S.14. In this setting, side information variables such as dm, which records the mass difference between the peptide and the spectrum, could be highly informative given that some differences will correspond to masses of common PTMs. Thus, it is not surprising that RESET, which can take advantage of such side information, does even better than FDP-SD in this context. Specifically, Figure 8 shows a greater improvement over FDP-SD across all confidence levels when analyzing open searching results.
Figure 8: Comparison of RESET ensemble and FDP-SD using the recently popular ‘open’ search. The left panel shows boxplots of the average number of discoveries using RESET divided by the average number of discoveries using FDP-SD for each of the PRIDE-20 datasets at a 1% FDP threshold and varying confidence parameters. We implement an ‘open’ search for each PRIDE-20 dataset (see Section S.14). The averages are taken over 10 applications, one for each of the 10 target-decoy databases used for each PRIDE-20 dataset. The right panel is a ‘zoomed’ in version of the left panel.
6 Experiments: p-value setting
6.1 Numerical experiments
In this section, we consider the p-value setting, comparing RESET with the FDR controlling procedures we reviewed in Section 3.1 on a set of simulations introduced by Lei and Fithian [33]. As in the competition setting, each setup is designed to satisfy Assumption 3, which is required for our theoretical guarantees. Details of each method used is given in Supplementary Section S.6-S.10.
6.1.1 Simulation 5: p-value based testing with geometric side information
In this section, we look at Simulation 1 of Lei and Fithian [33]. For each hypothesis, they obtain a p-value $p_{i}$ from a one-sided normal test so that $p_{i}=1-\Phi(z_{i})$ where $z_{i}\sim\mathcal{N}(\mu,1)$ with $\mu=2$ for a false null hypothesis and $\mu=0$ for a true null hypothesis. A total of $m=2500$ hypotheses are used. Similar to Simulation 3 of Section 5.1.3, each hypothesis’ side information is defined as unique pair of values $\mathbf{x}_{i}=\left(\mathbf{x}_{i1},\mathbf{x}_{i2}\right)$ from $50\times 50$ equally spaced values in the square $\left[-100,100\right]\times\left[-100,100\right]$. The false null hypotheses correspond to certain regions in the square. They considered three scenarios: (a) when the false nulls form a circle in the middle of the square, (b) when the false nulls form a circle in the top right of the square and (c) when the false nulls form an ellipse. Each of these scenarios is depicted in Figure 9. We used 100 independent runs where in each run, we randomly draw the 2500 p-values as described.
Figure 9: Location of false null hypotheses according to the pair of values $(x,y)$ in the square $\left[-100,100\right]\times\left[-100,100\right]$. The same picture is given in [33].
Figure 10: RESET vs. other methods using two-dimensional side information with p-values. The first row plots the power and the second row plots the estimated FDR for each FDR threshold ranging from 1% to 20%. Each color and shape combination corresponds to a unique method. The three columns correspond to the hypotheses’ relationship to the side-information according to Figure 9. The horizontal segments in the second row of panels mark the corresponding FDR thresholds on the y-axis. For readability, each point is jittered in the horizontal direction.
Figure 10 shows the results of RESET compared with several p-value based methods that we reviewed in Section 3.1. We find that in the first two simulations, (a) and (b), RESET Ensemble is roughly equal best in terms of power at the 5%, 10% and 20% FDR levels, along with AdaPT/AdaPTg GAM. However, in (c), clearly AdaPT/AdaPTg GAM are more powerful than the RESET methods, with RESET Ensemble coming arguably as third in the power ranking at the 5%, 10%, and 20% FDR levels. At 1% FDR, all RESET methods have negligible power, although this is consistent with many of the other methods. Interestingly, many of the finite-FDR controlling procedures such as the various RESET and AdaPT methods generally have better power than the asymptotic approaches, ZAP and AdaFDR, suggesting perhaps a misspecification of the fitted model used by these approaches. Figure 10 also shows the estimated FDR of each method. We find that in (a), AdaFDR’s estimated FDR at 20% (21.2% estimated FDR with 50% power) slightly exceeds the FDR threshold. This is possibly explained by the fact that AdaFDR only guarantees asymptotic FDP control. Consisitently, Chao and Fithian [8] reported inflated estimated FDRs at high thresholds in simulation studies for AdaFDR.
6.1.2 Simulation 6: p-value based testing with 100-dimensional side information
In Simulation 2 of Lei and Fithian [33], the side information for $m=2000$ hypotheses consists of 100 i.i.d draws from $U[0,1]$. This setup simulates ‘noisy’ side information since only the first 2 of the 100 draws are used in distinguishing between the true and false null hypotheses. Each p-value $p_{i}$ is drawn i.i.d from a two-group beta mixture model, where the true nulls have a uniform distribution and the false nulls have an alternative beta distribution. Specifically, the beta mixture model is given by:
$f(p_{i}\mid\mathbf{x}_{i})=1-\pi(\mathbf{x}_{i})+\pi(\mathbf{x}_{i})\cdot h(p_{i};\mu(\mathbf{x}_{i})),\quad h(p_{i};\mu(\mathbf{x}_{i}))=\left(\frac{1}{\mu(\mathbf{x}_{i})}p_{i}^{\frac{1}{\mu(\mathbf{x}_{i})}-1}\right),$ (6)
where $\mathbf{x}_{i}\in[0,1]^{100}$ denotes the side information of the $i$th hypothesis, $\pi(\mathbf{x}_{i})$ denotes proportion of the false nulls in the mixture, $h$ denotes the false null distribution of $Beta(1/\mu(\mathbf{x}_{i}),1)$. The $\pi(\mathbf{x}_{i})$’s and $\mu(\mathbf{x}_{i})$’s are determined by the following relationships with the 100-dimesional side information:
$\log\left(\frac{\pi}{1-\pi}\right)=\theta_{0}+\mathbf{x}_{i}^{T}\theta,\quad\mu_{i}=\max\{\mathbf{x}_{i}^{T}\beta,1\},$
where $\theta=(3,3,0,\dots,0)\in\mathbb{R}^{100}$, $\beta=(2,2,0,\dots,0)\in\mathbb{R}^{100}$ and $\theta_{0}$ is chosen to satisfy $\frac{1}{m}\sum_{i}\pi(\mathbf{x}_{i})=0.3$. Clearly, only the first two values in $\mathbf{x}_{i}\in[0,1]^{100}$ contribute to the beta mixture model, as intended. We used 100 independent runs, where in each we drew the 2000 p-values as described.
Figure 11: RESET vs. other methods using 100-dimensional side information with p-values. The left panel plots the power and the right panel plots the estimated FDR as a function of the FDR thresholds: 1%, 5%, 10% and 20%. Each color and shape combination corresponds to a unique method. The horizontal segments in the right panel mark the corresponding FDR thresholds on the y-axis. For readability, each point is jittered in the horizontal direction.
There were a couple of tools we had to omit from the analysis of this data. Suprisingly, we found that AdaFDR was too slow in this simuation. Similarly, we were not able to apply AdaPT GAM or AdaPTg GAM as we did in the previous simulation, since it would have required using a smoothing spline on all 100 side-information variables, which is impractical. Hence, we only applied the GLMnet variants of AdaPT.
Examining Figure 11, which graphically summarizes this analysis, it appears that RESET Ensemble, RESET NN, and AdaPT GLMnet deliver the most discoveries at the 5%, 10%, and 20% FDR levels. At 1% FDR, only AdaPTg GLMnet appears to have non-negligible power, albeit a fairly low number of discoveries. Finally, we point out that in this simulation, the model specification of AdaPT/AdaPTg perfectly matches the true model. In other words, the fitted mixture model used by AdaPT/AdaPTg coincides with Equation (6). Given this, we find it reassuring that RESET Ensemble is still comparable at 5% and higher thresholds.
Notably, the performance of AdaPT in these last two simulation setups is highly dependent on the choice of implementation. In Simulation 5, AdaPT GAM significantly outperforms AdaPT GLMnet while in Simulation 6, AdaPT GLMnet is superior (and AdaPT GAM is not practical). In contrast, RESET Ensemble selects from a collection of flexible machine learning models to adapt itself to the problem at hand.
6.2 Real data experiments with p-values
We next evaluate RESET and related methods on a collection of publically available datasets with p-values that were investigated by Zhang et al. [50] when introducing AdaFDR and by Lei and Fithian [33] when introducing AdaPT. Since AdaFDR and AdaPT operate under essentially the same assumptions as RESET’s, we expect these datasets to therefore satisfy Assumption 3 with the exception of two datasets, which we discuss at the end of this section. We provide a brief summary of each dataset below:
  • Three RNA-seq datasets (Airway [22], Bottomly [4], Pasilla [6]): the goal is to identify genes that are associated with varying gene expression levels in response to a change of conditions. In the Airway data, the differential expression analysis is conducted in response to a drug, dexamethasone; in the bottomly data, the analysis is conducted between two mouse strains; and in the Pasilla data, the analysis is conducted between normal and so-called ‘Pasilla knockdown’ conditions [23]. In all cases the side information consists of the log-normalized gene counts.
  • A proteomics dataset (Proteomics [14]): the goal is to identify proteins that have different protein abundances when treated with rapamycin versus DMSO in yeast cells. Each protein consists of side information equal to the natural log of the number of peptides “quantified” across all samples, i.e., the number of peptides belonging to the protein with their abundances measured.
  • Two microbiome datasets (Microbiome Enigma Ph and Microbiome Enigma Al [44, 31]): the goal is to identify microbiome organisms, recorded as OTUs (Operational Taxonomic Units), that are associated with certain conditions (pH and Al). Each hypothesis (OTU) is equipped with a two-dimensional side information — the ubiquity, which is defined as the percentage of samples detected with the OTU, and the mean nonzero abundance, which is defined as the average abundance of each OTU across the samples in which it was detected.
  • fMRI datasets (Auditory and Imagination) [47]: the goal is to identify voxels that are activated in response to different stimuli. In the first dataset, a single individual received auditory stimulus and in the second dataset, a single individual was instructed to imagine playing tennis. Each hypothesis (voxel) is equipped with four dimensional side information: the spatial position of the voxel (in three dimensions) and a categorical variable with the Brodmann label [5], which is used to delineate different areas of the brain with different functions.
  • Two eQTL datasets (Subcutaneous and Omentum [11, 10]): the goal is to identify single nucleotide polymorphisms (SNPs) that are associated with gene expression levels. Such an association is referred to as an expressive quantitative trait loci (eQTL). Each hypothesis (SNP) is equipped with the following four side information variables: (1) the distance between the SNP and the gene trascription’s start site (TSS), (2) the log gene expression level, (3) the alternative allele frequency (AAF) of the SNP and (4) the chromatin state of the SNP.
  • Gene-drug response data (Estrogen [12]): the goal is to identify genes that are associated with a change in gene expression levels in response to a low-dosage estrogen treatment applied to breast cancer cells. Each hypothesis (gene) is assigned a ‘rank’ that was generated using an analysis of the gene expression levels in response to (1) a high-dosage estrogen treatment and (2) a medium-dosage estrogen treatment. Here, the ranks are in terms of the strength of association between the expression level of each gene and the level of dosage in (1) and (2). Thus, we can analyze the data twice, once for each side information variable based on (1) and (2).
Figure 12: Comparison of RESET and other methods using a collection of publically available datasets. Each method was evaluated at the FDR thresholds of 1%, 5%, 10%, and 20%. The number of discoveries reported by each RESET method is averaged over 10 applications at FDR threshold and dataset. Each color and shape combination corresponds to a unique method. For readability, each point is jittered in the horizontal direction.
The results of each method applied to the collection of the above datasets is visually summarized in Figure 12. We can clearly see that RESET Ensemble is comparable to the other methods in terms of discoveries. Indeed, to argue this more accurately, we ranked each method across each dataset and at each FDR threshold considered (ties were averaged). Then we calculated the median rank for each method. RESET Ensemble and AdaPTg-GMM were tied with the highest median ranks.
There are a couple of implementation details we would like to highlight here (for full details, see Supplementary Sections S.5-S.10). First, we used the same default options of RESET Ensemble on all datasets, except for the two fMRI datasets which exhibit a spike of p-values close to 1. In this case, it is sensible to define asymmetric target and decoy regions, $[0,0.3)$ and $(0.3,0.9]$ respectively as Chao and Fithian [8] did in their application of AdaPTg-GMM to this dataset. We also point out that we implemented AdaPTg-GMM and ZAP using the p-value information, and not the test statistic information (see Section 3.1 for background on these methods). To do so, AdaPTg-GMM allows the user to directly input p-values instead of z-statistics. For ZAP, we converted the p-value information into z-values by defining $z_{i}=\pm\Phi^{-1}(p_{i}/2)$ of each p-value, $p_{i}$, where the signs are chosen i.i.d uniformly. In other words, we input plausible z-values that correspond to the same p-value information used by the other tools. As we mentioned in Section 3, our implementation of RESET focuses on the use of p-values, and we leave the extension to test statistic information as part of future work.
To evaulate the computation times for each method, we used the two fMRI datasets, which uses the most number of side information variables and has a large number of hypotheses. Figure 13(A-B) shows the computation times for each considered method where the fastest are the collection of RESET methods, ZAP and AdaFDR. RESET NN, RESET GAM and RESET RF are slightly faster than our ensemble approach which considers all of them.
Figure 13: Comparison of runtime and power. In (A) and (B), we compare the computation time of each method on the fMRI datasets. We averaged the times of each RESET method over 10 applications. In (C) and (D), we compare the number of discoveries using the two eQTL datasets. The number of discoveries reported by each RESET method is averaged over 10 applications at FDR threshold and dataset. In all panels, we evaluated each method at FDR thresholds of: 1%, 5%, 10%, 20%.
Noticeably, we have excluded the analysis of the two eQTL studies from Figure 12. This is because this dataset exhibits a high degree of p-value dependency, even among the true null hypotheses and thus violates Assumption 3 which guarantees our type-1 error control. Indeed, this phenomenon is due to linkage disequilibrium in which SNPs behave dependently. As a consequence, the dataset exhibits virtually identical copies of the same p-value, with similar side information, even if the p-value is large suggesting a low confident hypothesis. Figure 13(C-D) shows that RESET Ensemble reports substantially more discoveries on these eQTL datasets than all the non-RESET variants. This is because the underlying random forest in the ensemble is able to identify the same ‘copies’ of low confident p-values in both the training and test sets, and if these copies are mutually positively labelled, which many are, then they will be rescored higher by our RESET algorithm and ultimately discovered. Evidently, RESET NN is more robust to this problem: the number of discoveries is in the same ‘ballpark’ as other methods. Note that RESET GAM shares identical results with RESET RF since it fails to perform a spline basis expansion on one of the side information variables (the chromatin state of the SNP), and so RESET GAM switches to RESET RF in this case.
6.3 FDP control with p-values
Figure 14: Comparison of RESET and other methods with FDP control using a collection of publically available data. Each panel plots the number of discoveries obtained by RESET Ensemble, FDP-SD and GR-SD at a range of FDR thresholds (1%, 5%, 10%, 20%) with varying confidence $1-\gamma$ (0.5, 0.8, 0.9).
In this final section, we demonstrate RESET’s application of FDP control in the p-value setting. Here we define the following asymmetric target and decoy regions, $[0,0.3)$ and $(0.3,0.9]$ respectively, for all datasets. The rationale stems from our later analysis in the discussion, as well as our comments at the end of Section 5.2.3, which point out the fact that when the ratio of estimating decoys to true null targets is small, the subsequent FDP-SD step in RESET might incur a loss of power. Hence, to rectify this, we essentially double the number of estimating decoys under consideration by defining a decoy region that is twice as large as the corresponding target region.
As far as we can tell, there are no p-value based testing procedures that use side information while controlling the FDP in the sense of Section 2. Hence, we compare RESET to the following generic methods for FDP control. The first of these methods is the previously considered FDP-SD that we analyzed in the competition setting, but adapted for the p-value setup. Each of the labels and scores $(L,W)$ required for FDP-SD are constructed in the same way as in RESET’s Step 1 (Section 4), where we used the same asymmetric target and decoy regions described above. The second method we compare to is the p-value based analogue called GR-SD [20] (see Section 3).
Figure 14 displays the number of discoveries obtained by each method on the collection of real datasets we considered in the previous section, with varying FDR thresholds (1%, 5%, 10%, 20%) and confidence levels (50%, 80%, 90%). RESET Ensemble is much more powerful than the alternative non-side information approaches with few exceptions.
7 Discussion
We introduced RESET, a flexible wrapper for any semi-supervised learning algorithm. Alongside, we provide an implementation of RESET, called RESET Ensemble, that (1) is power-wise competitive with recently developed methods for competition as well as p-value based multiple testing problems with side information, (2) is fast compared to most tools, and (3) is able to achieve finite FDR/FDP control.
In evaluating the claim of (1), we highlight the fact that competing tools such as the variants of AdaPT or Adaptive Knockoffs, require an initial selection of one of those variants. Making a selection after observing the discovery lists of each tool will clearly cheat any guarantee of FDR control. Thus, we compared RESET Ensemble to the default competing method if one was available, which was the case of Adaptive Knockoff’s GAM filter, or by the variants separately, which was the case of AdaPT. Based on our comparisons, we conclude that RESET Ensemble is able to achieve comparable or slightly greater power to many of these tools. RESET Ensemble’s intended generality alleviates the user from having to choose the “right” method to begin with.
In the case of (2), we found that our collection of RESET methods were a faster alternative to all competing methods with finite FDR control in our considered experiments. The only case in which RESET was surpassed in terms of speed was ZAP, which demonstrated marginally faster runtimes than RESET Ensemble in one of the fMRI datasets (Figure 13B). The most striking difference in runtimes was in the peptide detection context using mass spectrometry data, in which the variants of Adaptive Knockoffs were generally too slow to be used as tools for this application.
Finally in addressing (3), we point out that RESET’s compatibility with FDR or FDP control is unique. In the competition setting, we note that it would be challenging to combine Adaptive Knockoffs with FDP-SD because FDP-SD is a stepdown procedure while Adaptive Knockoffs is essentially a step-up procedure. Katsevich et al. [27] provide a general method for adaptive ordering using their method for simultaneous FDP bounds. However, (a) they do not provide an exact implementation of this, only a theoretical discussion, and (b) their approach is conservative since it controls the FDP at every cutoff (which allows the analyst to explore the nested discovery lists without breaking FDP control). To the best of our knowledge, in the competition framework with side information, RESET is the only method of reordering that may be flexibly used with FDP-SD without the additional cost of (b). Similairly, we know of no method that allows finite sample FDP control in the p-value setting with side information either.
On the topic of RESET’s FDP control, we point out the following limitation. When the number of discoveries is relatively large, RESET yields more power over the generic FDP-SD. However, when the number of discoveries is small, or the confidence is too high, RESET may yield less power, even if the side information is reasonably informative. To illustrate this, we computed the bounds, $\delta_{i}$ on the number of decoy wins in the top $i$ scores from Algorithm S2, that are used to determine the reported list of discoveries in FDP-SD (with $c=1/2$) and RESET (using the default $c=2/3$). We considered a confidence level of $1-\gamma=90\%$ and an FDP threshold of $\alpha=1\%$. Since we removed approximately half the decoys for training purposes in RESET, the bounds need to be doubled to fairly compare between the two approaches.
Refer to caption Refer to caption
Figure 15: RESET’s and FDP-SD’s bounds on the number of decoy wins. We recorded, on the left y-axis, the bounds used to determine the list of FDP-SD discoveries and twice the bounds used to determine RESET’s discoveries at a range of indices at $\alpha=1\%$ and $1-\gamma=90\%$. On the right y-axis, we plot the ratio of these bounds (in black). The left panel looks at indices between 1K to 2K, while the right panel looks at indices from 10K to 20K. An index corresponds to the number of top scoring hypotheeses (regardless of their labels).
Figure 15 shows the bounds for FDP-SD, and double the bounds for RESET, along with the ratio of the two at a range of different indices. We can see that for smaller indices, this ratio is much smaller, initially at 50% at an index of 1K, while at 10K this ratio is much higher at 94%. If the order of the hypotheses is the same, RESET’s smaller bounds imply it will report fewer discoveries because RESET employs FDP-SD to control the FDP by searching for the largest index $i$ s.t. $D_{i_{0}}\leq\delta_{i},\dots,D_{i}\leq\delta_{i}$, and reports the top $i$ scoring positively labelled hypotheses. In other words, if RESET is to report roughly the same number of discoveries as FDP-SD, RESET needs to rearrange the targets and estimating decoys successfully enough so as to make up for a 50% difference in the bounds when considering the top 1K scores, while only 6% when considering the top 10K. If the side information is not rich enough, it is possible RESET will report few discoveries than FDP-SD in these cases. We intend on investigating this as part of future work.
Figure 16: Variability in the number of discoveries in the Airway data and in Simulation 5 In (A) We used a 10% FDR threshold and varied the internal seed of RESET Ensemble over 100 applications. In (B) we used a 10% FDR threshold and compared the variability in RESET’s and AdaPT’s power. There are two outliers in RESET’s case (which occurred when the signal was considerably low).
Notably, RESET’s discovery list is variable since it randomly splits the decoys into two sets: one set for training and the other for estimating the false discoveries. This variability is demonstrated in the histogram of the number of discoveries in Figure 16A using 100 applications of RESET Ensemble to the Airway dataset at an FDR threshold of 10%. The resulting number of discoveries has a mean of 5998 with a standard deviation of about 85. While this is problematic to an extent, keep in mind that the data distribution is random, and the added variance from RESET may be marginal at times. Indeed, in Figure 16(B), we find that in the case of Simulation 5 from Section 6.1.1, RESET is overall arguably less variable than AdaPT’s even though AdaPT does not employ any internal randomization. In practice, users may get around RESET’s randomization by setting the RNG as a function of the input data and parameters. We plan on adding this feature to a future update.
Lastly, while RESET is a flexible method for the competition and p-value based testing, we wish to further extend it to the test-statistic setting. There is more than one way of generalizing RESET to accommodate a test statistic $z_{i}$ for each hypothesis $H_{i}$. Moreover there are several ‘types’ of interactions between the side information $\mathbf{x}_{i}$ and the test statistic $z_{i}$ which are not preserved when switching to p-values, as outlined by Leung and Sun [35]. Thus it is not clear that RESET Ensemble will be equally applicable in this setting without substantial modifications.
Appendix A Supplement
S.1 Algorithms
Algorithm S1 Selective SeqStep / SeqStep+ (adopted from Selective Sequential Step+ of [1])
1:
$\bullet$ $(L_{i},W_{i})_{i=1}^{m}$ the list of paired winning labels and scores;
$\bullet$ $c\in(0,1)$ the probability of a null target/feature win;
$\bullet$ $\alpha\in(0,1)$ the FDR threshold;
2:A discovery list $R_{\alpha}$
3:sort the paired $(L_{i},W_{i})$ in decreasing order of $W_{i}$triangleright ties are randomly broken
4:$D_{k}\leftarrow\#\{i\leq k:L_{i}=-1\}$triangleright number of decoy wins in top $k$ scores
5:$T_{k}\leftarrow\#\{i\leq k:L_{i}=1\}$triangleright number of target/feate wins in top $k$ scores
6:if SeqStep then
7:     $k_{0}\leftarrow\max\{k:\frac{D_{k}}{T_{k}\vee 1}\cdot\frac{c}{1-c}\leq\alpha\}$
8:else if SeqStep+ then
9:     $k_{0}\leftarrow\max\{k:\frac{D_{k}+1}{T_{k}\vee 1}\cdot\frac{c}{1-c}\leq\alpha\}$
10:end if
11:return $R_{\alpha}\leftarrow\{i:\text{the (pre-sorted) $W_{i}$ is in the top $k_{0}$ ranks and $L_{i}=1$}\}$
Algorithm S2 FDP-SD (adopted from [39])
1:
$\bullet$ $(L_{i},W_{i})_{i=1}^{m}$ the list of paired winning labels and scores;
$\bullet$ $c\in(0,1)$ the probability of a null target/feature win;
$\bullet$ $\alpha\in(0,1)$ the FDP threshold;
$\bullet$ $\gamma\in(0,1)$ for a $1-\gamma$ confidence level;
2:A discovery list $R_{\alpha,\gamma}$
3:sort the paired $(L_{i},W_{i})$ in decreasing order of $W_{i}$triangleright ties are randomly broken
4:$\lambda\leftarrow 1-c$triangleright the probability of a null decoy win
5:$R\leftarrow(1-\lambda)/(c+1-\lambda)$
6:$i_{0}\leftarrow\max\{1,\lceil(\lceil\log_{1-R}(\gamma)\rceil)/\alpha\rceil\}$
7:for $i=i_{0},\dots,m$ do
8:     $D_{i}\leftarrow\#\{j\leq i:L_{j}=-1\}$triangleright number of decoy wins in top $i$ scores
9:     $\delta_{i}\leftarrow\max\{d\in\{0,1,\dots,i\}:F_{B(\lfloor(i-d)\alpha+1+d,R\rfloor)}(d)\leq\gamma\}$triangleright $F_{B(n,p)}$ denotes the CDF of a Binomial$(n,p)$ RV
10:end for
11:$i\leftarrow i_{0}$ and $\delta_{i_{0}-1}\leftarrow-1$ and $\bar{\delta}_{i_{0}-1}\leftarrow 0$
12:while $i\leq m$ do
13:     $k_{0}\leftarrow\lfloor(i-\delta_{i})\cdot\alpha\rfloor+1$
14:     $k_{1}\leftarrow\lfloor(i-\delta_{i}+1)\cdot\alpha\rfloor+1$
15:     $p_{0}\leftarrow F_{B(k_{0}+\delta_{i},R)}(\delta_{i})$
16:     $p_{1}\leftarrow F_{B(k_{1}+\delta_{i}+1,R)}(\delta_{i}+1)$
17:     $w_{i}\leftarrow(p_{1}-\gamma)/(p_{1}-p_{0})$
18:     if $\bar{\delta}_{i-1}=\delta_{i}+1$ then
19:         $\bar{\delta}_{i}\leftarrow\bar{\delta}_{i-1}$
20:     else
21:         if $\delta_{i}>\delta_{i-1}$ then
22:              $w^{\prime}\leftarrow w_{i}$
23:         else
24:              $w^{\prime}\leftarrow w_{i}/w_{i-1}$
25:         end if
26:     end if
27:     Randomly set $\bar{\delta}_{i}\leftarrow\delta_{i}$ or $\bar{\delta}_{i}\leftarrow\delta_{i}+1$ with probabilities $w^{\prime}$ and $1-w^{\prime}$ respectively
28:     if $D_{i}\leq\bar{\delta}_{i}$ then
29:         $i\leftarrow i+1$
30:     else
31:         break
32:     end if
33:end while
34:if $D_{i_{0}}\leq\bar{\delta}_{i_{0}}$ then
35:     $k_{FDP}\leftarrow i-1$
36:else
37:     $k_{FDP}\leftarrow 0$
38:end if
39:$R_{\alpha,\gamma}\leftarrow\{i:\text{the (pre-sorted) $W_{i}$ is in the top $k_{FDP}$ ranks and $L_{i}=1$}\}$
Algorithm S3 GR-SD (adopted from [20])
1:
$\bullet$ $(p_{i})_{i=1}^{m}$ the list of p-values;
$\bullet$ $\alpha\in(0,1)$ the FDP threshold;
$\bullet$ $\gamma\in(0,1)$ for a $1-\gamma$ confidence level;
2:A discovery list $R_{\alpha,\gamma}$
3:sort the p-values $(p_{i})$ in ascending order $\triangleright$ ties are randomly broken
4:for $i=1,\dots,m$ do
5:     $k_{i}\leftarrow\lfloor\alpha i\rfloor+1$
6:     $\delta_{i}\leftarrow F^{-1}[k_{i},m-i+1](\gamma)$triangleright $F^{-1}[\alpha,\beta](\cdot)$ denotes the CDF of a Beta$(\alpha,\beta)$ RV
7:end for
8:$k_{GR}\leftarrow\max\{i:\prod_{j=1}^{i}1_{p_{i}\leq\delta_{i}}=1\text{ or }i=0\}$
9:$R_{\alpha,\gamma}\leftarrow\{i\in[m]:i\leq k_{GR}\}$
Algorithm S4 RESET
1:
$\bullet$ $\{(L_{i},W_{i},\mathbf{x}_{i})\,:\,i=1,\dots,m\}$ each hypothesis’ (label, winning score, side information);
$\bullet$ $\alpha$ FDR threshold for the discovery list;
$\bullet$ $s$ the probability of assigning a decoy to the training set (default: $s=1/2$);
$\bullet$ $f$ a semi-supervised machine learning model;
$\bullet$ $c_{0}$ an upper bound probability for a true null label $\mathbb{P}(L_{i}=1)\leq c_{0}$ ($i$ is a true null);
$\bullet$ isFDR - a boolean which determines whether FDR or FDP control is desired;
$\bullet$ $\gamma$ a confidence parameter if FDP control is desired;
2:A discovery list $R$
3:$I\leftarrow$ a subset of the decoy win indices, $\{i:L_{i}=-1\}$, determined by randomly including each one with probability $s$
4:$\tilde{L}_{i}\leftarrow-1$ for $i\in I$triangleright training decoys
5:$\tilde{L}_{i}\leftarrow 1$ for $i\in J:=I^{c}$triangleright pseudo targets
6:$(\tilde{W}_{i})_{i=1}^{m}\leftarrow f\left((W_{i},U_{i},\tilde{L}_{i})_{i=1}^{m}\right)$triangleright where $f$ is any machine learning model
7:if isFDR then
8:     $R\leftarrow\text{SSS+}((\tilde{W}_{i},L_{i})_{i\in J},c=\frac{c_{0}}{1-s\cdot(1-c_{0})},\alpha)$triangleright Selective SeqStep+
9:else
10:     $R\leftarrow\text{FDP-SD}((\tilde{W}_{i},L_{i})_{i\in J},c=\frac{c_{0}}{1-s\cdot(1-c_{0})},\alpha,\gamma)$triangleright FDP-stepdown
11:end if
12:return $R$
S.2 Proof of Lemma 1 in the main text
See 1
Proof.
Let $W$ be all the scores, $\mathbf{x}$ be all the side information, and $L_{-i}$ be all the other labels that are not $L_{i}$. Then for a true null p-value, $p_{i}$:
$\mathbb{P}(L_{i}=1\mid W,\mathbf{x},L_{-i})$ $=\frac{\mathbb{P}(p_{i}\in[0,a)\mid W,\mathbf{x},L_{-i})}{\mathbb{P}(p_{i}\in[0,a)\mid W,\mathbf{x},L_{-i})+\mathbb{P}(p_{i}\in(b_{1},b_{2}]\mid W,\mathbf{x},L_{-i})}$
$\leq\frac{\mathbb{P}(p_{i}\in[0,a)\mid W,\mathbf{x},L_{-i})}{\mathbb{P}(p_{i}\in[0,a)\mid W,\mathbf{x},L_{-i})+\mathbb{P}(p_{i}\in[0,a)\mid W,\mathbf{x},L_{-i})\cdot\frac{b_{2}-b_{1}}{a}}$
$=\frac{a}{a+b_{2}-b_{1}},$
where the first line follows since the denominator sums to 1, and the second line clearly follows from the non-decreasing property. For a more formal argument for the second line, note that the non-decreasing property implies that:
$\mathbb{P}(p_{i}\in B)\leq\frac{a}{b_{2}-b_{1}}\cdot\mathbb{P}(p_{i}\in q(B)),$
where $B\in[0,a)$, $p_{i}$ is a true null p-value and $q$ denotes the inverse transformation of the mirrored p-values that map p-values in $(b_{1},b_{2}]$ onto $[0,a)$, that is $q(p):=b_{2}-\frac{b_{2}-b_{1}}{a}\cdot p$. Then consider for every measurable-$A$:
[…]
where $B$ is the corresponding region such that $\{W_{i}\in A\}=\{p_{i}\in B\}\cup\{p_{i}\in q(B)\}$. It follows that:
$\mathbb{P}(p_{i}\in[0,a)\mid W,\mathbf{x},L_{-i})$ $=\mathbb{P}(p_{i}\in[0,a)\mid W_{i})$
$\leq\frac{a}{b_{2}-b_{1}}\mathbb{P}(p_{i}\in(b_{1},b_{2}]\mid W_{i})=\frac{a}{b_{2}-b_{1}}\mathbb{P}(p_{i}\in(b_{1},b_{2}]\mid W,\mathbf{x},L_{-i})$
S.3 RESET controls the FDR or FDP
See 3
In this section we prove that given Assumption 3 RESET controls the FDR or FDP. We use the target-decoy terminology to refer to a positive label $L_{i}=1$ as a target win or simply a target, and a negative label $L_{i}=-1$ as a decoy win or decoy.
Lemma S1.
Let $\tilde{L}_{i}$ for $i\in[p]$ denote the training labels, where $\tilde{L}_{i}=-1$ for a training decoy and $\tilde{L}_{i}=1$ for a pseudo target. Let $\mathcal{F}$ be the $sigma$-algebra generated by the winning scores $W$, the side information $\mathbf{x}$, and the labels of the false nulls. Let $N$ be the indices of the true nulls, and $L_{-i},\tilde{L}_{-i}$ be the corresponding labels without the $i$th one. Then we have for $i\in N$:
$L_{i}$ =$=\begin{cases}+1,&\mathbb{P}(L_{i}=+1)\leq\frac{c_{0}}{1-s\cdot(1-c_{0})}\\ -1,&\mathbb{P}(L_{i}=-1)\geq\frac{(1-c_{0})\cdot(1-s)}{1-s\cdot(1-c_{0})}\\ \end{cases},$
independently of $\{\tilde{L}_{i}=1\},\mathcal{F},\tilde{L}_{-i},L_{-i}$.
Proof.
By Assumption 3, $\{L_{i}\,:\,i\in N\}$ are i.i.d with $P(L_{i}=1)\leq c_{0}$, independently of $\sigma(\mathcal{F},\tilde{L}_{-i},L_{-i})$. In addition, since each decoy is randomly assigned to the training decoy set independently of anything else with probability $s$, for each $i\in N$ we have:
$\mathbb{P}(L_{i}=-1,\tilde{L}_{i}=1\mid\mathcal{F},\tilde{L}_{-i},L_{-i})$ $=\mathbb{P}(L_{i}=-1\mid\mathcal{F},\tilde{L}_{-i},L_{-i})\cdot\mathbb{P}(\tilde{L}_{i}=1\mid L_{i}=-1,\mathcal{F},\tilde{L}_{-i},L_{-i})$
$\geq(1-c_{0})\cdot(1-s),$
and
$\mathbb{P}(L_{i}=1,\tilde{L}_{i}=1\mid\mathcal{F},\tilde{L}_{-i},L_{-i})$ $=\mathbb{P}(L_{i}=1\mid\mathcal{F},\tilde{L}_{-i},L_{-i})\cdot\mathbb{P}(\tilde{L}_{i}=1\mid L_{i}=1,\mathcal{F},\tilde{L}_{-i},L_{-i})$
$\leq c_{0}\cdot 1,$
Hence,
$\mathbb{P}(L_{i}=-1\mid\tilde{L}_{i}=1,\mathcal{F},\tilde{L}_{-i},L_{-i})$ $=\frac{\mathbb{P}(L_{i}=-1,\tilde{L}_{i}=1\mid\mathcal{F},\tilde{L}_{-i},L_{-i})}{P(\tilde{L}_{i}=1\mid\mathcal{F},\tilde{L}_{-i},L_{-i})}$
$\geq\frac{(1-c_{0})\cdot(1-s)}{(1-s)\cdot\mathbb{P}(L_{i}=-1\mid\mathcal{F},\tilde{L}_{-i},L_{-i})+1\cdot\mathbb{P}(L_{i}=1\mid\mathcal{F},\tilde{L}_{-i},L_{-i})}$
$=\frac{(1-c_{0})\cdot(1-s)}{1-s\cdot\mathbb{P}(L_{i}=-1\mid\mathcal{F},\tilde{L}_{-i},L_{-i})}$
$\geq\frac{(1-c_{0})\cdot(1-s)}{1-s\cdot(1-c_{0})}$
The proof is complete noting that $\frac{c_{0}}{1-s\cdot(1-c_{0})}+\frac{(1-c_{0})\cdot(1-s)}{1-s\cdot(1-c_{0})}=1$ and $L_{i}\in\{\pm 1\}$ a.s.. ∎
Remark 1.
Lemma S1 essentially states that the labels $L_{i}$ for $i\in N$ with $\tilde{L}_{i}=1$ are i.i.d random variables with $\mathbb{P}(L_{i}=+1)\leq\frac{c_{0}}{1-s\cdot(1-c_{0})}$ independently of all other information: the scores $W$, the side information $\mathbf{x}$, the labels of the false null discoveries, and all the labels $\tilde{L}$.
See 1
Proof.
We rely on Theorem 3 of Barber and Candès [1] applied to the set of pseudo-targets ($\tilde{L}_{i}=1$) ordered in decreasing order of the learned scores $\tilde{W}_{i}$. Specifically, we assign to each pseudo-target a one-bit p-value:
$\tilde{p}_{i}=\begin{cases}\frac{c_{0}}{1-s\cdot(1-c_{0})}&L_{i}=+1\\ 1&L_{i}=-1\end{cases}.$
Recall that the scores $\tilde{W}$ are themselves a function of the values $(\tilde{L},W,x)$ and clearly the one bit p-values are a function of the labels $L_{i}$. Then it follows from the last Lemma that given the scores $\tilde{W}$ and the one-bit p-values of the false nulls, the one-bit p-values of the true nulls hypotheses ($i\in N$ with $\tilde{L}_{i}=1$) are i.i.d with $\mathbb{P}(\tilde{p}_{i}\leq\frac{c_{0}}{1-s\cdot(1-c_{0})})=\mathbb{P}(L_{i}=1)\leq\frac{c_{0}}{1-s\cdot(1-c_{0})}$, i.e., they stochastically dominate the uniform distribution and hence are valid p-values. Therefore, the condition of Theorem 3 holds so applying, as RESET does, Selective SeqStep+ with $c=\frac{c_{0}}{1-s\cdot(1-c_{0})}$ to these p-values controls the FDR at level $\alpha$. Note that the original formulation of Selective SeqStep+ is equivalent to Algorithm S1 since we identify the event $\tilde{p}_{i}\leq\frac{c_{0}}{1-s\cdot(1-c_{0})}$ with $L_{i}=1$ and $\tilde{p}_{i}=1$ with $L_{i}=-1$. ∎
See 2
Proof.
The proof of Theorem 1 shows that conditional on the learned scores $\tilde{W}$ and the false null labels the labels $L_{i}$ for $i\in N$ with $\tilde{L}_{i}=1$ are distributed as i.i.d ±$\pm 1$ RVs, with $\mathbb{P}(L_{i}=1)\leq\frac{c_{0}}{1-s\cdot(1-c_{0})}$. Assuming for a moment that $\mathbb{P}(L_{i}=1)=\frac{c_{0}}{1-s\cdot(1-c_{0})}$, then RESET satisfies the assumption of the multiple decoy version of FDP-SD with $c=\lambda=\frac{c_{0}}{1-s\cdot(1-c_{0})}$ and therefore it controls the FDP at $\alpha$ with confidence $1-\gamma$ (Section 4 of [39]). In the case that $\mathbb{P}(L_{i}=1)<\frac{c_{0}}{1-s\cdot(1-c_{0})}$, then the procedure is conservative with $c=\lambda=\frac{c_{0}}{1-s\cdot(1-c_{0})}$ and the proof in [39] follows without any modification. ∎
S.4 Further implementation details of RESET
We point out some additional minor details regarding our implementation of RESET. We use the randomxForest package, nnet and mgcv to implement the random forest, two-layer neural network and generative additive model in R, respectively. We used the default parameters for the random forest, the default parameters for the two-layer neural network and the default parameters for the generative additive model (except we use drop.intercept = TRUE). If the number of side information variables used by RESET is $\leq 3$, a smoothing spline is fitted with mgcv::s otherwise a natural cubic spline is fitted using 5 degrees of freedom with splines::ns on each side information variable separately (for computational efficiency). For the use of GAM in RESET, we used tryCatch to switch to a random forest in the rare case that the GAM was not able to complete. Some real data sets contain p-values that are ‘zero’. Hence when we calculate $W_{i}=\lvert\Phi^{-1}(p_{i})\rvert$ for such p-values, we obtain Inf in R. We replace instances of Inf for the maximum real-valued score in the dataset. Lastly we used the default settings of RESET for all numerical and real data experiments unless otherwise stated in the main text.
S.5 Further implementation details of Adaptive Knockoffs
There were some small differences between the implementations in the adaptiveKnockoff package in R and the code used in the original manuscript. Accordingly, we used the updated package version. We also fixed a minor bug, which pruned the candidate set $\mathcal{S}$ before checking the estimated FDR is $\leq\alpha$ — possibly missing out on a marginally larger discovery list. There are several parameters that may be set before applying the Adaptive Knockoff filters, e.g. the initial proportion of hypotheses revealed, reveal_prop. In Simulations 1-3, we used the same parameters set by Ren and Candès from their paper. In Simulation 4, we used the same parameters in Simulation 3.
In the application to peptide detection, the primary scores $W$ may be negative. Since Adaptive Knockoffs encodes the labels $L$ by using the sign of the scores $W$, we shifted our peptide scores so that the minimum score was zero. For AdaKO EM, any peptide score $\leq 10^{-3}$ was subsequently assigned a zero score. Finally we used, reveal_prop = 10% which reveals the hypotheses that are less than or equal to the 10% of non-zero scoring hypotheses (which is the same setting used by the simulations). All other parameters were default.
S.6 Further implementation details of the AdaPT methods
We implemented AdaPT using the adaptMT package in R. In Simulation 5 and 6, we used the same settings considered by the authors [33], i.e., we looked at AdaPT GAM in Simulation 5 and AdaPT GLMnet in Simulation 6 with the same parameters. Moreover, we considered AdaPT GLMnet in Simulation 5 with the same settings from Simulation 6, demonstrating the importance of selecting the correct working model. In the real datasets looked at by Zhang et al. [50], we used AdaPT GAM to fit a natural cubic spline using splines::ns with 5 degrees of freedom (knots chosen by default) on each side information variable. The use of natural cubic splines is a suggested option from the package documentation. There was an exception to this for the GTEx datasets, which flagged an error when we tried to apply the spline basis transformation on the ‘chromatin state of the SNP’. In this case, we applied no transformation to the affected side information variable. Zhang et al. mention that they employed a ‘5-degree for each dimension’, but it is not clear to us if they mean the degree of the piece-wise polynomials that make the spline, or the degrees of freedom of the spline. In the latter case, it is not clear how the knots are chosen, or how they overcame the error associated with the ‘chromatin state of the SNP’. Regardless, our results using AdaPT appear to be essentially the same. Moreover, Zhang et al. omits the application of AdaPT to the two fMRI datasets on account of the categorical side information variable used in these datasets. However, Chao and Fithian appear to still apply AdaPTg without problem, and so analogously we applied AdaPT as well. Lastly, in the gene-drug response data, we followed the same implementation by Lei and Fithian [33], which used AdaPT GLM and a collection of candidate side information transformations using splines::ns with degrees of freedom ranging from 6 to 10.
S.7 Further implementation details of the AdaPTg methods
We implemented AdaPTg using the repository https://github.com/patrickrchao/adaptMT which has all the same implementations as AdaPT while offering the ability to define assymetric regions for mirroring the p-values. Accordingly, we followed the same implementation details as the previous section used by AdaPT and set the so-called ‘masking parameters’ as default.
S.8 Further implementation details of AdaPT-GMMg
We implemented AdaPT-GMMg from the repository https://github.com/patrickrchao/AdaPTGMM. In Simulations 5, we used the GAM filter (model_type = ‘mgcv’) to fit a smoothing spline using mgcv::s on the side information for the M-Step of the EM algorithm. In Simulation 6, we used a generalized linear model filter with regularization (model_type = ‘glmnet’) applied directly to the side information variables. For the real datasets used by Zhang et al. [50], we used the exact same implementation as Chao et al. from the file https://github.com/patrickrchao/AdaPTGMM_Experiments/blob/main/AdaFDR_experiments/run_all_exp.R. In the gene-drug response data, since only one-dimensional side information is used, we copied the parameters selected from the previous real datasets that also considered just one-dimensional side information.
S.9 Further implementation details of ZAP
We used the default parameters of zap_asymp from the ZAP package https://github.com/dmhleung/zap. In all simulations and real data experiments, we considered a natural cubic spline basis expansion on each side information variable with six degrees of freedom using splines::ns. This choice was based on their own analysis on two of the real data sets that we consider here, Airway and Bottomly. The exception was in Simulation 6, where we directly used all 100 side information variables rather than applying a spline basis expansion on each of them. In Simulation 5 and 6, we used the test statistics obtained as input, while in the real data experiments we converted the p-values to z-values by using the transformation ±$\pm\Phi^{-1}(p_{i}/2)$ on each p-value $p_{i}$ where the signs are chosen i.i.d uniformly.
S.10 Further implementation details of AdaFDR
We implemented AdaFDR using the method.adafdr_test function from the AdaFDR package https://github.com/martinjzhang/adafdr. We note that installation of the package required setting up a separate conda environment so that we could install an older version of Python (v3.6.13). This was because recent versions of Python do not support PyTorch v1.4.0, a requirement of the AdaFDR package setup.py file.
We used the exact same parameters in the vignettes https://github.com/martinjzhang/AdaFDRpaper/tree/master/vignettes to implement AdaFDR on the 10 datasets that Zhang et al. look at. In the gene-drug response data, we used AdaFDR with fast_mode = False. This is in accordance to the discussion from their paper that the fast version of AdaFDR is recommended when there are few discoveries to be expected or if the number of hypotheses are small (which is not the case for the gene-drug response data).
S.11 Computer specifications for computation times
Computation times were calculated using an M1 Mac Studio with 20-core CPU and 128GB RAM.
S.12 Data preparation and searching with HEK293 data
We downloaded the human proteome (UP000005640, downloaded on 2023/09/23 from UniProt) and prepared a target peptide database along with 5 randomly shuffled decoy peptide databases using Tide-index within the Crux Toolkit v4.1.6809338 [42, 29] with all options set to default. An output containing the target and decoy peptide pairs are conveniently provided using the --peptide-list T option in Tide-index. Each of the RAW 24 HEK293 spectrum files [9] were converted to mzML format using MSConvert 3.0.22314 with the vendor peak-picking filter using the default settings. For each of the 5 decoy databases, we searched each spectrum from the combined 24 mzML spectrum files against the combined target-decoy peptide database using Tide-search [15], using the options --top-match 1 --auto-precursor-window warn --auto-mz-bin-width warn --concat T. The resulting search files were then converted to so-called pin files using the make-pin function in Crux. Each row in the pin files correspond to the optimal peptide-spectrum match (PSM) for each spectrum, the primary score for quantifying this match, called XCorr scores, and a collection of auxilliary information regarding the PSM.
Side Information Description
deltCn The difference between the XCorr score of the two top ranked PSMs with respect to the combined database, divided by maximum of the top PSM XCorr score and 1
PepLen The length of the matched peptide, in residues
Charge The charge of the precursor ion (ranging from +1 to + 5)
lnNumSP The natural logarithm of the number of database peptides within the specified precursor range
dm The difference between the calculated and observed mass
absdM The absolute value of the difference between the calculated and observed mass
Table S4: List of described side information used by HEK293 data. The list of side information used by RESET and Adaptive Knockoffs in the HEK293 data, adapted from https://crux.ms/file-formats/features.html. Note that the Charge state is represented as a one-hot vector, where we left out a Charge state of +1 (due to linearity, since the one-hot vector sums to one).
Search algorithms like the one above are typically ‘spectrum-centric’, meaning that for each spectrum, we look for the best matching database peptide, as oppose to the other way round. Consequently, the same database peptide may be matched multiple times with different scores, or have no match at all. Hence for each target peptide, we define $Z_{i}$ as the maximum XCorr score of all the PSMs associated to that target peptide. Similarly for each decoy peptide, we define $\tilde{Z}_{i}$ as the maximum XCorr score of all the PSMs associated to that decoy peptide. For target or decoy peptides that do not appear in the pin file, they get a score of $\infty$- ∞. Then we compete each target and its corresponding paired decoy by recording the winning score $W_{i}=Z_{i}\vee\tilde{Z}_{i}$ along with the label $L_{i}=sign(Z_{i}-\tilde{Z}_{i})$, randomly breaking any ties. Peptides with a winning score of $\infty$- ∞, indicating that neither the target or its paired decoy are in the pin file, are thrown out. We assign each winning peptide the side information defined as the auxiliary information $x_{i}$ of the underlying PSM that had the XCorr score of $W_{i}$. The side information that was subsequently used by RESET and Adaptive Knockoffs is given in Table S4.
S.13 Estimating the computation times of Adaptive Knockoffs
Running Adaptive Knockoffs on all of the peptide datasets would be infeasible. Therefore, we resorted to estimating its runtimes by considering the time it took to complete a single iteration that determines which hypothesis to reveal next, and multiplying this value by the estimated number of iterations before the procedure stops. Since the time it takes a single iteration depends on the number of hypotheses that were already revealed, we took different such measurements as we varied this number. Specifically, we revealed the labels that were less than or equal to the 10% quantile of non-zero scoring hypotheses, up to a value of $10k$%, in 10% increments. We chose $k$ not to be too large, since hypothetically Adaptive Knockoffs might terminate at the 1% FDR level before it reaches $10k$of the revealed hypotheses. Of course, we are unable to compute when Adaptive Knockoffs might terminate, so we estimated this using the value $r$, the number of target and decoy peptides that scored above the 1% FDR cutoff using RESET ensemble. Then we defined $k=\lfloor(m-r)/m\cdot 10\rfloor$, where $m$ is the total number of hypotheses (peptides) considered by Adaptive Knockoffs. With $\bar{t}_{k}$ being the average of the $k$ measured single-iteration times, we estimated the runtime by $\bar{t}_{k}\cdot(m^{\prime}-r)$, where $m^{\prime}$ is the number of labels that are yet to be revealed (that are above the 10% quantile of non-zero scoring hypotheses) and $m^{\prime}-r$ is our estimate of the number of iterations until Adaptive Knockoffs stop. In Section 5.2.2, we considered 10 repeated applications of RESET ensemble, in which case $r$ is the average of the number of winning target and decoy peptides above the cutoff over each application.
S.14 Data preparation and searching with PRIDE-20 data
We followed the preparation and searching outlined in [18]. The only difference is that we used an updated version of the Crux toolkit. We provide the following brief description for the reader’s convenience. Twenty spectrum files and protein databases from separate projects, referred to as the PRIDE-20 dataset, were downloaded. Tide-index was used to prepare the target peptide database along with 10 randomly shuffled decoy peptide databases for each of the 20 spectrum files. Tide-search was used to search each spectrum file against the 10 combined target-decoy databases using two modes: ‘narrow’ and ‘open’ search options as they outline. Briefly a ‘narrow’ search, which is the default search mode, only searches each spectrum in the spectrum file against peptides in the target-decoy database that have a theoretical mass within a small tolerance of the sample peptide’s mass that generated the experimental spectrum. On the other hand, an ‘open’ search makes this mass-tolerance larger so that a spectrum may match with a peptide with a completely different mass. This is desirable at times, since sample peptides can undergo post-translational modifications which modify the mass of the peptide and therefore the experimental spectrum. Hence when searching in an ‘open’ mode, the modified spectrum may still be correctly matched. The search files were then converted to pin files using the make-pin function in Crux and subsequently filtered for the top 1 (the optimal) peptide-spectrum match for each spectrum.
For each spectrum file and search type, we obtained the triples $(L,W,\mathbf{x})$ in the following way. For each spectrum file that had no modifications (we explain this in more detail next), we proceeded essentially in the same way as in Section S.12. That is, for each peptide in the pin file, we define $Z_{i}$ (if a target peptide) or $\tilde{Z_{i}}$ (if a decoy peptide) as the maximum Tailor score of all the PSMs associated to the target peptide [46]. Here we are using the more sensitive Tailor scores and relegate the XCorr scores as one of the side information variables. Then we compete each target and its corresponding paired decoy in the same way as Section S.12, to obtain the triple $(L_{i},W_{i},\mathbf{x}_{i})$ where the side information $\mathbf{x}_{i}$ is obtained from the auxiliary information of the underlying PSM that had the Tailor score of $W_{i}$. A description of the side information used is given in Table S5.
Side Information Description
deltLCn The difference between the XCorr score of the top scoring PSM and the fifth/last ranked PSM with respect to the combined database, divided by the maximum of the top PSM’s XCorr score and 1
deltCn The difference between the XCorr score of the two top ranked PSMs with respect to the combined database, divided by maximum of the top PSM XCorr score and 1
Xcorr The SEQUEST cross-correlation PSM score
PepLen The length of the matched peptide, in residues
Charge The charge of the precursor ion (ranging from +1 to + 5)
lnNumSP The natural logarithm of the number of database peptides within the specified precursor range
dm The difference between the calculated and observed mass
absdM The absolute value of the difference between the calculated and observed mass
Table S5: List of described side information used by PRIDE-20 data. The list of side information used by RESET and Adaptive Knockoffs in the PRIDE-20 data, adapted from https://crux.ms/file-formats/features.html. Note that the Charge state is represented as a one-hot vector, where we left out a Charge state of +1 (due to linearity, since the one-hot vector sums to one). XCorr is used as side information since we define $W$ in terms of Tailor scores here instead.
We next consider spectrum files that were searched with variable modifications (---auto-modifications T). Here we need to make a slight adjustment to account for some dependency within the data. Specifically, using variable modifications creates several ‘copies’ of the peptides in the target-decoy database that are distinguished only by slight alterations to the mass of some of the amino acids. As an example, PEPTIDE may generate the following ‘copies’: PE[16]PTIDE, PEPTIDE[16] and PE[16]PTIDE[16], where [16] indicates an increase of 16 Daltons to the amino acid on the left. Consequently, each of these peptides are usually correlated in the data — i.e. if one of these peptides is scored high, than usually they all are. Hence in order to satisfy Assumption 3, we follow the protocol outlined in [18]. That is, we identify all the copies as being the same peptide, ‘PEPTIDE’, with each of the variable modifications ignored. Then we proceed with the same steps as the above paragraph, by determining the maximum score associated with PEPTIDE and recording it as $Z_{i}$ (or $\tilde{Z}_{i}$ if it is a decoy). We define the labels, winning scores and side information in the same way to obtain our triplets $(L,W,\mathbf{x}_{i})$.
Due to runtime considerations, we used 13 of the smallest narrow search files to assess Adaptive Knockoff’s power as described in Section 5.2.2 of the manuscript, and only one one of the 10 target-decoy databases. To fairly compare with Adaptive Knockoff, we used the same target-decoy databases for RESET ensemble. All 20 spectrum files and all 10 target-decoy databases for each spectrum file were used in the comparison of RESET ensemble and FDP-SD while controlling the FDP.
S.15 Brief justification of the assumptions used with HEK293 and PRIDE-20 datasets
A standard assumption of the mass spectrometry literature, which is also empirically validated, is that the label $L_{i}$, indicating whether an incorrect target or its decoy scored higher, is uniformly ±$\pm 1$ independently of all other target-decoy pairs and scores [21, 38]. RESET additionally requires this statement to be true if we further condition on the side information. This is obviously true for some side information variables which are constant between the target and corresponding decoys — e.g. PepLen, Charge, lnNumSP, dm, absdM. The other side information variables are score-related — e.g. Xcorr, deltCn and deltLCn, and so conditioning on them assumes no more than the “standard assumption”. Hence, we believe Assumption 3 with $P(L_{i}=1)=c_{0}=1/2$ to be reasonably satisfied.
S.16 Description and Data preparation using data from Zhang et al.
Each of the 10 datasets from Zhang et al. were downloaded from https://github.com/patrickrchao/AdaPTGMM_Experiments/tree/main/AdaFDR_experiments/data_files, containing the p-value and side information pairs, $(p,\mathbf{x})$. All pairs were then used to report a list of discoveries by each p-value based method with the following exceptions. In the RNA-seq datasets, Airway, Bottomly and Pasilla, there exists ‘spikes’ of high p-values. These ‘spikes’ concentrate in certain regions of the side information, i.e. when the log-normalized gene counts are low. Hence we removed those p-values with log-normalized gene counts that were less than zero. Figure S1 show the histograms of the p-values before and after this removal.
Figure S1: Histograms of p-values between in the RNA-seq data. The first row of panels plot the histograms before the removal of some of the p-values while the second row of panels plot the histograms after this removal.
S.17 Tables
Table S6: The number of discoveries and computations times for 13 of the PRIDE-20 dataset at the 1% FDR level. We calculated the number of discoveries at the 1% FDR using each method of Adaptive Knockoffs and TDC. We also computed computation times for Adaptive Knockoffs (in minutes). We used both the actual times and the estimated times according to Section S.13.
Project ID Method Estimated Time Actual Time Discoveries
PXD002470 AdaKO EM 0.58 1.17 300
AdaKO GAM 0.03 0.08 272
AdaKO LR 0.07 0.14 483
AdaKO RF 7.50 9.04 488
TDC - - 455
PXD006856 AdaKO EM 24.72 59.80 0
AdaKO GAM 0.80 2.17 379
AdaKO LR 3.67 4.63 428
AdaKO RF 427.01 486.70 421
TDC - - 425
PXD008920 AdaKO EM 33.81 48.63 8485
AdaKO GAM 0.95 2.39 8443
AdaKO LR 2.98 6.11 9143
AdaKO RF 250.66 270.67 9352
TDC - - 9086
PXD008996 AdaKO EM 119.73 188.54 7464
AdaKO GAM 3.39 7.55 7958
AdaKO LR 18.77 20.18 8481
AdaKO RF 1430.16 1626.96 8625
TDC - - 8413
PXD010504 AdaKO EM 134.07 185.59 459
AdaKO GAM 4.03 5.11 1358
AdaKO LR 9.25 11.37 1585
AdaKO RF 2134.38 2420.31 1626
TDC - - 1551
PXD012528 AdaKO EM 202.99 241.66 0
AdaKO GAM 6.37 8.34 239
AdaKO LR 13.57 16.34 314
AdaKO RF 3769.54 4143.21 405
TDC - - 342
PXD012611 AdaKO EM 4.79 7.59 2850
AdaKO GAM 0.15 0.26 3152
AdaKO LR 0.62 0.98 3250
AdaKO RF 43.53 47.01 3262
TDC - - 3260
PXD013274 AdaKO EM 34.65 49.52 14025
AdaKO GAM 0.81 1.70 14271
AdaKO LR 4.59 3.92 14791
AdaKO RF 205.10 248.82 14765
TDC - - 14769.0
PXD016724 AdaKO EM 499.32 598.31 3617
AdaKO GAM 13.76 24.01 3802
AdaKO LR 36.84 58.90 4536
AdaKO RF 8070.09 8882.65 4809
TDC - - 4745
PXD019186 AdaKO EM 247.15 389.51 0
AdaKO GAM 7.04 13.49 1020
AdaKO LR 34.77 30.35 942
AdaKO RF 4143.83 4427.20 1217
TDC - - 912
PXD022257 AdaKO EM 151.82 244.22 1619
AdaKO GAM 15.50 9.94 1918
AdaKO LR 12.34 23.81 2500
AdaKO RF 2475.32 2638.35 2501
TDC - - 2509
PXD023571 AdaKO EM 30.23 43.86 1030
AdaKO GAM 3.04 4.40 1743
AdaKO LR 2.92 10.34 0
AdaKO RF 437.75 474.17 2161
TDC - - 1505
PXD026895 AdaKO EM 62.89 85.62 3255
AdaKO GAM 8.94 4.94 3676
AdaKO LR 10.02 12.04 4044
AdaKO RF 847.23 835.02 4047
TDC - - 3980.0
S.18 Figures
Figure S2: Estimated FDR of each method in numerical simulations. Each panel plots the estimated FDR for each method at FDR thresholds ranging from 5% to 30%. The first row corresponds to Simulation 1 with three values of $k\in\{50,150,300\}$, the second row corresponds to Simulation 2 with three values of $k\in\{150,300,450\}$, and the last row corresponds to Simulation 3 and 4. For readability, the points are jittered in the horizontal direction.
Figure S3: Estimated power of each method in numerical simulations. Each panel plots the power for each method at FDR thresholds ranging from 5% to 30%. The first row corresponds to Simulation 1 with three values of $k\in\{50,150,300\}$, the second row corresponds to Simulation 2 with three values of $k\in\{150,300,450\}$, and the last row corresponds to Simulation 3 and 4. For readability, the points are jittered in the horizontal direction. A description of AdaKO GAM and KO can be found in Section 2.

### v2

@@
-A flexible and general semi-supervised approach to multiple hypothesis testing
+A semi-supervised framework for diverse multiple hypothesis testing scenarios
 Abstract
@@
 Recently there has been a growing interest in enhancing such procedures by combining additional information with the primary p-value or score.
-Specifically, such so-called “side information” can be leveraged to improve the separation between true and false nulls along additional “dimensions” thereby increasing the overall sensitivity.
-In line with this idea, we develop RESET (REScoring via Estimating and Training) which uses a unique data-splitting protocol that subsequently allows any semi-supervised learning approach to factor in the available side-information while maintaining finite-sample error rate control.
-Our practical implementation, RESET Ensemble, selects from an ensemble of classification algorithms so that it is compatible to a range of multiple testing scenarios without the need for the user to select the appropriate one.
-We apply RESET to both p-value and competition based multiple testing problems and show that RESET is (1) power-wise competitive, (2) fast compared to most tools and (3) is able to uniquely achieve finite sample FDR or FDP control, depending on the user’s preference.
-* Corresponding authors: J.Freestone@maths.usyd.edu.au, uri.keich@sydney.edu.au
+Specifically, such so-called side information can be leveraged to improve the separation between true and false nulls along additional dimensions thereby increasing the overall sensitivity.
+In line with this idea, we develop RESET (REScoring via Estimating and Training), which uses a unique data-splitting protocol that subsequently allows any semi-supervised learning approach to factor in the available side information while maintaining finite-sample error rate control.
+Our practical implementation, RESET Ensemble, selects from an ensemble of classification algorithms so that it is compatible with a range of multiple testing scenarios without the need for the user to select the appropriate one.
+We apply RESET to both p-value and competition based multiple testing problems and show that RESET is (1) power-wise competitive, (2) fast compared to most tools and (3) able to uniquely achieve finite sample false discovery rate or false discovery proportion control, depending on the user’s preference.
+Corresponding authors: J.Freestone@maths.usyd.edu.au, uri.keich@sydney.edu.au
 1 Introduction
@@
 As an example, a scientist may be interested in assessing which genes exhibit a change in gene expression levels in response to a particular drug treatment.
-In this case, the $i$th null hypothesis $H_{i}$ states that “the $i$th gene exhibits no change in gene expression level”.
+In this case, the $i$th null hypothesis $H_{i}$ states that ‘the $i$th gene exhibits no change in gene expression level’.
 This type of large-scale hypothesis testing is usually achieved by assigning a p-value, $p_{i}$, to each null hypothesis.
@@
 Given a confidence parameter $\gamma$ and a threshold $\alpha$, the FDP is said to be controlled if $\mathbb{P}(FDP>\alpha)\leq\gamma$.
-Evidently, such FDP control reduces the risk that the realized FDP exceeds $\alpha$ which is not guaranteed by FDR control [39, 27, 19, 37, 16].
-Determining a p-value for each null hypothesis can be challenging because (a) it requires knowledge of the underlying null distribution, and (b) it can be computationally intensive.
-A recently developed competition framework offers an alternative approach to FDP/FDR control that does not require computation of canonical p-values.
+Evidently, such FDP control reduces the risk that the realized FDP exceeds $\alpha$, which is not guaranteed by FDR control [40, 27, 19, 38, 16, 20].
+Determining a p-value for each null hypothesis can be challenging because (a) it requires knowledge of the underlying null distribution and (b) it can be computationally intensive.
+A recently developed competition framework offers an alternative approach to FDP/FDR control that does not require computation of canonical p-values [1, 21].
 Instead, all that these procedures require is a single draw, $\tilde{Z}_{i}$, from each null distribution.
 The null drawn $\tilde{Z}_{i}$ is then compared with $Z_{i}$, the originally observed score (test statistic) for $H_{i}$, thus conceptually providing a crude 1-bit p-value.
-Those p-values are too crude to be used in typical p-value based filters.
+These p-values are too crude to be used in typical p-value based filters.
 Instead, assuming that larger values of $Z_{i}$ provide more evidence against $H_{i}$, we compete each pair of scores $(Z_{i},\tilde{Z}_{i})$ by recording a label $L_{i}\in\{\pm 1\}$ indicating which was the largest of the two scores, as well as a winning score, $W_{i}=f(Z_{i},\tilde{Z}_{i})$.
-The function $f$ needs to be symmetric in its arguments and should be large if $Z_{i}$ is large, e.g., $W_{i}=Z_{i}\vee\tilde{Z}_{i}$, or if $Z_{i}$ is larger than its counterpart $\tilde{Z}_{i}$, e.g., $W_{i}=\lvert Z_{i}-\tilde{Z}_{i}\rvert$ [^1] In general, the functions can be hypothesis-dependent, i.e., $f_{i}$ instead of $f$..
+The function $f$ needs to be symmetric in its arguments and should increase with the values of $Z_{i}$ and $\tilde{Z}_{i}$, e.g., $W_{i}=Z_{i}\vee\tilde{Z}_{i}$, or with the discrepancy between the two, e.g., $W_{i}=\lvert Z_{i}-\tilde{Z}_{i}\rvert$.
+[^1] In general, the functions can be hypothesis-dependent, i.e., $f_{i}$ instead of $f$.
 For each true null hypothesis, $L_{i}$ is equally likely to be ±$\pm 1$ independent of all the scores $W=\{W_{i}\}$ and all other labels $L_{-i}$ because both scores were sampled from the same null distribution and $f$ is symmetric.
 Hence, the number of negative labels, those with $L_{i}=-1$, can be used to estimate the number of true nulls with positive labels among the top scoring hypotheses.
-Using this fact, the scores and labels $\{(W_{i},L_{i}):i=1,\dots,m\}$ are passed through a filter to reject a subset of the top scoring hypotheses, with either the FDR or FDP controlled depending on the user’s preference.
-More recently, research has been focused on improving these filtering methods by leveraging side-information $\mathbf{x}_{i}\in\mathcal{X}$ that complements each p-value $p_{i}$ or score $W_{i}$ to better discriminate between the true and false null hypotheses.
-These methods utilize a variety of strategies but ultimately operate with the same goal in mind: the side-information is used to weight or rescore the hypotheses according to how confident we believe each hypothesis is a false null.
-Indeed, there are several such tools that augment p-values with side-information [25, 34, 33, 36, 50, 35].
-When it comes to using side-information in the competition framework, we are aware of only one method with proper error-rate control along with a precise implementation, namely the recently published Adaptive Knockoffs [43] (AdaPT [33] offers a meta-procedure, but without a precise implementation).
-In this paper, we offer a novel approach for multiple testing with side-information.
-The idea is to randomly split a suspected subsets of the true nulls into two sets: one set is used for training a semi-supervised learning model to discriminate the true and false null hypotheses, while the second is reserved for estimating the number of false discoveries.
+Using this fact, the labels and scores $\{(L_{i},W_{i}):i=1,\dots,m\}$ are passed through a filter to reject a subset of the top scoring hypotheses, with either the FDR or FDP controlled depending on the user’s preference.
+More recently, research has been focused on improving these filtering methods by leveraging side information $\mathbf{x}_{i}\in\mathcal{X}$ that complements each p-value $p_{i}$ or score $W_{i}$ to better discriminate between the true and false null hypotheses.
+These methods utilize a variety of strategies but ultimately operate with the same goal in mind: the side information is used to weight or rescore the hypotheses according to how confidently we believe that each hypothesis is a false null.
+Indeed, there are several such tools that augment p-values with side information [25, 34, 33, 37, 51, 35].
+When it comes to using side information in the competition framework, we are aware of only one fully specified method with rigorous error-rate control, which is the recently published Adaptive Knockoffs [44] (AdaPT [33] only offers a meta-procedure).
+In this paper, we offer a novel approach for multiple testing with side information.
+The idea is to randomly split a set of suspected true nulls into two sets, where one set is used for training a semi-supervised learning model to discriminate between the true and false null hypotheses, and the other for estimating the number of false discoveries.
 Accordingly, we refer to our new approach as RESET—REScoring via Estimating and Training.
 Our method flexibly incorporates any semi-supervised approach to help distinguish between true and false null hypotheses.
-Moreover, it works the same in the competition and p-value context and allows the user to choose between the choice of FDR or FDP control.
+Moreover, it is equally applicable to the competition and p-value context and allows the user to choose between FDR or FDP control.
 We evaluate RESET in a range of simulations and real data applications to verify that it is a competitive alternative to existing methods in terms of statistical power and speed.
 2 Background: competition-based multiple testing
-Competition-based multiple testing was first established in the proteomics community where the goal is to infer which proteins, i.e.
-long chains of amino acids, are present in a biological sample [17].
-Direct protein identification is challenging and so a bottom-up approach is taken where the proteins are broken into small chains of amino acids, called peptides, and the objective is to identify these peptides instead.
+Competition-based multiple testing was first established in the proteomics community where the goal is to infer which proteins, i.e., long chains of amino acids, are present in a biological sample [17].
+Direct protein detection is challenging, and so a bottom-up approach is taken where the proteins are broken into small chains of amino acids, called peptides, and the objective is to identify these peptides instead.
 Then, a process called liquid chromatography tandem mass spectrometry (LC-MS/MS) is used to isolate each sample peptide and subsequently generate a fragmentation spectrum.
 Each such spectrum can be thought of as a fingerprint of the peptide that generated the spectrum, allowing us to hypothesize which peptides are present in the sample.
-This is done by searching each spectrum against a database of candidate peptides called targets to find its optimal peptide-spectrum match (PSM), along with a score indicating the strength of the match.
+This is done by searching each spectrum against a database of candidate peptides called targets to find the spectrum’s optimal peptide-spectrum match (PSM), along with a score indicating the strength of the match.
 Unfortunately, due to the incomplete nature of the database and the noise associated with generating the spectra, the PSM can be incorrect.
-Hence, we wish to employ a testing procedure for each null hypothesis $H_{i}$ that “the $i$th target does not belong to the sample”.
+Hence, we wish to employ a testing procedure for each null hypothesis $H_{i}$ that ‘the $i$th target is not present in the sample’.
 In order to decide which null hypotheses to reject, we pair each target peptide with a decoy peptide that is generated by randomly shuffling or reversing the target peptide’s sequence.
 Accordingly, a second search is performed for each spectrum against the database of decoys to obtain a second PSM and associated score, and only the highest scoring PSM out of the two is retained.
-As outlined in the introduction, we accompany $H_{i}$ with a score $Z_{i}$, defined as the maximum scoring PSM associated to the $i$th target peptide, and the higher the score, the more likely $H_{i}$ is false, i.e., the peptide is in the sample.
-Similarly we define $\tilde{Z}_{i}$ as the maximum scoring PSM associated to the $i$th target’s corresponding decoy peptide.
+As outlined in the introduction, we accompany $H_{i}$ with a score $Z_{i}$, defined as the maximum scoring PSM associated with the $i$th target peptide, and the higher the score, the more likely $H_{i}$ is false, i.e., the peptide is in the sample.
+Similarly, we define $\tilde{Z}_{i}$ as the maximum scoring PSM associated to the $i$th target’s corresponding decoy peptide.
 For each true null hypothesis, we expect $Z_{i}$ and $\tilde{Z}_{i}$ to be drawn from the same (hypothesis-specific) null distribution, independently of everything else.
 Finally, we obtain a winning score $W_{i}=Z_{i}\vee\tilde{Z}_{i}$ and a label $L_{i}$ indicating whether the winning score was from the $i$th target or its decoy [42].
-The statistics community independently formalized this idea in the context of variable selection where the goal is to identify relevant variables from $\{X_{i}:i=1,\dots,p\}$ that are associated with a response $y$.
+Subsequently, the statistics community independently formalized this idea in the context of variable selection, where the goal is to identify relevant variables from $\{X_{i}:i=1,\dots,p\}$ that are associated with a response $y$.
 The procedure that achieves this while controlling the FDR is called the knockoff filter, which begins by pairing each variable $X_{i}$ with an artificial variable $\tilde{X}_{i}$, called a knockoff.
@@
 First, the joint distribution of the variables, $X$, is assumed to be known.
-Second, each observation $(X_{i,1},\dots,X_{i,p},y_{i})$ is sampled i.i.d from the joint distribution $F_{XY}$.
-Because $X$ is random and its distribution known, the knockoffs $\tilde{X}$ are also constructed randomly based on $X$, requiring that they satisfy the following conditions:
+Second, each observation $(X_{i,1},\dots,X_{i,p},y_{i})$ is sampled i.i.d.
+from the joint distribution $F_{XY}$.
+Because $X$ is random and its distribution known, the knockoffs $\tilde{X}$ are also constructed randomly subject to the following conditions:
   1. 1.
@@
 The idea is that the label $L_{i}$ indicates whether the variable $X_{i}$ or the knockoff $\tilde{X}_{i}$ has greater evidence for being relevant to the model, and the score $W_{i}$ indicates by how much.
-Commonly, the score and label are combined by setting $sign(W_{i})=L_{i}$, however as it will become clear in our later presentation of RESET, it is instructive to separate them.
-In this paper, we consider the Lasso Coefficent Difference (LCD) score, which is the difference in the magnitude of the coefficients of the Lasso regression [48].
-Specifically, cross-validation is used to to determine the $\lambda$ penalty when regresson on the augmented design matrix $[X,\tilde{X}]$.
+Commonly, the score and label are combined by setting $sign(W_{i})=L_{i}$; however, as will become clear in our later presentation of RESET, it is instructive to separate them.
+In this paper, we consider the Lasso Coefficient Difference (LCD) score, which is the difference in the magnitude of the coefficients of the Lasso regression [49].
+Specifically, cross-validation is used to determine the $\lambda$ penalty when regressing on the augmented design matrix $[X,\tilde{X}]$.
 We denote the absolute values of the Lasso($\lambda$)-fitted coefficients as $Z_{i}$ for the variables and $\tilde{Z}_{i}$ for the knockoffs.
-The scores and labels are then defined as:
+The scores and labels are then defined as
 $W_{i}:=\big{|}Z_{i}-\tilde{Z_{i}}\big{|},\quad L_{i}:=sign(Z_{i}-\tilde{Z_{i}}).$ (1)
@@
 Let $N$ be the indices of the true null hypotheses.
-The labels $\{L_{i}:i\in N\}$ are i.i.d ±$\pm 1$ uniform random variables independent of all the scores $W$ and the labels of the false null hypotheses.
-For the knockoff filter, Assumption 1 follows from the knockoff construction while in the peptide detection problem, Assumption 1 is based on our previous discussion that the scores of a null target and its decoy, $Z_{i}$ and $\tilde{Z}_{i}$, are independent draws from the same null distribution, and morever, this happens independently of all other pairs of scores.
+The labels $\{L_{i}:i\in N\}$ are i.i.d.
+±$\pm 1$ uniform random variables independent of all the scores $W$ and the labels of the false null hypotheses.
+For the knockoff filter, Assumption 1 follows from the knockoff construction, whereas in the peptide detection problem, Assumption 1 is based on our previous discussion that the scores of a null target and its decoy, $Z_{i}$ and $\tilde{Z}_{i}$, are independent draws from the same null distribution, and moreover, this happens independently of all other pairs of scores.
 This assumption is commonly applied in the literature, with several analyses validating this claim empirically [39, 21].
-Given the above assumption, the number of hypotheses with $L_{i}=-1$ (i.e.
-decoy or knockoff wins) that score above a score cutoff of $\tau$ can be used to estimate the number of true null hypotheses that score above $\tau$ with $L_{i}=1$.
+Given the above assumption, the number of hypotheses with $L_{i}=-1$ (i.e., decoy or knockoff wins) that score above a score cutoff of $\tau$ can be used to estimate the number of true null hypotheses that score above $\tau$ with $L_{i}=1$.
 Using this strategy, we can report a list of discoveries while satisfying a type-1 error rate control in one of two ways.
-The first and by far the more common way is to apply Selective SeqStep+ (SSS+) [1] which controls the FDR at a prespecified threshold $\alpha$ by reporting the following list of discoveries $R_{\alpha}$:
-$R_{\alpha}:=\{i\in[p]:W_{i}\geq\tau,L_{i}=1\},\quad\text{where }\tau:=\min\{t\in\mathbb{R}:\frac{\#\{W_{i}\geq t,L_{i}=-1\}+1}{\#\{W_{i}\geq t,L_{i}=1\}\vee 1}\leq\alpha\}.$ (2)
+The first and by far the more common way is to apply Selective SeqStep+ (SSS+) [1], which controls the FDR at a prespecified threshold $\alpha$ by reporting the following list of discoveries $R_{\alpha}$:
+$R_{\alpha}:=\{i\in[p]:W_{i}\geq\tau,L_{i}=1\},\quad\text{where }\tau:=\min\left\{t\in\mathbb{R}:\frac{\#\{W_{i}\geq t,L_{i}=-1\}+1}{\#\{W_{i}\geq t,L_{i}=1\}\vee 1}\leq\alpha\right\}.$ (2)
 SSS+ is made precise in Algorithm S1 and works in a more general setting when the probabilities of the positive and negative labels of the true null hypotheses are no longer uniform, i.e., $\mathbb{P}(L_{i}=1)=c$ and $\mathbb{P}(L_{i}=-1)=1-c$.
 In the proteomics community, the application of Equation (2) in the peptide detection problem is referred to as target-decoy competition or TDC for short.
-As such, we will refer to the method as SSS+ in the variable selection setting, and TDC in the peptide detection setting.
+As such, we will refer to the method as TDC in the peptide detection setting, and SSS+ in the variable selection setting and more generally.
 Alternatively, we can control the FDP by using, for example, FDP-stepdown (FDP-SD, Algorithm S2[40].
 FDP-SD works by first reranking the labels $L$ according to the scores $W$ in descending order.
-Next, it compares the number of decoy (knockoff) wins in the first $i$ hypotheses, denoted as $D_{i}:=\#\{L_{j}=-1:j\leq i\}$, with a precomputed bound $\delta_{i}$ which is determined by $\alpha$ and $\gamma$.
-Each comparison is made in a ‘stepdown’ fashion, i.e, if the first $i$ hypotheses satisfy $D_{1}\leq\delta_{1},\dots,D_{i}\leq\delta_{i}$, then it moves onto the next comparison at $i+1$.
+Next, it compares the number of decoy (knockoff) wins in the first $i$ hypotheses, denoted as $D_{i}:=\#\{L_{j}=-1:j\leq i\}$, with a precomputed bound $\delta_{i}$, which is determined by $\alpha$ and $\gamma$.
+Each comparison is made in a ‘stepdown’ fashion, i.e., if the first $i$ hypotheses satisfy $D_{1}\leq\delta_{1},\dots,D_{i}\leq\delta_{i}$, then it moves on to the next comparison at $i+1$.
 For small $i$, it is impossible for $D_{i}\leq\delta_{i}$ and so FDP-SD begins this analysis at $i_{0}$, the smallest possible index for which $D_{i}\leq\delta_{i}$.
-Under Assumption 1, reporting the following list of discoveries $R_{\alpha,\gamma}$ controls the FDP.
+Under Assumption 1, reporting the following list of discoveries $R_{\alpha,\gamma}$ controls the FDP:
 $R_{\alpha,\gamma}:=\{i\in[p]:i\leq k_{FDP},L_{i}=1\},\quad k_{FDP}:=\max\{i:\Pi_{j=i_{0}}^{i}1_{D_{j}\leq\delta_{j}}=1\text{ or }i=0\}.$ (3)
 2.1 Extension to side information
-In proteomics, so-called post-processors such as PeptideProphet [28] and Percolator [26] have popularized the use of side-information to improve the number of detected peptides.
-These tools use machine learning procedures, leveraging features associated with the PSMs, such as the charge state associated with the spectrum, and the matched peptide’s length, to better distinguish correct PSMs, and hence to deliver more discoveries.
-These tools have no theoretical guarantees and having recently pointed out scenarious where such post-processors can fail to control the FDR [18], we will not analyze these tools in this paper.
-The only ‘intrinsic’ competition-based multiple testing procedure that uses arbitrary side information in a data-driven manner, that we are aware of, is the recent Adaptive Knockoffs (AdaKO) by Ren and Candès [43].
+In proteomics, so-called post-processors such as PeptideProphet [28] and Percolator [26] have popularized the use of side information to increase the number of detected peptides.
+These tools use machine learning procedures, leveraging features associated with the PSMs, such as the charge state associated with the spectrum and the matched peptide’s length, to better distinguish correct from incorrect PSMs, and hence to deliver more discoveries.
+These tools have no theoretical guarantees and, having recently pointed out scenarios where such post-processors can fail to control the FDR [18], we will not analyze these tools in this paper.
+To our knowledge, the only ‘intrinsic’ competition-based multiple testing procedure that uses arbitrary side information in a data-driven manner is the recent Adaptive Knockoffs (AdaKO) by Ren and Candès [44].
 AdaKO is a specialization of the p-value based procedure AdaPT [33] that we discuss in the next section.
@@
 Adaptive Knockoffs works by iteratively pruning $\mathcal{S}$, the candidate set of hypotheses, selecting each time the hypothesis it deems most likely to be a true null.
-To do that, it fits a classification algorithm using a strict set of information that partially masks the data: the scores $W$, the side information $\mathbf{x}$, the labels $L_{i}$ of the hypotheses that were pruned up to that point ($i\in[p]\setminus\mathcal{S}$), as well as #$\#\{L_{i}=1:i\in S\}$ and #$\#\{L_{i}=-1:i\in S\}$.
-In their paper, Ren and Candès investigate this prunning strategy using different types of “filters”: a logistic regression (LR), a generative additive model (GAM) which is the default filter in their R package, a random forest (RF) and a two-group probabilistic model which is fitted using the expectation-maximization algorithm (EM) [13].
+To do that, it fits a classification algorithm using a strict set of information that partially masks the data: the scores $W$, the side information $\mathbf{x}$, and the labels $L_{i}$ of the hypotheses that were pruned up to that point ($i\in[p]\setminus\mathcal{S}$), as well as #$\#\{L_{i}=1:i\in S\}$ and #$\#\{L_{i}=-1:i\in S\}$.
+In their paper, Ren and Candès investigate this pruning strategy using different types of ‘filters’: a logistic regression (LR), a generalized additive model (GAM) which is the default filter in their R package, a random forest (RF), and a two-group probabilistic model which is fitted using the expectation-maximization algorithm (EM) [13].
 The pruning process terminates when the estimated FDR is less than or equal to $\alpha$:
 $\widehat{FDR}_{\mathcal{S}}:=\frac{\#\{i\in\mathcal{S},L_{i}=-1\}+1}{\#\{i\in\mathcal{S},L_{i}=1\}\vee 1}\leq\alpha.$ (4)
-The remaining positively labelled hypotheses in $\mathcal{S}$ are then reported. It was shown that given the following natural extension of Assumption 1, Adaptive Knockoffs controls the FDR.
+The remaining positively labeled hypotheses in $\mathcal{S}$ are then reported. It was shown that, given the following natural extension of Assumption 1, Adaptive Knockoffs controls the FDR.
 Assumption 2.
 Let $N$ be the indices of the true null hypotheses.
-The labels $\{L_{i}:i\in N\}$ are i.i.d ±$\pm 1$ uniform random variables independent of all the scores $W$, the side information $\mathbf{x}$, and the labels of the false null hypotheses.
+The labels $\{L_{i}:i\in N\}$ are i.i.d.
+±$\pm 1$ uniform random variables independent of all the scores $W$, the side information $\mathbf{x}$, and the labels of the false null hypotheses.
 3 Background: p-value based multiple testing
-Multiple testing with FDR control was popularized by Benjamini and Hochberg when they developed the first testing procedure of this kind, which we refer to as the “BH procedure” [2].
+Multiple testing with FDR control was popularized by Benjamini and Hochberg when they developed the first testing procedure of this kind, which we refer to as the ‘BH procedure’ [2].
 Since then, many p-value based methods that control the FDR have been developed to improve statistical power.
-One such strategy for improving on BH is to estimate the fraction of true null hypotheses, denoted as $\pi_{0}$.
+One such strategy for improving on BH involves estimating the fraction of true null hypotheses, denoted as $\pi_{0}$.
 Indeed, the BH procedure controls the FDR conservatively at level $\pi_{0}\cdot\alpha$ instead of $\alpha$.
@@
 [3] estimate $\pi_{0}$, and essentially ‘plug-in’ the estimate $\hat{\pi}_{0}$, by applying the BH procedure at a threshold of $\alpha/\hat{\pi}_{0}$ instead of $\alpha$.
-Instead of the FDR, an alternative to BH that controls the FDP, is Guo and Romano’s stepdown approach [20], which we refer to as GR-SD.
+Instead of the FDR, Guo and Romano’s stepdown approach [20], which we refer to as GR-SD, controls the FDP.
 In fact, this stepdown approach inspired the development of FDP-SD in the competition setting.
-GR-SD works by first sorting the p-values in ascending order, i.e.
-$p_{1}\leq p_{2}\leq...\leq p_{m}$.
-Then it constructs a sequence of thresholds, $\delta_{i}$ for each p-value $p_{i}$ and searches for the largest $k$ such that $p_{1}\leq\delta_{1},\dots,p_{k}\leq\delta_{k}$ mutually holds, and then reports the $k$ smallest p-values.
+GR-SD works by first sorting the p-values in ascending order, i.e., $p_{1}\leq p_{2}\leq...\leq p_{m}$.
+Then it constructs a sequence of thresholds, $\delta_{i}$ for each p-value $p_{i}$ and searches for the largest $k$ such that $p_{1}\leq\delta_{1},\dots,p_{k}\leq\delta_{k}$ mutually holds, and then reports the corresponding $k$ smallest p-values.
 We provide the pseudocode for this method in Algorithm S3.
 3.1 Extension to side information
-Other methods utilize side-information in one of two ways: (a) they provide an improved ordering of the p-values based on how confident we believe each hypothesis is a false null, or (b) by assigning weights to the p-values, placing higher weights on hypotheses that we believe to be false nulls.
-However, the shortcoming of these side-information approaches is that the majority of them are not data-driven, or ‘adaptive’, that is, the ordering or weighting is fixed before the actual testing procedure takes place.
+Subsequent approaches were developed to take advantage of scenarios with available side information.
+Roughly speaking, these methods use side information to weight or rescore the hypotheses, placing higher weights or scores on hypotheses that are believed to be false nulls.
+Such procedures can be either ‘predetermined’, that is, the weighting or scoring is determined before the actual testing procedure takes place, or ‘adaptive’, in which the weighting or scoring is selected according to the data at hand and therefore often leads to greater statistical power.
 Method Error control Speed
@@
 AdaFDR (2019) Asymptotic FDP Fast
-Table 1: A table summarizing some popular and recent p-value based multiple testing procedures that use side information in terms of its type-1 error control and the computational speed.
-More recently, several methods have been developed that take an adaptive approach instead.
-Such methods include Adaptive P-value Thresholding (AdaPT) and its derivatives [33, 8], Z-value Adaptive Procedures (ZAP) [35] and AdaFDR [50].
+Table 1: Summary of some popular and recent p-value based multiple testing procedures that use side information, indicating each method’s type-1 error control and speed.
+Such adaptive methods include Adaptive P-value Thresholding (AdaPT) and its derivatives [33, 8], Z-value Adaptive Procedures (ZAP) [35] and AdaFDR [51].
 Table 1 provides a qualitative summary of each method in terms of its type-1 error control, its error control guarantees, and its computational speed.
-On balance, none of these mentioned tools enjoy both finite type-1 error control while also being fast, and as we will see, some of these methods have variable power when considering a range of simulated and real datasets.
+Notably, none of these mentioned tools offers a fast finite type-1 error control, and as we will see, some of these methods have variable power when considering a range of simulated and real datasets.
 We next provide some background on each of those methods.
-AdaFDR, splits the pairs $(p_{i},\mathbf{x}_{i})_{i=1}^{m}$ into two folds.
-For each fold, it fits a mixture model to optimally rescore the hypotheses, which is then applied to the other fold where a list of discoveries is obtained.
+AdaFDR splits the pairs $(p_{i},\mathbf{x}_{i})_{i=1}^{m}$ into two folds.
+For each fold, the method fits a mixture model to optimally rescore the hypotheses, and the model is then applied to the other fold where a list of discoveries is obtained.
 The two list of discoveries are joined together and then reported.
-AdaFDR’s mixture model is a composition of Gaussian components to model local ‘bumps’ of genuine signals (false nulls) within the data plus a generalized linear model to capture a monotonic relationship between the side information and the signals.
-AdaPT iteratively prunes a candidate set of hypotheses $\mathcal{S}$ similar to Adaptive Knockoffs.
-At each iteration, AdaPT fits a two group mixture model on a strict set of information: the p-values outside $\mathcal{S}$, the masked p-values in $\mathcal{S}$ given by $\{p_{i}\wedge(1-p_{i}):i\in\mathcal{S}\}$, the side information $\mathbf{x}$, as well as #$\#\{p_{i}>1/2:i\in\mathcal{S}\}$ and #$\#\{p_{i}<1/2:i\in\mathcal{S}\}$.
+AdaFDR’s mixture model is a composition of Gaussian components to model local ‘bumps’ of genuine signals (false nulls) within the data plus an additional component that captures the monotonic relationship between the side information and the signals.
+AdaPT iteratively prunes a candidate set of hypotheses $\mathcal{S}$, similar to Adaptive Knockoffs, which was derived from it.
+At each iteration, AdaPT fits a two-group mixture model on a strict set of information: the p-values outside $\mathcal{S}$, the masked p-values in $\mathcal{S}$ given by $\{p_{i}\wedge(1-p_{i}):i\in\mathcal{S}\}$ where $1-p_{i}$ is referred to as the mirrored p-value, the side information $\mathbf{x}$, as well as #$\#\{p_{i}>1/2:i\in\mathcal{S}\}$ and #$\#\{p_{i}<1/2:i\in\mathcal{S}\}$.
 It then removes the hypothesis from $\mathcal{S}$ that is deemed most likely to be a true null based on the fitted model.
 To fit the mixture model, the EM algorithm is used.
-The M-step of the EM algorithm results in a regression problem which allows the user to flexibly choose whichever regression method they like.
-They considered a generative additive model (GAM), a generalized linear model (GLM) and a generalized linear model with $L_{1}$ regularization (GLMnet).
-If we identify each p-value $p_{i}>1/2$ with a label $L_{i}=-1$ and each p-value $p_{i}<1/2$ with a label $L_{i}=1$, then AdaPT terminates this pruning procedure when the estimated FDR, like in Equation (4), is $\leq\alpha$.
+The M-step of the EM algorithm results in a regression problem which allows the user to flexibly use the regression method of their choice.
+The authors considered a generalized additive model (GAM), a generalized linear model (GLM), and a generalized linear model with $L_{1}$ regularization (GLMnet).
+If we identify each p-value $p_{i}>1/2$ with a label $L_{i}=-1$ and each p-value $p_{i}<1/2$ with a label $L_{i}=1$, then AdaPT terminates this pruning procedure when the estimated FDR, as in Equation (4), is $\leq\alpha$.
 The rationale is that for true null hypotheses, the p-values are usually uniformly distributed, so #$\#\{L_{i}=-1:i\in\mathcal{S}\}$ estimates the number of true null hypotheses with positive labels, #$\#\{L_{i}=1:i\in\mathcal{S}\}$.
-It then reports the remaining positively labelled hypotheses in $\mathcal{S}$ [^3] AdaPT and AdaPTg do not use labels explicitly – we introduce those here for notational convenience.
+It then reports the remaining positively labeled hypotheses in $\mathcal{S}$ [^3] AdaPT and AdaPTg do not use labels explicitly — we introduce those here for notational convenience.
 AdaPTg generalizes AdaPT by constructing asymmetric regions to define the positive and negative labels.
 For example, the p-values $p_{i}\in\left(0.3,0.9\right)$ can be identified with a label $L_{i}=-1$ and the p-values $p_{i}<0.3$ can be identified with a label $L_{i}=1$.
-One advantage of defining these asymmetric regions, is to avoid overestimating the number of true null hypotheses when there is a concentration of p-values towards 1.
+One advantage of defining these asymmetric regions, is to avoid overestimating the number of true null hypotheses in cases where their p-values are not uniformly distributed and there is a concentration of true null p-values near 1.
 Subsequently, the estimated FDR needs to be adjusted to account for the fact that we expect twice as many true null p-values with negative labels than positive labels.
-Moreover, the masked p-values are also no longer $p_{i}\wedge(1-p_{i})$ and also need adjustment.
+Moreover, the masked p-values are also no longer $p_{i}\wedge(1-p_{i})$, since $1-p_{i}$ no longer ‘correctly mirrors’ the values in $\left(0.3,0.9\right)$ onto $(0,0.3)$ and therefore needs adjustment.
 In essence, the rest of the procedure remains the same.
 Lastly, AdaPT-GMMg replaces the two group mixture model in AdaPT/AdaPTg for a Gaussian mixture model (GMM).
-Similar to AdaPT/AdaPTg, it also allows the user decide from a collection of regression methods to fit the model at the M-step of the EM algorithm.
-Lastly, ZAP-asymp directly fits a beta mixture model to the data.
-ZAP-asymp’s direct application of the model means that it is only able to offer asymptotic FDR control.
+Similar to AdaPT/AdaPTg, it also allows the user to choose from a collection of regression methods to fit the model at the M-step of the EM algorithm.
+Lastly, ZAP-asymp directly fits a beta mixture model to the data but is only able to offer asymptotic FDR control.
 Importantly, ZAP-asymp can operate directly on the test statistic instead of just the corresponding p-values to rerank the hypotheses.
@@
 In Section S.3, we prove that RESET controls the FDR/FDP in the finite-sample setting.
-For simplicity, we use the target-decoy terminology, i.e., we refer to a hypothesis with $L=1$ as a ‘target win’ or a target for short, and $L=-1$ as a ‘decoy win’ or a decoy for short.
+For simplicity, we use the target-decoy terminology, i.e., we refer to a hypothesis with $L_{i}=1$ as a ‘target win’ or a target for short, and $L_{i}=-1$ as a ‘decoy win’ or a decoy for short.
 4.1 RESET outline
   1. 1.
-    Intrinsically, RESET’s input consists of the labels, winning scores and side information $(L,W,\mathbf{x})$.
+    Intrinsically, RESET’s input consists of the labels, winning scores, and side information $(L,W,\mathbf{x})$.
 Hence, in the p-value setting, each $p_{i}$ is first converted to a pair $(L_{i},W_{i})$ in the following way:
@@
     where $0<a\leq 1/2\leq b_{1}<b_{2}\leq 1$ determine the cutoff regions for defining a positive and negative label, and $\Phi$ is the standard normal CDF.
-Hypotheses with p-values outside of $[0,a)\cup(b_{1},b_{2}]$ are thrown out and the default is $a=b_{1}=1/2$ and $b_{2}=1$.
+Hypotheses with p-values outside of $[0,a)\cup(b_{1},b_{2}]$ are thrown out.
+The default is $a=b_{1}=1/2$ and $b_{2}=1$.
   2. 2.
     RESET independently and randomly assigns each winning decoy as a training decoy with probability $s$ (we used $s=1/2$ throughout).
-The complementary set of decoy wins defines the estimating decoy set, and the set of pseudo targets is the set of all estimating decoys and target wins.
-The pseudo labels, $\tilde{L}_{i}$, denotes whether the $i$th hypothesis is a pseudo target ($\tilde{L}_{i}=1$) or a training decoy ($\tilde{L}_{i}=-1$).
+The complementary set of decoy wins defines the estimating decoy set, and the set of pseudo-targets is the set of all estimating decoys and target wins.
+The pseudo labels, $\tilde{L}_{i}$, denotes whether the $i$th hypothesis is a pseudo-target ($\tilde{L}_{i}=1$) or a training decoy ($\tilde{L}_{i}=-1$).
   3. 3.
-    Next, RESET applies a user-selected semi-supervised machine learning model which uses the pseudo labels, winning scores and side information $(\tilde{L},W,\mathbf{x})$.
-Note that this is a semi-supervised task as the positive pseudo labels will typically contain a mixture of true and false nulls, while the negative pseudo labels will indicate true null hypotheses.
-    The output of this step is a rescoring of the training decoys and pseudo targets, denoted as $\tilde{W}$.
-Ideally, $\tilde{W}$ scores many of the false nulls among the pseudo targets higher than they were scored originally.
+    Next, RESET applies a user-selected semi-supervised machine learning model which uses the pseudo labels, winning scores, and side information $(\tilde{L},W,\mathbf{x})$.
+Note that this is a semi-supervised task because the positive pseudo labels typically contain a mixture of true and false nulls, while the negative pseudo labels indicate true null hypotheses.
+    The output of this step is a rescoring of the training decoys and pseudo-targets, denoted as $\tilde{W}$.
+Ideally, $\tilde{W}$ scores many of the false nulls among the pseudo-targets higher than they were scored originally.
 We provide a specific framework that we developed for this crucial step in Section 4.3.
   4. 4.
-    With the training complete, the training decoys are thrown out, and the original labels $L$ of the remaining pseudo targets are revealed.
+    With the training complete, the training decoys are thrown out, and the original labels $L$ of the remaining pseudo-targets are revealed.
   5. 5.
-    RESET then determines its list of discoveries, by applying to the pseudo-targets using the original labels $L$ and the new scores $\tilde{W}$ either
+    RESET then determines its list of discoveries by applying to the pseudo-targets using the original labels $L$ and the new scores $\tilde{W}$ either
     • SSS+ (Algorithm S1) at the desired level $\alpha$ for FDR control, or
@@
 This parameter is used to define the expected ratio of the number of null targets to decoys, $\frac{c}{1-c}$.
-Typically, this is set to $c=1/2$ for an expected target-decoy ratio of 1, however, RESET uses approximately half of the decoys for training purposes.
+Typically, this is set to $c=1/2$ for an expected target-decoy ratio of 1; however, RESET uses approximately half of the decoys for training purposes (by default).
 Since those decoys are subsequently thrown out, this parameter needs to be adjusted.
-Indeed, assuming that the labels $L$ of the true nulls are i.i.d uniform ±$\pm 1$ RVs, and $s=1/2$, we set the parameter $c=2/3$, so that the expected target-decoy ratio is 2.
-More generally, if we know $\mathbb{P}(L_{i}=1)\leq c_{0}$, then we set $c=\frac{c_{0}}{1-s\cdot(1-c_{0})}$.
+Indeed, assuming that the labels $L$ of the true nulls are i.i.d.
+uniform ±$\pm 1$ RVs, and $s=1/2$, we set the parameter $c=2/3$, so that the expected target-decoy ratio is 2.
+More generally, if we know that for a true null $\mathbb{P}(L_{i}=1)\leq c_{0}$, then we set $c=\frac{c_{0}}{1-s\cdot(1-c_{0})}$.
 For example in the p-value setting, $c_{0}=\frac{a}{a+b_{2}-b_{1}}$.
@@
 We apply $\Phi^{-1}$ to the p-value $p_{i}$ or the mirrored p-value $(b_{2}-p_{i})\cdot\frac{a}{b_{2}-b_{1}}$ because we found that doing so makes the semi-supervised learning task of Step 3 to be generally less variable and more powerful (data not shown).
-The idea of mirrored p-value was established by Lei and Fithian [33] and later generalized by [8] as discussed in the previous section.
 Crucially in Step 4, the training decoys are thrown out.
-This is important since during the training phase of Step 3, the semi-supervised machine learning model will have presumably learned to discriminate between the training decoys and the pseudo targets.
-Hence, the training decoys are no longer suitable for estimating the true nulls among those pseudo targets, because the rescoring will tend to rank them lower than those true nulls.
+This is important since during the training phase of Step 3, the semi-supervised machine learning model will have presumably learned to discriminate between the training decoys and the pseudo-targets.
+Hence, true null targets will tend to score higher than the training decoys making the latter unsuitable for estimating the number of true nulls among the top scoring hypotheses.
 4.2 RESET controls the FDR or FDP
@@
 Let $N$ be the indices of the true null hypotheses.
-The labels $\{L_{i}:i\in N\}$ are i.i.d ±$\pm 1$ random variables with $\mathbb{P}(L_{i}=1)\leq c_{0}$ independently of all the scores $W$, the side information $\mathbf{x}$, and the labels of the false null hypotheses.
+The labels $\{L_{i}:i\in N\}$ are i.i.d.
+±$\pm 1$ random variables with $\mathbb{P}(L_{i}=1)\leq c_{0}$ and are independent of all the scores $W$, the side information $\mathbf{x}$, and the labels of the false null hypotheses.
 In the p-value setting, Assumption 3 is satisfied under some mild conditions regarding the distribution of the true null p-values.
@@
 Lemma 1 (Chao and Fithian).
-Assume that the true null p-values are mutually independent and independent of the false null p-values, all the side information, and assume the true null p-values have a non-decreasing density.
+Assume that the true null p-values have a non-decreasing density and that they are mutually independent and that they are independent of the false null p-values as well as of all the side information.
 Then Assumption 3 is satisfied with $c_{0}=\frac{a}{a+b_{2}-b_{1}}$.
@@
 In practice, Assumption 3 (or the conditions of Lemma 1) needs to be verified.
-We provide reasonable justification of this assumption in our real data applications in Sections 5.2 and 6.2.
+We provide reasonable heuristic justification of this assumption in our real data applications in Sections 5.2 and 6.2.
 We are now ready to state our two main theorems.
 Theorem 1.
-Under Assumption 3 RESET controls the FDR at the user-specified threshold of $\alpha$.
+Under Assumption 3 RESET controls the FDR at the user-specified threshold $\alpha$.
 Theorem 2.
-Under Assumption 3 RESET controls the FDP at the user-specified threshold of $\alpha$ with confidence $1-\gamma$.
-We defer the main proof of RESET’s FDR and FDP control to Supplementary Section S.3.
-The proof requires an intermediate Lemma (Lemma S1) which states that the labels $L$ of the true null pseudo targets in Step 4 remaim i.i.d.
+Under Assumption 3 RESET controls the FDP at the user-specified threshold $\alpha$ and confidence $1-\gamma$.
+We defer the proof of RESET’s FDR and FDP control to Supplementary Section S.3.
+The proof requires an intermediate lemma (Lemma S1), which states that the labels $L$ of the true null pseudo-targets in Step 4 remain i.i.d.
 ±$\pm 1$ RVs independently of everything else, even after sampling the training decoys in Step 2 and the application of the machine learning model in Step 3.
@@
 In this section we describe our specific framework for RESET’s Step 3, which as mentioned, can use any semi-supervised machine learning model.
-It consists of two man steps: (1) we apply a range of classification algorithms to estimate a high quality positive set from the data, containing presumably many false nulls, and (2) this positive set is then subsequently used for a second application of the classification algorithms to rescore the hypotheses.
+It consists of two main steps: (1) we apply a range of classification algorithms to estimate a high quality positive set from the data, containing presumably many false nulls, and (2) this positive set is then subsequently used for a second application of the classification algorithms to rescore the hypotheses.
   1. i.
-    Define the negative set as the set of all training decoys and the positive set as a subset of the top scoring pseudo targets (which we elaborate in Step III at the end of this section).
+    Define the negative set as the set of all training decoys and the positive set as a subset of the top scoring pseudo-targets (which we elaborate on in Step III at the end of this section).
 We define the features that are subsequently used by a collection of classification algorithms as the combined score and side information $(W,\mathbf{x})$.
@@
 For each $k\in\{1,2,\dots,K\}$, we train the classification algorithm using the positive and negative set on the features in the $K\setminus\{k\}$ folds and apply the trained model to all the hypotheses in the $k$th test fold.
-We consider a random forest (RF), a generative additive model (GAM) and a collection of two-layer neural networks (NN) with decay parameter $\lambda\in\{0,0.1,1\}$ and a number of hidden layer nodes $h\in\{2,5,10\}$.
+We consider a random forest (RF), a generalized additive model (GAM), and a collection of two-layer neural networks (NN) with decay parameter $\lambda\in\{0,0.1,1\}$ and a number of hidden layer nodes $h\in\{2,5,10\}$.
   3. iii.
@@
   4. iv.
-    To evaluate each classification algorithm, we apply Selective SeqStep (SSS) (Algorithm S1 which is the same as SSS+ of Equation (2) without the “+1”) at an FDR threshold $\alpha$ to each of the $rK$ test folds, using the scores obtained from the classification algorithm and the pseudo labels.
+    To evaluate each classification algorithm, we apply Selective SeqStep (SSS) (Algorithm S1, which is the similar to SSS+ of Equation (2) without the ‘+1’) at an FDR threshold $\alpha$ to each of the $rK$ test folds, using the scores obtained from the classification algorithm and the pseudo labels.
 If $\mathbb{P}(L_{i}=1)\leq c_{0}$ then we set $c=1-s\cdot(1-c_{0})$.
-This is to ensure that the ratio $\frac{c}{1-c}$ correctly accounts for the number of null pseudo targets to training decoys [^4] This is not a requirement for error control..
+This is to ensure that the ratio $\frac{c}{1-c}$ correctly accounts for the number of null pseudo-targets to training decoys [^4] This is not a requirement for error control..
   5. v.
@@
     Using the chosen classification algorithm, we assign new scores $\tilde{W}$ to the hypotheses.
-$\tilde{W}_{i}$ is defined as the average decision value for the $i$th hypothesis, taken over the $K$ test folds that the $i$th hypothesis appears in, and the $r$ repetitions executed in Step iii.
+$\tilde{W}_{i}$ is defined as the average decision value for the $i$th hypothesis in the test fold that it appears in, taken over the $r$ repetitions executed in Step iii.
   7. vii.
@@
   8. viii.
-    The positive set is redefined as those pseudo targets that are ‘discovered’ in Step vii.
+    The positive set is redefined as those pseudo-targets that are ‘discovered’ in Step vii.
 Steps ii-vi are then repeated with the new positive set.
-To ensure the positive set is not too small, we increase $\alpha$ by increments of 0.01 until the positive set reaches min_positive or until all the pseudo targets are in the positive set.
+To ensure that the positive set is not too small, we increase $\alpha$ by increments of 0.01 until the positive set reaches min_positive or until all the pseudo-targets are in the positive set.
 We used min_positive = 50.
@@
     To reduce the effect of ‘noisy’ side information that is unable to distinguish between true and false nulls, we implement an initial side information selection procedure.
-That is, we use a generative additive model to fit a smoothing spline using $W\cdot\tilde{L}$ as the response against each user-provided side information variable one at a time.
+That is, we use a generalized additive model to fit a smoothing spline using $W\cdot\tilde{L}$ as the response against each user-provided side information variable one at a time.
 We use tryCatch in case the smoothing spline fails, in which case we fit the response directly to the side information variable, that is, we use a linear fit.
-Again, we try to make use of the zero-scoring hypotheses where we can by setting $W\cdot\tilde{L}=0$ in these cases.
+Again, we try to make use of some of the zero-scoring hypotheses setting $W\cdot\tilde{L}=0$ in these cases.
 We keep only the side information variables with sufficiently small p-values in the fitted model ($p<0.01$).
   3. III.
-    Not all the pseudo targets need to be included in the initial positive set in Step (i).
-After all, we expect many of the pseudo targets to be true nulls.
+    Not all the pseudo-targets need to be included in the initial positive set in Step i.
+After all, we expect many of the pseudo-targets to be true nulls.
 Hence, to improve the performance of RESET, we consider the following strategy.
@@
 Then select the feature that maximizes the number of pseudo discoveries.
-Then using the chosen feature, we define the positive set as those pseudo targets discovered at $\alpha_{0}$ (we used $\alpha_{0}=50\%$ throughout), again using SSS with the pseudo labels $\tilde{L}$.
+Then using the chosen feature, we define the positive set as those pseudo-targets discovered at $\alpha_{0}$, again using SSS with the pseudo labels $\tilde{L}$ (we used $\alpha_{0}=50\%$ throughout).
 Usually the score is selected, but in some cases the side information is highly informative and is selected instead.
-    Users interested in a small FDR level, like $\alpha=1\%$, could probably benefit from reducing $\alpha_{0}$ to further improve the performance, and in particular the speed of RESET (though as we mentioned, we kept $\alpha_{0}=50\%$ across the board).
-To ensure the positive set is not too small, we increase $\alpha_{0}$ by increments of 0.01 until the positive set reaches min_positive or until all the pseudo targets are in the positive set.
-We refer to the general RESET wrapper in combination with our semi-supervised approach as RESET Ensemble.
-We refer to RESET RF/GAM/NN when the ensemble of machine learning methods in Step 2 is reduced to a random forest, generative additive model, and a collection of neural networks as described, respectively.
+    Users interested in a small FDR level, like $\alpha=1\%$, could probably benefit from reducing $\alpha_{0}$ to further improve the performance, and in particular the speed, of RESET (though as we mentioned, we kept $\alpha_{0}=50\%$ across the board).
+To ensure that the positive set is not too small, we increase $\alpha_{0}$ by increments of 0.01 until the positive set reaches min_positive or until all the pseudo-targets are in the positive set.
+We refer to the general RESET wrapper in combination with the above semi-supervised approach as RESET Ensemble.
+We refer to RESET RF/GAM/NN when the ensemble of machine learning methods in Step ii is reduced to a random forest, generalized additive model, and a collection of neural networks as described, respectively.
 4.4 Code availability
-Our code is available at https://github.com/freejstone/stat_RESET_paper_code. We intend on developing a package for RESET in R.
+Our code is available at https://github.com/freejstone/stat_RESET_paper_code with an Apache-2.0 license. We intend to develop a package for RESET in R.
 5 Experiments: competition setting
 5.1 Numerical experiments
-In this section, we conduct a wide range of simulations in the variable selection problem based on those introducted by Ren and Candès [43].
+In this section, we conduct a wide range of simulations in the variable selection problem based on those introduced by Ren and Candès [44].
 We compare RESET Ensemble to Adaptive Knockoffs (AdaKO) and the generic Knockoff Filter (KO) (see Section 2).
-Further implementation details for AdaKO are given in Supplementary Secton S.5.
+Further implementation details for AdaKO are given in Supplementary Section S.5.
 We provide a brief description of each simulation setup, referring to [44] for further details.
@@
 The labels and scores are obtained according to Equation (1).
-We focus on FDR control in this section, and study FDP control in Section 5.2.
+We focus on FDR control in this section, and we study FDP control in Section 5.2.
 5.1.1 Simulation 1: Linear model with one-dimensional side information
@@
 Then $k$ indices from $\{1,2,\dots,p\}$ are sampled without replacement with indices $i\in\{1,\dots,2k\}$ having probability proportional to $1/i^{2}$ and the indices $i\in\{2k+1,\dots,p\}$ having zero probability.
-The $k$ chosen indices are the nonzero $\beta$’s and are assigned a value of ±$\pm 3.5/\sqrt{n}$ where the signs are chosen i.i.d uniformly.
+The $k$ chosen indices are the nonzero $\beta$’s and are assigned a value of ±$\pm 3.5/\sqrt{n}$ where the signs are chosen i.i.d.
+uniformly.
 We consider $k\in\{50,150,300\}$ while the original simulation only considers $k=150$.
-Each hypothesis is supported by the side information $\mathbf{x}_{i}=i$, which should help in detecting the nonzero $\beta$’s as smaller values of $i$ are more likely to correspond to a nonzero $\beta_{i}$.
+Each hypothesis is supported by the side information $\mathbf{x}_{i}=i$, which should help in detecting the nonzero $\beta$’s, because smaller values of $i$ are more likely to correspond to a nonzero $\beta_{i}$.
 We used 100 independent runs of the above, where each run consisted of drawing those $n$ observations.
 The original simulation in [44] sampled the $\beta$’s once and fixed them across the 100 runs.
-Here we redrew the $\beta$’s for each run so that we may average our results over a range of models, as we noticed significant variability between the draws.
+Here we redrew the $\beta$’s for each run so that we may average our results over a range of models, as we noticed that all the methods exhibited significant variability between the draws.
 5.1.2 Simulation 2: Linear model with two-dimensional side information
@@
 Accordingly, we also adjusted the number of false nulls to $k\in\{150,300,450\}$.
+Note that unlike the previous simulation, it is not relevant how the $k$ variable indices are selected from $\{1,2,\dots,p\}$.
 Each hypothesis is equipped with the following two-dimensional side information.
-For the $i$th hypothesis, we draw a pair of observations $\mathbf{x}_{i}$ from a bivariate normal $\mathcal{N}(\boldsymbol{\mu},I)$ with $\boldsymbol{\mu}=(0,0)$ for a true null hypothesis and $\boldsymbol{\mu}=\pm(2,2)$ for a false null hypothesis, where the signs are chosen i.i.d uniformly.
-We used 20 independent runs of this larger simulation instead of 100 where in each run we draw those $n$ observations, $\beta$ coefficients and side information variables.
+For the $i$th hypothesis, we draw a pair of observations $\mathbf{x}_{i}$ from a bivariate normal $\mathcal{N}(\boldsymbol{\mu},I)$ with $\boldsymbol{\mu}=(0,0)$ for a true null hypothesis and $\boldsymbol{\mu}=\pm(2,2)$ for a false null hypothesis, where the signs are chosen i.i.d.
+uniformly.
+We used 20 independent runs of this larger simulation instead of 100, where in each run we draw those $n$ observations, $\beta$ coefficients, and side information variables.
 5.1.3 Simulation 3: Logistic model with two-dimensional side information
 Next, we consider Simulation 2 by Ren and Candès [44].
-They generate $n=1000$ observations with $p=1600$ variables sampled from the joint data distribution of $(X,Y)$ where $X$ is distributed according to a multivariate normal distribution and $Y|X\sim\text{Bernoulli}\left(e^{\beta\cdot X}/\left(1+e^{\beta\cdot X}\right)\right)$ is a logistic model (see [43] for further details).
+They generate $n=1000$ observations with $p=1600$ variables sampled from the joint data distribution of $(X,Y)$, where $X$ is distributed according to a multivariate normal distribution and $Y|X\sim\text{Bernoulli}\left(e^{\beta\cdot X}/\left(1+e^{\beta\cdot X}\right)\right)$ is a logistic model (see [44] for further details).
 Each hypothesis’ side information is a unique pair of values $\mathbf{x}_{i}=\left(\mathbf{x}_{i1},\mathbf{x}_{i2}\right)$ from the lattice $[-20,19]\times[-20,19]\subseteq\mathbb{Z}^{2}$.
@@
 An image of the nonzero $\beta$’s is given in Figure 1.
-Each nonzero $\beta$ is set to ±$\pm 25/\sqrt{n}$ where the signs are chosen i.i.d uniformly.
+Each nonzero $\beta$ is set to ±$\pm 25/\sqrt{n}$ where the signs are chosen i.i.d.
+uniformly.
 We used 98 runs of the above setup (since in 2 runs, AdaKO EM failed with an error), where each run consisted of drawing those $n$ observations.
-The original simulation in [43] sampled the sign of the $\beta$’s once, fixing them across the 98 runs.
-Instead, we redraw the sign of the $\beta$’s in each of the 98 runs since we found the results significantly varied with those draws.
+The original simulation in [44] sampled the sign of the $\beta$’s once, fixing them across the 100 runs.
+Instead, we redraw the sign of the $\beta$’s in each of the 98 runs since we found that the results of all the methods significantly varied with those draws.
 Figure 1: Location of nonzero $\beta$’s according to the pair of values $(x,y)$ in the lattice. The same image is given in [44].
@@
 Each variable is equipped with the following three-dimensional side information.
-The first $k$ variables are randomly assigned a drawn from $i\in\{1,\dots,2k\}$ taken without replacement and having probability proportional to $1/i^{2}$.
+The first $k$ variables are randomly assigned a value drawn from $i\in\{1,\dots,2k\}$ taken without replacement and having probability proportional to $1/i^{2}$.
 The remaining $p-k$ variables are uniquely assigned to one of the remaining values from $\{1,2,\dots,p\}$ uniformly at random.
-We expect that most of the nonzero $\beta$’s to have small values of $i$, but the effect of the random assignment makes the resulting side information variable less informative than in Simulation 1 as depicted in Figure 2.
+We expect most of the nonzero $\beta$’s to have small values of $i$, but the effect of the random assignment makes the resulting side information variable less informative than in Simulation 1, as depicted in Figure 2.
 This process is repeated three times, thus associating three side information variables with each hypothesis, which taken together, should be reasonably informative.
-We conducted 5 independent runs where in each run we drew the $n$ observations, $\beta$ coefficients and side information.
+We conducted 5 independent runs, where in each run we drew the $n$ observations, $\beta$ coefficients, and side information.
 5.1.5 Results
 We first compare RESET Ensemble with Adaptive Knockoff’s default method in R, AdaKO GAM, and the generic Knockoff Filter (KO).
-In Figure 3, we plot the power of each simulation using these methods as a function of the selected FDR thresholds of $\alpha\in\{0.05,0.1,0.2,0.3\}$.
+In Figure 3, for each simulation we plot the power of each method as a function of the selected FDR thresholds of $\alpha\in\{0.05,0.1,0.2,0.3\}$.
 The power is calculated as the proportion of relevant variables correctly discovered, averaged over the number of runs used in each simulation.
-Similarly, in Supplementary Figure S2, we plot the estimated FDR as the proportion of irrelevant variables in the discovery list, averaged over the runs.
+Similarly, in Supplementary Figure S2, we plot the estimated FDR: the proportion of irrelevant variables in the discovery list, averaged over the runs.
 Figure 3: RESET Ensemble vs.
@@
 A description of AdaKO GAM and KO can be found in Section 2.
-Unsurprisingly, RESET Ensemble uniformly dominates KO, demonstrating it successfully leverages the side information. At the same time, AdaKO GAM mostly improves on KO, with some exceptions.
+Unsurprisingly, RESET Ensemble uniformly dominates KO, demonstrating that it successfully leverages the side information. At the same time, AdaKO GAM mostly improves on KO, with some exceptions.
 In Simulation 1, we find that at higher FDR thresholds, RESET Ensemble is the preferred approach in terms of power.
 On the other hand, AdaKO GAM is more powerful than RESET at lower FDR thresholds, particularly in the case of $k=50$ and $\alpha=5\%$.
-This is not too surprising given the following two reasons.
+This is not too surprising for two reasons.
 First, RESET trains its classification algorithms on a relatively small positive set of false nulls.
-Second, RESET incurs a greater cost in terms of power at lower FDR thresholds if the classification algorithm overits and incorrectly scores a knockoff higher than it should.
+Second, RESET incurs a greater cost in terms of power at lower FDR thresholds if the classification algorithm overfits and incorrectly scores a knockoff higher than it should.
 Indeed, at an FDR threshold of $\alpha=5\%$, each knockoff may cost up to $402/\alpha=$ 40 discoveries.
 We find in Simulations 2 and 3 that RESET Ensemble is consistently more powerful than AdaKO GAM; in fact, AdaKO GAM appears to essentially have the same power as KO.
-The result in Simulation 3 is consistent with the findings of Ren and Candès in their analagous simulation.
-AdaKO GAM does not implement any smooth terms when the number of side information variables is two or more, and is therefore unable to leverage the side information if the true and false null hypotheses are not “easily” separable.
+The result in Simulation 3 is consistent with the findings of Ren and Candès in their analogous simulation.
+Indeed, AdaKO GAM does not implement any smooth terms when the number of side information variables is two or more, and is therefore unable to leverage the side information if the true and false null hypotheses are not ‘easily’ separable.
 Finally in Simulation 4, we see that RESET Ensemble is generally the most powerful, with essentially the same power as AdaKO GAM at $\alpha=5\%$.
@@
 Figure 4: Relative power and runtimes for each method.
-In (A), each boxplot describes the relative powers for each method, computed across the combination of simulations and FDR thresholds in Figure 3.
-(B) shows the log10 runtimes for each method applied at the 5% FDR level for 5 runs of Simulation 4.
+(A) Each boxplot describes the relative powers for each method, computed across the combination of simulations and FDR thresholds in Figure 3.
+(B) The log10 runtimes of each method applied at the 5% FDR level for 5 runs of Simulation 4.
 For readability, the points are jittered in the horizontal direction.
 We compared RESET Ensemble with AdaKO GAM because those are the respective default methods.
-We complement this comparision with Figure 4A where we summarize the performance of all RESET and Adaptive Knockoff methods.
+We complement this comparison with Figure 4A where we summarize the performance of all RESET and Adaptive Knockoff methods.
 Specifically, for each panel and FDR threshold in Figure 3, we determined the maximum power across all methods and computed the relative power of each method to this maximum.
@@
 On the other hand, RESET Ensemble’s performance is consistent across these simulations, and overall typically achieves the highest relative power when the results are aggregated.
-Interestingly, RESET NN seems to be be marginally more powerful than RESET Ensemble.
-While this may be the case, we still recommend RESET Ensemble as the default approach to alleviate the burden of selecting any singular classification algorithm.
+Interestingly, RESET NN seems to be marginally more powerful than RESET Ensemble.
+While this may be the case, we still recommend RESET Ensemble as the default approach to alleviate the burden of selecting any single classification algorithm.
 One advantage of RESET is its computational speed.
 Indeed, our implementation of RESET, as described in Section 4.3, has many steps that can be parallelized: the $r$ evaluations of each classification algorithm on the $K$ folds can all be done in parallel using multiple cores.
-In contrast, Adaptive Knockoffs must iteratively apply a classifiation algorithm for every nonzero scoring hypothesis in the candidate set $\mathcal{S}$ until the estimated FDR is $\leq\alpha$, or all hypotheses are revealed.
-Figure 4B demonstrates this advantage and shows the (log10) runtimes of RESET using 10-cores and Adaptive Knockoffs at the 5% FDR level across 5 runs of Simulation 4.
+In contrast, Adaptive Knockoffs must iteratively apply a classification algorithm for every nonzero scoring hypothesis in the candidate set $\mathcal{S}$ until the estimated FDR is $\leq\alpha$, or all hypotheses are revealed.
+Figure 4B demonstrates this advantage showing the (log10) runtimes of RESET (using 10 of a 20-cores CPU) and of Adaptive Knockoffs at the 5% FDR level across 5 runs of Simulation 4.
 We find that each version of RESET is significantly faster than each version of Adaptive Knockoffs.
 RESET Ensemble is around $10^{3}$ to $10^{4}$ times faster than AdaKO EM, $10^{2}$ to $10^{3}$ times faster than AdaKO RF, and about 3-4 times faster than AdaKO LR and AdaKO GAM.
-In scenarios where AdaKO EM and AdaKO RF are infeasible, but AdaKO GAM and AdaKO LR are unable to leverage the side information as in Simulations 3 and 4, RESET is arguably more suitable.
-We investigate such a scenario in the following application.
+In scenarios, as in Simulations 3 and 4, where AdaKO EM and AdaKO RF are infeasible, and AdaKO GAM and AdaKO LR are unable to leverage the side information RESET is clearly more suitable.
+We investigate such a commonly encountered scenario in the following application.
 5.2 Application to peptide detection
-In this section we compare RESET to Adaptive Knockoffs in the peptide detection context using mass spectrometry data and show that (a) the Adaptive Knockoff methods are often computationally infeasible, (b) feasibility aside, RESET is more powerful than Adaptive Knockoff’s default filter, AdaKO GAM, when controlling the FDR and (c) RESET is able to typically discover more peptides than the vanilla FDP-SD when controlling the FDP.
-We provide justification of Assumption 3 which is required for RESET’s theoretical guarantees in Section S.15.
+In this section we compare RESET to Adaptive Knockoffs in the peptide detection context using mass spectrometry data and show that (a) the Adaptive Knockoff methods are often computationally infeasible, (b) feasibility aside, RESET is more powerful than Adaptive Knockoff’s default filter, AdaKO GAM, when controlling the FDR, and (c) when controlling the FDP, RESET is able to typically discover more peptides than the vanilla FDP-SD.
+We provide a heuristic justification of Assumption 3, which is required for RESET’s theoretical guarantees, in Section S.15.
 5.2.1 HEK293 data
-We show that the application of Adaptive Knockoffs is often too slow to be used effectively in high throughput scenarios such as the peptide detection problem.
-We demonstrate this, using HEK293 data [9], a popular dataset used in benchmarking the development of new software [30, 49, 32] that we downloaded from the Proteomics Idenitification Database (PRIDE) [40].
+We show that the application of Adaptive Knockoffs is often too slow to be used effectively in scenarios such as the peptide detection problem that involve a large amount of data.
+We demonstrate this using HEK293 data [9], a popular dataset used in benchmarking the development of new software [30, 50, 32] that we downloaded from the Proteomics Identification Database (PRIDE) [41].
 After data preparation and searching the combined 24 HEK293 spectrum files, we determined the scores $Z_{i}$ for each target peptide and $\tilde{Z}_{i}$ for each decoy peptide using the XCorr score function (see Supplementary Section S.12 for details).
 Next, we competed each pair by recording the score $W_{i}=Z_{i}\vee\tilde{Z}_{i}$ and label $L_{i}=1$ if $Z_{i}>\tilde{Z}_{i}$ and $L_{i}=-1$ if $Z_{i}<\tilde{Z}_{i}$ (ties were randomly broken).
-In addition to the primary score associated with each peptide-spectrum match (PSM) is a set of features, such as the charge state or peptide length.
+In addition to its primary score we associate with each peptide-spectrum match (PSM) a set of features, such as the charge state or peptide length.
 Each winning peptide inherits these features, or side information $\mathbf{x}_{i}$, from its maximally associated PSM, allowing us to define the desired triples $(L,W,\mathbf{x})$.
@@
 Due to the aforementioned runtime problem, we had to use small datasets when evaluating the power of Adaptive Knockoffs in the peptide detection context.
-Specifically, we focused on thirteen spectrum files (each with $<$ 33K PSMs) from the PRIDE-20 dataset that we recently used in [18].
-In the following results, we only considered searching the spectrum files once against a single target-decoy database since unlike RESET we can only feasibly apply a single application of Adaptive Knockoffs.
-Data preparation, search settings and details on how each triple $(L,W,\mathbf{x})$ is obtained are given in Supplementary Section S.14 along with the list of side information used in Supplementary Table S5.
+Specifically, we focused on thirteen spectrum files (each with $<$ 33K PSMs) from the twenty PRIDE-20 datasets that we recently used in [18].
+In the following results, we only considered searching the spectrum files once against a single target-decoy database, since unlike RESET we can only feasibly apply a single application of Adaptive Knockoffs.
+Data preparation, search settings and the detailed definition of the triples $(L,W,\mathbf{x})$ are given in Supplementary Section S.14 along with the list of side information used in Supplementary Table S5.555The side information here slightly differs from above due to different search settings.
 Figure 5: Comparison of RESET Ensemble and AdaKO GAM.
 For each of the 13 PRIDE datasets, we computed the average number of discoveries using 10 applications of RESET Ensemble, varying RESET’s internal seed.
-The figure shows a boxplot of those average RESET Ensemble discoveries divided by the number of discoveries using AdaKO GAM (left boxplot) and TDC (right boxplot) using the 13 PRIDE-20 datasets (there are total of 13 points for each boxplot).
+The figure shows a boxplot of those RESET Ensemble averages divided by the number of discoveries using AdaKO GAM (left boxplot) and TDC (right boxplot) using the 13 PRIDE-20 datasets (there are total of 13 points for each boxplot).
 Figure 5 shows the average number of discoveries reported using 10 applications of RESET Ensemble, each applied at 1% FDR to the same dataset, divided by the number of discoveries reported by AdaKO GAM (left boxplot) and TDC (right boxplot).
 In each of the 10 applications of RESET, we varied the internal seed which randomly splits the decoys into their training and estimating sets.
-Clearly, RESET Ensemble is a more powerful alternative than AdaKO GAM and TDC.
-Interestingly, the relative ratios in the number of discoveries between AdaKO GAM and TDC suggests that AdaKO GAM is reporting less discoveries than the generic TDC.
-In a second comparision, we looked at the performance of RESET Ensemble versus AdaKO RF because in this setup it was the most powerful implementation of Adaptive Knockoffs.
+Clearly, RESET Ensemble is overall more powerful than AdaKO GAM and TDC.
+Interestingly, the relative ratios of the number of discoveries between AdaKO GAM and TDC suggest that AdaKO GAM is reporting fewer discoveries than the generic TDC.
+In a second comparison, we looked at the performance of RESET Ensemble versus AdaKO RF because in this setup it was the most powerful implementation of Adaptive Knockoffs.
 In Figure 6A, the left boxplot displays the average number of discoveries by RESET Ensemble, divided by the number of discoveries by AdaKO RF.
@@
 We also assessed the computational times of AdaKO RF in Figure 6B, using both the actual and estimated runtimes as outlined in Section S.13.
-Reassuringly, our estimated times were less than the actual times taken by each method to complete thus validating our estimated times in the previous section.
+Reassuringly, our estimated times were less than the actual times taken by each method to complete, thus validating the conservative nature of our estimated times in the previous section.
 Although we find that, when analyzing these smaller spectrum sets, AdaKO RF is the most powerful method of Adaptive Knockoffs, and is competitive to RESET Ensemble power-wise, it is again clear that AdaKO RF is generally too slow with some runtimes exceeding multiple days even for these small datasets.
@@
 5.2.3 FDP control
-Lastly, we evaluated RESET Ensemble’s FDP control using the HEK293 and PRIDE-20 spectrum files from the previous sections.
+Lastly, we evaluated RESET Ensemble’s FDP control using the HEK293 and all twenty PRIDE-20 spectrum files from the previous sections.
 As in Section 5.2.1, the joint HEK293 spectrum files were searched against 5 distinct target-decoy databases, where each target-decoy database was generated by concatenating the target database to a decoy database consisting of shuffled target peptides.
@@
 The averages are taken over 5 applications of each method, once for each of the 5 combined target-decoy databases.
-Analyzing the HEK293 results first, RESET Ensemble is clearly more powerful with an approximately 9-10% increase in the average number of discoveries over FDP-SD at all confidence levels.
-Analyzing the PRIDE-20 datasets, while RESET Ensemble appears to be generally more powerful, in a few datasets, it fails to make any discoveries at high confidence levels.
+Analyzing the HEK293 results first, RESET Ensemble is clearly more powerful, with an approximately 9–10% increase in the average number of discoveries over FDP-SD at all confidence levels.
+Analyzing the PRIDE-20 datasets, while RESET Ensemble appears to be generally more powerful, in a few datasets it fails to make any discoveries at high confidence levels.
 Upon a closer look, the numbers of discoveries made by FDP-SD in these datasets are relatively small, with the largest average number of discoveries among these datasets being 417.
-When the number of discoveries is expected to be small and the confidence parameter is high, RESET may yield less power, even if the side information is reasonably informative.
-We dwell on this fact in the Discussion.
-Regardless, more often than not RESET Ensemble produces more discoveries and if multiple spectrum files are analyzed jointly, as we do with the HEK293 spectrum files, then the risk of having too few discoveries is reduced.
+When the number of discoveries is expected to be small and the confidence parameter is high, RESET may yield less power, even if the side information is reasonably informative — a phenomenon we dwell on in the Discussion.
+Regardless, more often than not RESET Ensemble produces more discoveries, and if multiple spectrum files are analyzed jointly, as we do with the HEK293 spectrum files, then the risk of having too few discoveries is reduced.
 Figure 7: Comparison of RESET Ensemble and FDP-SD.
-The left panel shows boxplots of the average number of discoveries using RESET divided by the average number of discoveries using FDP-SD for each of the PRIDE-20 datasets at a 1% FDP threshold and varying confidence parameters.
+The left panel shows boxplots of the average number of discoveries using RESET divided by the average number of discoveries using FDP-SD for each of the twenty PRIDE-20 datasets at a 1% FDP threshold and varying confidence parameters.
 The averages are taken over 10 applications, one for each of the 10 target-decoy databases used for each PRIDE-20 dataset.
@@
 The results above consider a canonical ‘narrow’ search.
-Specfically, during the searching phase, each database peptide and spectrum are only allowed to match if the mass of the database peptide and the mass of the sample peptide generating the spectrum, the ‘spectrum mass’, differ by a small amount.
+Specifically, during the searching phase, each database peptide and spectrum are only allowed to match if the mass of the database peptide and the observed mass of the sample peptide generating the spectrum, the ‘spectrum mass’, differ by a small amount.
 Recently, ‘open’ searching, where each spectrum and peptide can match even if their masses differ by hundreds of Daltons, has become a popular method for detecting peptides [30].
-Such a search allows for the detection of peptides whose mass have been altered by post-translational modifications (PTMs).
+Such a search allows for the detection of peptides whose masses have been altered by post-translational modifications (PTMs).
 We provide further detail in Supplementary Section S.14.
-In this setting, side information variables such as dm, which records the mass difference between the peptide and the spectrum, could be highly informative given that some differences will correspond to masses of common PTMs.
-Thus, it is not surprising that RESET, which can take advantage of such side information, does even better than FDP-SD in this context.
-Specifically, Figure 8 shows a greater improvement over FDP-SD across all confidence levels when analyzing open searching results.
+In this setting, side information variables such as dm, which records the mass difference between the peptide and the spectrum, can be highly informative given that some differences will correspond to masses of common PTMs.
+Thus, it is not surprising that RESET, which can take advantage of such side information, improves on FDP-SD even further across all confidence levels in this context (Figure 8).
 Figure 8: Comparison of RESET ensemble and FDP-SD using the recently popular ‘open’ search.
@@
 As in the competition setting, each setup is designed to satisfy Assumption 3, which is required for our theoretical guarantees.
-Details of each method used is given in Supplementary Section S.6-S.10.
+Details of each method used are given in Supplementary Sections S.6S.10.
 6.1.1 Simulation 5: p-value based testing with geometric side information
@@
 A total of $m=2500$ hypotheses are used.
-Similar to Simulation 3 of Section 5.1.3, each hypothesis’ side information is defined as unique pair of values $\mathbf{x}_{i}=\left(\mathbf{x}_{i1},\mathbf{x}_{i2}\right)$ from $50\times 50$ equally spaced values in the square $\left[-100,100\right]\times\left[-100,100\right]$.
+Similar to Simulation 3 of Section 5.1.3, each hypothesis’ side information is defined as a unique pair of values $\mathbf{x}_{i}=\left(\mathbf{x}_{i1},\mathbf{x}_{i2}\right)$ from $50\times 50$ equally spaced values in the square $\left[-100,100\right]\times\left[-100,100\right]$.
 The false null hypotheses correspond to certain regions in the square.
-They considered three scenarios: (a) when the false nulls form a circle in the middle of the square, (b) when the false nulls form a circle in the top right of the square and (c) when the false nulls form an ellipse.
-Each of these scenarios is depicted in Figure 9.
-We used 100 independent runs where in each run, we randomly draw the 2500 p-values as described.
-Figure 9: Location of false null hypotheses according to the pair of values $(x,y)$ in the square $\left[-100,100\right]\times\left[-100,100\right]$. The same picture is given in [33].
+They considered three scenarios: (a) when the false nulls form a circle in the middle of the square, (b) when the false nulls form a circle in the top right of the square and (c) when the false nulls form an ellipse (Figure 9).
+We used 100 independent runs, where in each run we randomly draw the 2500 p-values as described.
+Figure 9: Location of false null hypotheses according to the pair of values $\left(\mathbf{x}_{i1},\mathbf{x}_{i2}\right)$ in the square $\left[-100,100\right]\times\left[-100,100\right]$.
+The same picture is given in [33].
 Figure 10: RESET vs.
@@
 Each color and shape combination corresponds to a unique method.
-The three columns correspond to the hypotheses’ relationship to the side-information according to Figure 9.
+The three columns correspond to the hypotheses’ relationship to the side information according to Figure 9.
 The horizontal segments in the second row of panels mark the corresponding FDR thresholds on the y-axis.
@@
 Figure 10 shows the results of RESET compared with several p-value based methods that we reviewed in Section 3.1.
-We find that in the first two simulations, (a) and (b), RESET Ensemble is roughly equal best in terms of power at the 5%, 10% and 20% FDR levels, along with AdaPT/AdaPTg GAM.
+We find that in the first two simulations, (a) and (b), RESET Ensemble is roughly equal to the best method in terms of power at the 5%, 10% and 20% FDR levels, along with AdaPT/AdaPTg GAM.
 However, in (c), clearly AdaPT/AdaPTg GAM are more powerful than the RESET methods, with RESET Ensemble coming arguably as third in the power ranking at the 5%, 10%, and 20% FDR levels.
@@
 This is possibly explained by the fact that AdaFDR only guarantees asymptotic FDP control.
-Consisitently, Chao and Fithian [8] reported inflated estimated FDRs at high thresholds in simulation studies for AdaFDR.
+Consistently, Chao and Fithian [8] reported inflated estimated FDRs at high thresholds in simulation studies for AdaFDR.
 6.1.2 Simulation 6: p-value based testing with 100-dimensional side information
-In Simulation 2 of Lei and Fithian [33], the side information for $m=2000$ hypotheses consists of 100 i.i.d draws from $U[0,1]$.
+In Simulation 2 of Lei and Fithian [33], the side information for $m=2000$ hypotheses consists of 100 i.i.d.
+draws from $U[0,1]$.
 This setup simulates ‘noisy’ side information since only the first 2 of the 100 draws are used in distinguishing between the true and false null hypotheses.
-Each p-value $p_{i}$ is drawn i.i.d from a two-group beta mixture model, where the true nulls have a uniform distribution and the false nulls have an alternative beta distribution.
+Each p-value $p_{i}$ is drawn i.i.d.
+from a two-group beta mixture model, where the true nulls have a uniform distribution and the false nulls have an alternative beta distribution.
 Specifically, the beta mixture model is given by:
@@
 where $\mathbf{x}_{i}\in[0,1]^{100}$ denotes the side information of the $i$th hypothesis, $\pi(\mathbf{x}_{i})$ denotes proportion of the false nulls in the mixture, $h$ denotes the false null distribution of $Beta(1/\mu(\mathbf{x}_{i}),1)$.
-The $\pi(\mathbf{x}_{i})$’s and $\mu(\mathbf{x}_{i})$’s are determined by the following relationships with the 100-dimesional side information:
+The $\pi(\mathbf{x}_{i})$’s and $\mu(\mathbf{x}_{i})$’s are determined by the following relationships with the 100-dimensional side information:
 $\log\left(\frac{\pi}{1-\pi}\right)=\theta_{0}+\mathbf{x}_{i}^{T}\theta,\quad\mu_{i}=\max\{\mathbf{x}_{i}^{T}\beta,1\},$
@@
 There were a couple of tools we had to omit from the analysis of this data.
-Suprisingly, we found that AdaFDR was too slow in this simuation.
+Surprisingly, we found that AdaFDR was too slow in this simulation.
 Similarly, we were not able to apply AdaPT GAM or AdaPTg GAM as we did in the previous simulation, since it would have required using a smoothing spline on all 100 side-information variables, which is impractical.
@@
 Notably, the performance of AdaPT in these last two simulation setups is highly dependent on the choice of implementation.
-In Simulation 5, AdaPT GAM significantly outperforms AdaPT GLMnet while in Simulation 6, AdaPT GLMnet is superior (and AdaPT GAM is not practical).
+In Simulation 5, AdaPT GAM significantly outperforms AdaPT GLMnet while in Simulation 6, AdaPT GLMnet is superior and AdaPT GAM is not even practical.
 In contrast, RESET Ensemble selects from a collection of flexible machine learning models to adapt itself to the problem at hand.
 6.2 Real data experiments with p-values
-We next evaluate RESET and related methods on a collection of publically available datasets with p-values that were investigated by Zhang et al.
+We next evaluate RESET and related methods on a collection of publicly available datasets with p-values that were investigated by Zhang et al.
 [51] when introducing AdaFDR and by Lei and Fithian [33] when introducing AdaPT.
 Since AdaFDR and AdaPT operate under essentially the same assumptions as RESET’s, we expect these datasets to therefore satisfy Assumption 3 with the exception of two datasets, which we discuss at the end of this section.
-We provide a brief summary of each dataset below:
+We provide a brief summary of each dataset below.
   • Three RNA-seq datasets (Airway [22], Bottomly [4], Pasilla [6]): The goal is to identify genes that are associated with varying gene expression levels in response to a change of conditions.
@@
   • A proteomics dataset (Proteomics [14]): The goal is to identify proteins that have different protein abundances when treated with rapamycin versus DMSO in yeast cells.
-Each protein consists of side information equal to the natural log of the number of peptides “quantified” across all samples, i.e., the number of peptides belonging to the protein with their abundances measured.
+Each protein consists of side information equal to the natural log of the number of peptides ‘quantified’ across all samples, i.e., the number of peptides belonging to the protein with their abundances measured.
   • Two microbiome datasets (Microbiome Enigma Ph and Microbiome Enigma Al [45, 31]): The goal is to identify microbiome organisms, recorded as OTUs (Operational Taxonomic Units), that are associated with certain conditions (pH and Al).
 Each hypothesis (OTU) is equipped with a two-dimensional side information — the ubiquity, which is defined as the percentage of samples detected with the OTU, and the mean nonzero abundance, which is defined as the average abundance of each OTU across the samples in which it was detected.
-  • fMRI datasets (Auditory and Imagination) [47]: the goal is to identify voxels that are activated in response to different stimuli.
+  • Two fMRI datasets (Auditory and Imagination) [48]: The goal is to identify voxels that are activated in response to different stimuli.
 In the first dataset, a single individual received auditory stimulus and in the second dataset, a single individual was instructed to imagine playing tennis.
 Each hypothesis (voxel) is equipped with four dimensional side information: the spatial position of the voxel (in three dimensions) and a categorical variable with the Brodmann label [5], which is used to delineate different areas of the brain with different functions.
-  • Two eQTL datasets (Subcutaneous and Omentum [11, 10]): the goal is to identify single nucleotide polymorphisms (SNPs) that are associated with gene expression levels.
-Such an association is referred to as an expressive quantitative trait loci (eQTL).
-Each hypothesis (SNP) is equipped with the following four side information variables: (1) the distance between the SNP and the gene trascription’s start site (TSS), (2) the log gene expression level, (3) the alternative allele frequency (AAF) of the SNP and (4) the chromatin state of the SNP.
-  • Gene-drug response data (Estrogen [12]): the goal is to identify genes that are associated with a change in gene expression levels in response to a low-dosage estrogen treatment applied to breast cancer cells.
+  • Two eQTL datasets (Subcutaneous and Omentum [11, 10]): The goal is to identify single nucleotide polymorphisms (SNPs) that are associated with changes in gene expression levels.
+Such an association is referred to as an expression quantitative trait locus (eQTL).
+Each hypothesis (SNP) is equipped with the following four side information variables: (1) the distance between the SNP and the gene transcription’s start site (TSS), (2) the log gene expression level, (3) the alternative allele frequency (AAF) of the SNP and (4) the chromatin state of the SNP.
+  • Gene-drug response data (Estrogen [12, 36]): The goal is to identify genes that are differentially expressed in response to a low-dosage estrogen treatment applied to breast cancer cells.
 Each hypothesis (gene) is assigned a ‘rank’ that was generated using an analysis of the gene expression levels in response to (1) a high-dosage estrogen treatment and (2) a medium-dosage estrogen treatment.
@@
 Thus, we can analyze the data twice, once for each side information variable based on (1) and (2).
-Figure 12: Comparison of RESET and other methods using a collection of publically available datasets.
+Figure 12: Comparison of RESET and other methods using a collection of publicly available datasets.
 Each method was evaluated at the FDR thresholds of 1%, 5%, 10%, and 20%.
@@
 Then we calculated the median rank for each method.
-RESET Ensemble and AdaPTg-GMM were tied with the highest median ranks.
+RESET Ensemble and AdaPTg-GMM were tied with the highest median rank.
+Moreover, for each dataset and each FDR threshold in Figure 12, we computed the maximum number of discoveries across all methods and determined the relative power of each method to this maximum.
+The boxplots in Figure 13 display the relative powers for each method.
+Visually, RESET Ensemble and AdaPTg-GMM have the highest medians, with RESET Ensemble at 95.9% and AdaPTg-GMM at 94.5% with RESET Ensemble’s boxplot edging higher overall.
+Figure 13: Relative power of each method. Each boxplot describes the relative power of each method computed for each dataset and FDR threshold in Figure 12.
 There are a couple of implementation details we would like to highlight here (for full details, see Supplementary Sections S.6-S.10).
@@
 In this case, it is sensible to define asymmetric target and decoy regions, $[0,0.3)$ and $(0.3,0.9]$ respectively as Chao and Fithian [8] did in their application of AdaPTg-GMM to this dataset.
-We also point out that we implemented AdaPTg-GMM and ZAP using the p-value information, and not the test statistic information (see Section 3.1 for background on these methods).
+We also point out that we applied AdaPTg-GMM and ZAP using the p-value information rather than the test statistic information (see Section 3.1 for background on these methods).
 To do so, AdaPTg-GMM allows the user to directly input p-values instead of z-statistics.
-For ZAP, we converted the p-value information into z-values by defining $z_{i}=\pm\Phi^{-1}(p_{i}/2)$ of each p-value, $p_{i}$, where the signs are chosen i.i.d uniformly.
+For ZAP, we converted the p-value information into z-values by defining $z_{i}=\pm\Phi^{-1}(p_{i}/2)$ of each p-value, $p_{i}$, where the signs are chosen i.i.d.
+uniformly.
 In other words, we input plausible z-values that correspond to the same p-value information used by the other tools.
 As we mentioned in Section 3, our implementation of RESET focuses on the use of p-values, and we leave the extension to test statistic information as part of future work.
-To evaulate the computation times for each method, we used the two fMRI datasets, which uses the most number of side information variables and has a large number of hypotheses.
-Figure 13(A-B) shows the computation times for each considered method where the fastest are the collection of RESET methods, ZAP and AdaFDR.
+To evaluate the computation times for each method, we used the two fMRI datasets, which uses the largest number of side information variables and has a large number of hypotheses.
+Figure 14(A-B) shows the computation times for each considered method, where the fastest are the collection of RESET methods, ZAP and AdaFDR.
 RESET NN, RESET GAM and RESET RF are slightly faster than our ensemble approach which considers all of them.
@@
 In all panels, we evaluated each method at FDR thresholds of: 1%, 5%, 10%, 20%.
-Noticeably, we have excluded the analysis of the two eQTL studies from Figure 12.
+Notably, we have excluded the analysis of the two eQTL studies from Figure 12.
 This is because this dataset exhibits a high degree of p-value dependency, even among the true null hypotheses and thus violates Assumption 3 which guarantees our type-1 error control.
 Indeed, this phenomenon is due to linkage disequilibrium in which SNPs behave dependently.
-As a consequence, the dataset exhibits virtually identical copies of the same p-value, with similar side information, even if the p-value is large suggesting a low confident hypothesis.
+As a consequence, the dataset exhibits virtually identical copies of the same p-value, with similar side information, even if the p-value is large, suggesting a true null hypothesis.
 Figure 14(C-D) shows that RESET Ensemble reports substantially more discoveries on these eQTL datasets than all the non-RESET variants.
-This is because the underlying random forest in the ensemble is able to identify the same ‘copies’ of low confident p-values in both the training and test sets, and if these copies are mutually positively labelled, which many are, then they will be rescored higher by our RESET algorithm and ultimately discovered.
+This is because many estimating and training decoys are virtually the same and the random forest is sensitive to this leakage of information.
 Evidently, RESET NN is more robust to this problem: the number of discoveries is in the same ‘ballpark’ as other methods.
@@
 6.3 FDP control with p-values
-Figure 14: Comparison of RESET and other methods with FDP control using a collection of publically available data.
+Figure 15: Comparison of RESET and other methods with FDP control using a collection of publicly available data.
 Each panel plots the number of discoveries obtained by RESET Ensemble, FDP-SD and GR-SD at a range of FDR thresholds (1%, 5%, 10%, 20%) with varying confidence $1-\gamma$ (0.5, 0.8, 0.9).
@@
 Here we define the following asymmetric target and decoy regions, $[0,0.3)$ and $(0.3,0.9]$ respectively, for all datasets.
-The rationale stems from our later analysis in the discussion, as well as our comments at the end of Section 5.2.3, which point out the fact that when the ratio of estimating decoys to true null targets is small, the subsequent FDP-SD step in RESET might incur a loss of power.
+The rationale stems from our later analysis in the Discussion, as well as our comments at the end of Section 5.2.3, which point out that when the ratio of estimating decoys to true null targets is small, the subsequent FDP-SD step in RESET might incur a loss of power.
 Hence, to rectify this, we essentially double the number of estimating decoys under consideration by defining a decoy region that is twice as large as the corresponding target region.
@@
 Figure 15 displays the number of discoveries obtained by each method on the collection of real datasets we considered in the previous section, with varying FDR thresholds (1%, 5%, 10%, 20%) and confidence levels (50%, 80%, 90%).
-RESET Ensemble is much more powerful than the alternative non-side information approaches with few exceptions.
+RESET Ensemble is much more powerful than the alternative non-side information approaches with a few exceptions.
 7 Discussion
 We introduced RESET, a flexible wrapper for any semi-supervised learning algorithm.
-Alongside, we provide an implementation of RESET, called RESET Ensemble, that (1) is power-wise competitive with recently developed methods for competition as well as p-value based multiple testing problems with side information, (2) is fast compared to most tools, and (3) is able to achieve finite FDR/FDP control.
-In evaluating the claim of (1), we highlight the fact that competing tools such as the variants of AdaPT or Adaptive Knockoffs, require an initial selection of one of those variants.
-Making a selection after observing the discovery lists of each tool will clearly cheat any guarantee of FDR control.
-Thus, we compared RESET Ensemble to the default competing method if one was available, which was the case of Adaptive Knockoff’s GAM filter, or by the variants separately, which was the case of AdaPT.
-Based on our comparisons, we conclude that RESET Ensemble is able to achieve comparable or slightly greater power to many of these tools.
-RESET Ensemble’s intended generality alleviates the user from having to choose the “right” method to begin with.
-In the case of (2), we found that our collection of RESET methods were a faster alternative to all competing methods with finite FDR control in our considered experiments.
-The only case in which RESET was surpassed in terms of speed was ZAP, which demonstrated marginally faster runtimes than RESET Ensemble in one of the fMRI datasets (Figure 13B).
-The most striking difference in runtimes was in the peptide detection context using mass spectrometry data, in which the variants of Adaptive Knockoffs were generally too slow to be used as tools for this application.
-Finally in addressing (3), we point out that RESET’s compatibility with FDR or FDP control is unique.
+Alongside, we provide an implementation of RESET, called RESET Ensemble, that is power-wise competitive with recently developed methods for competition as well as p-value based multiple testing problems with side information, fast compared to most tools, and flexible as it allows finite-sample of either FDR or FDP control.
+In terms of power, we highlight the fact that competing tools, such as the variants of AdaPT or Adaptive Knockoffs, require an initial selection of one of those variants.
+Making a selection after observing the discovery lists of each tool will generally compromise its FDR control.
+Thus, we compared RESET Ensemble to the default competing method if one was available, which was the case of Adaptive Knockoff’s GAM filter, or separately to each variant, which was the case of AdaPT.
+Based on our comparisons, we conclude that RESET Ensemble is able to achieve comparable or slightly greater power than many of these tools while freeing the user from having to choose the ‘right’ method to begin with.
+As far as computational runtime, we found that our collection of RESET methods provide a faster alternative to all competing methods with finite FDR control in our considered experiments.
+The only case in which RESET was surpassed in terms of speed was ZAP-asymp, which demonstrated marginally faster runtimes than RESET Ensemble in one of the fMRI datasets (Figure 14B).
+However, ZAP-asymp’s fast runtimes come at the cost of asymptotic FDR control.
+The most striking difference in runtimes was in the peptide detection context using mass spectrometry data, in which all variants of Adaptive Knockoffs were generally too slow to be practical.
+Finally, we point out that RESET’s compatibility with FDR or FDP control is unique.
 In the competition setting, we note that it would be challenging to combine Adaptive Knockoffs with FDP-SD because FDP-SD is a stepdown procedure while Adaptive Knockoffs is essentially a step-up procedure.
-Katsevich et al.
-[27] provide a general method for adaptive ordering using their method for simultaneous FDP bounds.
+Katsevich and Ramdas [27] provide a general method for adaptive ordering using their method for simultaneous FDP bounds.
 However, (a) they do not provide an exact implementation of this, only a theoretical discussion, and (b) their approach is conservative since it controls the FDP at every cutoff (which allows the analyst to explore the nested discovery lists without breaking FDP control).
 To the best of our knowledge, in the competition framework with side information, RESET is the only method of reordering that may be flexibly used with FDP-SD without the additional cost of (b).
-Similairly, we know of no method that allows finite sample FDP control in the p-value setting with side information either.
+Similarly, we know of no method that allows finite sample FDP control in the p-value setting with side information.
 On the topic of RESET’s FDP control, we point out the following limitation.
-When the number of discoveries is relatively large, RESET yields more power over the generic FDP-SD.
-However, when the number of discoveries is small, or the confidence is too high, RESET may yield less power, even if the side information is reasonably informative.
-To illustrate this, we computed the bounds, $\delta_{i}$ on the number of decoy wins in the top $i$ scores from Algorithm S2, that are used to determine the reported list of discoveries in FDP-SD (with $c=1/2$) and RESET (using the default $c=2/3$).
+When the number of discoveries is small, e.g., when the confidence is too high, RESET may yield less power, even if the side information is reasonably informative.
+To illustrate this phenomenon, we computed the bounds, $\delta_{i}$ on the number of decoy wins in the top $i$ scores from Algorithm S2, which are used to determine the reported list of discoveries in FDP-SD (with $c=1/2$) and RESET (using the default $c=2/3$).
 We considered a confidence level of $1-\gamma=90\%$ and an FDP threshold of $\alpha=1\%$.
@@
 On the right y-axis, we plot the ratio of these bounds (in black).
-The left panel looks at indices between 1K to 2K, while the right panel looks at indices from 10K to 20K.
-An index corresponds to the number of top scoring hypotheeses (regardless of their labels).
+The left panel looks at indices between 501 to 2K, while the right panel looks at indices from 10K to 20K.
+An index corresponds to the number of top scoring hypotheses (regardless of their labels).
 Figure 16 shows the bounds for FDP-SD, and double the bounds for RESET, along with the ratio of the two at a range of different indices.
-We can see that for smaller indices, this ratio is much smaller, initially at 50% at an index of 1K, while at 10K this ratio is much higher at 94%.
-If the order of the hypotheses is the same, RESET’s smaller bounds imply it will report fewer discoveries because RESET employs FDP-SD to control the FDP by searching for the largest index $i$ s.t.
-$D_{i_{0}}\leq\delta_{i},\dots,D_{i}\leq\delta_{i}$, and reports the top $i$ scoring positively labelled hypotheses.
-In other words, if RESET is to report roughly the same number of discoveries as FDP-SD, RESET needs to rearrange the targets and estimating decoys successfully enough so as to make up for a 50% difference in the bounds when considering the top 1K scores, while only 6% when considering the top 10K.
-If the side information is not rich enough, it is possible RESET will report few discoveries than FDP-SD in these cases.
-We intend on investigating this as part of future work.
-Figure 16: Variability in the number of discoveries in the Airway data and in Simulation 5 In (A) We used a 10% FDR threshold and varied the internal seed of RESET Ensemble over 100 applications.
-In (B) we used a 10% FDR threshold and compared the variability in RESET’s and AdaPT’s power.
+We can see that for smaller indices, this ratio is much smaller, initially at 0% at an index of 501 and 50% at an index of 1K, while at 10K this ratio is much higher at 94%.
+If the order of the hypotheses is the same, then RESET’s smaller bounds imply it will report fewer discoveries because RESET employs FDP-SD to control the FDP by searching for the largest index $i$ s.t.
+$D_{i_{0}}\leq\delta_{i},\dots,D_{i}\leq\delta_{i}$, and reports the top $i$ scoring positively labeled hypotheses.
+In other words, if RESET is to report roughly the same number of discoveries as FDP-SD, then RESET needs to rearrange the targets and estimating decoys successfully enough so as to make up for a 50% difference in the bounds when considering the top 1K scores, while only 6% when considering the top 10K.
+If the side information is not rich enough, then it is possible RESET will report few discoveries than FDP-SD in these cases.
+We intend to investigate this as part of future work.
+Figure 17: Variability in the number of discoveries.
+in the Airway data and in Simulation 5 (A) Histogram of the number of RESET Ensemble discoveries applied 100 times to the Airway data with a 10% FDR threshold and varying its internal seed.
+(B) Boxplots of RESET’s and AdaPT’s power in 100 applications of the ‘circle in the middle’ setup of Simulation 5 using a 10% FDR threshold.
 There are two outliers in RESET’s case (which occurred when the signal was considerably low).
@@
 While this is problematic to an extent, keep in mind that the data distribution is random, and the added variance from RESET may be marginal at times.
-Indeed, in Figure 16(B), we find that in the case of Simulation 5 from Section 6.1.1, RESET is overall arguably less variable than AdaPT’s even though AdaPT does not employ any internal randomization.
-In practice, users may get around RESET’s randomization by setting the RNG as a function of the input data and parameters.
+Indeed, in Figure 17(B), we find that in the case of Simulation 5 from Section 6.1.1, RESET is overall arguably less variable than AdaPT, even though AdaPT does not employ any internal randomization.
+In practice, users may get around RESET’s randomization by setting the random number generator to be a function of the input data and parameters.
 We plan on adding this feature to a future update.
@@
 There is more than one way of generalizing RESET to accommodate a test statistic $z_{i}$ for each hypothesis $H_{i}$.
-Moreover there are several ‘types’ of interactions between the side information $\mathbf{x}_{i}$ and the test statistic $z_{i}$ which are not preserved when switching to p-values, as outlined by Leung and Sun [35].
+Moreover, there are several ‘types’ of interactions between the side information $\mathbf{x}_{i}$ and the test statistic $z_{i}$ which are not preserved when switching to p-values, as outlined by Leung and Sun [35].
 Thus it is not clear that RESET Ensemble will be equally applicable in this setting without substantial modifications.
-Appendix A Supplement
+Supplement
 S.1 Algorithms
@@
 3:sort the paired $(L_{i},W_{i})$ in decreasing order of $W_{i}$triangleright ties are randomly broken
-4:$D_{k}\leftarrow\#\{i\leq k:L_{i}=-1\}$triangleright number of decoy wins in top $k$ scores
-5:$T_{k}\leftarrow\#\{i\leq k:L_{i}=1\}$triangleright number of target/feate wins in top $k$ scores
+4:$D_{k}\leftarrow\#\{i\leq k:L_{i}=-1\}$triangleright number of decoy/knockoffs wins in top $k$ scores
+5:$T_{k}\leftarrow\#\{i\leq k:L_{i}=1\}$triangleright number of target/original feature wins in top $k$ scores
 6:if SeqStep then
@@
 4:$\tilde{L}_{i}\leftarrow-1$ for $i\in I$triangleright training decoys
-5:$\tilde{L}_{i}\leftarrow 1$ for $i\in J:=I^{c}$triangleright pseudo targets
+5:$\tilde{L}_{i}\leftarrow 1$ for $i\in J:=I^{c}$triangleright pseudo-targets
 6:$(\tilde{W}_{i})_{i=1}^{m}\leftarrow f\left((W_{i},U_{i},\tilde{L}_{i})_{i=1}^{m}\right)$triangleright where $f$ is any machine learning model
@@
 S.2 Proof of Lemma 1 in the main text
+We first remind the reader of the Lemma’s statement.
+See 3
 See 1
-Proof.
+Proof of Lemma 4.1.
 Let $W$ be all the scores, $\mathbf{x}$ be all the side information, and $L_{-i}$ be all the other labels that are not $L_{i}$. Then for a true null p-value, $p_{i}$:
@@
 $=\frac{a}{a+b_{2}-b_{1}},$
-where the first line follows since the denominator sums to 1, and the second line clearly follows from the non-decreasing property.
-For a more formal argument for the second line, note that the non-decreasing property implies that:
+where the first line follows since the denominator sums to 1, and the second line clearly follows from the non-decreasing property as explained next.
+First, note that the non-decreasing density property implies that:
 $\mathbb{P}(p_{i}\in B)\leq\frac{a}{b_{2}-b_{1}}\cdot\mathbb{P}(p_{i}\in q(B)),$
-where $B\in[0,a)$, $p_{i}$ is a true null p-value and $q$ denotes the inverse transformation of the mirrored p-values that map p-values in $(b_{1},b_{2}]$ onto $[0,a)$, that is $q(p):=b_{2}-\frac{b_{2}-b_{1}}{a}\cdot p$.
-Then consider for every measurable-$A$:
+where $B\subseteq[0,a)$ is measurable, $p_{i}$ is a true null p-value and $q$ denotes the inverse transformation of the mirrored p-values that map p-values in $(b_{1},b_{2}]$ onto $[0,a)$, that is $q(p):=b_{2}-\frac{b_{2}-b_{1}}{a}\cdot p$.
+Let $A\subseteq\mathbb{R}$ be measurable and let $B\subseteq[0,a)$ be such that $\{W_{i}\in A\}=\{p_{i}\in B\}\cup\{p_{i}\in q(B)\}$. Then,
 […]
-where $B$ is the corresponding region such that $\{W_{i}\in A\}=\{p_{i}\in B\}\cup\{p_{i}\in q(B)\}$. It follows that:
+$\leq\frac{a}{b_{2}-b_{1}}\mathbb{P}(p_{i}\in q(B))$
+[…]
+[…]
+It follows that:
 $\mathbb{P}(p_{i}\in[0,a)\mid W,\mathbf{x},L_{-i})$ $=\mathbb{P}(p_{i}\in[0,a)\mid W_{i})$
-$\leq\frac{a}{b_{2}-b_{1}}\mathbb{P}(p_{i}\in(b_{1},b_{2}]\mid W_{i})=\frac{a}{b_{2}-b_{1}}\mathbb{P}(p_{i}\in(b_{1},b_{2}]\mid W,\mathbf{x},L_{-i})$
+$\leq\frac{a}{b_{2}-b_{1}}\mathbb{P}(p_{i}\in(b_{1},b_{2}]\mid W_{i})=\frac{a}{b_{2}-b_{1}}\mathbb{P}(p_{i}\in(b_{1},b_{2}]\mid W,\mathbf{x},L_{-i}).$
 S.3 RESET controls the FDR or FDP
-See 3
-In this section we prove that given Assumption 3 RESET controls the FDR or FDP.
+In this section we prove that given Assumption 3 RESET controls the FDR or the FDP.
 We use the target-decoy terminology to refer to a positive label $L_{i}=1$ as a target win or simply a target, and a negative label $L_{i}=-1$ as a decoy win or decoy.
 Lemma S1.
-Let $\tilde{L}_{i}$ for $i\in[p]$ denote the training labels, where $\tilde{L}_{i}=-1$ for a training decoy and $\tilde{L}_{i}=1$ for a pseudo target.
+Let $\tilde{L}_{i}$ for $i\in[p]$ denote the pseudo labels, where $\tilde{L}_{i}=-1$ for a training decoy and $\tilde{L}_{i}=1$ for a pseudo-target.
 Let $\mathcal{F}$ be the $sigma$-algebra generated by the winning scores $W$, the side information $\mathbf{x}$, and the labels of the false nulls.
 Let $N$ be the indices of the true nulls, and $L_{-i},\tilde{L}_{-i}$ be the corresponding labels without the $i$th one.
-Then we have for $i\in N$:
-$L_{i}$ =$=\begin{cases}+1,&\mathbb{P}(L_{i}=+1)\leq\frac{c_{0}}{1-s\cdot(1-c_{0})}\\ -1,&\mathbb{P}(L_{i}=-1)\geq\frac{(1-c_{0})\cdot(1-s)}{1-s\cdot(1-c_{0})}\\ \end{cases},$
-independently of $\{\tilde{L}_{i}=1\},\mathcal{F},\tilde{L}_{-i},L_{-i}$.
+Then under Assumption 3, we have for $i\in N$, $L_{i}=\pm 1$ with
+$\mathbb{P}(L_{i}=+1\mid\tilde{L}_{i}=1)\leq\frac{c_{0}}{1-s\cdot(1-c_{0})},\quad\mathbb{P}(L_{i}=-1\mid\tilde{L}_{i}=1)\geq\frac{(1-c_{0})\cdot(1-s)}{1-s\cdot(1-c_{0})},$
+independently of $\mathcal{F},\tilde{L}_{-i},L_{-i}$.
 Proof.
-By Assumption 3, $\{L_{i}\,:\,i\in N\}$ are i.i.d with $P(L_{i}=1)\leq c_{0}$, independently of $\sigma(\mathcal{F},\tilde{L}_{-i},L_{-i})$.
+By Assumption 3, we know $\{L_{i}\,:\,i\in N\}$ are i.i.d with $P(L_{i}=1)\leq c_{0}$, independently of $\mathcal{G}:=\sigma(\mathcal{F},\tilde{L}_{-i},L_{-i})$.
+Hence, the following probability statements are independent of $\mathcal{G}$.
 In addition, since each decoy is randomly assigned to the training decoy set independently of anything else with probability $s$, for each $i\in N$ we have:
-$\mathbb{P}(L_{i}=-1,\tilde{L}_{i}=1\mid\mathcal{F},\tilde{L}_{-i},L_{-i})$ $=\mathbb{P}(L_{i}=-1\mid\mathcal{F},\tilde{L}_{-i},L_{-i})\cdot\mathbb{P}(\tilde{L}_{i}=1\mid L_{i}=-1,\mathcal{F},\tilde{L}_{-i},L_{-i})$
+$\mathbb{P}(L_{i}=-1,\tilde{L}_{i}=1)$ $=\mathbb{P}(L_{i}=-1)\cdot\mathbb{P}(\tilde{L}_{i}=1\mid L_{i}=-1)$
 $\geq(1-c_{0})\cdot(1-s),$
-and
-$\mathbb{P}(L_{i}=1,\tilde{L}_{i}=1\mid\mathcal{F},\tilde{L}_{-i},L_{-i})$ $=\mathbb{P}(L_{i}=1\mid\mathcal{F},\tilde{L}_{-i},L_{-i})\cdot\mathbb{P}(\tilde{L}_{i}=1\mid L_{i}=1,\mathcal{F},\tilde{L}_{-i},L_{-i})$
-$\leq c_{0}\cdot 1,$
 Hence,
-$\mathbb{P}(L_{i}=-1\mid\tilde{L}_{i}=1,\mathcal{F},\tilde{L}_{-i},L_{-i})$ $=\frac{\mathbb{P}(L_{i}=-1,\tilde{L}_{i}=1\mid\mathcal{F},\tilde{L}_{-i},L_{-i})}{P(\tilde{L}_{i}=1\mid\mathcal{F},\tilde{L}_{-i},L_{-i})}$
-$\geq\frac{(1-c_{0})\cdot(1-s)}{(1-s)\cdot\mathbb{P}(L_{i}=-1\mid\mathcal{F},\tilde{L}_{-i},L_{-i})+1\cdot\mathbb{P}(L_{i}=1\mid\mathcal{F},\tilde{L}_{-i},L_{-i})}$
-$=\frac{(1-c_{0})\cdot(1-s)}{1-s\cdot\mathbb{P}(L_{i}=-1\mid\mathcal{F},\tilde{L}_{-i},L_{-i})}$
-$\geq\frac{(1-c_{0})\cdot(1-s)}{1-s\cdot(1-c_{0})}$
+$\mathbb{P}(L_{i}=-1\mid\tilde{L}_{i}=1)$ $=\frac{\mathbb{P}(L_{i}=-1,\tilde{L}_{i}=1)}{P(\tilde{L}_{i}=1)}$
+$\geq\frac{(1-c_{0})\cdot(1-s)}{(1-s)\cdot\mathbb{P}(L_{i}=-1)+1\cdot\mathbb{P}(L_{i}=1)}$
+$=\frac{(1-c_{0})\cdot(1-s)}{1-s\cdot\mathbb{P}(L_{i}=-1)}$
+$\geq\frac{(1-c_{0})\cdot(1-s)}{1-s\cdot(1-c_{0})}.$
 The proof is complete noting that $\frac{c_{0}}{1-s\cdot(1-c_{0})}+\frac{(1-c_{0})\cdot(1-s)}{1-s\cdot(1-c_{0})}=1$ and $L_{i}\in\{\pm 1\}$ a.s.. ∎
 Remark 1.
-Lemma S1 essentially states that the labels $L_{i}$ for $i\in N$ with $\tilde{L}_{i}=1$ are i.i.d random variables with $\mathbb{P}(L_{i}=+1)\leq\frac{c_{0}}{1-s\cdot(1-c_{0})}$ independently of all other information: the scores $W$, the side information $\mathbf{x}$, the labels of the false null discoveries, and all the labels $\tilde{L}$.
+Lemma S1 essentially states that the labels $L_{i}$ for $i\in N$ with $\tilde{L}_{i}=1$ are i.i.d.
+random variables with $\mathbb{P}(L_{i}=+1)\leq\frac{c_{0}}{1-s\cdot(1-c_{0})}$ independently of all other information: the scores $W$, the side information $\mathbf{x}$, the labels of the false null discoveries, and all the labels $\tilde{L}$.
 See 1
@@
 Recall that the scores $\tilde{W}$ are themselves a function of the values $(\tilde{L},W,x)$ and clearly the one bit p-values are a function of the labels $L_{i}$.
-Then it follows from the last Lemma that given the scores $\tilde{W}$ and the one-bit p-values of the false nulls, the one-bit p-values of the true nulls hypotheses ($i\in N$ with $\tilde{L}_{i}=1$) are i.i.d with $\mathbb{P}(\tilde{p}_{i}\leq\frac{c_{0}}{1-s\cdot(1-c_{0})})=\mathbb{P}(L_{i}=1)\leq\frac{c_{0}}{1-s\cdot(1-c_{0})}$, i.e., they stochastically dominate the uniform distribution and hence are valid p-values.
+Then it follows from the last Lemma that given the scores $\tilde{W}$ and the one-bit p-values of the false nulls, the one-bit p-values of the true nulls hypotheses ($i\in N$ with $\tilde{L}_{i}=1$) are i.i.d.
+with $\mathbb{P}(\tilde{p}_{i}\leq\frac{c_{0}}{1-s\cdot(1-c_{0})})=\mathbb{P}(L_{i}=1)\leq\frac{c_{0}}{1-s\cdot(1-c_{0})}$, i.e., they stochastically dominate the uniform distribution and hence are valid p-values.
 Therefore, the condition of Theorem 3 holds so applying, as RESET does, Selective SeqStep+ with $c=\frac{c_{0}}{1-s\cdot(1-c_{0})}$ to these p-values controls the FDR at level $\alpha$.
@@
 Proof.
-The proof of Theorem 1 shows that conditional on the learned scores $\tilde{W}$ and the false null labels the labels $L_{i}$ for $i\in N$ with $\tilde{L}_{i}=1$ are distributed as i.i.d ±$\pm 1$ RVs, with $\mathbb{P}(L_{i}=1)\leq\frac{c_{0}}{1-s\cdot(1-c_{0})}$.
+The proof of Theorem 1 shows that conditional on the learned scores $\tilde{W}$ and the false null labels the labels $L_{i}$ for $i\in N$ with $\tilde{L}_{i}=1$ are distributed as i.i.d.
+±$\pm 1$ RVs, with $\mathbb{P}(L_{i}=1)\leq\frac{c_{0}}{1-s\cdot(1-c_{0})}$.
 Assuming for a moment that $\mathbb{P}(L_{i}=1)=\frac{c_{0}}{1-s\cdot(1-c_{0})}$, then RESET satisfies the assumption of the multiple decoy version of FDP-SD with $c=\lambda=\frac{c_{0}}{1-s\cdot(1-c_{0})}$ and therefore it controls the FDP at $\alpha$ with confidence $1-\gamma$ (Section 4 of [40]).
@@
 We point out some additional minor details regarding our implementation of RESET.
-We use the randomxForest package, nnet and mgcv to implement the random forest, two-layer neural network and generative additive model in R, respectively.
-We used the default parameters for the random forest, the default parameters for the two-layer neural network and the default parameters for the generative additive model (except we use drop.intercept = TRUE).
+We use the randomxForest package, nnet and mgcv to implement the random forest, two-layer neural network and generalized additive model in R, respectively.
+We used the default parameters for the random forest, the default parameters for the two-layer neural network and the default parameters for the generalized additive model (except we use drop.intercept = TRUE).
 If the number of side information variables used by RESET is $\leq 3$, a smoothing spline is fitted with mgcv::s otherwise a natural cubic spline is fitted using 5 degrees of freedom with splines::ns on each side information variable separately (for computational efficiency).
@@
 We also fixed a minor bug, which pruned the candidate set $\mathcal{S}$ before checking the estimated FDR is $\leq\alpha$ — possibly missing out on a marginally larger discovery list.
-There are several parameters that may be set before applying the Adaptive Knockoff filters, e.g.
-the initial proportion of hypotheses revealed, reveal_prop.
-In Simulations 1-3, we used the same parameters set by Ren and Candès from their paper.
-In Simulation 4, we used the same parameters in Simulation 3.
+There are several parameters that may be set before applying the Adaptive Knockoff filters, e.g., the initial proportion of hypotheses revealed, reveal_prop.
+We used the same parameter settings in Simulation 1 as Simulation 1 in [44], and the same parameter settings in Simulation 2-4 as their Simulation 2 [44].
 In the application to peptide detection, the primary scores $W$ may be negative.
@@
 For AdaKO EM, any peptide score $\leq 10^{-3}$ was subsequently assigned a zero score.
-Finally we used, reveal_prop = 10% which reveals the hypotheses that are less than or equal to the 10% of non-zero scoring hypotheses (which is the same setting used by the simulations).
+Finally we used, reveal_prop = 10% which reveals the hypotheses that are less than or equal to the 10% of positive scoring hypotheses (which is the same setting used by the simulations).
 All other parameters were default.
@@
 S.7 Further implementation details of the AdaPTg methods
-We implemented AdaPTg using the repository https://github.com/patrickrchao/adaptMT which has all the same implementations as AdaPT while offering the ability to define assymetric regions for mirroring the p-values.
+We implemented AdaPTg using the repository https://github.com/patrickrchao/adaptMT which has all the same implementations as AdaPT while offering the ability to define asymmetric regions for mirroring the p-values.
 Accordingly, we followed the same implementation details as the previous section used by AdaPT and set the so-called ‘masking parameters’ as default.
@@
 The exception was in Simulation 6, where we directly used all 100 side information variables rather than applying a spline basis expansion on each of them.
-In Simulation 5 and 6, we used the test statistics obtained as input, while in the real data experiments we converted the p-values to z-values by using the transformation ±$\pm\Phi^{-1}(p_{i}/2)$ on each p-value $p_{i}$ where the signs are chosen i.i.d uniformly.
+In Simulation 5 and 6, we used the test statistics obtained as input, while in the real data experiments we converted the p-values to z-values by using the transformation ±$\pm\Phi^{-1}(p_{i}/2)$ on each p-value $p_{i}$ where the signs are chosen i.i.d.
+uniformly.
 S.10 Further implementation details of AdaFDR
-We implemented AdaFDR using the method.adafdr_test function from the AdaFDR package https://github.com/martinjzhang/adafdr.
+We applied AdaFDR using the method.adafdr_test function from the AdaFDR package https://github.com/martinjzhang/adafdr.
 We note that installation of the package required setting up a separate conda environment so that we could install an older version of Python (v3.6.13).
@@
 look at.
+Interestingly, the two fMRI datasets only consider the Brodmann label as the side information.
+We tried using all the available side information, but the statistical power was worse (data not shown).
 In the gene-drug response data, we used AdaFDR with fast_mode = False.
-This is in accordance to the discussion from their paper that the fast version of AdaFDR is recommended when there are few discoveries to be expected or if the number of hypotheses are small (which is not the case for the gene-drug response data).
+This is in accordance to the discussion from their paper that the fast version of AdaFDR is recommended when few discoveries are expected or if the number of hypotheses is small (which is not the case for both of the gene-drug response datasets).
 S.11 Computer specifications for computation times
@@
 The resulting search files were then converted to so-called pin files using the make-pin function in Crux.
-Each row in the pin files correspond to the optimal peptide-spectrum match (PSM) for each spectrum, the primary score for quantifying this match, called XCorr scores, and a collection of auxilliary information regarding the PSM.
+Each row in the pin files correspond to the optimal peptide-spectrum match (PSM) for each spectrum, the primary score for quantifying this match, called XCorr scores, and a collection of auxiliary information regarding the PSM.
 Side Information Description
@@
 Peptides with a winning score of $\infty$- ∞, indicating that neither the target or its paired decoy are in the pin file, are thrown out.
-We assign each winning peptide the side information defined as the auxiliary information $x_{i}$ of the underlying PSM that had the XCorr score of $W_{i}$.
+We assign each winning peptide the side information defined as the auxiliary information $\mathbf{x}_{i}$ of the underlying PSM that had the XCorr score of $W_{i}$.
 The side information that was subsequently used by RESET and Adaptive Knockoffs is given in Table S4.
 S.13 Estimating the computation times of Adaptive Knockoffs
-Running Adaptive Knockoffs on all of the peptide datasets would be infeasible.
+Running Adaptive Knockoffs (AdaKO) on all peptide datasets would be infeasible.
 Therefore, we resorted to estimating its runtimes by considering the time it took to complete a single iteration that determines which hypothesis to reveal next, and multiplying this value by the estimated number of iterations before the procedure stops.
-Since the time it takes a single iteration depends on the number of hypotheses that were already revealed, we took different such measurements as we varied this number.
-Specifically, we revealed the labels that were less than or equal to the 10% quantile of non-zero scoring hypotheses, up to a value of $10k$%, in 10% increments.
-We chose $k$ not to be too large, since hypothetically Adaptive Knockoffs might terminate at the 1% FDR level before it reaches $10k$of the revealed hypotheses.
-Of course, we are unable to compute when Adaptive Knockoffs might terminate, so we estimated this using the value $r$, the number of target and decoy peptides that scored above the 1% FDR cutoff using RESET ensemble.
-Then we defined $k=\lfloor(m-r)/m\cdot 10\rfloor$, where $m$ is the total number of hypotheses (peptides) considered by Adaptive Knockoffs.
-With $\bar{t}_{k}$ being the average of the $k$ measured single-iteration times, we estimated the runtime by $\bar{t}_{k}\cdot(m^{\prime}-r)$, where $m^{\prime}$ is the number of labels that are yet to be revealed (that are above the 10% quantile of non-zero scoring hypotheses) and $m^{\prime}-r$ is our estimate of the number of iterations until Adaptive Knockoffs stop.
-In Section 5.2.2, we considered 10 repeated applications of RESET ensemble, in which case $r$ is the average of the number of winning target and decoy peptides above the cutoff over each application.
+Since the time it takes a single iteration depends on the number of hypotheses that were already revealed, we took different such measurements as we varied this number, and averaged them.
+Specifically, as AdaKO does, we first revealed the labels that were scored less than or equal to the 10% quantile of non-zero scoring hypotheses, and recorded the time it takes for a single interation from this point.
+To be clear, the revealed labels thus includes those zero-scoring hypotheses since they will be less than this score cutoff — it is only the 10% quantile that considers the non-zero scoring hypotheses.
+Then we took similar measurements by increasing the ‘10%’ value, in 10% increments, up to a value of ‘10k%’ where we define $k$ next.
+We chose the value of $k$ so that the proportion of revealed labels, $10k\%$, estimates the number of revealed labels when AdaKO terminates at the 1% FDR level.
+That way, the $k$ times should provide an accurate sample of the number single iteration times across all of AdaKO’s iterations until it terminates.
+Of course, we are unable to compute when Adaptive Knockoffs might terminate, so we estimated this using the value $q$, the number of target and decoy peptides that scored above the 1% FDR cutoff using RESET Ensemble.
+Then we defined $k=\lfloor(m-q)/m\cdot 10\rfloor$, where $m$ is the total number of hypotheses (peptides) considered by Adaptive Knockoffs.
+We estimated the runtime by
+$\bar{t}_{k}\cdot(m^{\prime}-q),$
+where $\bar{t}_{k}$ is the average of the $k$ measured single-iteration times, and $m^{\prime}$ is the number of labels that are yet to be revealed at the start of AdaKO’s iterations, i.e., the number of labels with scores above the 10% quantile of non-zero scoring hypotheses.
+Thus the value $m^{\prime}-q$ is our estimate of the number of iterations until Adaptive Knockoffs stop.
+In Section 5.2.2, we considered 10 repeated applications of RESET Ensemble, varying RESET Ensemble’s internal seed.
+In this case, we define $\hat{\ell}$ as the average of the number of winning target and decoy peptides above the cutoff over each application.
 S.14 Data preparation and searching with PRIDE-20 data
@@
 As an example, PEPTIDE may generate the following ‘copies’: PE[16]PTIDE, PEPTIDE[16] and PE[16]PTIDE[16], where [16] indicates an increase of 16 Daltons to the amino acid on the left.
-Consequently, each of these peptides are usually correlated in the data — i.e.
-if one of these peptides is scored high, than usually they all are.
+Consequently, each of these peptides are usually correlated in the data — i.e., if one of these peptides is scored high, than usually they all are.
 Hence in order to satisfy Assumption 3, we follow the protocol outlined in [18].
@@
 We define the labels, winning scores and side information in the same way to obtain our triplets $(L,W,\mathbf{x}_{i})$.
-Due to runtime considerations, we used 13 of the smallest narrow search files to assess Adaptive Knockoff’s power as described in Section 5.2.2 of the manuscript, and only one one of the 10 target-decoy databases.
-To fairly compare with Adaptive Knockoff, we used the same target-decoy databases for RESET ensemble.
+Due to runtime considerations, we used 13 of the smallest narrow search files to assess Adaptive Knockoff’s power as described in Section 5.2.2 of the manuscript, and only one of the 10 target-decoy databases.
+To fairly compare with Adaptive Knockoffs, we used the same target-decoy databases for RESET ensemble.
 All 20 spectrum files and all 10 target-decoy databases for each spectrum file were used in the comparison of RESET ensemble and FDP-SD while controlling the FDP.
@@
 RESET additionally requires this statement to be true if we further condition on the side information.
-This is obviously true for some side information variables which are constant between the target and corresponding decoys — e.g.
-PepLen, Charge, lnNumSP, dm, absdM.
-The other side information variables are score-related — e.g.
-Xcorr, deltCn and deltLCn, and so conditioning on them assumes no more than the “standard assumption”.
+This is obviously true for some side information variables which are constant between the target and corresponding decoys — e.g., PepLen, Charge, lnNumSP, dm, absdM.
+The other side information variables are score-related — e.g., Xcorr, deltCn and deltLCn, and so conditioning on them assumes no more than the ‘standard assumption’.
 Hence, we believe Assumption 3 with $P(L_{i}=1)=c_{0}=1/2$ to be reasonably satisfied.
-S.16 Description and Data preparation using data from Zhang et al.
+S.16 Preparation of data with p-values
 Each of the 10 datasets from Zhang et al.
@@
 In the RNA-seq datasets, Airway, Bottomly and Pasilla, there exists ‘spikes’ of high p-values.
-These ‘spikes’ concentrate in certain regions of the side information, i.e.
-when the log-normalized gene counts are low.
+These ‘spikes’ concentrate in certain regions of the side information, i.e., when the log-normalized gene counts are low.
 Hence we removed those p-values with log-normalized gene counts that were less than zero.
 Figure S1 show the histograms of the p-values before and after this removal.
-Figure S1: Histograms of p-values between in the RNA-seq data.
-The first row of panels plot the histograms before the removal of some of the p-values while the second row of panels plot the histograms after this removal.
+The gene-drug response data was obtained from the adaptMT package.
+Figure S1: Histograms of p-values in the RNA-seq data before and after p-value removal.
+Top row: the original histograms of the p-values.
+Bottom row: the corresponding histograms after removal of the p-values associated with low counts.
 S.17 Tables
@@
 For readability, the points are jittered in the horizontal direction.
-A description of AdaKO GAM and KO can be found in Section 2.
+A description of the AdaKO methods and KO can be found in Section 2.

### v3

@@
 Recently there has been a growing interest in enhancing such procedures by combining additional information with the primary p-value or score.
-Specifically, such so-called side information can be leveraged to improve the separation between true and false nulls along additional dimensions thereby increasing the overall sensitivity.
-In line with this idea, we develop RESET (REScoring via Estimating and Training), which uses a unique data-splitting protocol that subsequently allows any semi-supervised learning approach to factor in the available side information while maintaining finite-sample error rate control.
+In line with this idea, we develop RESET (REScoring via Estimating and Training), which uses a unique data-splitting protocol that subsequently allows any semi-supervised learning approach to factor in the available side information while maintaining finite sample error rate control.
 Our practical implementation, RESET Ensemble, selects from an ensemble of classification algorithms so that it is compatible with a range of multiple testing scenarios without the need for the user to select the appropriate one.
-We apply RESET to both p-value and competition based multiple testing problems and show that RESET is (1) power-wise competitive, (2) fast compared to most tools and (3) able to uniquely achieve finite sample false discovery rate or false discovery proportion control, depending on the user’s preference.
-Corresponding authors: J.Freestone@maths.usyd.edu.au, uri.keich@sydney.edu.au
+We apply RESET to both p-value and competition based multiple testing problems and show that RESET is (1) power-wise competitive, (2) fast compared to most tools and (3) able to uniquely achieve finite sample false discovery rate or false discovery exceedance control, depending on the user’s preference.
+Corresponding authors: jack.freestone@mq.edu.au, uri.keich@sydney.edu.au
 1 Introduction
 Scientists are often interested in testing many null hypotheses, $\{H_{i}:i=1,\dots,m\}$.
-As an example, a scientist may be interested in assessing which genes exhibit a change in gene expression levels in response to a particular drug treatment.
-In this case, the $i$th null hypothesis $H_{i}$ states that ‘the $i$th gene exhibits no change in gene expression level’.
-This type of large-scale hypothesis testing is usually achieved by assigning a p-value, $p_{i}$, to each null hypothesis.
-The collection of p-values then undergoes a filtering process that rejects a subset of these null hypotheses that have sufficiently small p-values subject to a type-1 error rate control to minimize the number of misreported true null hypotheses.
-The most common choice of error rate in this setup is the false discovery rate (FDR), which is defined as the expectation of the false discovery proportion (FDP) — the fraction of true null hypotheses in the reported list of rejections, or discoveries [2].
+Large-scale hypothesis testing is usually achieved by assigning a p-value, $p_{i}$, or a score, $Z_{i}$, to each null hypothesis.
+The collection of p-values, or scores, then undergoes a filtering process that rejects a subset of these null hypotheses subject to type-1 error rate control to minimize the number of misreported true null hypotheses.
+The most common choice of error rate is the false discovery rate (FDR), which is defined as the expectation of the false discovery proportion (FDP) — the fraction of true null hypotheses in the reported list of rejections, or discoveries (Benjamini and Hochberg (1995)).
 In this case, the filtering procedure ensures that the FDR is $\leq\alpha$ where $\alpha\in\left(0,1\right)$ is a prespecified threshold.
-Alternatively, we might be interested in controlling the FDP rather than its expectation in the following sense.
-Given a confidence parameter $\gamma$ and a threshold $\alpha$, the FDP is said to be controlled if $\mathbb{P}(FDP>\alpha)\leq\gamma$.
-Evidently, such FDP control reduces the risk that the realized FDP exceeds $\alpha$, which is not guaranteed by FDR control [40, 27, 19, 38, 16, 20].
-Determining a p-value for each null hypothesis can be challenging because (a) it requires knowledge of the underlying null distribution and (b) it can be computationally intensive.
-A recently developed competition framework offers an alternative approach to FDP/FDR control that does not require computation of canonical p-values [1, 21].
-Instead, all that these procedures require is a single draw, $\tilde{Z}_{i}$, from each null distribution.
-The null drawn $\tilde{Z}_{i}$ is then compared with $Z_{i}$, the originally observed score (test statistic) for $H_{i}$, thus conceptually providing a crude 1-bit p-value.
-These p-values are too crude to be used in typical p-value based filters.
-Instead, assuming that larger values of $Z_{i}$ provide more evidence against $H_{i}$, we compete each pair of scores $(Z_{i},\tilde{Z}_{i})$ by recording a label $L_{i}\in\{\pm 1\}$ indicating which was the largest of the two scores, as well as a winning score, $W_{i}=f(Z_{i},\tilde{Z}_{i})$.
-The function $f$ needs to be symmetric in its arguments and should increase with the values of $Z_{i}$ and $\tilde{Z}_{i}$, e.g., $W_{i}=Z_{i}\vee\tilde{Z}_{i}$, or with the discrepancy between the two, e.g., $W_{i}=\lvert Z_{i}-\tilde{Z}_{i}\rvert$.
-[^1] In general, the functions can be hypothesis-dependent, i.e., $f_{i}$ instead of $f$.
-For each true null hypothesis, $L_{i}$ is equally likely to be ±$\pm 1$ independent of all the scores $W=\{W_{i}\}$ and all other labels $L_{-i}$ because both scores were sampled from the same null distribution and $f$ is symmetric.
-Hence, the number of negative labels, those with $L_{i}=-1$, can be used to estimate the number of true nulls with positive labels among the top scoring hypotheses.
-Using this fact, the labels and scores $\{(L_{i},W_{i}):i=1,\dots,m\}$ are passed through a filter to reject a subset of the top scoring hypotheses, with either the FDR or FDP controlled depending on the user’s preference.
-More recently, research has been focused on improving these filtering methods by leveraging side information $\mathbf{x}_{i}\in\mathcal{X}$ that complements each p-value $p_{i}$ or score $W_{i}$ to better discriminate between the true and false null hypotheses.
+Alternatively, we might be interested in controlling false discovery exceedance (FDX), which is the upper tail probability of the FDP.
+The FDX is said to be controlled with threshold $\alpha$ at confidence $1-\gamma$ if $\mathbb{P}(FDP>\alpha)\leq\gamma$.
+Evidently, such FDX control reduces the chance that the realized FDP exceeds $\alpha$, which is not guaranteed by FDR control (Guo and Romano (2007); Katsevich and Ramdas (2020); Goeman et al.
+(2021); Li et al.
+(2022); Luo et al.
+(2023); Ebadi et al.
+(2023)).
+Recently, researchers have focused on developing methods that leverage side information $\mathbf{x}_{i}\in\mathcal{X}$, which complements each p-value, or score, to better discriminate between true and false null hypotheses.
 These methods utilize a variety of strategies but ultimately operate with the same goal in mind: the side information is used to weight or rescore the hypotheses according to how confidently we believe that each hypothesis is a false null.
-Indeed, there are several such tools that augment p-values with side information [25, 34, 33, 37, 51, 35].
-When it comes to using side information in the competition framework, we are aware of only one fully specified method with rigorous error-rate control, which is the recently published Adaptive Knockoffs [44] (AdaPT [33] only offers a meta-procedure).
 In this paper, we offer a novel approach for multiple testing with side information.
@@
 Our method flexibly incorporates any semi-supervised approach to help distinguish between true and false null hypotheses.
-Moreover, it is equally applicable to the competition and p-value context and allows the user to choose between FDR or FDP control.
+Moreover, it is applicable to both competition- and p-value-based multiple testing scenarios, and allows the user to choose between FDR or FDX control.
 We evaluate RESET in a range of simulations and real data applications to verify that it is a competitive alternative to existing methods in terms of statistical power and speed.
 2 Background: competition-based multiple testing
-Competition-based multiple testing was first established in the proteomics community where the goal is to infer which proteins, i.e., long chains of amino acids, are present in a biological sample [17].
-Direct protein detection is challenging, and so a bottom-up approach is taken where the proteins are broken into small chains of amino acids, called peptides, and the objective is to identify these peptides instead.
-Then, a process called liquid chromatography tandem mass spectrometry (LC-MS/MS) is used to isolate each sample peptide and subsequently generate a fragmentation spectrum.
-Each such spectrum can be thought of as a fingerprint of the peptide that generated the spectrum, allowing us to hypothesize which peptides are present in the sample.
-This is done by searching each spectrum against a database of candidate peptides called targets to find the spectrum’s optimal peptide-spectrum match (PSM), along with a score indicating the strength of the match.
-Unfortunately, due to the incomplete nature of the database and the noise associated with generating the spectra, the PSM can be incorrect.
-Hence, we wish to employ a testing procedure for each null hypothesis $H_{i}$ that ‘the $i$th target is not present in the sample’.
-In order to decide which null hypotheses to reject, we pair each target peptide with a decoy peptide that is generated by randomly shuffling or reversing the target peptide’s sequence.
-Accordingly, a second search is performed for each spectrum against the database of decoys to obtain a second PSM and associated score, and only the highest scoring PSM out of the two is retained.
-As outlined in the introduction, we accompany $H_{i}$ with a score $Z_{i}$, defined as the maximum scoring PSM associated with the $i$th target peptide, and the higher the score, the more likely $H_{i}$ is false, i.e., the peptide is in the sample.
-Similarly, we define $\tilde{Z}_{i}$ as the maximum scoring PSM associated to the $i$th target’s corresponding decoy peptide.
-For each true null hypothesis, we expect $Z_{i}$ and $\tilde{Z}_{i}$ to be drawn from the same (hypothesis-specific) null distribution, independently of everything else.
-Finally, we obtain a winning score $W_{i}=Z_{i}\vee\tilde{Z}_{i}$ and a label $L_{i}$ indicating whether the winning score was from the $i$th target or its decoy [42].
-Subsequently, the statistics community independently formalized this idea in the context of variable selection, where the goal is to identify relevant variables from $\{X_{i}:i=1,\dots,p\}$ that are associated with a response $y$.
-The procedure that achieves this while controlling the FDR is called the knockoff filter, which begins by pairing each variable $X_{i}$ with an artificial variable $\tilde{X}_{i}$, called a knockoff.
-In this paper, we consider the model-X version of the knockoff filter [7] which makes the following assumptions [^2] However, our results are equally applicable to the fixed-X model as well [1]..
-First, the joint distribution of the variables, $X$, is assumed to be known.
-Second, each observation $(X_{i,1},\dots,X_{i,p},y_{i})$ is sampled i.i.d.
-from the joint distribution $F_{XY}$.
-Because $X$ is random and its distribution known, the knockoffs $\tilde{X}$ are also constructed randomly subject to the following conditions:
-  1. 1.
-[…]
-  2. 2.
-    We construct $\tilde{X}$ independently of $y$ by only looking at $X$, i.e., $\tilde{X}\perp\!\!\!\perp y\mid X$.
-Once the knockoffs have been constructed, each variable is given a user-defined score $W_{i}:=W_{i}([X,\tilde{X}],y)\in\mathbb{R}$ and a label $L_{i}([X,\tilde{X}],y)\in\{\pm 1\}$.
-The idea is that the label $L_{i}$ indicates whether the variable $X_{i}$ or the knockoff $\tilde{X}_{i}$ has greater evidence for being relevant to the model, and the score $W_{i}$ indicates by how much.
-Commonly, the score and label are combined by setting $sign(W_{i})=L_{i}$; however, as will become clear in our later presentation of RESET, it is instructive to separate them.
-In this paper, we consider the Lasso Coefficient Difference (LCD) score, which is the difference in the magnitude of the coefficients of the Lasso regression [49].
-Specifically, cross-validation is used to determine the $\lambda$ penalty when regressing on the augmented design matrix $[X,\tilde{X}]$.
-We denote the absolute values of the Lasso($\lambda$)-fitted coefficients as $Z_{i}$ for the variables and $\tilde{Z}_{i}$ for the knockoffs.
-The scores and labels are then defined as
-$W_{i}:=\big{|}Z_{i}-\tilde{Z_{i}}\big{|},\quad L_{i}:=sign(Z_{i}-\tilde{Z_{i}}).$ (1)
-Hypotheses with zero scores are typically ignored.
-In either context, peptide inference or variable selection, we expect that the labels of the true null hypotheses are independently and equally likely to be ±$\pm 1$, independently of everything else.
-This is formalized in the next assumption.
+In competition-based multiple testing, each hypothesis $H_{i}$ is associated with two scores: a test or target score, $Z_{i}$, where larger values indicate greater evidence against the null, and a decoy score, $\tilde{Z}_{i}$, a random draw from the null distribution.
+Each pair $(Z_{i},\tilde{Z}_{i})$ is evaluated in a head-to-head competition yielding a label $L_{i}\in\{\pm 1\}$ that indicates which score is higher and a winning score $W_{i}$:
+$L_{i}:=\begin{cases}+1&Z_{i}>\tilde{Z}_{i}\\ -1&Z_{i}<\tilde{Z}_{i}\end{cases},\quad W_{i}:=f(Z_{i},\tilde{Z}_{i}),$
+where $f$ is symmetric in its arguments and increases with the values of $Z_{i}$ and $\tilde{Z}_{i}$, or with the difference between them, e.g., $f(x,y)=\max\{x,y\}$ or $f(x,y)=\lvert x-y\rvert$.
+Hypotheses with ties (i.e., $Z_{i}=\tilde{Z}_{i}$) are either discarded or the ties are randomly broken.
+We refer to a hypothesis with $L_{i}=1$ as a target win or a target for short, and $L_{i}=-1$ as a decoy win or a decoy for short.
+The winning scores and labels have two properties.
+First, the false nulls with large $Z_{i}$ will likely win their competition, resulting in a target win ($L_{i}=1$) with large $W_{i}$.
+Second, the true null labels are independent and equally likely to be $\pm 1$, independently of everything else.
+This is formalized in the following assumption.
 Assumption 1.
 Let $N$ be the indices of the true null hypotheses.
-The labels $\{L_{i}:i\in N\}$ are i.i.d.
-±$\pm 1$ uniform random variables independent of all the scores $W$ and the labels of the false null hypotheses.
-For the knockoff filter, Assumption 1 follows from the knockoff construction, whereas in the peptide detection problem, Assumption 1 is based on our previous discussion that the scores of a null target and its decoy, $Z_{i}$ and $\tilde{Z}_{i}$, are independent draws from the same null distribution, and moreover, this happens independently of all other pairs of scores.
-This assumption is commonly applied in the literature, with several analyses validating this claim empirically [39, 21].
-Given the above assumption, the number of hypotheses with $L_{i}=-1$ (i.e., decoy or knockoff wins) that score above a score cutoff of $\tau$ can be used to estimate the number of true null hypotheses that score above $\tau$ with $L_{i}=1$.
-Using this strategy, we can report a list of discoveries while satisfying a type-1 error rate control in one of two ways.
-The first and by far the more common way is to apply Selective SeqStep+ (SSS+) [1], which controls the FDR at a prespecified threshold $\alpha$ by reporting the following list of discoveries $R_{\alpha}$:
-$R_{\alpha}:=\{i\in[p]:W_{i}\geq\tau,L_{i}=1\},\quad\text{where }\tau:=\min\left\{t\in\mathbb{R}:\frac{\#\{W_{i}\geq t,L_{i}=-1\}+1}{\#\{W_{i}\geq t,L_{i}=1\}\vee 1}\leq\alpha\right\}.$ (2)
-SSS+ is made precise in Algorithm S1 and works in a more general setting when the probabilities of the positive and negative labels of the true null hypotheses are no longer uniform, i.e., $\mathbb{P}(L_{i}=1)=c$ and $\mathbb{P}(L_{i}=-1)=1-c$.
-In the proteomics community, the application of Equation (2) in the peptide detection problem is referred to as target-decoy competition or TDC for short.
-As such, we will refer to the method as TDC in the peptide detection setting, and SSS+ in the variable selection setting and more generally.
-Alternatively, we can control the FDP by using, for example, FDP-stepdown (FDP-SD, Algorithm S2[40].
+Conditioned on all the scores $W$, and the labels of the false null hypotheses, the true null labels $\{L_{i}:i\in N\}$ are i.i.d.
+$\pm 1$ uniform random variables.
+Given Assumption 1 holds, the number of decoy wins ($L_{i}=-1$) that score above a cutoff $\tau$ can be used to conservatively estimate the number of true null hypotheses with $L_{i}=1$ that are also above $\tau$.
+Using this strategy, we can report a list of discoveries in one of two ways.
+The first and by far the more common way is to apply Selective SeqStep+ (SSS+, Algorithm A1) (Barber and Candès (2015)), which controls the FDR under Assumption 1 at a prespecified threshold $\alpha$.
+Assuming the labels $L$ are reranked according to the scores $W$ in descending order, SSS+ reports the following list of discoveries, $R_{\alpha}$:
+$R_{\alpha}:=\{i\leq\tau:L_{i}=1\},\text{ where }\tau:=\max\left\{t\in[m]:\frac{\#\{i\leq t,L_{i}=-1\}+1}{\#\{i\leq t,L_{i}=1\}\vee 1}\leq\alpha\right\},$ (1)
+where $m]:=\{1,2,\dots,m\}$.
+Notably, SSS+ can be adjusted to handle a more general setting when the probabilities of the positive and negative labels of the true null hypotheses are no longer uniform (Barber and Candès (2015); Lei and Fithian (2016)).
+Alternatively, we can control the FDX by using, for example, FDP-stepdown (FDP-SD, Algorithm A2) (Luo et al.
+(2023)).
 FDP-SD works by first reranking the labels $L$ according to the scores $W$ in descending order.
-Next, it compares the number of decoy (knockoff) wins in the first $i$ hypotheses, denoted as $D_{i}:=\#\{L_{j}=-1:j\leq i\}$, with a precomputed bound $\delta_{i}$, which is determined by $\alpha$ and $\gamma$.
-Each comparison is made in a ‘stepdown’ fashion, i.e., if the first $i$ hypotheses satisfy $D_{1}\leq\delta_{1},\dots,D_{i}\leq\delta_{i}$, then it moves on to the next comparison at $i+1$.
-For small $i$, it is impossible for $D_{i}\leq\delta_{i}$ and so FDP-SD begins this analysis at $i_{0}$, the smallest possible index for which $D_{i}\leq\delta_{i}$.
-Under Assumption 1, reporting the following list of discoveries $R_{\alpha,\gamma}$ controls the FDP:
-$R_{\alpha,\gamma}:=\{i\in[p]:i\leq k_{FDP},L_{i}=1\},\quad k_{FDP}:=\max\{i:\Pi_{j=i_{0}}^{i}1_{D_{j}\leq\delta_{j}}=1\text{ or }i=0\}.$ (3)
+Next, it compares the number of decoy wins in the first $i$ hypotheses, denoted as $D_{i}:=\#\{L_{j}=-1:j\leq i\}$, with a precomputed bound $\delta_{i}$, which is determined by $\alpha$ and $\gamma$.
+The comparisons are made in a ‘stepdown’ fashion: we test whether $D_{i}\leq\delta_{i}$ only if $D_{j}\leq\delta_{j}$ for $j<i$.
+For small $i$, it is impossible for $D_{i}\leq\delta_{i}$ and so FDP-SD begins this analysis at $i_{0}$, the smallest possible index for which $D_{i}\leq\delta_{i}$ can occur.
+Under Assumption 1, reporting the following list of discoveries $R_{\alpha,\gamma}$ controls the FDX:
+$R_{\alpha,\gamma}:=\{i\in[m]:i\leq k_{FDP},L_{i}=1\},\quad k_{FDP}:=\max\{i:\Pi_{j=i_{0}}^{i}1_{D_{j}\leq\delta_{j}}=1\text{ or }i=0\}.$ (2)
 2.1 Extension to side information
-In proteomics, so-called post-processors such as PeptideProphet [28] and Percolator [26] have popularized the use of side information to increase the number of detected peptides.
-These tools use machine learning procedures, leveraging features associated with the PSMs, such as the charge state associated with the spectrum and the matched peptide’s length, to better distinguish correct from incorrect PSMs, and hence to deliver more discoveries.
-These tools have no theoretical guarantees and, having recently pointed out scenarios where such post-processors can fail to control the FDR [18], we will not analyze these tools in this paper.
-To our knowledge, the only ‘intrinsic’ competition-based multiple testing procedure that uses arbitrary side information in a data-driven manner is the recent Adaptive Knockoffs (AdaKO) by Ren and Candès [44].
-AdaKO is a specialization of the p-value based procedure AdaPT [33] that we discuss in the next section.
-By ‘intrinsic’, we mean that while other p-value based approaches may be modified to be used in the competition setting, they are significantly less powerful and do not really qualify in this sense, as demonstrated in [44].
-Adaptive Knockoffs works by iteratively pruning $\mathcal{S}$, the candidate set of hypotheses, selecting each time the hypothesis it deems most likely to be a true null.
-To do that, it fits a classification algorithm using a strict set of information that partially masks the data: the scores $W$, the side information $\mathbf{x}$, and the labels $L_{i}$ of the hypotheses that were pruned up to that point ($i\in[p]\setminus\mathcal{S}$), as well as #$\#\{L_{i}=1:i\in S\}$ and #$\#\{L_{i}=-1:i\in S\}$.
-In their paper, Ren and Candès investigate this pruning strategy using different types of ‘filters’: a logistic regression (LR), a generalized additive model (GAM) which is the default filter in their R package, a random forest (RF), and a two-group probabilistic model which is fitted using the expectation-maximization algorithm (EM) [13].
-The pruning process terminates when the estimated FDR is less than or equal to $\alpha$:
-$\widehat{FDR}_{\mathcal{S}}:=\frac{\#\{i\in\mathcal{S},L_{i}=-1\}+1}{\#\{i\in\mathcal{S},L_{i}=1\}\vee 1}\leq\alpha.$ (4)
-The remaining positively labeled hypotheses in $\mathcal{S}$ are then reported. It was shown that, given the following natural extension of Assumption 1, Adaptive Knockoffs controls the FDR.
+To our knowledge, the only ‘intrinsic’ competition-based multiple testing procedure that uses arbitrary side information in a data-driven manner is the recent Adaptive Knockoffs (AdaKO) by Ren and Candès (2023).
+AdaKO is a specialization of the p-value based procedure AdaPT (Lei and Fithian (2018)) that we introduce in the next section with further details in Section C.1.
+By ‘intrinsic’, we mean that while other p-value based approaches may be modified to be used in the competition setting, Ren and Candès (2023) demonstrate they are substantially less powerful, and are not viable in this sense.
+Adaptive Knockoffs works by iteratively pruning $\mathcal{\hat{S}}_{t}$, the candidate set of hypotheses at each step $t$, by selecting the hypothesis it deems most likely to be a true null.
+To do that, it fits a classification algorithm using a strict subset of information that partially masks the data: the scores $W$, the side information $\mathbf{x}$, and the labels $(L_{i})_{i\in[m]\setminus\mathcal{\hat{S}}_{t}}$ of the hypotheses that were pruned up to that point, as well as $\{i\in\mathcal{\hat{S}}_{t}:L_{i}=1\}$ and $\{i\in\mathcal{\hat{S}}_{t}:L_{i}=-1\}$.
+In their paper, Ren and Candès (2023) investigate this pruning strategy using different types of filters: a logistic regression (LR), a generalized additive model (GAM) which is the default filter in their R package, a random forest (RF), and a two-group model which is fitted using the expectation-maximization algorithm (EM) (Dempster et al.
+(1977)).
+The pruning process terminates at step $\tau$ when the estimated FDR is less than or equal to $\alpha$:
+$\widehat{FDR}_{\mathcal{\hat{S}}_{\tau}}:=\frac{\#\{i\in\mathcal{\hat{S}}_{\tau}:L_{i}=-1\}+1}{\#\{i\in\mathcal{\hat{S}}_{\tau}:L_{i}=1\}\vee 1}\leq\alpha.$ (3)
+The remaining positively labelled hypotheses in $\mathcal{\hat{S}}_{\tau}$ are then reported.
+It was shown that, given the following natural extension of Assumption 1, Adaptive Knockoffs controls the FDR.
 Assumption 2.
 Let $N$ be the indices of the true null hypotheses.
-The labels $\{L_{i}:i\in N\}$ are i.i.d.
-±$\pm 1$ uniform random variables independent of all the scores $W$, the side information $\mathbf{x}$, and the labels of the false null hypotheses.
+Conditioned on all the scores $W$, the side information $\mathbf{x}$, and the labels of the false null hypotheses, the true null labels $\{L_{i}:i\in N\}$ are i.i.d.
+$\pm 1$ uniform random variables.
 3 Background: p-value based multiple testing
-Multiple testing with FDR control was popularized by Benjamini and Hochberg when they developed the first testing procedure of this kind, which we refer to as the ‘BH procedure’ [2].
-Since then, many p-value based methods that control the FDR have been developed to improve statistical power.
-One such strategy for improving on BH involves estimating the fraction of true null hypotheses, denoted as $\pi_{0}$.
-Indeed, the BH procedure controls the FDR conservatively at level $\pi_{0}\cdot\alpha$ instead of $\alpha$.
-Hence, in cases when $\pi_{0}\ll 1$, the BH procedure loses a considerable amount of power.
-Methods like those by Storey et al.
-[46] and Benjamini et al.
-[3] estimate $\pi_{0}$, and essentially ‘plug-in’ the estimate $\hat{\pi}_{0}$, by applying the BH procedure at a threshold of $\alpha/\hat{\pi}_{0}$ instead of $\alpha$.
-Instead of the FDR, Guo and Romano’s stepdown approach [20], which we refer to as GR-SD, controls the FDP.
-In fact, this stepdown approach inspired the development of FDP-SD in the competition setting.
-GR-SD works by first sorting the p-values in ascending order, i.e., $p_{1}\leq p_{2}\leq...\leq p_{m}$.
-Then it constructs a sequence of thresholds, $\delta_{i}$ for each p-value $p_{i}$ and searches for the largest $k$ such that $p_{1}\leq\delta_{1},\dots,p_{k}\leq\delta_{k}$ mutually holds, and then reports the corresponding $k$ smallest p-values.
-We provide the pseudocode for this method in Algorithm S3.
+Multiple testing with FDR control was popularized by the BH procedure (Benjamini and Hochberg (1995)).
+Since then, numerous p-value-based methods have been developed to address multiple testing in two main ways.
+First, many methods build on the BH framework to improve statistical power by estimating the fraction of true null hypotheses denoted as $\pi_{0}$ (Storey et al.
+(2004); Benjamini et al.
+(2006); Hu et al.
+(2010)), reweighting p-values (Genovese et al.
+(2006); Blanchard and Roquain (2008); Roquain and Van De Wiel (2009); Habiger (2017); Ramdas et al.
+(2019)), or by reordering them (Barber and Candès (2015); Lei and Fithian (2016); G’Sell et al.
+(2016); Li and Barber (2017)).
+A second approach focuses on the stricter criterion of FDX control.
+One example of this second direction is the stepdown procedure proposed by Guo and Romano (2007), which we will refer to as GR-SD in our later analyses.
 3.1 Extension to side information
-Subsequent approaches were developed to take advantage of scenarios with available side information.
-Roughly speaking, these methods use side information to weight or rescore the hypotheses, placing higher weights or scores on hypotheses that are believed to be false nulls.
-Such procedures can be either ‘predetermined’, that is, the weighting or scoring is determined before the actual testing procedure takes place, or ‘adaptive’, in which the weighting or scoring is selected according to the data at hand and therefore often leads to greater statistical power.
+Table 1: Summary of some popular recent p-value based multiple testing procedures that use side information.
 Method Error control Speed
@@
 ZAP-asymp (2022) Asymptotic FDR Fast
-AdaFDR (2019) Asymptotic FDP Fast
-Table 1: Summary of some popular and recent p-value based multiple testing procedures that use side information, indicating each method’s type-1 error control and speed.
-Such adaptive methods include Adaptive P-value Thresholding (AdaPT) and its derivatives [33, 8], Z-value Adaptive Procedures (ZAP) [35] and AdaFDR [51].
+AdaFDR (2019) Asymptotic FDX Fast
+More recent works harness side information in a data-adaptive way to further boost statistical power.
+Such adaptive methods include Adaptive P-value Thresholding (AdaPT) and its derivatives (Lei and Fithian (2018); Chao and Fithian (2021)), Z-value Adaptive Procedures (ZAP) (Leung and Sun (2022)) and AdaFDR (Zhang et al.
+(2019)).
 Table 1 provides a qualitative summary of each method in terms of its type-1 error control, its error control guarantees, and its computational speed.
-Notably, none of these mentioned tools offers a fast finite type-1 error control, and as we will see, some of these methods have variable power when considering a range of simulated and real datasets.
-We next provide some background on each of those methods.
-AdaFDR splits the pairs $(p_{i},\mathbf{x}_{i})_{i=1}^{m}$ into two folds.
-For each fold, the method fits a mixture model to optimally rescore the hypotheses, and the model is then applied to the other fold where a list of discoveries is obtained.
-The two list of discoveries are joined together and then reported.
-AdaFDR’s mixture model is a composition of Gaussian components to model local ‘bumps’ of genuine signals (false nulls) within the data plus an additional component that captures the monotonic relationship between the side information and the signals.
-AdaPT iteratively prunes a candidate set of hypotheses $\mathcal{S}$, similar to Adaptive Knockoffs, which was derived from it.
-At each iteration, AdaPT fits a two-group mixture model on a strict set of information: the p-values outside $\mathcal{S}$, the masked p-values in $\mathcal{S}$ given by $\{p_{i}\wedge(1-p_{i}):i\in\mathcal{S}\}$ where $1-p_{i}$ is referred to as the mirrored p-value, the side information $\mathbf{x}$, as well as #$\#\{p_{i}>1/2:i\in\mathcal{S}\}$ and #$\#\{p_{i}<1/2:i\in\mathcal{S}\}$.
-It then removes the hypothesis from $\mathcal{S}$ that is deemed most likely to be a true null based on the fitted model.
-To fit the mixture model, the EM algorithm is used.
-The M-step of the EM algorithm results in a regression problem which allows the user to flexibly use the regression method of their choice.
-The authors considered a generalized additive model (GAM), a generalized linear model (GLM), and a generalized linear model with $L_{1}$ regularization (GLMnet).
-If we identify each p-value $p_{i}>1/2$ with a label $L_{i}=-1$ and each p-value $p_{i}<1/2$ with a label $L_{i}=1$, then AdaPT terminates this pruning procedure when the estimated FDR, as in Equation (4), is $\leq\alpha$.
-The rationale is that for true null hypotheses, the p-values are usually uniformly distributed, so #$\#\{L_{i}=-1:i\in\mathcal{S}\}$ estimates the number of true null hypotheses with positive labels, #$\#\{L_{i}=1:i\in\mathcal{S}\}$.
-It then reports the remaining positively labeled hypotheses in $\mathcal{S}$ [^3] AdaPT and AdaPTg do not use labels explicitly — we introduce those here for notational convenience.
-AdaPTg generalizes AdaPT by constructing asymmetric regions to define the positive and negative labels.
-For example, the p-values $p_{i}\in\left(0.3,0.9\right)$ can be identified with a label $L_{i}=-1$ and the p-values $p_{i}<0.3$ can be identified with a label $L_{i}=1$.
-One advantage of defining these asymmetric regions, is to avoid overestimating the number of true null hypotheses in cases where their p-values are not uniformly distributed and there is a concentration of true null p-values near 1.
-Subsequently, the estimated FDR needs to be adjusted to account for the fact that we expect twice as many true null p-values with negative labels than positive labels.
-Moreover, the masked p-values are also no longer $p_{i}\wedge(1-p_{i})$, since $1-p_{i}$ no longer ‘correctly mirrors’ the values in $\left(0.3,0.9\right)$ onto $(0,0.3)$ and therefore needs adjustment.
-In essence, the rest of the procedure remains the same.
-Lastly, AdaPT-GMMg replaces the two group mixture model in AdaPT/AdaPTg for a Gaussian mixture model (GMM).
-Similar to AdaPT/AdaPTg, it also allows the user to choose from a collection of regression methods to fit the model at the M-step of the EM algorithm.
-Lastly, ZAP-asymp directly fits a beta mixture model to the data but is only able to offer asymptotic FDR control.
-Importantly, ZAP-asymp can operate directly on the test statistic instead of just the corresponding p-values to rerank the hypotheses.
-This is particularly advantageous for two-sided tests because the test statistic is mapped to its p-value in a non-bijective manner, losing information regarding the sign of the original test statistic along the way.
-Note that AdaPT-GMMg can also operate on test statistics, but in this paper, we focus on the p-value information.
+ZAP-asymp and AdaPT-GMMg also have the advantage of operating directly on the test statistic instead of the corresponding p-values, which can boost power.
+In this paper, we will focus our comparisons based on the p-value information.
+Notably, none of these mentioned tools offer a fast finite type-1 error control (FDR or FDX), and as we will see, some of these methods have variable power when considering a range of simulated and real datasets.
+We provide a description of these methods in Section C.1.
 The list of methods we examine here is incomplete.
-Others include Independent Hypothesis Weighting [25, 24] which splits the p-values and side information into several folds and learns a weight for each hypothesis within that fold using out-of-fold information.
-This was shown to control the FDR asymptotically, and with some later adjustments to the method, finite sample FDR control was achieved.
-Another is Structure Adaptive Benjamini Hochberg Algorithm (SABHA) [37], which performs a reweighting of the p-values (similar to IHW) with finite sample FDR control.
-In addition, ZAP-finite is the analogue of ZAP-asymp which offers finite-sample FDR control.
+Others include Independent Hypothesis Weighting (Ignatiadis et al.
+(2016); Ignatiadis and Huber (2021)) and Structure Adaptive Benjamini Hochberg Algorithm (SABHA) (Li and Barber (2019)) which both adaptively weight the p-values using side information.
+In addition, ZAP-finite is the analogue of ZAP-asymp which offers finite sample FDR control.
 It uses the same principles of the AdaPT methods where each hypothesis is pruned one at a time.
-However, in previous simulations and real data experiments, IHW and SABHA appeared to be consistently outperformed by some of the methods in Table 1.
+However, in previous simulations and real data experiments, IHW and SABHA appeared to be consistently outperformed by some of the methods in Table 1 of the main text.
 Because we consider the same simulations and real data experiments, we did not include those methods in our comparisons.
@@
 Given our comparisons to ZAP-asymp and the AdaPT methods, we decided to exclude ZAP-finite from our comparisons.
-4 RESET (REScoring via Estimating and Training)
-RESET is designed to control the FDR or the FDP in either the competition or p-value based setting.
-We next describe RESET’s main steps, with further details given in Algorithm S4.
-In Section S.3, we prove that RESET controls the FDR/FDP in the finite-sample setting.
-For simplicity, we use the target-decoy terminology, i.e., we refer to a hypothesis with $L_{i}=1$ as a ‘target win’ or a target for short, and $L_{i}=-1$ as a ‘decoy win’ or a decoy for short.
-4.1 RESET outline
+4 Methods
+RESET is designed to control the FDR or the FDX in either the competition or p-value based setting. We next describe RESET’s main steps, with further details given in Algorithm A4.
+4.1 RESET
   1. 1.
-    Intrinsically, RESET’s input consists of the labels, winning scores, and side information $(L,W,\mathbf{x})$.
+    Intrinsically, RESET’s input consists of the labels, $L=(L_{i})_{i=1}^{m}$, the winning scores, $W=(W_{i})_{i=1}^{m}$, and side information, $\mathbf{x}=(\mathbf{x}_{i})_{i=1}^{m}$.
 Hence, in the p-value setting, each $p_{i}$ is first converted to a pair $(L_{i},W_{i})$ in the following way:
     $L_{i}=\begin{cases}+1&p_{i}\in[0,a)\\ -1&p_{i}\in(b_{1},b_{2}]\\ \end{cases},\quad W_{i}=\begin{cases}\lvert\Phi^{-1}(p_{i})\rvert&p_{i}\in[0,a)\\ \lvert\Phi^{-1}((b_{2}-p_{i})\cdot\frac{a}{b_{2}-b_{1}})\rvert&p_{i}\in(b_{1},b_{2}]\\ \end{cases},$ (4)
-    where $0<a\leq 1/2\leq b_{1}<b_{2}\leq 1$ determine the cutoff regions for defining a positive and negative label, and $\Phi$ is the standard normal CDF.
-Hypotheses with p-values outside of $[0,a)\cup(b_{1},b_{2}]$ are thrown out.
+    where $0<a\leq b_{1}<b_{2}\leq 1$ and $a\leq 1/2$ determine the cutoff regions for defining a positive and negative label, and $\Phi$ is the standard normal CDF (the motivation for the introduction of $\Phi$ is discussed below).
+Hypotheses with p-values outside of $0,a)\cup(b_{1},b_{2}]$ are thrown out.
 The default is $a=b_{1}=1/2$ and $b_{2}=1$.
@@
     RESET independently and randomly assigns each winning decoy as a training decoy with probability $s$ (we used $s=1/2$ throughout).
-The complementary set of decoy wins defines the estimating decoy set, and the set of pseudo-targets is the set of all estimating decoys and target wins.
-The pseudo labels, $\tilde{L}_{i}$, denotes whether the $i$th hypothesis is a pseudo-target ($\tilde{L}_{i}=1$) or a training decoy ($\tilde{L}_{i}=-1$).
+The complementary set of decoy wins defines the estimating decoy set, and the set of pseudo targets is the set of all estimating decoys and target wins.
+The pseudo labels, $\tilde{L}_{i}$, denotes whether the $i$th hypothesis is a pseudo target ($\tilde{L}_{i}=1$) or a training decoy ($\tilde{L}_{i}=-1$).
   3. 3.
     Next, RESET applies a user-selected semi-supervised machine learning model which uses the pseudo labels, winning scores, and side information $(\tilde{L},W,\mathbf{x})$.
-Note that this is a semi-supervised task because the positive pseudo labels typically contain a mixture of true and false nulls, while the negative pseudo labels indicate true null hypotheses.
-    The output of this step is a rescoring of the training decoys and pseudo-targets, denoted as $\tilde{W}$.
-Ideally, $\tilde{W}$ scores many of the false nulls among the pseudo-targets higher than they were scored originally.
+Note that this is a semi-supervised task because the positive pseudo labels typically contain a mixture of true and false nulls, while the negative pseudo labels are mostly comprised of null hypotheses.
+    The output of this step is a rescoring of the training decoys and pseudo targets, denoted as $\widetilde{W}$.
+Ideally, $\widetilde{W}$ scores many of the false nulls among the pseudo targets higher than they were scored originally.
 We provide a specific framework that we developed for this crucial step in Section 4.4.
   4. 4.
-    With the training complete, the training decoys are thrown out, and the original labels $L$ of the remaining pseudo-targets are revealed.
+    With the training complete, the training decoys are thrown out, and the original labels $L$ of the remaining pseudo targets, $J:=\{i\in[m]:\tilde{L}_{i}=1\}$, are revealed.
   5. 5.
-    RESET then determines its list of discoveries by applying to the pseudo-targets using the original labels $L$ and the new scores $\tilde{W}$ either
-    • SSS+ (Algorithm S1) at the desired level $\alpha$ for FDR control, or
-    • FDP-SD (Algorithm S2) at the desired level $\alpha$ and confidence parameter $\gamma$ for FDP control.
+    Using the original labels $(L_{i})_{i\in J}$ and the new scores $(\widetilde{W}_{i})_{i\in J}$ RESET then determines its list of discoveries by applying either
+    1. (a)
+      SSS+ (Algorithm A1) at the desired level $\alpha$ for FDR control, or
+    2. (b)
+      FDP-SD (Algorithm A2) at the desired level $\alpha$ and confidence parameter $\gamma$ for FDX control.
     The two algorithms, SSS+ and FDP-SD, require an additional parameter $c$.
@@
 Indeed, assuming that the labels $L$ of the true nulls are i.i.d.
-uniform ±$\pm 1$ RVs, and $s=1/2$, we set the parameter $c=2/3$, so that the expected target-decoy ratio is 2.
-More generally, if we know that for a true null $\mathbb{P}(L_{i}=1)\leq c_{0}$, then we set $c=\frac{c_{0}}{1-s\cdot(1-c_{0})}$.
-For example in the p-value setting, $c_{0}=\frac{a}{a+b_{2}-b_{1}}$.
-This choice is made rigorous in Lemma S1.
+uniform $\pm 1$ RVs, and $s=1/2$, we set the parameter to $c=c_{e}:=2/3$, so that the expected target-decoy ratio is $\frac{c}{1-c}=2$.
+More generally, if we know that for a true null $\mathbb{P}(L_{i}=1)\leq c_{0}$, then we set $c=c_{e}:=\frac{c_{0}}{1-s\cdot(1-c_{0})}$.
+For example, in the p-value setting, $c_{0}=\frac{a}{a+b_{2}-b_{1}}$.
+This choice is made rigorous in Lemma 1.
 We note that the above application of $\Phi^{-1}$ in Equation (4) is not strictly necessary for type-1 error rate control.
-We apply $\Phi^{-1}$ to the p-value $p_{i}$ or the mirrored p-value $(b_{2}-p_{i})\cdot\frac{a}{b_{2}-b_{1}}$ because we found that doing so makes the semi-supervised learning task of Step 3 to be generally less variable and more powerful (data not shown).
-Crucially in Step 4, the training decoys are thrown out.
-This is important since during the training phase of Step 3, the semi-supervised machine learning model will have presumably learned to discriminate between the training decoys and the pseudo-targets.
+We apply $\Phi^{-1}$ to the p-value $p_{i}$ or the mirrored p-value, $(b_{2}-p_{i})\cdot\frac{a}{b_{2}-b_{1}}$, because we found that doing so makes the semi-supervised learning task of Step (3) to be generally less variable and more powerful (data not shown).
+Crucially, in Step (5) the training decoys are thrown out.
+This is important since during the training phase of Step (3), the semi-supervised machine learning model will have presumably learned to discriminate between the training decoys and the pseudo targets.
 Hence, true null targets will tend to score higher than the training decoys making the latter unsuitable for estimating the number of true nulls among the top scoring hypotheses.
-4.2 RESET controls the FDR or FDP
-To establish RESET’s control of the FDR or FDP in both the competition and p-value settings, we generalize Assumption 2 when the distribution of the true null labels $\{L_{i}:i\in N\}$ is non-uniform.
-This is given below:
+4.2 Other related work and contributions
+Recently, Marandon et al.
+(2024) developed a method called AdaDetect in the application of novelty detection.
+Given a sample of null observations $(Y_{1},\dots,Y_{n})$ with common marginal distribution $P_{0}$, the objective is to detect which observations in a test set $(X_{1},\dots,X_{m})$ are drawn from a different marginal distribution, i.e., $X_{i}\not\sim P_{0}$.
+AdaDetect uses a similar strategy to that of RESET by splitting the null observations into two parts, $(Y_{1},\dots,Y_{k})$ and $(Y_{k+1},\dots,Y_{n})$.
+The first part is used to assist in ranking the joint observations $(Y_{k+1},\dots,Y_{n},X_{1},\dots,X_{m})$ while the second part is used to construct empirical p-values (also called conformal p-values (Bates et al.
+(2023))) for each test set observation.
+Then, the BH procedure (or one of its $\pi_{0}$ derivatives) is applied.
+The critical assumption that guarantees AdaDetect’s FDR control is that the null observations $(Y_{1},\dots,Y_{n})$ and the null test set observations $(X_{i}:X_{i}\sim P_{0})$ are exchangeable, conditioned on $(X_{i}:X_{i}\not\sim P_{0})$.
+A natural question is whether a similar set of assumptions can be applied to the multiple hypothesis setting, so that an AdaDetect-like approach can be used to control the FDR.
+The answer is no, and we show in Section F.1 that applying AdaDetect’s idea can break FDR control.
+We recently developed Percolator-RESET (Freestone et al.
+(2025)) that combined the RESET approach with Percolator, a popular semi-supervised algorithm designed specifically for FDR control in peptide detection from mass spectrometry data (Käll et al.
+(2007)).
+This manuscript includes several novel contributions over Freestone et al.
+(2025).
+First, we introduce the general RESET algorithm, which is applicable to both p-value and competition-based multiple testing scenarios.
+Second, we also consider FDX control and provide theoretical justification of RESET’s finite sample FDR or FDX control (thereby also providing the previously-missing theoretical proof of Percolator-RESET’s FDR control).
+Third, we introduce “RESET Ensemble” as the new semi-supervised engine of RESET’s Step (c).
+Fourth, we conduct extensive simulated and real data experiments comparing RESET Ensemble with competing methods (including a comparison of the new RESET Ensemble to Adaptive Knockoffs in the peptide detection context).
+Fifth, we explore FDX control with side information for peptide detection.
+4.3 RESET controls the FDR or FDX
+To establish RESET’s control of the FDR or FDX in both the competition and p-value settings, we relax Assumption 2 so that it can be applied to the case where the true null labels are no longer necessarily uniformly or even identically distributed:
 Assumption 3.
 Let $N$ be the indices of the true null hypotheses.
-The labels $\{L_{i}:i\in N\}$ are i.i.d.
-±$\pm 1$ random variables with $\mathbb{P}(L_{i}=1)\leq c_{0}$ and are independent of all the scores $W$, the side information $\mathbf{x}$, and the labels of the false null hypotheses.
-In the p-value setting, Assumption 3 is satisfied under some mild conditions regarding the distribution of the true null p-values.
-The value of $c_{0}$ is determined by the cutoff regions $[0,a)$ and $(b_{1},b_{2}]$ which are used to assign the positive or negative labels (Equation (5)).
-This is given by the following Lemma originally established by Chao and Fithian [8].
-Lemma 1 (Chao and Fithian).
-Assume that the true null p-values have a non-decreasing density and that they are mutually independent and that they are independent of the false null p-values as well as of all the side information.
-Then Assumption 3 is satisfied with $c_{0}=\frac{a}{a+b_{2}-b_{1}}$.
-The above lemma is phrased differently by Chao and Fithian and is presented as Lemma A.1 in their paper, but it is materially the same.
-Most of the time, we will consider symmetric cutoff regions $[0,1/2)$ and $(1/2,1]$.
-For completeness, we provide a proof of this lemma in Supplementary Section S.2 using our notation.
-In practice, Assumption 3 (or the conditions of Lemma 1) needs to be verified.
-We provide reasonable heuristic justification of this assumption in our real data applications in Sections 5.2 and 6.2.
+Conditioned on $W$, the side information $\mathbf{x}$, and the labels of the false null hypotheses, the true null labels $(L_{i})_{i\in N}$ are independent $\pm 1$ random variables with $\mathbb{P}(L_{i}=1)=:c_{i}\leq c_{0}$.
+We discuss the applicability of this assumption in Sections 5 and 6 where we apply RESET in both the competition and p-value settings.
+RESET’s FDR/FDX control hinges on the following Lemma, which essentially states that the labels $L$ of the true null pseudo targets in Step (5) are conditionally independent $\pm 1$ RVs—even after we sample the training decoys in Step (2) and the application of the machine learning model in Step (3).
+Lemma 1.
+Let $\mathcal{G}:=\sigma\left(W,\mathbf{x},(L_{i})_{i\not\in N},(\tilde{L}_{i})_{i\in[m]}\right)$, $N$ be the indices of the true nulls, and $J:=\{i\in[m]:\tilde{L}_{i}=1\}$.
+Then, under Assumption 3, conditioned on $\mathcal{G}$, $(L_{i})_{i\in J\cap N}$ are independent $\pm 1$ RVs with
+$\mathbb{P}(L_{i}=-1\mid\mathcal{G})\geq 1-c_{e},$
+where $c_{e}:=\frac{c_{0}}{1-s\cdot(1-c_{0})}$.
+Proof.
+Note that we can express $\tilde{L}_{i}$ as $\tilde{L}_{i}=f(L_{i},B_{i})$ where $f$ is a deterministic function and $B_{i}$ are Bernoulli(s)(s) RVs that are independent given everything else.
+It follows from $\tilde{L}_{i}=f(L_{i},B_{i})$ and Assumption 3 that, conditioned on $\mathcal{G}_{0}:=\sigma\left(W,\mathbf{x},(L_{i})_{i\not\in N}\right)$, the pairs in $\{(L_{i},\tilde{L}_{i}):i\in N\}$ are independent of one another.
+Moreover, from $\tilde{L}_{i}=f(L_{i},B_{i})$ and $(L_{i})_{i\not\in N}\in\mathcal{G}_{0}$ we find that these pairs are still independent conditioned on $\sigma\left(\mathcal{G}_{0},(\tilde{L}_{i})_{i\not\in N}\right)$.
+This independence in turn implies that $(L_{i})_{i\in N}$ are independent if we further condition on all the pseudo labels, i.e., on $\mathcal{G}=\sigma\left(\mathcal{G}_{0},(\tilde{L}_{i})_{i\in[m]}\right)$.
+Since $J$ is fixed conditioned on $\mathcal{G}$, the labels $(L_{i})_{i\in J\cap N}$ are independent conditioned on $\mathcal{G}$, establishing the first part.
+It follows from Assumption 3 and the above discussion, that with $\mathcal{G}_{-i}:=\sigma\left(\mathcal{G}_{0},(\tilde{L}_{j})_{j\neq i}\right)$, $\mathbb{P}(L_{i}=-1\mid\mathcal{G}_{-i})\geq 1-c_{0}$.
+Consider $i\in N$ and an atom of $\mathcal{G}_{-i}$ where $W$, $\mathbf{x}$, $(L_{j})_{j\not\in N}$ and $(\tilde{L}_{j})_{j\neq i}$ are all fixed.
+Then construing the following as conditional on that atom we have
+$\mathbb{P}(L_{i}=-1,\tilde{L}_{i}=1)$ $=\mathbb{P}(L_{i}=-1)\cdot\mathbb{P}(\tilde{L}_{i}=1\mid L_{i}=-1)$
+$\geq(1-c_{0})\cdot(1-s).$
+Hence, under the same assumptions
+$\mathbb{P}(L_{i}=-1\mid\tilde{L}_{i}=1)$ $=\frac{\mathbb{P}(L_{i}=-1,\tilde{L}_{i}=1)}{P(\tilde{L}_{i}=1)}$
+$\geq\frac{(1-c_{0})\cdot(1-s)}{(1-s)\cdot\mathbb{P}(L_{i}=-1)+1\cdot\mathbb{P}(L_{i}=1)}$
+$=\frac{(1-c_{0})\cdot(1-s)}{1-s\cdot\mathbb{P}(L_{i}=-1)}$
+$\geq\frac{(1-c_{0})\cdot(1-s)}{1-s\cdot(1-c_{0})}$
+$=1-c_{e}.$
+Additionally, clearly
+$\mathbb{P}(L_{i}=-1\mid\tilde{L}_{i}=-1)=1\geq 1-c_{e}.$
+Therefore, under the same assumptions, the following holds a.s.
+$\mathbb{P}(L_{i}=-1\mid\mathcal{G})=\mathbb{P}(L_{i}=-1\mid\tilde{L}_{i},\mathcal{G}_{-i})\geq 1-c_{e}.$
+It is worth noting that Lemma 1 is not true for an arbitrary splitting of the decoys into the training and estimating sets, e.g., splitting the decoys into two halves of equal size.
+To see this, suppose that all hypotheses are true nulls, i.e., $N=[m]$.
+Then by observing the number of training decoys, we know that the number of true null decoys in $J$ must be the same, and the labels $(L_{i})_{i\in J\cap N}$ are no longer independent.
 We are now ready to state our two main theorems.
@@
 Theorem 2.
-Under Assumption 3 RESET controls the FDP at the user-specified threshold $\alpha$ and confidence $1-\gamma$.
-We defer the proof of RESET’s FDR and FDP control to Supplementary Section S.3.
-The proof requires an intermediate lemma (Lemma S1), which states that the labels $L$ of the true null pseudo-targets in Step 4 remain i.i.d.
-±$\pm 1$ RVs independently of everything else, even after sampling the training decoys in Step 2 and the application of the machine learning model in Step 3.
+Under Assumption 3 RESET controls the FDX at the user-specified threshold $\alpha$ and confidence $1-\gamma$.
+The proofs of Theorem 1 and 2 follow from Lemma 1 and rely on a technical discussion when the labels $(L_{i})_{i\in J\cap N}$, conditioned on $\mathcal{G}$, are non-identical, which we give in Section B.1.
+To briefly see the connection of Lemma 1 to Theorems 1 and 2, note that the labels $(L_{i})_{i\in J\cap N}$ are independent $\pm 1$ RVs even after observing the machine learning scores $\widetilde{W}$ in Step (3) and the false null labels $(L_{i})_{i\notin N}$.
+If the labels are identical, i.e., $\mathbb{P}(L_{i}=-1\mid\mathcal{G})=1-c_{e}$, then the assumptions of SSS+ in Barber and Candès (2015) and FDP-SD in Luo et al.
+(2023) are satisfied, and RESET controls the FDR or FDX, respectively.
+Intuitively, in the non-identical case, i.e., $\mathbb{P}(L_{i}=-1\mid\mathcal{G})\geq 1-c_{e}$, FDR or FDX control should be more conservative using these procedures.
 4.4 Semi-supervised approach
-In this section we describe our specific framework for RESET’s Step 3, which as mentioned, can use any semi-supervised machine learning model.
-It consists of two main steps: (1) we apply a range of classification algorithms to estimate a high quality positive set from the data, containing presumably many false nulls, and (2) this positive set is then subsequently used for a second application of the classification algorithms to rescore the hypotheses.
-  1. i.
-    Define the negative set as the set of all training decoys and the positive set as a subset of the top scoring pseudo-targets (which we elaborate on in Step III at the end of this section).
-We define the features that are subsequently used by a collection of classification algorithms as the combined score and side information $(W,\mathbf{x})$.
-  2. ii.
-    For each considered classification algorithm, we randomly split all the hypotheses into $K$ folds (we used $K=3$ throughout).
-For each $k\in\{1,2,\dots,K\}$, we train the classification algorithm using the positive and negative set on the features in the $K\setminus\{k\}$ folds and apply the trained model to all the hypotheses in the $k$th test fold.
+In this section we describe our specific framework for RESET’s Step (3), which can use any semi-supervised machine learning model.
+It consists of two main steps: (1) we apply a range of classification algorithms to estimate a high quality positive set from the data, presumably containing many false nulls, and (2) this positive set is then subsequently used for a second application of the classification algorithms to rescore the hypotheses.
+In Section C.2, we outline some additional heuristics that we employ to improve the procedure’s speed and power, as well as its applicability to dependent data.
+  1. (i)
+    We first define the negative and positive sets, as well as the features that are subsequently used by a collection of classification algorithms: the negative set is the set of all training decoys, the positive set is a subset of top scoring pseudo targets (Section C.2.3), and the features are the combined score and side information $\left(W,\mathbf{x}\right)$.
+  2. (ii)
+    We randomly split all the hypotheses into $K$ folds (we used $K=3$ throughout).
+For each considered classification algorithm, and for each $k\in[K]$, we train the classification algorithm using the features of the hypotheses in the subsets of the positive and negative sets that are in the $K]\setminus\{k\}$ training folds, and then we apply the trained model to all the hypotheses in the $k$th test fold.
 We consider a random forest (RF), a generalized additive model (GAM), and a collection of two-layer neural networks (NN) with decay parameter $\lambda\in\{0,0.1,1\}$ and a number of hidden layer nodes $h\in\{2,5,10\}$.
-  3. iii.
-    To reduce variability, we repeat Step ii $r$ times (we used $r=10$).
-  4. iv.
-    To evaluate each classification algorithm, we apply Selective SeqStep (SSS) (Algorithm S1, which is the similar to SSS+ of Equation (2) without the ‘+1’) at an FDR threshold $\alpha$ to each of the $rK$ test folds, using the scores obtained from the classification algorithm and the pseudo labels.
-If $\mathbb{P}(L_{i}=1)\leq c_{0}$ then we set $c=1-s\cdot(1-c_{0})$.
-This is to ensure that the ratio $\frac{c}{1-c}$ correctly accounts for the number of null pseudo-targets to training decoys [^4] This is not a requirement for error control..
-  5. v.
+  3. (iii)
+    To reduce variability, we repeat Step (ii) $r$ times (we used $r=10$).
+  4. (iv)
+    To evaluate each classification algorithm, we apply Selective SeqStep (SSS) (Algorithm A1, which is similar to SSS+ of Equation (1) but without the ‘+1’) at an FDR threshold $\alpha$ to each of the $rK$ test folds, using the scores obtained from the classification algorithm and the pseudo labels.
+We set $c=c_{t}:=1-s\cdot(1-c_{0})$ (recall that $\mathbb{P}(L_{i}=1)\leq c_{0}$ under Assumption 3).
+This is to ensure that the ratio $\frac{c}{1-c}$ correctly accounts for the number of null pseudo targets to training decoys (note, this is not a requirement of FDR/FDX control).
+  5. (v)
     We record the total number of ‘discoveries’ produced by SSS over the $rK$ test folds from the previous step and select the classification algorithm that maximizes this value.
-  6. vi.
-    Using the chosen classification algorithm, we assign new scores $\tilde{W}$ to the hypotheses.
-$\tilde{W}_{i}$ is defined as the average decision value for the $i$th hypothesis in the test fold that it appears in, taken over the $r$ repetitions executed in Step iii.
-  7. vii.
-    We reapply Selective SeqStep (SSS) at $\alpha$ using the pseudo labels and the updated scores, $(\tilde{L},\tilde{W})$ with $c=1-s\cdot(1-c_{0})$.
-  8. viii.
-    The positive set is redefined as those pseudo-targets that are ‘discovered’ in Step vii.
-Steps ii-vi are then repeated with the new positive set.
-To ensure that the positive set is not too small, we increase $\alpha$ by increments of 0.01 until the positive set reaches min_positive or until all the pseudo-targets are in the positive set.
-We used min_positive = 50.
-  9. ix.
-    The final scores $\tilde{W}$ are then reported.
-We apply the following additional heuristics to our implementation that occur prior to Step 1 and leave some minor details in Section S.4.
-  1. I.
-    RESET throws out the hypotheses with undefined labels, i.e., those hypotheses with $W_{i}=0$.
-This effect is negligible when using p-values.
-However, in variable selection, the LCD score in the knockoff filter forces many of the winning scores to be zero ($W_{i}=0$), leaving typically a few hypotheses to be considered by RESET and consequently losing some information.
-We try to recover some of this information in the following way.
-For each hypothesis with $W_{i}\neq 0$, we identify the k-nearest neighbors in terms of the side information (zero-scoring hypotheses included).
-Then we define an extra side information variable as the number of zero-scoring hypotheses among these identified neighbors (we used the 20-nearest neighbors).
-  2. II.
-    To reduce the effect of ‘noisy’ side information that is unable to distinguish between true and false nulls, we implement an initial side information selection procedure.
-That is, we use a generalized additive model to fit a smoothing spline using $W\cdot\tilde{L}$ as the response against each user-provided side information variable one at a time.
-We use tryCatch in case the smoothing spline fails, in which case we fit the response directly to the side information variable, that is, we use a linear fit.
-Again, we try to make use of some of the zero-scoring hypotheses setting $W\cdot\tilde{L}=0$ in these cases.
-We keep only the side information variables with sufficiently small p-values in the fitted model ($p<0.01$).
-  3. III.
-    Not all the pseudo-targets need to be included in the initial positive set in Step i.
-After all, we expect many of the pseudo-targets to be true nulls.
-Hence, to improve the performance of RESET, we consider the following strategy.
-For each feature in $(W,\mathbf{x})$, reorder the hypotheses according to that feature, and apply SSS to the pseudo labels $\tilde{L}$ at the FDR level $\alpha$.
-Then select the feature that maximizes the number of pseudo discoveries.
-Then using the chosen feature, we define the positive set as those pseudo-targets discovered at $\alpha_{0}$, again using SSS with the pseudo labels $\tilde{L}$ (we used $\alpha_{0}=50\%$ throughout).
-Usually the score is selected, but in some cases the side information is highly informative and is selected instead.
-    Users interested in a small FDR level, like $\alpha=1\%$, could probably benefit from reducing $\alpha_{0}$ to further improve the performance, and in particular the speed, of RESET (though as we mentioned, we kept $\alpha_{0}=50\%$ across the board).
-To ensure that the positive set is not too small, we increase $\alpha_{0}$ by increments of 0.01 until the positive set reaches min_positive or until all the pseudo-targets are in the positive set.
+  6. (vi)
+    Using the chosen classification algorithm, we assign new scores $\widetilde{W}$ to the hypotheses.
+$\widetilde{W}_{i}$ is defined as the average decision value for the $i$th hypothesis in the test fold that it appears in, taken over the $r$ repetitions executed in Step (iii).
+  7. (vii)
+[…]
+  8. (viii)
+    The positive set is redefined as those pseudo targets that are ‘discovered’ in Step (vii). Steps (ii)-(vi) are then repeated with the new positive set.
+  9. (ix)
+    The final scores $\widetilde{W}$ are then reported.
 We refer to the general RESET wrapper in combination with the above semi-supervised approach as RESET Ensemble.
-We refer to RESET RF/GAM/NN when the ensemble of machine learning methods in Step ii is reduced to a random forest, generalized additive model, and a collection of neural networks as described, respectively.
+We refer to RESET RF/GAM/NN when the ensemble of machine learning methods in Step (ii) is reduced to a random forest, generalized additive model, and a collection of neural networks as described, respectively.
 4.5 Code availability
-Our code is available at https://github.com/freejstone/stat_RESET_paper_code with an Apache-2.0 license. We intend to develop a package for RESET in R.
+Our code is available at https://github.com/freejstone/stat_RESET_paper_code with an Apache-2.0 licence. We intend on developing a package for RESET in R.
 5 Experiments: competition setting
-5.1 Numerical experiments
-In this section, we conduct a wide range of simulations in the variable selection problem based on those introduced by Ren and Candès [44].
-We compare RESET Ensemble to Adaptive Knockoffs (AdaKO) and the generic Knockoff Filter (KO) (see Section 2).
-Further implementation details for AdaKO are given in Supplementary Section S.5.
-We provide a brief description of each simulation setup, referring to [44] for further details.
-Each setup is designed to satisfy Assumption 3, which is required for our theoretical guarantees.
-The labels and scores are obtained according to Equation (1).
-We focus on FDR control in this section, and we study FDP control in Section 5.2.
-5.1.1 Simulation 1: Linear model with one-dimensional side information
-Our first simulation is based on Simulation 1 by Ren and Candès [44].
-They used $n=1000$ independent observations from the joint data distribution of $(X,Y)$, where $X$ is drawn from a hidden markov model (HMM) with $p=900$ variables and $Y|X\sim\mathcal{N}(X\beta,1)$ is a linear model.
-The $\beta$’s are selected in the following way.
-Let $k$ denote the number of nonzero $\beta$’s.
-Then $k$ indices from $\{1,2,\dots,p\}$ are sampled without replacement with indices $i\in\{1,\dots,2k\}$ having probability proportional to $1/i^{2}$ and the indices $i\in\{2k+1,\dots,p\}$ having zero probability.
-The $k$ chosen indices are the nonzero $\beta$’s and are assigned a value of ±$\pm 3.5/\sqrt{n}$ where the signs are chosen i.i.d.
-uniformly.
-We consider $k\in\{50,150,300\}$ while the original simulation only considers $k=150$.
-Each hypothesis is supported by the side information $\mathbf{x}_{i}=i$, which should help in detecting the nonzero $\beta$’s, because smaller values of $i$ are more likely to correspond to a nonzero $\beta_{i}$.
-We used 100 independent runs of the above, where each run consisted of drawing those $n$ observations.
-The original simulation in [44] sampled the $\beta$’s once and fixed them across the 100 runs.
-Here we redrew the $\beta$’s for each run so that we may average our results over a range of models, as we noticed that all the methods exhibited significant variability between the draws.
-5.1.2 Simulation 2: Linear model with two-dimensional side information
-In a second simulation, we considered a larger linear model where $n=2000$ observations are taken from the joint distribution $(X,Y)$ as described in Section 5.1.1, with $p=1800$.
-Accordingly, we also adjusted the number of false nulls to $k\in\{150,300,450\}$.
-Note that unlike the previous simulation, it is not relevant how the $k$ variable indices are selected from $\{1,2,\dots,p\}$.
-Each hypothesis is equipped with the following two-dimensional side information.
-For the $i$th hypothesis, we draw a pair of observations $\mathbf{x}_{i}$ from a bivariate normal $\mathcal{N}(\boldsymbol{\mu},I)$ with $\boldsymbol{\mu}=(0,0)$ for a true null hypothesis and $\boldsymbol{\mu}=\pm(2,2)$ for a false null hypothesis, where the signs are chosen i.i.d.
-uniformly.
-We used 20 independent runs of this larger simulation instead of 100, where in each run we draw those $n$ observations, $\beta$ coefficients, and side information variables.
-5.1.3 Simulation 3: Logistic model with two-dimensional side information
-Next, we consider Simulation 2 by Ren and Candès [44].
-They generate $n=1000$ observations with $p=1600$ variables sampled from the joint data distribution of $(X,Y)$, where $X$ is distributed according to a multivariate normal distribution and $Y|X\sim\text{Bernoulli}\left(e^{\beta\cdot X}/\left(1+e^{\beta\cdot X}\right)\right)$ is a logistic model (see [44] for further details).
-Each hypothesis’ side information is a unique pair of values $\mathbf{x}_{i}=\left(\mathbf{x}_{i1},\mathbf{x}_{i2}\right)$ from the lattice $[-20,19]\times[-20,19]\subseteq\mathbb{Z}^{2}$.
-The nonzero $\beta$’s are chosen according to the position of $\left(\mathbf{x}_{i1},\mathbf{x}_{i2}\right)$ in the lattice.
-An image of the nonzero $\beta$’s is given in Figure 1.
-Each nonzero $\beta$ is set to ±$\pm 25/\sqrt{n}$ where the signs are chosen i.i.d.
-uniformly.
-We used 98 runs of the above setup (since in 2 runs, AdaKO EM failed with an error), where each run consisted of drawing those $n$ observations.
-The original simulation in [44] sampled the sign of the $\beta$’s once, fixing them across the 100 runs.
-Instead, we redraw the sign of the $\beta$’s in each of the 98 runs since we found that the results of all the methods significantly varied with those draws.
-Figure 1: Location of nonzero $\beta$’s according to the pair of values $(x,y)$ in the lattice. The same image is given in [44].
-Figure 2: Scatterplot of each hypothesis according to the LCD score denoted as $W$ (x-axis) and side information variable (y-axis).
-Each hypothesis is colored according to whether it is a true (red) or false null hypothesis (blue).
-In (A), the side information variable is the index of the hypothesis in Simulation 1 (Section 5.1.1).
-In (B), we use one of the three side information variables as described in Simulation 4 (Section 5.1.4).
-5.1.4 Simulation 4: Large $p$ and $n$ and 3-d side information
-Finally, in the last simulation, $n=10K$ observations with $p=10K$ variables and $k=0.15p$ relevant variables are drawn from the joint data distribution $(X,Y)$ exactly as described in Simulation 1.
-Each variable is equipped with the following three-dimensional side information.
-The first $k$ variables are randomly assigned a value drawn from $i\in\{1,\dots,2k\}$ taken without replacement and having probability proportional to $1/i^{2}$.
-The remaining $p-k$ variables are uniquely assigned to one of the remaining values from $\{1,2,\dots,p\}$ uniformly at random.
-We expect most of the nonzero $\beta$’s to have small values of $i$, but the effect of the random assignment makes the resulting side information variable less informative than in Simulation 1, as depicted in Figure 2.
-This process is repeated three times, thus associating three side information variables with each hypothesis, which taken together, should be reasonably informative.
-We conducted 5 independent runs, where in each run we drew the $n$ observations, $\beta$ coefficients, and side information.
+In the following sections, we explore two applications of competition-based multiple testing.
+The first is in selecting relevant variables that are conditionally associated with a response.
+The second is in the detection of peptides from mass spectrometry data.
+5.1 Variable selection
+We conducted a wide range of simulations in the variable selection problem, where the goal is to identify relevant variables from $\{X_{i}:i=1,\dots,p\}$ that are conditionally associated to the response $y$.
+The variable selection problem uses a competition framework called the knockoff filter (Barber and Candès (2015); Candès et al.
+(2018)).
+Briefly, each variable $X_{i}$ is paired with an artificial knockoff variable, denoted as $\tilde{X}_{i}$.
+Each variable and knockoff is scored according to how well it correlates with the response $y$, producing $Z_{i}$ and $\tilde{Z}_{i}$, respectively.
+Then following Section 2, the pairs of scores can be converted into a winning score $W_{i}$ and a label $L_{i}$, and a competition-based multiple testing procedure can be applied (see Section C.3 for details).
+The following simulations, detailed in Section C.5, are based on those introduced by Ren and Candès (2023).
+Each setup, is designed to satisfy Assumption 3, and in each we compare RESET Ensemble with Adaptive Knockoffs (AdaKO, with its default GAM filter, see Section C.4 for details), as well as with the generic Knockoff Filter (KO) focusing on FDR control.
+  • In simulation 1 we consider a linear model with $n=1000$ observations with $p=900$ variables.
+We use one dimensional side-information $\mathbf{x}_{i}$ where lower values correspond to a greater chance of being a relevant feature ($\beta_{i}\neq 0$).
+We vary the number of relevant variables $k\in\{50,150,300\}$ and generate and analyze 100 datasets for each.
+  • In simulation 2 we consider a larger linear model with $n=2000$ and $p=1800$ with two-dimensional side information.
+The side information is drawn from a joint normal distribution, with different mean vectors for the relevant and irrelevant variables respectively.
+We vary the number of relevant variables $k\in\{150,300,450\}$ and generate and analyze 20 datasets for each.
+  • In simulation 3 we consider a high-dimensional logistic model with $n=1000$ observations and $p=1600$ variables.
+Relevant and irrelevant variables are determined according to the spatially dependent two-dimensional side information $\mathbf{x}_{i}=(\mathbf{x}_{i1},\mathbf{x}_{i2})$ (see Figure C9).
+We generate and analyze 100 datasets.
+  • In simulation 4 we consider a linear model with $n=10K$ observations and $p=10K$ variables with $k=0.15p$ relevant variables and three dimensional side information $\mathbf{x_{i}}=(\mathbf{x}_{i1},\mathbf{x}_{i2},\mathbf{x}_{i3})$.
+Lower values of the $\mathbf{x}_{ij}$’s correspond to a greater chance of being a relevant feature.
+We generate and analyze 5 datasets.
 5.1.1 Results
-We first compare RESET Ensemble with Adaptive Knockoff’s default method in R, AdaKO GAM, and the generic Knockoff Filter (KO).
 In Figure 1, for each simulation we plot the power of each method as a function of the selected FDR thresholds of $\alpha\in\{0.05,0.1,0.2,0.3\}$.
-The power is calculated as the proportion of relevant variables correctly discovered, averaged over the number of runs used in each simulation.
-Similarly, in Supplementary Figure S2, we plot the estimated FDR: the proportion of irrelevant variables in the discovery list, averaged over the runs.
-Figure 3: RESET Ensemble vs.
-AdaKO GAM vs.
-KO.
-Each panel plots the power for each method at FDR thresholds ranging from 5% to 30%.
+The power is calculated as the proportion of relevant variables correctly discovered, averaged over the number of datasets drawn for each simulation.
+Unsurprisingly, RESET Ensemble uniformly dominates KO, demonstrating that it successfully leverages the side information.
+At the same time, AdaKO GAM mostly improves on KO, with some exceptions.
+Figure 1: Comparing the power of RESET Ensemble, AdaKO GAM and KO.
+Each panel shows the power of each method at FDR thresholds ranging from 5% to 30%.
 The first row corresponds to Simulation 1 with three values of $k\in\{50,150,300\}$, the second row corresponds to Simulation 2 with three values of $k\in\{150,300,450\}$, and the last row corresponds to Simulation 3 and 4.
 For readability, the points are jittered in the horizontal direction.
-A description of AdaKO GAM and KO can be found in Section 2.
-Unsurprisingly, RESET Ensemble uniformly dominates KO, demonstrating that it successfully leverages the side information. At the same time, AdaKO GAM mostly improves on KO, with some exceptions.
 In Simulation 1, we find that at higher FDR thresholds, RESET Ensemble is the preferred approach in terms of power.
@@
 First, RESET trains its classification algorithms on a relatively small positive set of false nulls.
-Second, RESET incurs a greater cost in terms of power at lower FDR thresholds if the classification algorithm overfits and incorrectly scores a knockoff higher than it should.
-Indeed, at an FDR threshold of $\alpha=5\%$, each knockoff may cost up to $402/\alpha=$ 40 discoveries.
+Second, RESET incurs a greater cost in terms of power at lower FDR thresholds if the classification algorithm overfits and incorrectly scores a knockoff win higher than it should.
+Indeed, at an FDR threshold of $\alpha=5\%$, each knockoff win may cost up to $2/\alpha=40$ discoveries.
 We find in Simulations 2 and 3 that RESET Ensemble is consistently more powerful than AdaKO GAM; in fact, AdaKO GAM appears to essentially have the same power as KO.
-The result in Simulation 3 is consistent with the findings of Ren and Candès in their analogous simulation.
-Indeed, AdaKO GAM does not implement any smooth terms when the number of side information variables is two or more, and is therefore unable to leverage the side information if the true and false null hypotheses are not ‘easily’ separable.
+Indeed, The result in Simulation 3 is consistent with the findings of Ren and Candès in their analogous simulation.
+AdaKO GAM does not implement any smooth terms when the number of side information variables is two or more, effectively reducing to a linear model, and is therefore unable to leverage the side information if the true and false null hypotheses are not ‘easily’ separable.
 Finally in Simulation 4, we see that RESET Ensemble is generally the most powerful, with essentially the same power as AdaKO GAM at $\alpha=5\%$.
 We hypothesize that the greater number of variables in this example, particularly relevant variables, appears to overcome the relatively poor performance of RESET Ensemble in Simulation 1 at low FDR thresholds.
-Figure 4: Relative power and runtimes for each method.
-(A) Each boxplot describes the relative powers for each method, computed across the combination of simulations and FDR thresholds in Figure 3.
-(B) The log10 runtimes of each method applied at the 5% FDR level for 5 runs of Simulation 4.
-For readability, the points are jittered in the horizontal direction.
 We compared RESET Ensemble with AdaKO GAM because those are the respective default methods.
-We complement this comparison with Figure 4A where we summarize the performance of all RESET and Adaptive Knockoff methods.
-Specifically, for each panel and FDR threshold in Figure 3, we determined the maximum power across all methods and computed the relative power of each method to this maximum.
-The resulting relative powers are reported as boxplots, giving an indication of the performance of each method across the simulations as a whole.
-A plot of the actual powers for each of the methods is given in Supplementary Figure S3.
-While in Supplementary Figure S3 we often find a version of AdaKO that is more powerful, or at least comparable to RESET Ensemble, this varies between the versions.
-On the other hand, RESET Ensemble’s performance is consistent across these simulations, and overall typically achieves the highest relative power when the results are aggregated.
-Interestingly, RESET NN seems to be marginally more powerful than RESET Ensemble.
-While this may be the case, we still recommend RESET Ensemble as the default approach to alleviate the burden of selecting any single classification algorithm.
-One advantage of RESET is its computational speed.
-Indeed, our implementation of RESET, as described in Section 4.3, has many steps that can be parallelized: the $r$ evaluations of each classification algorithm on the $K$ folds can all be done in parallel using multiple cores.
-In contrast, Adaptive Knockoffs must iteratively apply a classification algorithm for every nonzero scoring hypothesis in the candidate set $\mathcal{S}$ until the estimated FDR is $\leq\alpha$, or all hypotheses are revealed.
-Figure 4B demonstrates this advantage showing the (log10) runtimes of RESET (using 10 of a 20-cores CPU) and of Adaptive Knockoffs at the 5% FDR level across 5 runs of Simulation 4.
-We find that each version of RESET is significantly faster than each version of Adaptive Knockoffs.
-RESET Ensemble is around $10^{3}$ to $10^{4}$ times faster than AdaKO EM, $10^{2}$ to $10^{3}$ times faster than AdaKO RF, and about 3-4 times faster than AdaKO LR and AdaKO GAM.
-In scenarios, as in Simulations 3 and 4, where AdaKO EM and AdaKO RF are infeasible, and AdaKO GAM and AdaKO LR are unable to leverage the side information RESET is clearly more suitable.
-We investigate such a commonly encountered scenario in the following application.
+In Section E.1, we complement this comparison where we summarize the performance of all RESET and AdaKO methods, and also analyze the runtimes of those methods.
 5.2 Application to peptide detection
-In this section we compare RESET to Adaptive Knockoffs in the peptide detection context using mass spectrometry data and show that (a) the Adaptive Knockoff methods are often computationally infeasible, (b) feasibility aside, RESET is more powerful than Adaptive Knockoff’s default filter, AdaKO GAM, when controlling the FDR, and (c) when controlling the FDP, RESET is able to typically discover more peptides than the vanilla FDP-SD.
-We provide a heuristic justification of Assumption 3, which is required for RESET’s theoretical guarantees, in Section S.15.
-5.2.1 HEK293 data
-We show that the application of Adaptive Knockoffs is often too slow to be used effectively in scenarios such as the peptide detection problem that involve a large amount of data.
-We demonstrate this using HEK293 data [9], a popular dataset used in benchmarking the development of new software [30, 50, 32] that we downloaded from the Proteomics Identification Database (PRIDE) [41].
-After data preparation and searching the combined 24 HEK293 spectrum files, we determined the scores $Z_{i}$ for each target peptide and $\tilde{Z}_{i}$ for each decoy peptide using the XCorr score function (see Supplementary Section S.12 for details).
-Next, we competed each pair by recording the score $W_{i}=Z_{i}\vee\tilde{Z}_{i}$ and label $L_{i}=1$ if $Z_{i}>\tilde{Z}_{i}$ and $L_{i}=-1$ if $Z_{i}<\tilde{Z}_{i}$ (ties were randomly broken).
-In addition to its primary score we associate with each peptide-spectrum match (PSM) a set of features, such as the charge state or peptide length.
-Each winning peptide inherits these features, or side information $\mathbf{x}_{i}$, from its maximally associated PSM, allowing us to define the desired triples $(L,W,\mathbf{x})$.
-A complete description of the side information used is given in Table S4 of the Supplement.
+The end-to-end strategy for identifying peptides in an unknown sample can be broken into three steps: data acquisition, search, and competition-based multiple testing.
+We next provide a brief summary on each of these steps with details given in Section C.6.
+During data acquisition, peptides are isolated and selected from a sample where each peptide generates a spectrum using a technique called liquid chromatography mass spectrometry (LC-MS/MS).
+In the search phase, each spectrum is searched against a database of peptides, called target peptides, for a possible candidate peptide that generated the spectrum.
+This search phase allows us to provide each target peptide with a score, $Z_{i}$, which quantifies evidence against the null hypothesis $H_{i}$ that ‘the target peptide is not in the sample’.
+Since not all target peptides are in the sample, each spectrum is searched against a database of artificially constructed peptides, called decoy peptides.
+The decoy peptides are obtained by randomly shuffling or reversing the target peptides.
+Each decoy peptide is therefore paired with a corresponding target peptide, and acquires a similar score from the search, $\tilde{Z}_{i}$.
+In the final testing phase, we obtain a winning score, $W_{i}=Z_{i}\vee\tilde{Z}_{i}$ and a label $L_{i}$ indicating whether the $i$th target peptide ($L_{i}=1$) or its corresponding decoy peptide ($L_{i}=-1$) scored higher.
+We can then apply the SSS+ procedure — with $c=1/2$ and $\alpha$ typically set to the $1\%$ FDR threshold — to $L$ and $W$ to obtain a list of target peptides believed to be in the sample.
+Note that the SSS+ procedure (with $c=1/2$) is commonly referred to as target-decoy competition or TDC in the proteomics community.
+In the following results, we show that (a) the Adaptive Knockoff methods are often computationally infeasible in this context, and (b) feasibility aside, RESET is more powerful than Adaptive Knockoff’s default filter, AdaKO GAM, when controlling the FDR.
+In Section F.2, we provide a heuristic justification of Assumption 3, which is required for RESET’s theoretical guarantees.
+In terms of FDX control, RESET typically discovers more peptides than the side-information-oblivious FDP-SD (Section E.3).
+As outlined in Section 4.2, these comparative analyses of RESET Ensemble are new and were not undertaken in Freestone et al.
+(2025).
+5.2.1 Runtime comparison
+We argue that the application of Adaptive Knockoffs is often too slow to be used effectively in scenarios such as the peptide detection problem that involve a large amount of data.
+We demonstrate this using HEK293 (Chick et al.
+(2015)), a popular dataset used in benchmarking the development of new software (Kong et al.
+(2017); Yu et al.
+(2020); Lazear (2023)) that we downloaded from the Proteomics Identification Database (PRIDE) (Martens et al.
+(2005)).
+Details regarding the HEK293 dataset and the search phase analysis are given in Section D.1.
+Table 2: The average runtimes for each method at 1% FDR on the HEK293 dataset.
 Method Time
-RESET Ensemble 27.5 minutes
-AdaKO GAM 1.06 days
-AdaKO EM 33.71 days
-AdaKO LR 3.35 days
-AdaKO RF 836.03 days
-Table 2: The average runtimes for each method at 1% FDR on the HEK293 dataset.
-The average runtimes reported for each Adaptive Knockoffs method is estimated (as described in Section S.13), while we computed the actual average runtime for RESET Ensemble.
-The average is taken over 5 randomly constructed decoy databases.
-We used 20 cores for RESET Ensemble.
-Finally, using the method described in Supplementary Section S.13, we estimated the time it would take for each method of Adaptive Knockoffs to report discoveries at the 1% FDR level (which is the typical FDR threshold in the peptide detection context).
-Table 2 shows the average estimated times taken over 5 distinct target-decoy databases, with decoys varying each time by randomly shuffling the target peptides.
-It is clear that running any implementation of AdaKO is generally too expensive, with the worst being AdaKO RF at an estimate of 836 days, and the default method, AdaKO GAM, taking an estimate of 1.06 days.
-In contrast, we were able to successfully run RESET Ensemble with FDR control in 27.5 minutes.
-5.2.2 PRIDE data
-Due to the aforementioned runtime problem, we had to use small datasets when evaluating the power of Adaptive Knockoffs in the peptide detection context.
-Specifically, we focused on thirteen spectrum files (each with $<$ 33K PSMs) from the twenty PRIDE-20 datasets that we recently used in [18].
-In the following results, we only considered searching the spectrum files once against a single target-decoy database, since unlike RESET we can only feasibly apply a single application of Adaptive Knockoffs.
-Data preparation, search settings and the detailed definition of the triples $(L,W,\mathbf{x})$ are given in Supplementary Section S.14 along with the list of side information used in Supplementary Table S5.555The side information here slightly differs from above due to different search settings.
-Figure 5: Comparison of RESET Ensemble and AdaKO GAM.
-For each of the 13 PRIDE datasets, we computed the average number of discoveries using 10 applications of RESET Ensemble, varying RESET’s internal seed.
-The figure shows a boxplot of those RESET Ensemble averages divided by the number of discoveries using AdaKO GAM (left boxplot) and TDC (right boxplot) using the 13 PRIDE-20 datasets (there are total of 13 points for each boxplot).
-Figure 5 shows the average number of discoveries reported using 10 applications of RESET Ensemble, each applied at 1% FDR to the same dataset, divided by the number of discoveries reported by AdaKO GAM (left boxplot) and TDC (right boxplot).
-In each of the 10 applications of RESET, we varied the internal seed which randomly splits the decoys into their training and estimating sets.
+RESET Ensemble 17.7 minutes
+AdaKO GAM 1.07 days
+AdaKO EM 34.17 days
+AdaKO LR 3.50 days
+AdaKO RF 846.33 days
+We estimated the time it would take for each method of Adaptive Knockoffs to report discoveries at the 1% FDR level (see Section C.14).
+Table 2 shows the estimated times averaged over 5 different constructions of the decoy database by randomly shuffling the target peptides.
+All the implementations of AdaKO are not practical for analyzing this data, with the worst being AdaKO RF at an estimated 846.33 days, and the default method, AdaKO GAM, taking an estimate of 1.07 days.
+In contrast, we were able to run RESET Ensemble in 17.7 minutes on average using 20 cores (Section C.13).
+5.2.2 Power comparison
+General applicability aside, we resorted to using small datasets to evaluate the power of Adaptive Knockoffs in the peptide detection context.
+Specifically, we focused on the thirteen smallest datasets from PRIDE-20, a collection of twenty datasets that we recently used (Freestone et al.
+(2024)).
+For each of the thirteen spectrum files in PRIDE-20, we performed two types of searches: a narrow and an open search.
+Briefly, in a narrow search, spectra can only match database peptides that are within a small tolerance of the machine-reported mass of the molecules that generated those spectra.
+On the other hand, in open searches the same tolerance is several orders of magnitude larger.
+Details regarding PRIDE-20 and the search settings are given in Section D.2.
+Figure 2: Comparison of RESET Ensemble to AdaKO GAM and TDC.
+For each of the PRIDE-20 datasets, we computed the average number of discoveries using 10 applications of RESET Ensemble, varying RESET’s internal seed.
+The figure shows a boxplot of those RESET Ensemble averages divided by the number of discoveries using AdaKO GAM (left boxplots) and TDC (right boxplots) applied to thirteen PRIDE-20 datasets and two different search modes (there are total of 13 points in each boxplot).
+Figure 2 shows the average number of discoveries reported using RESET Ensemble divided by the number of discoveries reported by AdaKO with its default GAM filter (left boxplots) and TDC (right boxplots), using the two different search settings.
+We averaged RESET Ensemble over 10 applications by varying the internal seed which randomly splits the decoys into their training and estimating sets.
 Clearly, RESET Ensemble is overall more powerful than AdaKO GAM and TDC.
-Interestingly, the relative ratios of the number of discoveries between AdaKO GAM and TDC suggest that AdaKO GAM is reporting fewer discoveries than the generic TDC.
-In a second comparison, we looked at the performance of RESET Ensemble versus AdaKO RF because in this setup it was the most powerful implementation of Adaptive Knockoffs.
-In Figure 6A, the left boxplot displays the average number of discoveries by RESET Ensemble, divided by the number of discoveries by AdaKO RF.
-In six of the thirteen PRIDE spectrum files, we found RESET Ensemble to be the superior method, while in the remaining seven spectrum files, AdaKO RF produced more discoveries.
-In one dataset, AdaKO RF discovered 1217 peptides while RESET Ensemble only discovered 1104 peptides, about a 10% increase over RESET Ensemble.
-We also assessed the computational times of AdaKO RF in Figure 6B, using both the actual and estimated runtimes as outlined in Section S.13.
-Reassuringly, our estimated times were less than the actual times taken by each method to complete, thus validating the conservative nature of our estimated times in the previous section.
-Although we find that, when analyzing these smaller spectrum sets, AdaKO RF is the most powerful method of Adaptive Knockoffs, and is competitive to RESET Ensemble power-wise, it is again clear that AdaKO RF is generally too slow with some runtimes exceeding multiple days even for these small datasets.
-In fact, on the dataset where AdaKO RF outperforms RESET Ensemble by about 10%, it takes 3.07 days for AdaKO RF to discover its peptides while RESET Ensemble takes about half a minute to complete.
-Figure 6C shows that in the worst case, RESET takes roughly 1 minute.
-The issue of runtimes is especially problematic considering the fact that multiple spectrum files are often analyzed jointly, as in our HEK293 analysis in the previous section.
-Figure 6: Comparison of RESET Ensemble and AdaKO RF.
-For each of the 13 PRIDE datasets, we computed the average number of discoveries using 10 applications of RESET Ensemble, varying RESET’s internal seed.
-The left panel shows a boxplot of those average RESET Ensemble discoveries divided by the number of discoveries using AdaKO RF (left boxplot) and TDC (right boxplot) using the 13 PRIDE-20 datasets (there are total of 13 points for each boxplot).
-In the centre panel, we compare the estimated and actual runtimes for AdaKO RF.
-Lastly, in the right panel, we compare the actual runtimes of RESET Ensemble and AdaKO RF (in log10 scale).
-5.2.3 FDP control
-Lastly, we evaluated RESET Ensemble’s FDP control using the HEK293 and all twenty PRIDE-20 spectrum files from the previous sections.
-As in Section 5.2.1, the joint HEK293 spectrum files were searched against 5 distinct target-decoy databases, where each target-decoy database was generated by concatenating the target database to a decoy database consisting of shuffled target peptides.
-Similarly, each PRIDE-20 spectrum file was searched against 10 distinct target-decoy databases.
-We used the coinflip version of FDP-SD [40], which provides a uniform increase in the number of discoveries (at the cost of some randomness in the discovery list, see Supplementary Algorithm S2).
-Our results using RESET Ensemble and FDP-SD are given in Table 3 and Figure 7.
-Confidence $1-\gamma$
-Method 0.5 0.8 0.9
-RESET Ensemble 82987 82729 82558
-FDP-SD 75977 75464 75254
-Table 3: The average number of discoveries at 1% FDP threshold at varying confidence parameters using the HEK293 data.
-The average number of discoveries reported for RESET Ensemble and FDP-SD on the combined HEK293 spectrum files at different confidence values $0.91-\gamma=0.5,0.8,$ 0.9.
-The averages are taken over 5 applications of each method, once for each of the 5 combined target-decoy databases.
-Analyzing the HEK293 results first, RESET Ensemble is clearly more powerful, with an approximately 9–10% increase in the average number of discoveries over FDP-SD at all confidence levels.
-Analyzing the PRIDE-20 datasets, while RESET Ensemble appears to be generally more powerful, in a few datasets it fails to make any discoveries at high confidence levels.
-Upon a closer look, the numbers of discoveries made by FDP-SD in these datasets are relatively small, with the largest average number of discoveries among these datasets being 417.
-When the number of discoveries is expected to be small and the confidence parameter is high, RESET may yield less power, even if the side information is reasonably informative — a phenomenon we dwell on in the Discussion.
-Regardless, more often than not RESET Ensemble produces more discoveries, and if multiple spectrum files are analyzed jointly, as we do with the HEK293 spectrum files, then the risk of having too few discoveries is reduced.
-Figure 7: Comparison of RESET Ensemble and FDP-SD.
-The left panel shows boxplots of the average number of discoveries using RESET divided by the average number of discoveries using FDP-SD for each of the twenty PRIDE-20 datasets at a 1% FDP threshold and varying confidence parameters.
-The averages are taken over 10 applications, one for each of the 10 target-decoy databases used for each PRIDE-20 dataset.
-The right panel is a ‘zoomed’ in version of the left panel.
-The results above consider a canonical ‘narrow’ search.
-Specifically, during the searching phase, each database peptide and spectrum are only allowed to match if the mass of the database peptide and the observed mass of the sample peptide generating the spectrum, the ‘spectrum mass’, differ by a small amount.
-Recently, ‘open’ searching, where each spectrum and peptide can match even if their masses differ by hundreds of Daltons, has become a popular method for detecting peptides [30].
-Such a search allows for the detection of peptides whose masses have been altered by post-translational modifications (PTMs).
-We provide further detail in Supplementary Section S.14.
-In this setting, side information variables such as dm, which records the mass difference between the peptide and the spectrum, can be highly informative given that some differences will correspond to masses of common PTMs.
-Thus, it is not surprising that RESET, which can take advantage of such side information, improves on FDP-SD even further across all confidence levels in this context (Figure 8).
-Figure 8: Comparison of RESET ensemble and FDP-SD using the recently popular ‘open’ search.
-The left panel shows boxplots of the average number of discoveries using RESET divided by the average number of discoveries using FDP-SD for each of the PRIDE-20 datasets at a 1% FDP threshold and varying confidence parameters.
-We implement an ‘open’ search for each PRIDE-20 dataset (see Section S.14).
-The averages are taken over 10 applications, one for each of the 10 target-decoy databases used for each PRIDE-20 dataset.
-The right panel is a ‘zoomed’ in version of the left panel.
+Interestingly, the relative ratios of the number of discoveries between AdaKO GAM and TDC suggest that AdaKO GAM is at times reporting fewer discoveries than the generic TDC.
+In Section E.2, we compare RESET Ensemble with AdaKO RF since we found AdaKO RF to be the most powerful variant of Adaptive Knockoffs in this setting.
+The results show that RESET Ensemble is comparable to AdaKO RF in terms of power, with the latter still being generally too slow to be practical.
 6 Experiments: p-value setting
-6.1 Numerical experiments
-In this section, we consider the p-value setting, comparing RESET with the FDR controlling procedures we reviewed in Section 3.1 on a set of simulations introduced by Lei and Fithian [33].
-As in the competition setting, each setup is designed to satisfy Assumption 3, which is required for our theoretical guarantees.
-Details of each method used are given in Supplementary Sections S.6S.10.
+6.1 Assumption guaranteeing finite-sample FDR control
+Recall that RESET and Assumption 3 that guarantees its finite-sample error control are formulated in terms of the labels and winning score $(L,W)$.
+To apply RESET in the p-value setting, we first convert each $p_{i}$ into a pair $(L_{i},W_{i})$ as in Equation (4).
+Similarly, instead of Assumption 3 we adopt the following assumption formulated by Chao and Fithian (2021) who also showed it implies the former assumption (the next lemma).
+Assumption 4.
+Let $N$ be the indices of the true null hypotheses.
+Conditioned on the side information $\mathbf{x}$, and the false null p-values $(p_{i})_{i\not\in N}$, the true null p-values $(p_{i})_{i\in N}$ are independent and have a non-decreasing density.
+Lemma 2 (Chao and Fithian).
+Assumption 4 implies Assumption 3 with $c_{0}=\frac{a}{a+b_{2}-b_{1}}$, where $[0,a)$ and $(b_{1},b_{2}]$ are the regions corresponding to the positive and negative labels as described in Equation (4).
+The above lemma is phrased differently by Chao and Fithian (2021) and is presented as Lemma A.1 in their paper, but it is materially the same.
+Most of the time, we will consider symmetric cutoff regions $[0,1/2)$ and $(1/2,1]$.
+For completeness, we provide a proof of this lemma in Section B.2 using our notation.
+6.2 Simulated data
+In this section, we consider the p-value setting, comparing RESET with the procedures we introduced in Section 3.1 on a set of simulations introduced by Lei and Fithian (2018).
+Each setup, detailed through Sections C.7C.11, is designed to satisfy Assumption 4 and hence Assumption 3, which is required for our theoretical guarantees.
 6.2.1 Simulation 5: p-value based testing with geometric side information
-In this section, we look at Simulation 1 of Lei and Fithian [33].
+In this section, we look at Simulation 1 of Lei and Fithian (2018).
 For each hypothesis, they obtain a p-value $p_{i}$ from a one-sided normal test so that $p_{i}=1-\Phi(z_{i})$ where $z_{i}\sim\mathcal{N}(\mu,1)$ with $\mu=2$ for a false null hypothesis and $\mu=0$ for a true null hypothesis.
 A total of $m=2500$ hypotheses are used.
-Similar to Simulation 3 of Section 5.1.3, each hypothesis’ side information is defined as a unique pair of values $\mathbf{x}_{i}=\left(\mathbf{x}_{i1},\mathbf{x}_{i2}\right)$ from $50\times 50$ equally spaced values in the square $\left[-100,100\right]\times\left[-100,100\right]$.
+Each hypothesis’ side information is defined as a unique pair of values $\mathbf{x}_{i}=\left(\mathbf{x}_{i1},\mathbf{x}_{i2}\right)$ from $50\times 50$ equally spaced values in the square $\left[-100,100\right]\times\left[-100,100\right]$.
 The false null hypotheses correspond to certain regions in the square.
-They considered three scenarios: (a) when the false nulls form a circle in the middle of the square, (b) when the false nulls form a circle in the top right of the square and (c) when the false nulls form an ellipse (Figure 9).
-We used 100 independent runs, where in each run we randomly draw the 2500 p-values as described.
-Figure 9: Location of false null hypotheses according to the pair of values $\left(\mathbf{x}_{i1},\mathbf{x}_{i2}\right)$ in the square $\left[-100,100\right]\times\left[-100,100\right]$.
-The same picture is given in [33].
+They considered three scenarios: (a) when the false nulls form a circle in the middle of the square, (b) when the false nulls form a circle in the top right of the square and (c) when the false nulls form an ellipse (Figure C12).
+We generated and analyzed 100 independent datasets, where in each we randomly draw the 2500 p-values as described.
 Figure 3: RESET vs.
-other methods using two-dimensional side information with p-values.
-The first row plots the power and the second row plots the estimated FDR for each FDR threshold ranging from 1% to 20%.
+other methods using p-values with two-dimensional side information.
+The power (1st row) and the estimated FDR (2nd row) are plotted against FDR thresholds ranging from 1% to 20%.
 Each color and shape combination corresponds to a unique method.
-The three columns correspond to the hypotheses’ relationship to the side information according to Figure 9.
-The horizontal segments in the second row of panels mark the corresponding FDR thresholds on the y-axis.
+The three columns correspond to the hypotheses’ relationship to the side information according to Figure C12.
+The black horizontal segments in the second row of panels mark the corresponding FDR thresholds on the y-axis.
 For readability, each point is jittered in the horizontal direction.
@@
 We find that in (a), AdaFDR’s estimated FDR at 20% (21.2% estimated FDR with 50% power) slightly exceeds the FDR threshold.
-This is possibly explained by the fact that AdaFDR only guarantees asymptotic FDP control.
-Consistently, Chao and Fithian [8] reported inflated estimated FDRs at high thresholds in simulation studies for AdaFDR.
+This is possibly explained by the fact that AdaFDR only guarantees asymptotic FDX control.
+Consistently, Chao and Fithian (2021) reported inflated estimated FDRs at high thresholds in simulation studies for AdaFDR.
 6.2.2 Simulation 6: p-value based testing with 100-dimensional side information
-In Simulation 2 of Lei and Fithian [33], the side information for $m=2000$ hypotheses consists of 100 i.i.d.
-draws from $U[0,1]$.
-This setup simulates ‘noisy’ side information since only the first 2 of the 100 draws are used in distinguishing between the true and false null hypotheses.
-Each p-value $p_{i}$ is drawn i.i.d.
-from a two-group beta mixture model, where the true nulls have a uniform distribution and the false nulls have an alternative beta distribution.
-Specifically, the beta mixture model is given by:
-$f(p_{i}\mid\mathbf{x}_{i})=1-\pi(\mathbf{x}_{i})+\pi(\mathbf{x}_{i})\cdot h(p_{i};\mu(\mathbf{x}_{i})),\quad h(p_{i};\mu(\mathbf{x}_{i}))=\left(\frac{1}{\mu(\mathbf{x}_{i})}p_{i}^{\frac{1}{\mu(\mathbf{x}_{i})}-1}\right),$ (6)
-where $\mathbf{x}_{i}\in[0,1]^{100}$ denotes the side information of the $i$th hypothesis, $\pi(\mathbf{x}_{i})$ denotes proportion of the false nulls in the mixture, $h$ denotes the false null distribution of $Beta(1/\mu(\mathbf{x}_{i}),1)$.
-The $\pi(\mathbf{x}_{i})$’s and $\mu(\mathbf{x}_{i})$’s are determined by the following relationships with the 100-dimensional side information:
-$\log\left(\frac{\pi}{1-\pi}\right)=\theta_{0}+\mathbf{x}_{i}^{T}\theta,\quad\mu_{i}=\max\{\mathbf{x}_{i}^{T}\beta,1\},$
-where $\theta=(3,3,0,\dots,0)\in\mathbb{R}^{100}$, $\beta=(2,2,0,\dots,0)\in\mathbb{R}^{100}$ and $\theta_{0}$ is chosen to satisfy $\frac{1}{m}\sum_{i}\pi(\mathbf{x}_{i})=0.3$.
-Clearly, only the first two values in $\mathbf{x}_{i}\in[0,1]^{100}$ contribute to the beta mixture model, as intended.
-We used 100 independent runs, where in each we drew the 2000 p-values as described.
+We next adopted Simulation 2 of Lei and Fithian (2018).
+The side information for $m=2000$ hypotheses consists of 100 i.i.d.
+draws from $U[0,1]$ in which only the first 2 of the 100 draws are used in distinguishing between the true and false null hypotheses.
+Details are given in Section C.12.
+We conducted 100 independent runs, where in each run we randomly draw the $m=2000$ p-values.
 Figure 4: RESET vs.
-other methods using 100-dimensional side information with p-values.
-The left panel plots the power and the right panel plots the estimated FDR as a function of the FDR thresholds: 1%, 5%, 10% and 20%.
+other methods using p-values with 100-dimensional side information.
+The power (left panel) and the estimated FDR (right panel) are plotted as a function of the FDR thresholds: 1%, 5%, 10% and 20%.
 Each color and shape combination corresponds to a unique method.
-The horizontal segments in the right panel mark the corresponding FDR thresholds on the y-axis.
+The black horizontal segments in the right panel mark the corresponding FDR thresholds on the y-axis.
 For readability, each point is jittered in the horizontal direction.
@@
 At 1% FDR, only AdaPTg GLMnet appears to have non-negligible power, albeit a fairly low number of discoveries.
-Finally, we point out that in this simulation, the model specification of AdaPT/AdaPTg perfectly matches the true model.
-In other words, the fitted mixture model used by AdaPT/AdaPTg coincides with Equation (6).
+Finally, we point out that in this simulation, the model specification of AdaPT/AdaPTg perfectly matches the underlying data generating model: the fitted mixture model used by AdaPT/AdaPTg coincides with Equation (7) of Section C.12.
 Given this, we find it reassuring that RESET Ensemble is still comparable at 5% and higher thresholds.
@@
 In contrast, RESET Ensemble selects from a collection of flexible machine learning models to adapt itself to the problem at hand.
+6.3 Robustness to dependent data
+In reality, real datasets often exhibit varying degrees of dependency, therefore failing to satisfy Assumptions 3 and 4.
+RESET Ensemble is equipped with a heuristic to account for such data (Section C.2.4).
+The following simulations adapted from Zhang et al.
+(2019) empirically demonstrate that RESET Ensemble exhibits some robustness to both weak and strong forms of local dependency.
+6.3.1 Simulation 7: Weak dependent p-values
+This simulation is based on Data 4 of Zhang et al.
+(2019).
+We generated $m=20K$ hypotheses with i.i.d side information drawn from $\mathbf{x}_{i}\sim U[0,1]$.
+Each hypothesis was declared a false null with probability $\pi_{i}:=\pi(\mathbf{x}_{i})$ which is dependent on the drawn side-information.
+The function $\pi$ is a convex combination of (truncated) normal densities and an exponential function (Section C.12.2).
+For every 10 consecutive hypotheses, starting with $i=1,\dots,10$, we drew $z_{i}$ jointly from a $\mathcal{N}(0,\Sigma)$ where $\Sigma$ is a covariance matrix with upper triangular matrix given by $\Sigma_{jj}=1$, $\Sigma_{jk}=0.25$ for $1\leq j<k\leq 5$, $\Sigma_{jk}=-0.25$ for $1\leq j\leq 5<k\leq 10$ and $\Sigma_{jk}=0$ for all other entries in the upper triangle.
+Hypotheses which are false nulls had their z-scores shifted by 2.
+A one-sided p-value $p_{i}=1-\Phi(z_{i})$ is then obtained.
+Notably, the covariance matrix $\Sigma$ induces a local weak dependency among the p-values, effecting both true and false null hypotheses.
+6.3.2 Simulation 8: Strong dependent p-values
+This simulation is based on Data 5 of Zhang et al.
+(2019).
+We used $m=20K$ hypotheses with side information $\mathbf{x}_{i}$ and false null probability $\pi_{i}$ defined as in Simulation 7.
+For every 5 consecutive hypotheses, starting with $i=1,\dots,5$, $z_{i}=z$ where $z\sim\mathcal{N}(0,1)$, creating 5 identical z-scores.
+False null hypotheses are shifted by 2 and one-sided p-values $p_{i}=1-\Phi(z_{i})$ were obtained.
+In this case, the data exhibits a local strong dependency effecting both true and false null p-values.
+6.3.3 Results
+Figure 5 shows the estimated FDR calculated from 100 runs of the above simulations.
+In the weakly dependent case (A), the majority of the methods seem to empirically control the FDR, with the exception of AdaFDR and RESET GAM which appears to marginally exceed the 5% threshold.
+In the strongly dependent data (B), we see several more methods exceeding their respective thresholds, albeit relatively mildly.
+RESET Ensemble seems to be robust and appears to approximately control the FDR in these experiments.
+Figure 5: Estimated FDR using RESET vs.
+other methods on dependent p-values.
+The estimated FDR is plotted as a function of the FDR thresholds: 1%, 5%, 10% and 20% for weak (A) and strong dependence (B).
+Each color and shape combination corresponds to a unique method.
+The black horizontal segments mark the corresponding FDR thresholds on the y-axis.
+For readability, each point is jittered in the horizontal direction.
 6.4 Real data experiments with p-values
-We next evaluate RESET and related methods on a collection of publicly available datasets with p-values that were investigated by Zhang et al.
-[51] when introducing AdaFDR and by Lei and Fithian [33] when introducing AdaPT.
-Since AdaFDR and AdaPT operate under essentially the same assumptions as RESET’s, we expect these datasets to therefore satisfy Assumption 3 with the exception of two datasets, which we discuss at the end of this section.
-We provide a brief summary of each dataset below.
-  • Three RNA-seq datasets (Airway [22], Bottomly [4], Pasilla [6]): The goal is to identify genes that are associated with varying gene expression levels in response to a change of conditions.
-In the Airway data, the differential expression analysis is conducted in response to a drug, dexamethasone; in the Bottomly data, the analysis is conducted between two mouse strains; and in the Pasilla data, the analysis is conducted between normal and so-called ‘Pasilla knockdown’ conditions [23].
-In all cases the side information consists of the log-normalized gene counts.
-  • A proteomics dataset (Proteomics [14]): The goal is to identify proteins that have different protein abundances when treated with rapamycin versus DMSO in yeast cells.
-Each protein consists of side information equal to the natural log of the number of peptides ‘quantified’ across all samples, i.e., the number of peptides belonging to the protein with their abundances measured.
-  • Two microbiome datasets (Microbiome Enigma Ph and Microbiome Enigma Al [45, 31]): The goal is to identify microbiome organisms, recorded as OTUs (Operational Taxonomic Units), that are associated with certain conditions (pH and Al).
-Each hypothesis (OTU) is equipped with a two-dimensional side information — the ubiquity, which is defined as the percentage of samples detected with the OTU, and the mean nonzero abundance, which is defined as the average abundance of each OTU across the samples in which it was detected.
-  • Two fMRI datasets (Auditory and Imagination) [48]: The goal is to identify voxels that are activated in response to different stimuli.
-In the first dataset, a single individual received auditory stimulus and in the second dataset, a single individual was instructed to imagine playing tennis.
-Each hypothesis (voxel) is equipped with four dimensional side information: the spatial position of the voxel (in three dimensions) and a categorical variable with the Brodmann label [5], which is used to delineate different areas of the brain with different functions.
-  • Two eQTL datasets (Subcutaneous and Omentum [11, 10]): The goal is to identify single nucleotide polymorphisms (SNPs) that are associated with changes in gene expression levels.
-Such an association is referred to as an expression quantitative trait locus (eQTL).
-Each hypothesis (SNP) is equipped with the following four side information variables: (1) the distance between the SNP and the gene transcription’s start site (TSS), (2) the log gene expression level, (3) the alternative allele frequency (AAF) of the SNP and (4) the chromatin state of the SNP.
-  • Gene-drug response data (Estrogen [12, 36]): The goal is to identify genes that are differentially expressed in response to a low-dosage estrogen treatment applied to breast cancer cells.
-Each hypothesis (gene) is assigned a ‘rank’ that was generated using an analysis of the gene expression levels in response to (1) a high-dosage estrogen treatment and (2) a medium-dosage estrogen treatment.
-Here, the ranks are in terms of the strength of association between the expression level of each gene and the level of dosage in (1) and (2).
-Thus, we can analyze the data twice, once for each side information variable based on (1) and (2).
-Figure 12: Comparison of RESET and other methods using a collection of publicly available datasets.
-Each method was evaluated at the FDR thresholds of 1%, 5%, 10%, and 20%.
-The number of discoveries reported by each RESET method is averaged over 10 applications at FDR threshold and dataset.
-Each color and shape combination corresponds to a unique method.
-For readability, each point is jittered in the horizontal direction.
-The results of each method applied to the collection of the above datasets is visually summarized in Figure 12.
-We can clearly see that RESET Ensemble is comparable to the other methods in terms of discoveries.
-Indeed, to argue this more accurately, we ranked each method across each dataset and at each FDR threshold considered (ties were averaged).
-Then we calculated the median rank for each method.
-RESET Ensemble and AdaPTg-GMM were tied with the highest median rank.
-Moreover, for each dataset and each FDR threshold in Figure 12, we computed the maximum number of discoveries across all methods and determined the relative power of each method to this maximum.
-The boxplots in Figure 13 display the relative powers for each method.
-Visually, RESET Ensemble and AdaPTg-GMM have the highest medians, with RESET Ensemble at 95.9% and AdaPTg-GMM at 94.5% with RESET Ensemble’s boxplot edging higher overall.
-Figure 13: Relative power of each method. Each boxplot describes the relative power of each method computed for each dataset and FDR threshold in Figure 12.
-There are a couple of implementation details we would like to highlight here (for full details, see Supplementary Sections S.6-S.10).
-First, we used the same default options of RESET Ensemble on all datasets, except for the two fMRI datasets which exhibit a spike of p-values close to 1.
-In this case, it is sensible to define asymmetric target and decoy regions, $[0,0.3)$ and $(0.3,0.9]$ respectively as Chao and Fithian [8] did in their application of AdaPTg-GMM to this dataset.
-We also point out that we applied AdaPTg-GMM and ZAP using the p-value information rather than the test statistic information (see Section 3.1 for background on these methods).
+We next evaluate RESET and related methods on a collection of publicly available datasets with p-values.
+These were investigated by Zhang et al.
+(2019), when introducing AdaFDR, and by Lei and Fithian (2018) when introducing AdaPT.
+A description of each dataset can be found in Section D.3.
+As we saw above, RESET Ensemble and the other tools we analyze here exhibit some robustness to deviations from the independence assumption that is built into Assumptions 3 and 4.
+However, despite this fact, we sought to remove the two Adipose datasets from this analysis because RESET’s heuristics, described in Section C.2.4, flagged these datasets for high dependency (cf.
+Section F.3).
+Figure 6: Power comparison.
+Each boxplot is made of 40 data points showing the relative power of a method on all combinations of dataset and FDR threshold.
+There are ten dataset as described in Section D.3 (excluding the two Adipose datasets) and four FDR thresholds: 1%, 5%, 10% and 20%.
+RESET Ensemble’s power and its variants were averaged over 10 applications.
+For each dataset (excluding the two Adipose datasets) and each FDR threshold in $\{1\%,5\%,10\%,20\%\}$, we computed the maximum number of discoveries across all methods and determined the power of each method relative to this maximum (Figure 6).
+RESET Ensemble (94.4%), RESET NN (93.6%) and AdaPTg-GMM (91.5%) have the highest medians, with RESET Ensemble also arguably having the tightest boxplot.
+Figure E6 shows the number of discoveries for each method on each of the 40 combinations of dataset and FDR threshold.
+There are a couple of implementation comments we highlight here (see Sections C.7-C.11 for details).
+First, we used the same default options of RESET Ensemble for all datasets, except for the two fMRI datasets which exhibit a spike of p-values close to 1.
+To address this, we followed Chao and Fithian (2021)’s suggestion of defining asymmetric target and decoy regions of $[0,0.3)$ and $(0.3,0.9]$ respectively.
+Second, we applied AdaPTg-GMM and ZAP using the p-value information rather than the test statistic information.
 To do so, AdaPTg-GMM allows the user to directly input p-values instead of z-statistics.
-For ZAP, we converted the p-value information into z-values by defining $z_{i}=\pm\Phi^{-1}(p_{i}/2)$ of each p-value, $p_{i}$, where the signs are chosen i.i.d.
+For ZAP, we converted the p-value information into z-values by defining $z_{i}=\pm\Phi^{-1}(p_{i}/2)$ for each p-value, $p_{i}$, where the signs are chosen i.i.d.
 uniformly.
@@
 As we mentioned in Section 3, our implementation of RESET focuses on the use of p-values, and we leave the extension to test statistic information as part of future work.
-To evaluate the computation times for each method, we used the two fMRI datasets, which uses the largest number of side information variables and has a large number of hypotheses.
-Figure 14(A-B) shows the computation times for each considered method, where the fastest are the collection of RESET methods, ZAP and AdaFDR.
+To evaluate the runtime of each method, we used the two fMRI datasets, which use the largest number of side information variables and has a large number of hypotheses.
+Figure 7(A-B) shows the runtimes for each considered method, where the fastest are the collection of RESET methods, ZAP and AdaFDR.
 RESET NN, RESET GAM and RESET RF are slightly faster than our ensemble approach which considers all of them.
-Figure 14: Comparison of runtime and power.
-In (A) and (B), we compare the computation time of each method on the fMRI datasets.
-We averaged the times of each RESET method over 10 applications.
-In (C) and (D), we compare the number of discoveries using the two eQTL datasets.
-The number of discoveries reported by each RESET method is averaged over 10 applications at FDR threshold and dataset.
-In all panels, we evaluated each method at FDR thresholds of: 1%, 5%, 10%, 20%.
-Notably, we have excluded the analysis of the two eQTL studies from Figure 12.
-This is because this dataset exhibits a high degree of p-value dependency, even among the true null hypotheses and thus violates Assumption 3 which guarantees our type-1 error control.
-Indeed, this phenomenon is due to linkage disequilibrium in which SNPs behave dependently.
-As a consequence, the dataset exhibits virtually identical copies of the same p-value, with similar side information, even if the p-value is large, suggesting a true null hypothesis.
-Figure 14(C-D) shows that RESET Ensemble reports substantially more discoveries on these eQTL datasets than all the non-RESET variants.
-This is because many estimating and training decoys are virtually the same and the random forest is sensitive to this leakage of information.
-Evidently, RESET NN is more robust to this problem: the number of discoveries is in the same ‘ballpark’ as other methods.
-Note that RESET GAM shares identical results with RESET RF since it fails to perform a spline basis expansion on one of the side information variables (the chromatin state of the SNP), and so RESET GAM switches to RESET RF in this case.
-6.3 FDP control with p-values
-Figure 15: Comparison of RESET and other methods with FDP control using a collection of publicly available data.
-Each panel plots the number of discoveries obtained by RESET Ensemble, FDP-SD and GR-SD at a range of FDR thresholds (1%, 5%, 10%, 20%) with varying confidence $1-\gamma$ (0.5, 0.8, 0.9).
-In this final section, we demonstrate RESET’s application of FDP control in the p-value setting.
-Here we define the following asymmetric target and decoy regions, $[0,0.3)$ and $(0.3,0.9]$ respectively, for all datasets.
-The rationale stems from our later analysis in the Discussion, as well as our comments at the end of Section 5.2.3, which point out that when the ratio of estimating decoys to true null targets is small, the subsequent FDP-SD step in RESET might incur a loss of power.
-Hence, to rectify this, we essentially double the number of estimating decoys under consideration by defining a decoy region that is twice as large as the corresponding target region.
-As far as we can tell, there are no p-value based testing procedures that use side information while controlling the FDP in the sense of Section 2.
-Hence, we compare RESET to the following generic methods for FDP control.
-The first of these methods is the previously considered FDP-SD that we analyzed in the competition setting, but adapted for the p-value setup.
-Each of the labels and scores $(L,W)$ required for FDP-SD are constructed in the same way as in RESET’s Step 1 (Section 4), where we used the same asymmetric target and decoy regions described above.
-The second method we compare to is the p-value based analogue called GR-SD [20] (see Section 3).
-Figure 15 displays the number of discoveries obtained by each method on the collection of real datasets we considered in the previous section, with varying FDR thresholds (1%, 5%, 10%, 20%) and confidence levels (50%, 80%, 90%).
-RESET Ensemble is much more powerful than the alternative non-side information approaches with a few exceptions.
+Figure 7: Runtime comparison. In (A) and (B), we compare the runtime of each method on the fMRI datasets. We averaged the times of each RESET method over 10 applications.
+6.5 FDX control with p-values
+Figure 8: Comparison of RESET and other methods with FDX control using a collection of publicly available data.
+Each panel shows the number of discoveries obtained by RESET Ensemble, FDP-SD and GR-SD at a range of FDR thresholds (1%, 5%, 10%, 20%) with varying confidence $1-\gamma$ (0.5, 0.8, 0.9).
+RESET Ensemble and FDP-SD were averaged over 10 applications.
+In this final section, we demonstrate RESET’s FDX control in the p-value setting.
+To the best of our knowledge, there are no existing p-value-based testing procedures that offer finite sample FDX control while utilizing side information.
+Therefore, we compare RESET Ensemble against FDP-SD (Algorithm A2) and GR-SD (Algorithm A3).
+The labels and scores $(L,W)$ used in RESET Ensemble and FDP-SD are defined using the asymmetric target and decoy regions—$[0,0.3)$ and $(0.3,0.9]$, respectively—for all datasets.
+The rationale stems from our analysis in Section F.4, which points out that when the ratio of estimating decoys to true null targets is small, RESET might incur a loss of power when controlling the FDX.
+To rectify this, we define a decoy region that is twice as large as the corresponding target region.
+Figure 8 displays the number of discoveries obtained by each method on the collection of real datasets we considered in the previous section.
+The numbers are averaged over 10 runs while varying the FDR thresholds (1%, 5%, 10%, 20%) and the confidence levels (50%, 80%, 90%).
+With a few exceptions RESET Ensemble is generally much more powerful than the alternative approaches that make no use of the side information.
 7 Discussion
-We introduced RESET, a flexible wrapper for any semi-supervised learning algorithm.
-Alongside, we provide an implementation of RESET, called RESET Ensemble, that is power-wise competitive with recently developed methods for competition as well as p-value based multiple testing problems with side information, fast compared to most tools, and flexible as it allows finite-sample of either FDR or FDP control.
+RESET provides a general framework for FDR and FDX control through a flexible wrapper for any semi-supervised learning algorithm.
+Alongside, we introduced an implementation of RESET, called RESET Ensemble, that is power-wise competitive with recently developed methods for competition as well as p-value based multiple testing problems with side information, is fast compared to most tools, and is flexible allowing finite sample FDR and FDX control.
 In terms of power, we highlight the fact that competing tools, such as the variants of AdaPT or Adaptive Knockoffs, require an initial selection of one of those variants.
 Making a selection after observing the discovery lists of each tool will generally compromise its FDR control.
-Thus, we compared RESET Ensemble to the default competing method if one was available, which was the case of Adaptive Knockoff’s GAM filter, or separately to each variant, which was the case of AdaPT.
+Thus, we compared RESET Ensemble to the default competing method if one was available, which is the case with Adaptive Knockoff’s GAM filter, or separately to each variant, which is the case for AdaPT.
 Based on our comparisons, we conclude that RESET Ensemble is able to achieve comparable or slightly greater power than many of these tools while freeing the user from having to choose the ‘right’ method to begin with.
-As far as computational runtime, we found that our collection of RESET methods provide a faster alternative to all competing methods with finite FDR control in our considered experiments.
-The only case in which RESET was surpassed in terms of speed was ZAP-asymp, which demonstrated marginally faster runtimes than RESET Ensemble in one of the fMRI datasets (Figure 14B).
+We found that each of our RESET methods was consistently faster than any of the competing methods for finite FDR control with side-information we considered here.
+More generally, the only case in which RESET was surpassed in terms of speed was ZAP-asymp, which demonstrated marginally faster runtimes than RESET Ensemble in one of the fMRI datasets (Figure 7B).
 However, ZAP-asymp’s fast runtimes come at the cost of asymptotic FDR control.
 The most striking difference in runtimes was in the peptide detection context using mass spectrometry data, in which all variants of Adaptive Knockoffs were generally too slow to be practical.
-Finally, we point out that RESET’s compatibility with FDR or FDP control is unique.
+RESET’s impressive speed compared to Adaptive Knockoffs and the AdaPT variants is due to the fact that these methods work by iterating between training a classification algorithm, and deciding which hypothesis to remove from the candidate set $\mathcal{\hat{S}}_{t}$ as explained in Section 2.1.
+In contrast, RESET Ensemble can be heavily parallelized, and only performs two iterations.
+Finally, we point out that RESET’s compatibility with FDR as well as FDX control is unique.
 In the competition setting, we note that it would be challenging to combine Adaptive Knockoffs with FDP-SD because FDP-SD is a stepdown procedure while Adaptive Knockoffs is essentially a step-up procedure.
-Katsevich and Ramdas [27] provide a general method for adaptive ordering using their method for simultaneous FDP bounds.
-However, (a) they do not provide an exact implementation of this, only a theoretical discussion, and (b) their approach is conservative since it controls the FDP at every cutoff (which allows the analyst to explore the nested discovery lists without breaking FDP control).
+Katsevich and Ramdas (2020) provide a general method for adaptive ordering using their method for simultaneous FDP bounds.
+However, (a) they do not provide an exact implementation of this, only a theoretical discussion, and (b) their approach is conservative since it controls the FDX at every cutoff (which allows the analyst to explore the nested discovery lists without breaking FDX control).
 To the best of our knowledge, in the competition framework with side information, RESET is the only method of reordering that may be flexibly used with FDP-SD without the additional cost of (b).
-Similarly, we know of no method that allows finite sample FDP control in the p-value setting with side information.
-On the topic of RESET’s FDP control, we point out the following limitation.
-When the number of discoveries is small, e.g., when the confidence is too high, RESET may yield less power, even if the side information is reasonably informative.
-To illustrate this phenomenon, we computed the bounds, $\delta_{i}$ on the number of decoy wins in the top $i$ scores from Algorithm S2, which are used to determine the reported list of discoveries in FDP-SD (with $c=1/2$) and RESET (using the default $c=2/3$).
-We considered a confidence level of $1-\gamma=90\%$ and an FDP threshold of $\alpha=1\%$.
-Since we removed approximately half the decoys for training purposes in RESET, the bounds need to be doubled to fairly compare between the two approaches.
-Refer to caption Refer to caption
-Figure 16: RESET’s and FDP-SD’s bounds on the number of decoy wins.
-We recorded, on the left y-axis, the bounds used to determine the list of FDP-SD discoveries and twice the bounds used to determine RESET’s discoveries at a range of indices at $\alpha=1\%$ and $1-\gamma=90\%$.
-On the right y-axis, we plot the ratio of these bounds (in black).
-The left panel looks at indices between 501 to 2K, while the right panel looks at indices from 10K to 20K.
-An index corresponds to the number of top scoring hypotheses (regardless of their labels).
-Figure 16 shows the bounds for FDP-SD, and double the bounds for RESET, along with the ratio of the two at a range of different indices.
-We can see that for smaller indices, this ratio is much smaller, initially at 0% at an index of 501 and 50% at an index of 1K, while at 10K this ratio is much higher at 94%.
-If the order of the hypotheses is the same, then RESET’s smaller bounds imply it will report fewer discoveries because RESET employs FDP-SD to control the FDP by searching for the largest index $i$ s.t.
-$D_{i_{0}}\leq\delta_{i},\dots,D_{i}\leq\delta_{i}$, and reports the top $i$ scoring positively labeled hypotheses.
-In other words, if RESET is to report roughly the same number of discoveries as FDP-SD, then RESET needs to rearrange the targets and estimating decoys successfully enough so as to make up for a 50% difference in the bounds when considering the top 1K scores, while only 6% when considering the top 10K.
-If the side information is not rich enough, then it is possible RESET will report few discoveries than FDP-SD in these cases.
-We intend to investigate this as part of future work.
-Figure 17: Variability in the number of discoveries.
-in the Airway data and in Simulation 5 (A) Histogram of the number of RESET Ensemble discoveries applied 100 times to the Airway data with a 10% FDR threshold and varying its internal seed.
-(B) Boxplots of RESET’s and AdaPT’s power in 100 applications of the ‘circle in the middle’ setup of Simulation 5 using a 10% FDR threshold.
-There are two outliers in RESET’s case (which occurred when the signal was considerably low).
+Similarly, we know of no method that allows finite sample FDX control in the p-value setting with side information.
 Notably, RESET’s discovery list is variable since it randomly splits the decoys into two sets: one set for training and the other for estimating the false discoveries.
-This variability is demonstrated in the histogram of the number of discoveries in Figure 17A using 100 applications of RESET Ensemble to the Airway dataset at an FDR threshold of 10%.
-The resulting number of discoveries has a mean of 5998 with a standard deviation of about 85.
-While this is problematic to an extent, keep in mind that the data distribution is random, and the added variance from RESET may be marginal at times.
-Indeed, in Figure 17(B), we find that in the case of Simulation 5 from Section 6.1.1, RESET is overall arguably less variable than AdaPT, even though AdaPT does not employ any internal randomization.
-In practice, users may get around RESET’s randomization by setting the random number generator to be a function of the input data and parameters.
-We plan on adding this feature to a future update.
-Lastly, while RESET is a flexible method for the competition and p-value based testing, we wish to further extend it to the test-statistic setting.
+While this is problematic to an extent, keep in mind that the data distribution is random, and the added variance from RESET may be marginal at times (cf.
+Section F.5).
+Lastly, we point out a few avenues for further research.
+While RESET is a flexible method for the competition and p-value based testing, we wish to further extend it to the test-statistic setting.
 There is more than one way of generalizing RESET to accommodate a test statistic $z_{i}$ for each hypothesis $H_{i}$.
-Moreover, there are several ‘types’ of interactions between the side information $\mathbf{x}_{i}$ and the test statistic $z_{i}$ which are not preserved when switching to p-values, as outlined by Leung and Sun [35].
+Moreover, there are several ‘types’ of interactions between the side information $\mathbf{x}_{i}$ and the test statistic $z_{i}$ which are not preserved when switching to p-values, as outlined by Leung and Sun (2022).
 Thus it is not clear that RESET Ensemble will be equally applicable in this setting without substantial modifications.
+Second, RESET relies on a handful of prespecified quantities: the probability $s$ of assigning decoys to the estimation set, and the definitions of the target and decoy regions as $[0,a)$ and $(b_{1},b_{2}]$, respectively.
+Moreover, in Step (1) we employ a mirroring function that converts larger $p$-values in the decoy region into smaller values in the target region.
+Given that the true null p-values are possibly non-uniform, we could consider other masking functions as suggested by Chao and Fithian (2021).
+The selection of these parameters and the masking function would ideally be data-driven which we leave as part of future work.
+Finally, we may consider alternatives to our semi-supervised learning engine, RESET Ensemble.
+One direction is to pair RESET’s data-splitting scheme with AdaPT’s iterative fitting procedure, which could further increase power (at the expense of additional computation time).
+Another is to broaden RESET Ensemble’s library of classifiers to include the two-group models that are used by many of the competing methods.
 Supplement
-S.1 Algorithms
-Algorithm S1 Selective SeqStep / SeqStep+ (adopted from Selective Sequential Step+ of [1])
+Appendix A Algorithms
+Algorithm A1 Selective SeqStep / SeqStep+ (adopted from Selective Sequential Step+ of Barber and Candès (2015))
 1:
 $\bullet$ $(L_{i},W_{i})_{i=1}^{m}$ the list of paired winning labels and scores;
-$\bullet$ $c\in(0,1)$ the probability of a null target/feature win;
-$\bullet$ $\alpha\in(0,1)$ the FDR threshold;
+$\bullet$ $c\in(0,1)$ - the probability of a null target/feature win;
+$\bullet$ $\alpha\in(0,1)$ - the FDR threshold;
 2:A discovery list $R_{\alpha}$
-3:sort the paired $(L_{i},W_{i})$ in decreasing order of $W_{i}$triangleright ties are randomly broken
-4:$D_{k}\leftarrow\#\{i\leq k:L_{i}=-1\}$triangleright number of decoy/knockoffs wins in top $k$ scores
-5:$T_{k}\leftarrow\#\{i\leq k:L_{i}=1\}$triangleright number of target/original feature wins in top $k$ scores
+3:sort the paired $(L_{i},W_{i})$ in decreasing order of $W_{i}$ $\triangleright$ ties are randomly broken
+[…]
+[…]
 6:if SeqStep then
-7:     $k_{0}\leftarrow\max\{k:\frac{D_{k}}{T_{k}\vee 1}\cdot\frac{c}{1-c}\leq\alpha\}$
+7:  $\tau\leftarrow\max\{t:\frac{A_{t}}{R_{t}\vee 1}\cdot\frac{c}{1-c}\leq\alpha\}$
 8:else if SeqStep+ then
-9:     $k_{0}\leftarrow\max\{k:\frac{D_{k}+1}{T_{k}\vee 1}\cdot\frac{c}{1-c}\leq\alpha\}$
+9:  $\tau\leftarrow\max\{t:\frac{A_{t}+1}{R_{t}\vee 1}\cdot\frac{c}{1-c}\leq\alpha\}$
 10:end if
-11:return $R_{\alpha}\leftarrow\{i:\text{the (pre-sorted) $W_{i}$ is in the top $k_{0}$ ranks and $L_{i}=1$}\}$
-Algorithm S2 FDP-SD (adopted from [40])
+[…]
+Algorithm A2 FDP-SD (adopted from Luo et al. (2023))
 1:
 $\bullet$ $(L_{i},W_{i})_{i=1}^{m}$ the list of paired winning labels and scores;
-$\bullet$ $c\in(0,1)$ the probability of a null target/feature win;
-$\bullet$ $\alpha\in(0,1)$ the FDP threshold;
-$\bullet$ $\gamma\in(0,1)$ for a $1-\gamma$ confidence level;
+$\bullet$ $c\in(0,1)$ - the probability of a null target/feature win;
+$\bullet$ $\alpha\in(0,1)$ - the FDP threshold;
+$\bullet$ $\gamma\in(0,1)$ - for a $1-\gamma$ confidence level;
 2:A discovery list $R_{\alpha,\gamma}$
-3:sort the paired $(L_{i},W_{i})$ in decreasing order of $W_{i}$triangleright ties are randomly broken
-4:$\lambda\leftarrow 1-c$triangleright the probability of a null decoy win
-5:$R\leftarrow(1-\lambda)/(c+1-\lambda)$
-6:$i_{0}\leftarrow\max\{1,\lceil(\lceil\log_{1-R}(\gamma)\rceil)/\alpha\rceil\}$
+3:sort the paired $(L_{i},W_{i})$ in decreasing order of $W_{i}$ $\triangleright$ ties are randomly broken
+[…]
+[…]
+[…]
 7:for $i=i_{0},\dots,m$ do
-8:     $D_{i}\leftarrow\#\{j\leq i:L_{j}=-1\}$triangleright number of decoy wins in top $i$ scores
-9:     $\delta_{i}\leftarrow\max\{d\in\{0,1,\dots,i\}:F_{B(\lfloor(i-d)\alpha+1+d,R\rfloor)}(d)\leq\gamma\}$triangleright $F_{B(n,p)}$ denotes the CDF of a Binomial$(n,p)$ RV
+8:  $A_{i}\leftarrow\#\{j\leq i:L_{j}=-1\}$ $\triangleright$ number of decoy wins in top $i$ scores
+9:  $\delta_{i}\leftarrow\max\{d\in\{0,1,\dots,i\}:F_{B(\lfloor(i-d)\alpha+1+d,R\rfloor)}(d)\leq\gamma\}$ $\triangleright$ $F_{B(n,p)}$ denotes the CDF of a Binomial(n,p)(n,p) RV
 10:end for
-11:$i\leftarrow i_{0}$ and $\delta_{i_{0}-1}\leftarrow-1$ and $\bar{\delta}_{i_{0}-1}\leftarrow 0$
+[…]
 12:while $i\leq m$ do
@@
 27:  Randomly set $\bar{\delta}_{i}\leftarrow\delta_{i}$ or $\bar{\delta}_{i}\leftarrow\delta_{i}+1$ with probabilities $w^{\prime}$ and $1-w^{\prime}$ respectively
-28:     if $D_{i}\leq\bar{\delta}_{i}$ then
+28:  if $A_{i}\leq\bar{\delta}_{i}$ then
 29:   $i\leftarrow i+1$
@@
 33:end while
-34:if $D_{i_{0}}\leq\bar{\delta}_{i_{0}}$ then
+34:if $A_{i_{0}}\leq\bar{\delta}_{i_{0}}$ then
 35:  $k_{FDP}\leftarrow i-1$
@@
 38:end if
-39:$R_{\alpha,\gamma}\leftarrow\{i:\text{the (pre-sorted) $W_{i}$ is in the top $k_{FDP}$ ranks and $L_{i}=1$}\}$
-Algorithm S3 GR-SD (adopted from [20])
+[…]
+Algorithm A3 GR-SD (adopted from Guo and Romano (2007))
 1:
 $\bullet$ $(p_{i})_{i=1}^{m}$ the list of p-values;
-$\bullet$ $\alpha\in(0,1)$ the FDP threshold;
-$\bullet$ $\gamma\in(0,1)$ for a $1-\gamma$ confidence level;
+$\bullet$ $\alpha\in(0,1)$ - the FDP threshold;
+$\bullet$ $\gamma\in(0,1)$ - for a $1-\gamma$ confidence level;
 2:A discovery list $R_{\alpha,\gamma}$
@@
 5:  $k_{i}\leftarrow\lfloor\alpha i\rfloor+1$
-6:     $\delta_{i}\leftarrow F^{-1}[k_{i},m-i+1](\gamma)$triangleright $F^{-1}[\alpha,\beta](\cdot)$ denotes the CDF of a Beta$(\alpha,\beta)$ RV
+[…]
 7:end for
-8:$k_{GR}\leftarrow\max\{i:\prod_{j=1}^{i}1_{p_{i}\leq\delta_{i}}=1\text{ or }i=0\}$
-9:$R_{\alpha,\gamma}\leftarrow\{i\in[m]:i\leq k_{GR}\}$
-Algorithm S4 RESET
+[…]
+[…]
+Algorithm A4 RESET
 1:
-$\bullet$ $\{(L_{i},W_{i},\mathbf{x}_{i})\,:\,i=1,\dots,m\}$ each hypothesis’ (label, winning score, side information);
-$\bullet$ $\alpha$ FDR threshold for the discovery list;
-$\bullet$ $s$ the probability of assigning a decoy to the training set (default: $s=1/2$);
-$\bullet$ $f$ a semi-supervised machine learning model;
-$\bullet$ $c_{0}$ an upper bound probability for a true null label $\mathbb{P}(L_{i}=1)\leq c_{0}$ ($i$ is a true null);
-$\bullet$ isFDR - a boolean which determines whether FDR or FDP control is desired;
-$\bullet$ $\gamma$ a confidence parameter if FDP control is desired;
+$\bullet$ $\{(L_{i},W_{i},\mathbf{x}_{i})\,:\,i=1,\dots,m\}$ - each hypothesis’ (label, winning score, side information);
+$\bullet$ $\alpha$ - FDR threshold for the discovery list;
+$\bullet$ $s$ - the probability of assigning a decoy to the training set (default: $s=1/2$);
+$\bullet$ $f$ - a semi-supervised machine learning model;
+$\bullet$ $c_{0}$ - an upper bound probability for a true null label $\mathbb{P}(L_{i}=1)\leq c_{0}$ (ii is a true null);
+$\bullet$ isFDR - a boolean which determines whether FDR or FDX control is desired;
+$\bullet$ $\gamma$ - a confidence parameter if FDX control is desired;
 2:A discovery list $R$
-3:$I\leftarrow$ a subset of the decoy win indices, $\{i:L_{i}=-1\}$, determined by randomly including each one with probability $s$
-4:$\tilde{L}_{i}\leftarrow-1$ for $i\in I$triangleright training decoys
-5:$\tilde{L}_{i}\leftarrow 1$ for $i\in J:=I^{c}$triangleright pseudo-targets
-6:$(\tilde{W}_{i})_{i=1}^{m}\leftarrow f\left((W_{i},U_{i},\tilde{L}_{i})_{i=1}^{m}\right)$triangleright where $f$ is any machine learning model
+[…]
+[…]
+[…]
+[…]
 7:if isFDR then
-8:     $R\leftarrow\text{SSS+}((\tilde{W}_{i},L_{i})_{i\in J},c=\frac{c_{0}}{1-s\cdot(1-c_{0})},\alpha)$triangleright Selective SeqStep+
+8:  $R\leftarrow\text{SSS+}((\widetilde{W}_{i},L_{i})_{i\in J},c=c_{e}:=\frac{c_{0}}{1-s\cdot(1-c_{0})},\alpha)$ $\triangleright$ Selective SeqStep+
 9:else
-10:     $R\leftarrow\text{FDP-SD}((\tilde{W}_{i},L_{i})_{i\in J},c=\frac{c_{0}}{1-s\cdot(1-c_{0})},\alpha,\gamma)$triangleright FDP-stepdown
+10:  $R\leftarrow\text{FDP-SD}((\widetilde{W}_{i},L_{i})_{i\in J},c=c_{e}=\frac{c_{0}}{1-s\cdot(1-c_{0})},\alpha,\gamma)$ $\triangleright$ FDP-stepdown
 11:end if
 12:return $R$
-S.2 Proof of Lemma 1 in the main text
-We first remind the reader of the Lemma’s statement.
-See 3
-See 1
-Proof of Lemma 4.1.
-Let $W$ be all the scores, $\mathbf{x}$ be all the side information, and $L_{-i}$ be all the other labels that are not $L_{i}$. Then for a true null p-value, $p_{i}$:
-$\mathbb{P}(L_{i}=1\mid W,\mathbf{x},L_{-i})$ $=\frac{\mathbb{P}(p_{i}\in[0,a)\mid W,\mathbf{x},L_{-i})}{\mathbb{P}(p_{i}\in[0,a)\mid W,\mathbf{x},L_{-i})+\mathbb{P}(p_{i}\in(b_{1},b_{2}]\mid W,\mathbf{x},L_{-i})}$
-$\leq\frac{\mathbb{P}(p_{i}\in[0,a)\mid W,\mathbf{x},L_{-i})}{\mathbb{P}(p_{i}\in[0,a)\mid W,\mathbf{x},L_{-i})+\mathbb{P}(p_{i}\in[0,a)\mid W,\mathbf{x},L_{-i})\cdot\frac{b_{2}-b_{1}}{a}}$
-$=\frac{a}{a+b_{2}-b_{1}},$
-where the first line follows since the denominator sums to 1, and the second line clearly follows from the non-decreasing property as explained next.
-First, note that the non-decreasing density property implies that:
-$\mathbb{P}(p_{i}\in B)\leq\frac{a}{b_{2}-b_{1}}\cdot\mathbb{P}(p_{i}\in q(B)),$
-where $B\subseteq[0,a)$ is measurable, $p_{i}$ is a true null p-value and $q$ denotes the inverse transformation of the mirrored p-values that map p-values in $(b_{1},b_{2}]$ onto $[0,a)$, that is $q(p):=b_{2}-\frac{b_{2}-b_{1}}{a}\cdot p$.
+Appendix B Proofs
+B.1 RESET controls the FDR or FDX
+By Lemma 1 in the main text, conditional on $\mathcal{G}$ the true null labels among the pseudo-targets are independent $\pm 1$ random variables, though not necessarily identically distributed.
+From this and either Lemma 1 of Barber and Candès (2015) or Lemma 2 of Lei and Fithian (2018), it can be shown that RESET controls the FDR.
+On the other hand, FDX control does not immediately follow from existing lemmas; one must separately check that the FDP-SD proof still holds when the true null labels are non-identical.
+We therefore introduce a new two-step proof that simultaneously addresses RESET’s FDR and FDX control in the non-identical setting.
+Our approach may be of independent interest, since the extension to the non-identical setting holds for a generic multiple testing procedure.
+Specifically, we show that any procedure which controls FDR or FDX with true null probabilities $\mathbb{P}(L_{i}=1)=c_{i}$ also controls them whenever $\mathbb{P}(L_{i}=1)\leq c_{i}$.
+Second, we verify RESET’s control in the special case of Lemma 1 where the true null labels are identically distributed conditioned on $\mathcal{G}$, i.e., $\mathbb{P}(L_{i}=1\mid\mathcal{G})=c_{e}$, and then invoke the first step to cover the general case, $\mathbb{P}(L_{i}=1\mid\mathcal{G})\leq c_{e}$.
+Definition 1.
+Let $\left(n_{i}\right)_{i=1}^{m_{0}}$ and $\left(a_{i}\right)_{i=1}^{m_{1}}$ be subsequences that partition $[m]$.
+Let $\left(c_{i}\right)_{i=1}^{m}$ be a sequence of probabilities in $[0,1]$.
+An input template $\mathcal{I}$ is any configuration of the data such that:
+  1. 1.
+    $\left(H_{n_{i}}\right)_{i=1}^{m_{0}}$ are true null hypotheses and $\left(H_{a_{i}}\right)_{i=1}^{m_{1}}$ are false null hypotheses;
+  2. 2.
+    all the scores $\left(W_{i}\right)_{i=1}^{m}$ as well as the labels $\left(L_{a_{i}}\right)_{i=1}^{m_{1}}$ of the false null hypotheses are fixed;
+  3. 3.
+    conditioned on the false null labels $\left(L_{a_{i}}\right)_{i=1}^{m_{1}}$ and on all the scores $\left(W_{i}\right)_{i=1}^{m}$ the labels of the true null hypotheses are independent with $P(L_{n_{i}}=1)=c_{n_{i}}$.
+Lemma B3.
+Suppose $R$ is a multiple hypothesis selection procedure (filter) that:
+  1. 1.
+    Accepts as input label-score pairs $\left(L_{i},W_{i}\right)_{i=1}^{m}$;
+  2. 2.
+    Reports only a subset of the target wins $\left(L_{i}=1\right)$;
+  3. 3.
+    For fixed $\left(c_{i}\right)_{i=1}^{m}$, $R$ controls the FDR/FDX for any input template $\mathcal{I}$.
+Then $R$ controls the FDR/FDX for a fixed $\left(c^{*}_{i}\right)_{i=1}^{m}$ in the above sense (3), where $c^{*}_{i}=c_{i}$ for $i\neq j$ and $c^{*}_{j}=0$.
+Proof.
+Let $\mathcal{I}^{*}$ be any input template for the sequence $\left(c^{*}_{i}\right)_{i=1}^{m}$.
+Suppose $H_{j}$ is a false null hypothesis.
+Then FDR/FDX control trivially follows from the original assumptions since $\mathcal{I}^{*}$ only depends on $\left(c^{*}_{i}\right)_{i=1}^{m}$ only through $(c^{*}_{n_{i}})_{i=1}^{m_{0}}=(c_{n_{i}})_{i=1}^{m_{0}}$.
+Otherwise, there exists $n_{i}=j$ for some $i\in[m_{0}]$.
+Define $\mathcal{I}^{\prime}$ to be the input template that is identical to $\mathcal{I}^{*}$ except that $H_{j}$ is a false null hypothesis and $L_{j}\coloneqq-1$.
+Again, $R$ controls the FDR/FDX on $\mathcal{I}^{\prime}$ by the original assumptions since $\mathcal{I}^{\prime}$ depends on $\left(c^{*}_{i}\right)_{i=1}^{m}$ only through $(c_{n_{i}}^{*})_{i=1,n_{i}\neq j}^{m_{0}}=(c_{n_{i}})_{i=1,n_{i}\neq j}^{m_{0}}$.
+Notably, since for $\mathcal{I}^{*}$ we have $P\left(L_{j}=1\right)=0$ and since we set $L_{j}\coloneqq-1$ for $\mathcal{I}^{\prime}$, from the perspective of $R$ the two input templates are identical.
+Moreover, because $R$ cannot report decoy wins, the FDP in its list of discoveries is the same for $\mathcal{I}^{*}$ and $\mathcal{I}^{\prime}$, hence $R$ also controls the FDR/FDX on $\mathcal{I}^{*}$ with $\left(c^{*}_{i}\right)_{i=1}^{m}$.
+∎
+Lemma B4.
+Under the assumptions of Lemma B3, $R$ controls the FDR/FDX for a fixed $(c_{i}^{*})_{i=1}^{m}$ where $c_{i}^{*}=c_{i}$ for $i\neq j$ and $c_{j}^{*}\in[0,c_{j}]$.
+Proof.
+If $c_{j}=0$, then the Lemma is trivial.
+Otherwise, consider any input template $\mathcal{I}^{*}$ with $c^{*}_{j}\in[0,c_{j}]$.
+Define $\mathcal{I}_{1}$ to be the input template that is identical to $\mathcal{I}^{*}$ except that $c^{*}_{j}$ is set to $c_{j}$.
+Similarly, define $\mathcal{I}_{2}$ to be the input template that is identical to $\mathcal{I}^{*}$ except that $c^{*}_{j}$ is set to 0.
+We can generate data from $\mathcal{I}^{*}$ by stochastically sampling from the two related templates $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$.
+First, independently of everything else we draw a value $B$ from a Bernoulli distribution with $P(B=1)=c/c_{j}$.
+Next, if $B=1$ we sample from $\mathcal{I}_{1}$, otherwise we sample from $\mathcal{I}_{2}$.
+Note that $R$ controls the FDR/FDX on samples generated according to the $\mathcal{I}_{1}$ template by the original assumptions, and it similarly controls the FDR/FDX on $\mathcal{I}_{2}$ by Lemma B3.
+It therefore follows from the law of total expectation/probability that $S$ controls the FDR/FDX on the template $\mathcal{I}^{*}$.
+∎
+Corollary B5.
+Under the assumptions of Lemma B3, $R$ controls the FDR/FDX for a fixed $(c_{i}^{*})_{i=1}^{m}$ where $c^{*}_{j}\in[0,c_{j}]$ for all $j\in[m]$.
+Proof.
+The result follows by induction on $m$ using Lemma B4. ∎
+We are now ready to prove that RESET controls the FDR or FDX where we first address the special case where the true null labels are identically distributed conditioned on $\mathcal{G}$.
+We use Corollary B5 to extend our proof to the general case.
+Proof of Theorem 1.
+For each $i\in J$ where $J=\{i\in[m]:\tilde{L}_{i}=1\}$, we assign a one-bit p-value:
+$\tilde{p}_{i}=\begin{cases}c_{e}&L_{i}=+1\\ 1&L_{i}=-1.\end{cases}$
+Recall that the scores $\widetilde{W}$ are themselves a function of $\mathcal{G}$ and clearly each one bit p-value $\tilde{p}_{i}$ is a function of the labels $L_{i}$.
+It follows from Lemma 1 of the main text that conditioned on the scores $\widetilde{W}$ and the one-bit p-values of the false nulls, the one-bit p-values of the true nulls $(\tilde{p}_{i})_{i\in J\cap N}$ are independent with $\mathbb{P}(\tilde{p}_{i}\leq c_{e})=\mathbb{P}(L_{i}=1)\leq c_{e}$, i.e., they stochastically dominate the uniform distribution.
+Suppose the labels are identically distributed so that $\mathbb{P}(\tilde{p}_{i}\leq c_{e})=\mathbb{P}(L_{i}=1)=c_{e}$ for $i\in J\cap N$.
+Then the conditions of Theorem 3 in Barber and Candès (2015) hold and applying, as RESET does, Selective SeqStep+ to the labels $\left(L_{i}\right)_{i\in J}$ and scores $\left(W_{i}\right)_{i\in J}$ with $c=c_{e}$ controls the FDR at level $\alpha$.
+Now suppose the labels are non-identical so that $\mathbb{P}(L_{i}=1)\leq c_{e}$ for $i\in J\cap N$.
+Selective SeqStep+ is a procedure that satisfies the conditions of Lemma B3 and so it follows from Corollary B5 that RESET also controls the FDR at level $\alpha$.
+∎
+Remark 1.
+It is worth noting that the original formulation of Selective SeqStep+ in Barber and Candès (2015) uses one-bit p-values instead of labels.
+However, these are equivalent since we identify the event $\{\tilde{p}_{i}=c_{e}\}$ with $\{L_{i}=1\}$ and $\{\tilde{p}_{i}=1\}$ with $\{L_{i}=-1\}$.
+Proof of Theorem 2.
+It follows from Lemma 1 of the main text that conditioned on the scores $\widetilde{W}$ and the labels of the false nulls, the labels of the true nulls $(L_{i})_{i\in J\cap N}$ are independent with with $\mathbb{P}(L_{i}=1)\leq c_{e}$.
+Suppose the labels are identically distributed so that $\mathbb{P}(L_{i}=1)=c_{e}$ for $i\in J\cap N$.
+Then this satisfies the assumption of FDP-SD with $c=\lambda=c_{e}$ (see Section 4 of Luo et al.
+(2023)) and applying, as RESET does, FDP-SD to the labels $\left(L_{i}\right)_{i\in J}$ and scores $\left(W_{i}\right)_{i\in J}$ controls the FDX at $\alpha$ with confidence $1-\gamma$.
+Now suppose the labels are non-identical so that $\mathbb{P}(L_{i}=1)\leq c_{e}$ for $i\in J\cap N$.
+FDP-SD is a procedure that satisfies the conditions of Lemma B3 and so it follows from Corollary B5 that RESET also controls the FDX at level $\alpha$ with confidence $1-\gamma$.
+∎
+B.2 Proof of Lemma 2 in the main text
+We provide a proof of Lemma 2 which is essentially the same as in Chao and Fithian (2021) but has been adapted using our notation.
+Proof of Lemma 2.
+By Assumption 4, since the pairs $(L_{i},W_{i})$ are determined by $p_{i}$ (and vice versa), the null pairs $(L_{i},W_{i})_{i\in N}$ are independent conditioned on the side information $\mathbf{x}$ and the false null pairs $(L_{i},W_{i})_{i\not\in N}$.
+Since each $L_{i}$ only possibly depends on $W_{i}$, the labels $(L_{i})_{i\in N}$ are independent if we further condition on all of $W$, the side information $\mathbf{x}$, and the labels $(L_{i})_{i\not\in N}$.
+This proves the first part of Assumption 3.
+Suppose $\mathbf{x}$ and $(p_{i})_{i\not\in N}$ are fixed in the remaining probability statements. Then for $i\in N$, since $L_{i}$ only possibly depends on $W_{i}$,
+$\mathbb{P}(L_{i}=1\mid W_{i})$ $=\frac{\mathbb{P}(p_{i}\in[0,a)\mid W_{i})}{\mathbb{P}(p_{i}\in[0,a)\mid W_{i})+\mathbb{P}(p_{i}\in(b_{1},b_{2}]\mid W_{i})}$
+$\leq\frac{\mathbb{P}(p_{i}\in[0,a)\mid W_{i})}{\mathbb{P}(p_{i}\in[0,a)\mid W_{i})+\mathbb{P}(p_{i}\in[0,a)\mid W_{i})\cdot\frac{b_{2}-b_{1}}{a}}$
+$=\frac{a}{a+b_{2}-b_{1}}$
+$=c_{0},$
+where the first line follows since the denominator sums to 1, and the second line follows from the non-decreasing property as explained next.
+First, note that the non-decreasing density property from Assumption 4 implies that for a true null p-value $p_{i}$ ($i\in N$) and a measurable $B\subseteq[0,a)$
+$\mathbb{P}(p_{i}\in B)\leq\frac{a}{b_{2}-b_{1}}\cdot\mathbb{P}(p_{i}\in q(B)),$ (5)
+where $q$ denotes the inverse transformation of the mirroring that maps p-values in $(b_{1},b_{2}]$ onto $[0,a)$, that is $q(p):=b_{2}-\frac{b_{2}-b_{1}}{a}\cdot p$.
 Let $A\subseteq\mathbb{R}$ be measurable and let $B\subseteq[0,a)$ be such that $\{W_{i}\in A\}=\{p_{i}\in B\}\cup\{p_{i}\in q(B)\}$. Then,
-[…]
+$\int_{W_{i}\in A}1_{p_{i}\in[0,a)}d\mathbb{P}=\int 1_{\{p_{i}\in[0,a)\cap B\}}d\mathbb{P}$ $=\mathbb{P}(p_{i}\in B)$
 $\leq\frac{a}{b_{2}-b_{1}}\mathbb{P}(p_{i}\in q(B))$
-[…]
-[…]
-It follows that:
-$\mathbb{P}(p_{i}\in[0,a)\mid W,\mathbf{x},L_{-i})$ $=\mathbb{P}(p_{i}\in[0,a)\mid W_{i})$
-$\leq\frac{a}{b_{2}-b_{1}}\mathbb{P}(p_{i}\in(b_{1},b_{2}]\mid W_{i})=\frac{a}{b_{2}-b_{1}}\mathbb{P}(p_{i}\in(b_{1},b_{2}]\mid W,\mathbf{x},L_{-i}).$
-S.3 RESET controls the FDR or FDP
-In this section we prove that given Assumption 3 RESET controls the FDR or the FDP.
-We use the target-decoy terminology to refer to a positive label $L_{i}=1$ as a target win or simply a target, and a negative label $L_{i}=-1$ as a decoy win or decoy.
-Lemma S1.
-Let $\tilde{L}_{i}$ for $i\in[p]$ denote the pseudo labels, where $\tilde{L}_{i}=-1$ for a training decoy and $\tilde{L}_{i}=1$ for a pseudo-target.
-Let $\mathcal{F}$ be the $sigma$-algebra generated by the winning scores $W$, the side information $\mathbf{x}$, and the labels of the false nulls.
-Let $N$ be the indices of the true nulls, and $L_{-i},\tilde{L}_{-i}$ be the corresponding labels without the $i$th one.
-Then under Assumption 3, we have for $i\in N$, $L_{i}=\pm 1$ with
-$\mathbb{P}(L_{i}=+1\mid\tilde{L}_{i}=1)\leq\frac{c_{0}}{1-s\cdot(1-c_{0})},\quad\mathbb{P}(L_{i}=-1\mid\tilde{L}_{i}=1)\geq\frac{(1-c_{0})\cdot(1-s)}{1-s\cdot(1-c_{0})},$
-independently of $\mathcal{F},\tilde{L}_{-i},L_{-i}$.
-Proof.
-By Assumption 3, we know $\{L_{i}\,:\,i\in N\}$ are i.i.d with $P(L_{i}=1)\leq c_{0}$, independently of $\mathcal{G}:=\sigma(\mathcal{F},\tilde{L}_{-i},L_{-i})$.
-Hence, the following probability statements are independent of $\mathcal{G}$.
-In addition, since each decoy is randomly assigned to the training decoy set independently of anything else with probability $s$, for each $i\in N$ we have:
-$\mathbb{P}(L_{i}=-1,\tilde{L}_{i}=1)$ $=\mathbb{P}(L_{i}=-1)\cdot\mathbb{P}(\tilde{L}_{i}=1\mid L_{i}=-1)$
-$\geq(1-c_{0})\cdot(1-s),$
-Hence,
-$\mathbb{P}(L_{i}=-1\mid\tilde{L}_{i}=1)$ $=\frac{\mathbb{P}(L_{i}=-1,\tilde{L}_{i}=1)}{P(\tilde{L}_{i}=1)}$
-$\geq\frac{(1-c_{0})\cdot(1-s)}{(1-s)\cdot\mathbb{P}(L_{i}=-1)+1\cdot\mathbb{P}(L_{i}=1)}$
-$=\frac{(1-c_{0})\cdot(1-s)}{1-s\cdot\mathbb{P}(L_{i}=-1)}$
-$\geq\frac{(1-c_{0})\cdot(1-s)}{1-s\cdot(1-c_{0})}.$
-The proof is complete noting that $\frac{c_{0}}{1-s\cdot(1-c_{0})}+\frac{(1-c_{0})\cdot(1-s)}{1-s\cdot(1-c_{0})}=1$ and $L_{i}\in\{\pm 1\}$ a.s.. ∎
+$=\frac{a}{b_{2}-b_{1}}\int 1_{\{p_{i}\in(b_{1},b_{2}]\cap q(B)\}}d\mathbb{P}$
+$=\frac{a}{b_{2}-b_{1}}\int_{W_{i}\in A}1_{p_{i}\in(b_{1},b_{2}]}d\mathbb{P},$
+Since $A$ was arbitrary, it follows that:
+$\mathbb{P}(p_{i}\in[0,a)\mid W_{i})\leq\frac{a}{b_{2}-b_{1}}\mathbb{P}(p_{i}\in(b_{1},b_{2}]\mid W_{i}).$
 Remark 2.
-Lemma S1 essentially states that the labels $L_{i}$ for $i\in N$ with $\tilde{L}_{i}=1$ are i.i.d.
-random variables with $\mathbb{P}(L_{i}=+1)\leq\frac{c_{0}}{1-s\cdot(1-c_{0})}$ independently of all other information: the scores $W$, the side information $\mathbf{x}$, the labels of the false null discoveries, and all the labels $\tilde{L}$.
-See 1
-Proof.
-We rely on Theorem 3 of Barber and Candès [1] applied to the set of pseudo-targets ($\tilde{L}_{i}=1$) ordered in decreasing order of the learned scores $\tilde{W}_{i}$.
-Specifically, we assign to each pseudo-target a one-bit p-value:
-$\tilde{p}_{i}=\begin{cases}\frac{c_{0}}{1-s\cdot(1-c_{0})}&L_{i}=+1\\ 1&L_{i}=-1\end{cases}.$
-Recall that the scores $\tilde{W}$ are themselves a function of the values $(\tilde{L},W,x)$ and clearly the one bit p-values are a function of the labels $L_{i}$.
-Then it follows from the last Lemma that given the scores $\tilde{W}$ and the one-bit p-values of the false nulls, the one-bit p-values of the true nulls hypotheses ($i\in N$ with $\tilde{L}_{i}=1$) are i.i.d.
-with $\mathbb{P}(\tilde{p}_{i}\leq\frac{c_{0}}{1-s\cdot(1-c_{0})})=\mathbb{P}(L_{i}=1)\leq\frac{c_{0}}{1-s\cdot(1-c_{0})}$, i.e., they stochastically dominate the uniform distribution and hence are valid p-values.
-Therefore, the condition of Theorem 3 holds so applying, as RESET does, Selective SeqStep+ with $c=\frac{c_{0}}{1-s\cdot(1-c_{0})}$ to these p-values controls the FDR at level $\alpha$.
-Note that the original formulation of Selective SeqStep+ is equivalent to Algorithm S1 since we identify the event $\tilde{p}_{i}\leq\frac{c_{0}}{1-s\cdot(1-c_{0})}$ with $L_{i}=1$ and $\tilde{p}_{i}=1$ with $L_{i}=-1$.
-∎
-See 2
-Proof.
-The proof of Theorem 1 shows that conditional on the learned scores $\tilde{W}$ and the false null labels the labels $L_{i}$ for $i\in N$ with $\tilde{L}_{i}=1$ are distributed as i.i.d.
-±$\pm 1$ RVs, with $\mathbb{P}(L_{i}=1)\leq\frac{c_{0}}{1-s\cdot(1-c_{0})}$.
-Assuming for a moment that $\mathbb{P}(L_{i}=1)=\frac{c_{0}}{1-s\cdot(1-c_{0})}$, then RESET satisfies the assumption of the multiple decoy version of FDP-SD with $c=\lambda=\frac{c_{0}}{1-s\cdot(1-c_{0})}$ and therefore it controls the FDP at $\alpha$ with confidence $1-\gamma$ (Section 4 of [40]).
-In the case that $\mathbb{P}(L_{i}=1)<\frac{c_{0}}{1-s\cdot(1-c_{0})}$, then the procedure is conservative with $c=\lambda=\frac{c_{0}}{1-s\cdot(1-c_{0})}$ and the proof in [40] follows without any modification.
-∎
-S.4 Further implementation details of RESET
+As part of AdaPT’s assumptions, Lei and Fithian (2018) directly assume Equation (5), which they call mirror-conservatism.
+Lemma 2 is therefore true under a broader class of true null p-values that satisfy mirror-conservatism, not just those that are non-decreasing as prescribed by AdaPTg’s assumptions (Chao and Fithian (2021)) and our Assumption 4.
+Appendix C Supplementary Methods
+C.1 Summary of methods
+  1. 1.
+    AdaFDR: AdaFDR splits the pairs $(p_{i},\mathbf{x}_{i})_{i=1}^{m}$ into two folds.
+For each fold, the method fits a mixture model to optimally rescore the hypotheses, and the model is then applied to the other fold where a list of discoveries is obtained.
+The two list of discoveries are joined together and then reported.
+AdaFDR’s mixture model is a composition of Gaussian components to model local ‘bumps’ of genuine signals (false nulls) within the data plus an additional component that captures the monotonic relationship between the side information and the signals.
+  2. 2.
+    AdaPT: AdaPT works by iteratively pruning a candidate set of hypotheses $\mathcal{\hat{S}}_{t}$, similar to Adaptive Knockoffs, which was derived from it.
+At each iteration, AdaPT fits a two-group mixture model on a strict subset of information: the p-values outside $\mathcal{\hat{S}}_{t}$, the masked p-values in $\mathcal{\hat{S}}_{t}$ given by $\left(p_{i}\wedge(1-p_{i})\right)_{i\in\mathcal{\hat{S}}_{t}}$ where $1-p_{i}$ is referred to as the mirrored p-value, the side information $\mathbf{x}$, as well as $\{i\in\mathcal{\hat{S}}_{t}:p_{i}>1/2\}$ and $\{i\in\mathcal{\hat{S}}_{t}:p_{i}<1/2\}$.
+It then removes the hypothesis from $\mathcal{\hat{S}}_{t}$ that is deemed most likely to be a true null based on the fitted model.
+To fit the mixture model, the EM algorithm is used.
+The M-step of the EM algorithm results in a regression problem which allows the user to flexibly use the regression method of their choice.
+The authors considered a generalized additive model (GAM), a generalized linear model (GLM), and a generalized linear model with $L_{1}$ regularization (GLMnet).
+If we identify each p-value $p_{i}>1/2$ with a label $L_{i}=-1$ and each p-value $p_{i}<1/2$ with a label $L_{i}=1$, then AdaPT terminates this pruning procedure at step $\tau$ when the estimated FDR, as in Equation (3) of the main text, is $\leq\alpha$.
+The rationale is that for true null hypotheses, the p-values are usually uniformly distributed, so $\{i\in\mathcal{\hat{S}}_{t}:L_{i}=-1\}$ estimates the number of true null hypotheses with positive labels, $\{i\in\mathcal{\hat{S}}_{t}:L_{i}=1\}$.
+It then reports the remaining positively labelled hypotheses in $\mathcal{\hat{S}}_{\tau}$.
+Note, AdaPT and its variants do not use labels explicitly — we introduce those here for notational convenience.
+  3. 3.
+    AdaPTg: AdaPTg generalizes AdaPT by constructing asymmetric regions to define the positive and negative labels.
+For example, the p-values $p_{i}\in\left(0.3,0.9\right)$ can be identified with a label $L_{i}=-1$ and the p-values $p_{i}<0.3$ can be identified with a label $L_{i}=1$.
+One advantage of defining these asymmetric regions, is to avoid overestimating the number of true null hypotheses in cases where their p-values are not uniformly distributed and there is a concentration of true null p-values near 1.
+Subsequently, the estimated FDR needs to be adjusted to account for the fact that we expect twice as many true null p-values with negative labels than positive labels.
+  4. 4.
+    AdaPT-GMMg: AdaPT-GMMg replaces the two group mixture model in AdaPTg for a Gaussian mixture model (GMM).
+Similar to AdaPT/AdaPTg, it also allows the user to choose from a collection of regression methods to fit the model at the M-step of the EM algorithm.
+  5. 5.
+    ZAP-asymp: Lastly, ZAP-asymp directly fits a beta mixture model to the data but is only able to offer asymptotic FDR control.
+Importantly, ZAP-asymp can operate directly on the test statistic instead of just the corresponding p-values to rerank the hypotheses.
+This is particularly advantageous for two-sided tests because the test statistic is mapped to its p-value in a non-bijective manner, losing information regarding the sign of the original test statistic along the way.
+Note that AdaPT-GMMg can also operate on test statistics, but in this paper, we focus on the p-value information.
+C.2 Further implementation details of RESET
+C.2.1 Handling hypotheses with a score of zero
+RESET throws out the hypotheses with undefined labels, i.e., those hypotheses with $W_{i}=0$.
+We try to recover some of this information in the following way.
+If the proportion of zero-scoring hypotheses is at least $0.01$, then for each hypothesis with $W_{i}\neq 0$, we identify the $k$-nearest neighbours in terms of the side information (zero-scoring hypotheses included).
+Then we define an extra side information variable as the number of zero-scoring hypotheses among these identified neighbours (we used $k=20$).
+C.2.2 Handling noisy side information
+To reduce the effect of ‘noisy’ side information that is unable to distinguish between true and false nulls, we implement an initial side information selection procedure.
+Specifically, we use a generalized additive model to fit a smoothing spline using $W\cdot\tilde{L}$ as the response against each user-provided side information variable one at a time.
+We use tryCatch in case the smoothing spline fails, in which case we fit the response directly to the side information variable, that is, we use a linear fit.
+Again, we try to make use of some of the zero-scoring hypotheses setting $W\cdot\tilde{L}=0$ in these cases.
+We keep only the side information variables with sufficiently small p-values in the fitted model ($p<0.01$).
+C.2.3 Positive set for training
+Not all the pseudo-targets need to be included in the initial positive set in Step (i).
+After all, we expect many of the pseudo-targets to be true nulls.
+Hence, to improve the performance of RESET, we consider the following strategy.
+For each feature in $(W,\mathbf{x})$, reorder the hypotheses according to that feature, and apply SSS to the pseudo labels $\tilde{L}$ at the FDR level $\alpha$.
+Then select the feature that maximizes the number of pseudo discoveries.
+Then using the chosen feature, we define the positive set as those pseudo-targets discovered at $\alpha_{0}$, again using SSS with the pseudo labels $\tilde{L}$ (we used $\alpha_{0}=50\%$ throughout).
+Usually the score is selected, but in some cases the side information is highly informative and is selected instead.
+Users interested in a small FDR level, like $\alpha=1\%$, could probably benefit from reducing $\alpha_{0}$ to further improve the performance, and in particular the speed, of RESET (though as we mentioned, we kept $\alpha_{0}=50\%$ across the board).
+To ensure that the positive set is always large enough, we increase $\alpha_{0}$ by increments of 0.01 until the positive set reaches min_positive = 50 or until all the pseudo-targets are in the positive set.
+This is applied to both the initial positive set and the second round of training.
+C.2.4 Handling dependent data
+Real data often exhibits dependency to some degree, failing to satisfy Assumption 3, which guarantees RESET’s finite sample FDR/FDX control.
+To overcome this challenge, we implement the following two modifications when applied to p-value based data in RESET Ensemble and its variants.
+The first modification is in Step (ii).
+For each fold $k\in[K]$, we replace each score $W_{i}$ from the negative set in the training folds, $K]\setminus\{k\}$, with scores drawn i.i.d from $\lvert\Phi^{-1}((b_{2}-p_{i})\cdot\frac{a}{b_{2}-b_{1}})\rvert$ where $p_{i}\sim U[b_{1},b_{2}]$.
+In other words, we replace the scores of the negative set for those obtained from uniform p-values before each training phase.
+Note that we only modify the p-values of the negative set in the training folds: all the p-values in the test folds remain unchanged.
+This reduces any leakage of information from the training folds into the test folds and in the final stage when the estimating decoys are used to report a list of discoveries.
+The following second modification is motivated by the observation that the Random Forest is particularly sensitive to local dependencies.
+Specifically, if there are several hypotheses that exhibit essentially the same p-value and side information, then the same observation is essentially split across all folds.
+The flexibility of Random Forest enables it to overfit to these local regions of hypotheses and significantly overshoot the FDR threshold.
+In the second modification we apply $k$-nearest neighbours to the side information of the set of training decoys.
+Next, we sum over the squared differences between each training decoy’s score $W_{i}$ and its $k$-neighbours, denoted as $ss_{\text{obs}}$.
+We perform a bootstrap-like procedure in which we shuffle the scores of the training decoys $B=1000$ times and subsequently compute the sum of squares as previously described for each shuffle, producing $\left(ss_{i}\right)_{i=1}^{B}$.
+We compute a one-sided p-value by calculating the fraction of bootstrapped sum-of-squares smaller than $ss_{\text{obs}}$.
+If the p-value is $\leq 0.01$, the user is notified that the data may exhibit local dependency, and we remove the Random Forest from the ensemble of classification algorithms.
+If RESET RF is implemented and the p-value is $\leq 0.01$, then RESET RF switches to RESET NN with default parameters.
+Our detection method above compares how close the scores are, through $ss_{\text{obs}}$, versus the collection of $\left(ss_{i}\right)_{i=1}^{B}$ in which the scores are no longer dependent on the side information.
+Thus a p-value that is $\leq 0.01$ is likely to catch data exhibiting local dependency.
+To make the k-nearest neighbours more efficient, and to be certain that the procedure is only applied to true null hypotheses, we consider only the training decoys that have a small enough mirrored p-value, i.e., $(b_{2}-p_{i})\cdot\frac{a}{b_{2}-b_{1}}\leq 0.1$ and we restrict number of neighbours to $k=100\wedge\lceil 0.01\cdot m\rceil$, where $m$ is the number of considered hypotheses used in RESET.
+We demonstrate the robustness of our modifications in Section 6.3 in which we show that RESET Ensemble appears to empirically control the FDR using dependent p-values, along with our analysis of real data in Section E.2.
+In cases when the data satisfies Assumption 3, our above modifications do not violate finite sample FDR/FDX control.
+C.2.5 Other technical details
 We point out some additional minor details regarding our implementation of RESET.
-We use the randomxForest package, nnet and mgcv to implement the random forest, two-layer neural network and generalized additive model in R, respectively.
-We used the default parameters for the random forest, the default parameters for the two-layer neural network and the default parameters for the generalized additive model (except we use drop.intercept = TRUE).
+We use the randomForest package, nnet and mgcv to implement the random forest, two-layer neural network and generalized additive model in R, respectively.
+We used the default parameters for the random forest except with $1000$ trees, the default parameters for the two-layer neural network and the default parameters for the generalized additive model (except we use drop.intercept = TRUE).
 If the number of side information variables used by RESET is $\leq 3$, a smoothing spline is fitted with mgcv::s otherwise a natural cubic spline is fitted using 5 degrees of freedom with splines::ns on each side information variable separately (for computational efficiency).
-For the use of GAM in RESET, we used tryCatch to switch to a random forest in the rare case that the GAM was not able to complete.
+If GAM fails to fit, we used tryCatch to remove the smooth terms and therefore becoming just a linear model.
 Some real datasets contain p-values that are ‘zero’.
@@
 We replace instances of Inf for the maximum real-valued score in the dataset.
-Lastly we used the default settings of RESET for all numerical and real data experiments unless otherwise stated in the main text.
-S.5 Further implementation details of Adaptive Knockoffs
+Unless otherwise stated, we used the same default settings of RESET for all simulated and real data experiments, which includes all of the heuristics described in this section (Section C.2).
+C.3 Competition-based multiple testing: the knockoff filter
+The competition framework was popularized in the statistics community in the context of variable selection, where the goal is to identify relevant variables from $\{X_{i}:i=1,\dots,p\}$ that are conditionally associated with a response $y$.
+The procedure that achieves this while controlling the FDR is called the knockoff filter, which begins by pairing each variable $X_{i}$ with an artificial variable $\tilde{X}_{i}$, called a knockoff.
+In this paper, we consider the model-X version of the knockoff filter (Candès et al.
+(2018)) which makes the following assumptions.
+First, the joint distribution of the variables, $X$, is assumed to be known.
+Second, each observation $(X_{i,1},\dots,X_{i,p},y_{i})$ is sampled i.i.d.
+from the joint distribution $F_{XY}$.
+Because $X$ is random and its distribution known, the knockoffs $\tilde{X}$ are also constructed randomly subject to the following conditions:
+  1. 1.
+    The joint distribution of $\hat{X}:=[X,\tilde{X}$] is the same if we swap a subset of the variables and their corresponding knockoffs, i.e., $\hat{X}\circ\Pi\overset{d}{=}\hat{X}$, where $\Pi$ is a permutation that swaps the subset of variable indices for their corresponding knockoff indices.
+  2. 2.
+[…]
+Once the knockoffs have been constructed, each variable is given a user-defined score $W_{i}([X,\tilde{X}],y)\in\mathbb{R}$ and a label $L_{i}([X,\tilde{X}],y)\in\{\pm 1\}$.
+As in the competition framework described in Section 2, the idea is that the label $L_{i}$ indicates whether the variable $X_{i}$ or the knockoff $\tilde{X}_{i}$ has greater evidence for being relevant to the model, and the score $W_{i}$ indicates by how much.
+Commonly, the score and label are combined by setting $sign(W_{i})=L_{i}$; however, given our presentation of RESET, it is instructive to separate them.
+In this paper, we consider the Lasso Coefficient Difference (LCD) score, which is the difference in the magnitude of the coefficients of the Lasso regression (Tibshirani (1996)).
+Specifically, cross-validation is used to determine the $\lambda$ penalty when regressing on the augmented design matrix $X,\tilde{X}$].
+We denote the absolute values of the Lasso($\lambda$)-fitted coefficients as $Z_{i}$ for the variables and $\tilde{Z}_{i}$ for the knockoffs.
+The scores and labels are then defined as
+[…]
+where hypotheses with zero scores are thrown out.
+C.4 Further implementation details of Adaptive Knockoffs
 There were some small differences between the implementations in the adaptiveKnockoff package in R and the code used in the original manuscript.
 Accordingly, we used the updated package version.
-We also fixed a minor bug, which pruned the candidate set $\mathcal{S}$ before checking the estimated FDR is $\leq\alpha$ — possibly missing out on a marginally larger discovery list.
+We also fixed a minor bug, which pruned the candidate set $\mathcal{\hat{S}}_{t}$ before checking the estimated FDR is $\leq\alpha$ — possibly missing out on a marginally larger discovery list.
 There are several parameters that may be set before applying the Adaptive Knockoff filters, e.g., the initial proportion of hypotheses revealed, reveal_prop.
-We used the same parameter settings in Simulation 1 as Simulation 1 in [44], and the same parameter settings in Simulation 2-4 as their Simulation 2 [44].
-In the application to peptide detection, the primary scores $W$ may be negative.
-Since Adaptive Knockoffs encodes the labels $L$ by using the sign of the scores $W$, we shifted our peptide scores so that the minimum score was zero.
+We used the same parameter settings in Simulation 1 here as in Simulation 1 in Ren and Candès (2023) (which uses one-dimensional side information), and the same parameter settings in Simulations 2-4 here as Simulation 2 in Ren and Candès (2023) (which uses two-dimensional side information or more).
+In the application to peptide detection, the primary score $W$ can be negative.
+Since Adaptive Knockoffs encodes the labels $L$ by using the sign of the scores $W$, we shifted our peptide scores so that the minimum score was zero in those cases.
 For AdaKO EM, any peptide score $\leq 10^{-3}$ was subsequently assigned a zero score.
@@
 All other parameters were default.
-S.6 Further implementation details of the AdaPT methods
-We implemented AdaPT using the adaptMT package in R.
-In Simulation 5 and 6, we used the same settings considered by the authors [33], i.e., we looked at AdaPT GAM in Simulation 5 and AdaPT GLMnet in Simulation 6 with the same parameters.
+C.5 Simulation details for the competition setting
+C.5.1 Simulation 1: Linear model with one-dimensional side information
+Our first simulation is based on Simulation 1 of Ren and Candès (2023).
+Specifically, $n=1000$ independent observations from the joint data distribution of $(X,Y)$ are generated by drawing $X$ from a hidden markov model (HMM) with $p=900$ variables and $Y|X\sim\mathcal{N}(X\beta,I)$ (linear model).
+The $\beta$’s are selected in the following way.
+Let $k$ denote the number of nonzero $\beta$’s.
+Then $k$ indices from $\{1,2,\dots,2k\}$ are sampled without replacement, each with probability proportional to $1/i^{2}$.
+The $k$ chosen indices are the nonzero $\beta$’s and are assigned a value of $\pm 3.5/\sqrt{n}$, where the signs are chosen i.i.d.
+uniformly.
+We consider $k\in\{50,150,300\}$ while the original simulation only considers $k=150$.
+We generate HMM knockoffs from Sesia et al.
+(2019).
+Each hypothesis is supported by the side information $\mathbf{x}_{i}=i$, which should help in detecting the nonzero $\beta$’s, because smaller values of $i$ are more likely to correspond to a nonzero $\beta_{i}$.
+We generated 100 independent datasets each consisting of drawing $n$ observations as described above.
+The original simulation in Ren and Candès (2023) sampled the $\beta$’s once and fixed them across the 100 datasets.
+Here we redrew the $\beta$’s for each dataset so that we may average our results over a range of models, as we noticed that all the methods exhibited significant variability between the draws.
+C.5.2 Simulation 2: Linear model with two-dimensional side information
+In a second simulation, we considered a larger linear model where $n=2000$ observations are taken from the joint distribution $(X,Y)$ as described in Section C.5.1, with $p=1800$.
+Accordingly, we also adjusted the number of false nulls to $k\in\{150,300,450\}$.
+Note that unlike the previous simulation, it is not relevant how the $k$ variable indices are selected from $\{1,2,\dots,p\}$.
+Each hypothesis is equipped with the following two-dimensional side information.
+For the $i$th hypothesis, we draw a pair of observations $\mathbf{x}_{i}$ from a bivariate normal $\mathcal{N}(\boldsymbol{\mu},I)$ with $\boldsymbol{\mu}=(0,0)$ for a true null hypothesis and $\boldsymbol{\mu}=\pm(2,2)$ for a false null hypothesis, where the signs are chosen i.i.d.
+uniformly.
+We generated 20 independent datasets of this larger simulation instead of 100, where each involved drawing those $n$ observations, $\beta$ coefficients, and side information variables.
+C.5.3 Simulation 3: Logistic model with two-dimensional side information
+Next, we consider Simulation 2 by Ren and Candès (2023).
+They generate $n=1000$ observations with $p=1600$ variables sampled from the joint data distribution of $(X,Y)$, where $X$ is distributed according to a multivariate normal distribution and $Y|X\sim\text{Bernoulli}\left(e^{\beta\cdot X}/\left(1+e^{\beta\cdot X}\right)\right)$ is a logistic model (see Ren and Candès (2023) for further details).
+Each hypothesis’ side information is a unique pair of values $\mathbf{x}_{i}=\left(\mathbf{x}_{i1},\mathbf{x}_{i2}\right)$ from the lattice $20,19]\times[-20,19]\subseteq\mathbb{Z}^{2}$.
+The nonzero $\beta$’s are chosen according to the position of $\left(\mathbf{x}_{i1},\mathbf{x}_{i2}\right)$ in the lattice.
+An image of the nonzero $\beta$’s is given in Figure C9.
+Each nonzero $\beta$ is set to $\pm 25/\sqrt{n}$, where the signs are chosen i.i.d.
+uniformly.
+We generate approximate SDP Gaussian knockoffs as described in Candès et al.
+(2018).
+We generated 100 independent datasets drawing $n$ observations as above in each but for 2 of those AdaKO EM failed with an error so we used only 98 of the datasets.
+The original simulation in Ren and Candès (2023) sampled the sign of the $\beta$’s once, fixing them across the 100 datasets.
+Instead, we redraw the sign of the $\beta$’s in each of the 98 datasets since we found that the results of all the methods significantly varied with those draws.
+Figure C9: Location of nonzero $\beta$’s according to the pair of values $(x,y)$ in the lattice. The same image is given in Ren and Candès (2023).
+C.5.4 Simulation 4: Large $p$ and $n$ and 3-d side information
+Finally, in the last simulation, $n=10K$ observations with $p=10K$ variables and $k=0.15p$ relevant variables are drawn from the joint data distribution $(X,Y)$ exactly as described in Simulation 1.
+Each variable is equipped with the following three-dimensional side information.
+The first $k$ variables are randomly assigned a value drawn from $i\in\{1,\dots,2k\}$ taken without replacement and having probability proportional to $1/i^{2}$.
+The remaining pkp-k variables are uniquely assigned to one of the remaining values from $i\in\{1,2,\dots,p\}$ uniformly at random.
+Overall, the final assigment is a permutation of the integers in $\{1,\dots,p\}$, subject to the constraint that the first $k$ variables take values exclusively from $\{1,\dots,2k\}$.
+We expect most of the nonzero $\beta$’s to have small values of $i$, but the effect of the random assignment makes the resulting side information variable less informative than in Simulation 1, as depicted in Figure C10.
+This process is repeated three times, thus associating three side information variables with each hypothesis, which taken together, should be reasonably informative.
+We generated 5 independent samples, drawing the $n$ observations, $\beta$ coefficients, and side information in each.
+Figure C10: Scatterplot of each hypothesis according to the LCD score denoted as $W$ (x-axis) and side information variable (y-axis).
+Each hypothesis is coloured according to whether it is a true (red) or false null hypothesis (blue).
+In (A), the side information variable is the index of the hypothesis in Simulation 1 (Section C.5.1).
+In (B), we use one of the three side information variables as described in Simulation 4 (Section C.5.4).
+C.6 Competition-based multiple testing: peptide detection
+Competition-based multiple testing was first established in the proteomics community where the goal is to infer which proteins, i.e., long chains of amino acids, are present in a biological sample (Elias and Gygi (2007)).
+Direct protein detection is challenging, and so a bottom-up approach is taken where the proteins are digested into small chains of amino acids, called peptides, and the objective is to identify these peptides instead.
+Then, a process called liquid chromatography tandem mass spectrometry (LC-MS/MS) is used in which a first round of mass spectrometry isolates each sample peptide and reports the abundance or intensity of each sample peptide.
+In a second round of mass spectrometry, each sample peptide that reports a high enough intensity is selected and further fragmented.
+The intensity of each fragment is then reported in a corresponding fragmentation spectrum.
+Each such spectrum can be thought of as a fingerprint of the peptide that generated the spectrum, allowing us to hypothesize which peptides are present in the sample.
+This is done by searching each spectrum against a database of candidate peptides called targets to find the spectrum’s optimal peptide-spectrum match (PSM), along with a score indicating the strength of the match.
+Unfortunately, due to the incomplete nature of the database and the noise associated with generating the spectra, the PSM can be incorrect.
+Hence, we wish to employ a testing procedure for each null hypothesis $H_{i}$ that ‘the $i$th target is not present in the sample’.
+In order to decide which null hypotheses to reject, we pair each target peptide with a decoy peptide that is generated by randomly shuffling or reversing the target peptide’s sequence.
+Accordingly, a second search is performed for each spectrum against the database of decoys to obtain a second PSM and associated score, and only the highest scoring PSM out of the two is retained.
+The search phase described above is dictated by a range of user-specified parameters.
+The most common configuration of these parameters correspond to two broad types of searches, namely narrow and open searches, which we consider in this paper.
+Briefly a ‘narrow’ search, which is the default search mode, only searches each spectrum in the spectrum file against peptides in the target-decoy database that have a theoretical mass within a small tolerance of the mass of the sample peptide that generated the experimental spectrum.
+On the other hand, an ‘open’ search makes this mass-tolerance larger so that a spectrum may match with a peptide with a completely different mass.
+This is desirable at times, since sample peptides can undergo post-translational modifications which modify the mass of the peptide so it no longer matches its theoretical mass.
+However, when searching in an ‘open’ mode, the modified spectrum may still be correctly matched.
+The same database peptide, target or decoy, may be matched multiple times with different scores, or have no match at all.
+Hence, we associate with $H_{i}$ the score $Z_{i}$, defined as the maximum scoring PSM associated with the $i$th target peptide, and by its design the higher the score, the more likely $H_{i}$ is false, i.e., the peptide is in the sample.
+Similarly, we define the null drawn $\tilde{Z}_{i}$ as the maximum scoring PSM associated to the $i$th target’s corresponding decoy peptide.
+Target or decoy peptides that have no matching spectra are assigned a score of -$\infty$.
+Finally, we obtain a winning score $W_{i}=Z_{i}\vee\tilde{Z}_{i}$ and a label $L_{i}$ indicating whether the winning score was from the $i$th target or its decoy (where ties are randomly broken).
+Peptides with a winning score of -$\infty$, indicating that neither the target nor its paired decoy have a matching spectra, are thrown out.
+Finally, a competition-based multiple testing procedure like SSS+ or FDP-SD may be applied to these scores and labels (Nesvizhskii (2010)).
+C.7 Further implementation details of the AdaPT methods
+We applied AdaPT using the adaptMT package in R.
+In Simulations 5 and 6, we used the same settings considered by Lei and Fithian (2018), i.e., we used AdaPT GAM in Simulation 5 and AdaPT GLMnet in Simulation 6 with the same parameters.
 Moreover, we considered AdaPT GLMnet in Simulation 5 with the same settings from Simulation 6, demonstrating the importance of selecting the correct working model.
+In Simulations 7 and 8, we used the same settings as Simulation 5.
 In the real datasets looked at by Zhang et al.
-[51], we used AdaPT GAM to fit a natural cubic spline using splines::ns with 5 degrees of freedom (knots chosen by default) on each side information variable.
+(2019), we used AdaPT GAM to fit a natural cubic spline using splines::ns with 5 degrees of freedom (knots chosen by default) on each side information variable.
 The use of natural cubic splines is a suggested option from the package documentation.
@@
 Zhang et al.
-mention that they employed a ‘5-degree for each dimension’, but it is not clear to us if they mean the degree of the piece-wise polynomials that make the spline, or the degrees of freedom of the spline.
+mention that they employed a “5-degree for each dimension”, but it is not clear to us if they mean the degree of the piece-wise polynomials that make the spline, or the degrees of freedom of the spline.
 In the latter case, it is not clear how the knots are chosen, or how they overcame the error associated with the ‘chromatin state of the SNP’.
@@
 Moreover, Zhang et al.
-omits the application of AdaPT to the two fMRI datasets on account of the categorical side information variable used in these datasets.
+omit the application of AdaPT to the two fMRI datasets on account of the categorical side information variable used in these datasets.
 However, Chao and Fithian appear to still apply AdaPTg without problem, and so analogously we applied AdaPT as well.
-Lastly, in the gene-drug response data, we followed the same implementation by Lei and Fithian [33], which used AdaPT GLM and a collection of candidate side information transformations using splines::ns with degrees of freedom ranging from 6 to 10.
-S.7 Further implementation details of the AdaPTg methods
-We implemented AdaPTg using the repository https://github.com/patrickrchao/adaptMT which has all the same implementations as AdaPT while offering the ability to define asymmetric regions for mirroring the p-values.
+Lastly, in the gene-drug response data, we followed the same implementation by Lei and Fithian (2018), which used AdaPT GLM and a collection of candidate side information transformations using splines::ns with degrees of freedom ranging from 6 to 10.
+C.8 Further implementation details of the AdaPTg methods
+We downloaded AdaPTg from the repository https://github.com/patrickrchao/adaptMT which has all the same implementations as AdaPT while offering the ability to define asymmetric regions for mirroring the p-values.
 Accordingly, we followed the same implementation details as the previous section used by AdaPT and set the so-called ‘masking parameters’ as default.
-S.8 Further implementation details of AdaPT-GMMg
-We implemented AdaPT-GMMg from the repository https://github.com/patrickrchao/AdaPTGMM.
-In Simulations 5, we used the GAM filter (model_type = ‘mgcv’) to fit a smoothing spline using mgcv::s on the side information for the M-Step of the EM algorithm.
+C.9 Further implementation details of AdaPT-GMMg
+We downloaded AdaPT-GMMg from the repository https://github.com/patrickrchao/AdaPTGMM.
+In Simulations 5, 7 and 8, we used the GAM filter (model_type = ‘mgcv’) to fit a smoothing spline using mgcv::s on the side information for the M-Step of the EM algorithm.
 In Simulation 6, we used a generalized linear model filter with regularization (model_type = ‘glmnet’) applied directly to the side information variables.
 For the real datasets used by Zhang et al.
-[51], we used the exact same implementation as Chao et al.
+(2019), we used the exact same implementation as Chao et al.
 from the file https://github.com/patrickrchao/AdaPTGMM_Experiments/blob/main/AdaFDR_experiments/run_all_exp.R.
 In the gene-drug response data, since only one-dimensional side information is used, we copied the parameters selected from the previous real datasets that also considered just one-dimensional side information.
-S.9 Further implementation details of ZAP
+C.10 Further implementation details of ZAP
 We used the default parameters of zap_asymp from the ZAP package https://github.com/dmhleung/zap.
@@
 The exception was in Simulation 6, where we directly used all 100 side information variables rather than applying a spline basis expansion on each of them.
-In Simulation 5 and 6, we used the test statistics obtained as input, while in the real data experiments we converted the p-values to z-values by using the transformation ±$\pm\Phi^{-1}(p_{i}/2)$ on each p-value $p_{i}$ where the signs are chosen i.i.d.
+ZAP requires explicit test-statistics, specifically z-scores, as input.
+In Simulations 5, 7 and 8, we used the test statistics as input.
+Since the p-values are obtained from a one-sided test, there is a one-to-one correspondence between the p-value information and test statistic information.
+Thus our comparisons in Simulations 5, 7, and 8, are still effectively based on the p-values as we addressed in Section 3.1.
+In Simulation 6, explicit z-score test statistics were not generated by the simulation.
+Hence, we used the inverse normal CDF $\Phi^{-1}$ to convert the p-values to a z-score test statistic.
+In the real data experiments, instead of directly using the $\Phi^{-1}(p_{i})$, we converted the p-values to z-values using the transformation $\pm\Phi^{-1}(p_{i}/2)$ on each p-value $p_{i}$ where the signs are chosen i.i.d.
 uniformly.
-S.10 Further implementation details of AdaFDR
+Otherwise, applying the transformation $\Phi^{-1}(p_{i})$ to true null p-values with a non-decreasing density, or a spike near 1, will produce consistently large z-scores and will be incorrectly discovered by ZAP as an artefact of the data transformation.
+C.11 Further implementation details of AdaFDR
 We applied AdaFDR using the method.adafdr_test function from the AdaFDR package https://github.com/martinjzhang/adafdr.
@@
 This is in accordance to the discussion from their paper that the fast version of AdaFDR is recommended when few discoveries are expected or if the number of hypotheses is small (which is not the case for both of the gene-drug response datasets).
-S.11 Computer specifications for computation times
+For the same reasons, we used AdaFDR with fast_mode = True on Simulation 5 when the number of hypotheses was small, and fast_mode = False on Simulations 7 and 8 when the number of hypotheses were large.
+C.12 Simulation details for p-value setting
+C.12.1 Simulation 6
+Simulation 6 in Section 6.2.2 models “noisy” side information since only the first 2 of the 100 draws are used to distinguish between the true and false null hypotheses.
+Each p-value $p_{i}$ is drawn i.i.d.
+from a two-group beta mixture model, where the true nulls have a uniform distribution and the false nulls have an alternative beta distribution.
+Specifically, the beta mixture model is given by:
+$f(p_{i}\mid\mathbf{x}_{i})=1-\pi(\mathbf{x}_{i})+\pi(\mathbf{x}_{i})\cdot h(p_{i};\mu(\mathbf{x}_{i})),\quad h(p_{i};\mu(\mathbf{x}_{i}))=\left(\frac{1}{\mu(\mathbf{x}_{i})}p_{i}^{\frac{1}{\mu(\mathbf{x}_{i})}-1}\right),$ (7)
+where $\mathbf{x}_{i}\in[0,1]^{100}$ denotes the side information of the $i$th hypothesis, $\pi(\mathbf{x}_{i})$ denotes proportion of the false nulls in the mixture, $h$ denotes the false null distribution of $\text{Beta}(1/\mu(\mathbf{x}_{i}),1)$.
+The $\pi(\mathbf{x}_{i})$’s and $\mu(\mathbf{x}_{i})$’s are determined by the following relationships with the 100-dimensional side information:
+$\log\left(\frac{\pi}{1-\pi}\right)=\theta_{0}+\mathbf{x}_{i}^{T}\theta,\quad\mu_{i}=\max\{\mathbf{x}_{i}^{T}\beta,1\},$
+where $\theta=(3,3,0,\dots,0)\in\mathbb{R}^{100}$, $\beta=(2,2,0,\dots,0)\in\mathbb{R}^{100}$ and $\theta_{0}$ is chosen to satisfy $\frac{1}{m}\sum_{i}\pi(\mathbf{x}_{i})=0.3$.
+Clearly, only the first two values in $\mathbf{x}_{i}\in[0,1]^{100}$ contribute to the beta mixture model, as intended.
+We generated 100 independent datasets, where in each we drew the 2000 p-values as described.
+C.12.2 Simulation 7 (false null probability)
+The $i$th hypothesis is randomly identified as a false null dependent on its side information $\mathbf{x}_{i})$ with probability (Figure C11)
+$\pi_{i}:=\pi(\mathbf{x}_{i}):=0.1\cdot\sum_{j=0}^{2}w_{j}\cdot f_{j}(\mathbf{x}_{i}),$
+where $(w_{0},w_{1},w_{2})=(0.5,0.25,0.25)$, $f_{1}$ and $f_{2}$ are normal densities with parameters $(\mu_{1},\sigma_{1})=(0.25,0.05)$ and $(\mu_{2},\sigma_{2})=(0.75,0.05)$ respectively, truncated between 0 and 1 and normalized such that $\int_{[0,1]}f_{1}dx=\int_{[0,1]}f_{2}dx=1$, and $f_{0}(x):=a\cdot\exp(ax)/(\exp(a)-1)$ where $a=0.5$.
+Figure C11: The probabilities of a hypothesis being a false null, $\pi$, as a function of the hypothesis’ side information, $\mathbf{x}$.
+C.13 Computer specifications for computation times
 Computation times were calculated using an M1 Mac Studio with 20-core CPU and 128GB RAM.
-S.12 Data preparation and searching with HEK293 data
-We downloaded the human proteome (UP000005640, downloaded on 2023/09/23 from UniProt) and prepared a target peptide database along with 5 randomly shuffled decoy peptide databases using Tide-index within the Crux Toolkit v4.1.6809338 [43, 29] with all options set to default.
+C.14 Estimating the computation times of Adaptive Knockoffs
+Running Adaptive Knockoffs (AdaKO) on each of the HEK293 and PRIDE-20 spectrum files is impractical.
+We therefore resorted to estimate its runtime, which requires dealing with two challenges.
+The first is that the runtime of each iteration of AdaKO, where it determines which hypothesis to reveal next, varies with the number of remaining hypotheses.
+The second is that we cannot precisely predict the number of iterations AdaKO will take.
+Addressing the second challenge first, we note that AdaKO starts by revealing the label of any hypothesis $H_{i}$ for which $W_{i}\leq q^{*}_{0.1}$, where $q^{*}_{c}$ is the $c$-quantile of all positive scoring hypotheses.
+Let $m^{\prime}$ be the number of remaining hypotheses after this initial reveal.
+The number of iterations that AdaKO will make before stopping is then $m^{\prime}$-$\ell$, where $\ell$ is the number of hypotheses (target or decoy peptides) that remain when it stops.
+We can estimate $\ell$ with $\hat{\ell}$, the number of target and decoy peptides that score above the 1% FDR cutoff using RESET Ensemble.
+Thus we estimate the number of AdaKO iterations by $m^{\prime}$-$\hat{\ell}$.
+To estimate the runtime of each iteration we resorted to taking $k$ measurements.
+Indeed, for each $i\in\{1,\dots,k\}$, we revealed the labels of any hypotheses $H_{i}$ for which $W_{i}\leq q_{(0.1\cdot i)}^{*}$ and determined the time it took for AdaKO to reveal the next label.
+We chose $k$ to be the largest positive integer such that the number of remaining labels, after revealing the labels $L_{i}$ for which $W_{i}\leq q_{(0.1\cdot k)}^{*}$, is still more than $\hat{\ell}$.
+Thus, we obtain the runtime of $k$ single iterations that are evenly spaced, starting from AdaKO’s first iteration and ending at approximately AdaKO’s last iteration before it terminates with $\hat{\ell}$ remaining hypotheses.
+Next, we computed $\bar{t}_{k}$, the average of those $k$ measurements.
+Thus our overall estimate is:
+$\bar{t}_{k}\cdot(m^{\prime}-\hat{\ell}).$
+In the analysis of the PRIDE-20 data in Section E.2, we considered 10 repeated applications of RESET Ensemble, varying RESET Ensemble’s internal seed.
+In this case, we define $\hat{\ell}$ as the average of the number of winning target and decoy peptides above the cutoff over each application.
+C.15 Method-related figures
+Figure C12: Location of false null hypotheses according to the pair of values $\left(\mathbf{x}_{i1},\mathbf{x}_{i2}\right)$ in the square $\left[-100,100\right]\times\left[-100,100\right]$.
+The same figure is given in Lei and Fithian (2018).
+Appendix D Data preparation
+D.1 Data preparation and searching with HEK293 data
+We downloaded the human proteome (UP000005640, downloaded on 2023/09/23 from UniProt) and prepared a target peptide database along with 5 randomly shuffled decoy peptide databases using Tide-index within the Crux Toolkit v4.1.6809338 (Park et al.
+(2008); Kertesz-Farkas et al.
+(2023)) with all options set to default.
 An output containing the target and decoy peptide pairs are conveniently provided using the --peptide-list T option in Tide-index.
-Each of the RAW 24 HEK293 spectrum files [9] were converted to mzML format using MSConvert 3.0.22314 with the vendor peak-picking filter using the default settings.
-For each of the 5 decoy databases, we searched each spectrum from the combined 24 mzML spectrum files against the combined target-decoy peptide database using Tide-search [15], using the options --top-match 1 --auto-precursor-window warn --auto-mz-bin-width warn --concat T.
+Each of the RAW 24 HEK293 spectrum files (Chick et al.
+(2015)) were converted to mzML format using MSConvert 3.0.22314 with the vendor peak-picking filter using the default settings.
+For each of the 5 decoy databases, we performed a ‘narrow’ search on each spectrum from the combined 24 mzML spectrum files against the combined target-decoy peptide database using Tide-search (Diament and Noble (2011)), using the options --top-match 1 --auto-precursor-window warn --auto-mz-bin-width warn --concat T.
 The resulting search files were then converted to so-called pin files using the make-pin function in Crux.
@@
 absdM The absolute value of the difference between the calculated and observed mass
-Table S4: List of described side information used by HEK293 data.
+Table D3: List of described side information used by HEK293 data.
 The list of side information used by RESET and Adaptive Knockoffs in the HEK293 data, adapted from https://crux.ms/file-formats/features.html.
 Note that the Charge state is represented as a one-hot vector, where we left out a Charge state of +1 (due to linearity, since the one-hot vector sums to one).
-Search algorithms like the one above are typically ‘spectrum-centric’, meaning that for each spectrum, we look for the best matching database peptide, as oppose to the other way round.
-Consequently, the same database peptide may be matched multiple times with different scores, or have no match at all.
-Hence for each target peptide, we define $Z_{i}$ as the maximum XCorr score of all the PSMs associated to that target peptide.
-Similarly for each decoy peptide, we define $\tilde{Z}_{i}$ as the maximum XCorr score of all the PSMs associated to that decoy peptide.
-For target or decoy peptides that do not appear in the pin file, they get a score of $\infty$- ∞.
-Then we compete each target and its corresponding paired decoy by recording the winning score $W_{i}=Z_{i}\vee\tilde{Z}_{i}$ along with the label $L_{i}=sign(Z_{i}-\tilde{Z}_{i})$, randomly breaking any ties.
-Peptides with a winning score of $\infty$- ∞, indicating that neither the target or its paired decoy are in the pin file, are thrown out.
-We assign each winning peptide the side information defined as the auxiliary information $\mathbf{x}_{i}$ of the underlying PSM that had the XCorr score of $W_{i}$.
-The side information that was subsequently used by RESET and Adaptive Knockoffs is given in Table S4.
-S.13 Estimating the computation times of Adaptive Knockoffs
-Running Adaptive Knockoffs (AdaKO) on all peptide datasets would be infeasible.
-Therefore, we resorted to estimating its runtimes by considering the time it took to complete a single iteration that determines which hypothesis to reveal next, and multiplying this value by the estimated number of iterations before the procedure stops.
-Since the time it takes a single iteration depends on the number of hypotheses that were already revealed, we took different such measurements as we varied this number, and averaged them.
-Specifically, as AdaKO does, we first revealed the labels that were scored less than or equal to the 10% quantile of non-zero scoring hypotheses, and recorded the time it takes for a single interation from this point.
-To be clear, the revealed labels thus includes those zero-scoring hypotheses since they will be less than this score cutoff — it is only the 10% quantile that considers the non-zero scoring hypotheses.
-Then we took similar measurements by increasing the ‘10%’ value, in 10% increments, up to a value of ‘10k%’ where we define $k$ next.
-We chose the value of $k$ so that the proportion of revealed labels, $10k\%$, estimates the number of revealed labels when AdaKO terminates at the 1% FDR level.
-That way, the $k$ times should provide an accurate sample of the number single iteration times across all of AdaKO’s iterations until it terminates.
-Of course, we are unable to compute when Adaptive Knockoffs might terminate, so we estimated this using the value $q$, the number of target and decoy peptides that scored above the 1% FDR cutoff using RESET Ensemble.
-Then we defined $k=\lfloor(m-q)/m\cdot 10\rfloor$, where $m$ is the total number of hypotheses (peptides) considered by Adaptive Knockoffs.
-We estimated the runtime by
-$\bar{t}_{k}\cdot(m^{\prime}-q),$
-where $\bar{t}_{k}$ is the average of the $k$ measured single-iteration times, and $m^{\prime}$ is the number of labels that are yet to be revealed at the start of AdaKO’s iterations, i.e., the number of labels with scores above the 10% quantile of non-zero scoring hypotheses.
-Thus the value $m^{\prime}-q$ is our estimate of the number of iterations until Adaptive Knockoffs stop.
-In Section 5.2.2, we considered 10 repeated applications of RESET Ensemble, varying RESET Ensemble’s internal seed.
-In this case, we define $\hat{\ell}$ as the average of the number of winning target and decoy peptides above the cutoff over each application.
-S.14 Data preparation and searching with PRIDE-20 data
-We followed the preparation and searching outlined in [18].
+We constructed the target and decoy peptide scores, $Z_{i}$ and $\tilde{Z}_{i}$ respectively, according to our description in Section C.6 using the XCorr score as the PSM scoring function.
+Subsequently we record the winning score $W_{i}=Z_{i}\vee\tilde{Z}_{i}$ along with the label $L_{i}=sign(Z_{i}-\tilde{Z}_{i})$.
+In the case of ties ($Z_{i}=\tilde{Z}_{i}$), we randomly set $L_{i}$ to be +1+1 or 1-1 with equal probability.
+To each pair $(W_{i},L_{i})$ we attribute the side information defined as the auxiliary information $\mathbf{x}_{i}$ of the underlying PSM that had the XCorr score of $W_{i}$.
+In the case there is more than one PSM with the same XCorr score of $W_{i}$, we randomly select one of the target PSMs if $L_{i}=+1$ or one of the decoy PSMs if $L_{i}=-1$.
+The side information that was subsequently used by RESET and Adaptive Knockoffs is given in Table D3.
+D.2 Data preparation and searching with PRIDE-20 data
+We followed the preparation and searching outlined in Freestone et al.
+(2024).
 The only difference is that we used an updated version of the Crux toolkit.
 We provide the following brief description for the reader’s convenience.
-Twenty spectrum files and protein databases from separate projects, referred to as the PRIDE-20 dataset, were downloaded.
+Twenty spectrum files and protein databases from separate projects, referred to as the PRIDE-20 dataset, were downloaded (see Table 2 in Freestone et al.
+(2024)).
 Tide-index was used to prepare the target peptide database along with 10 randomly shuffled decoy peptide databases for each of the 20 spectrum files.
-Tide-search was used to search each spectrum file against the 10 combined target-decoy databases using two modes: ‘narrow’ and ‘open’ search options as they outline.
-Briefly a ‘narrow’ search, which is the default search mode, only searches each spectrum in the spectrum file against peptides in the target-decoy database that have a theoretical mass within a small tolerance of the sample peptide’s mass that generated the experimental spectrum.
-On the other hand, an ‘open’ search makes this mass-tolerance larger so that a spectrum may match with a peptide with a completely different mass.
-This is desirable at times, since sample peptides can undergo post-translational modifications which modify the mass of the peptide and therefore the experimental spectrum.
-Hence when searching in an ‘open’ mode, the modified spectrum may still be correctly matched.
+Tide-search was used to search each spectrum file against the 10 combined target-decoy databases using two modes: ‘narrow’ and ‘open’ search options as outlined in Freestone et al.
+(2024).
 The search files were then converted to pin files using the make-pin function in Crux and subsequently filtered for the top 1 (the optimal) peptide-spectrum match for each spectrum.
 For each spectrum file and search type, we obtained the triples $(L,W,\mathbf{x})$ in the following way.
-For each spectrum file that had no modifications (we explain this in more detail next), we proceeded essentially in the same way as in Section S.12.
-That is, for each peptide in the pin file, we define $Z_{i}$ (if a target peptide) or $\tilde{Z_{i}}$ (if a decoy peptide) as the maximum Tailor score of all the PSMs associated to the target peptide [47].
-Here we are using the more sensitive Tailor scores and relegate the XCorr scores as one of the side information variables.
-Then we compete each target and its corresponding paired decoy in the same way as Section S.12, to obtain the triple $(L_{i},W_{i},\mathbf{x}_{i})$ where the side information $\mathbf{x}_{i}$ is obtained from the auxiliary information of the underlying PSM that had the Tailor score of $W_{i}$.
-A description of the side information used is given in Table S5.
+For each spectrum file that had no modifications (we explain this in more detail next), we proceeded essentially in the same way as in Section D.1.
+That is, for each peptide in the pin file, we define $Z_{i}$ (if a target peptide) or $\tilde{Z_{i}}$ (if a decoy peptide) as the maximum Tailor score of all the PSMs associated to the target peptide (Sulimov and Kertész-Farkas (2020)).
+Here we are using the more sensitive Tailor score and relegate the XCorr score as one of the side information variables.
+Then we compete each target and its corresponding paired decoy in the same way as Section D.1, to obtain the triple $(L_{i},W_{i},\mathbf{x}_{i})$ where the side information $\mathbf{x}_{i}$ is obtained from the auxiliary information of the underlying PSM that had the Tailor score of $W_{i}$.
+A description of the side information used is given in Table D4.
 Side Information Description
@@
 absdM The absolute value of the difference between the calculated and observed mass
-Table S5: List of described side information used by PRIDE-20 data.
+Table D4: List of described side information used by PRIDE-20 data.
 The list of side information used by RESET and Adaptive Knockoffs in the PRIDE-20 data, adapted from https://crux.ms/file-formats/features.html.
@@
 As an example, PEPTIDE may generate the following ‘copies’: PE[16]PTIDE, PEPTIDE[16] and PE[16]PTIDE[16], where [16] indicates an increase of 16 Daltons to the amino acid on the left.
-Consequently, each of these peptides are usually correlated in the data — i.e., if one of these peptides is scored high, than usually they all are.
-Hence in order to satisfy Assumption 3, we follow the protocol outlined in [18].
+Consequently, the scores of all these peptides are usually correlated: if one of these is scored high, than often they all are.
+Hence in order to satisfy Assumption 3, we follow the protocol outlined in Freestone et al.
+(2024).
 That is, we identify all the copies as being the same peptide, ‘PEPTIDE’, with each of the variable modifications ignored.
 Then we proceed with the same steps as the above paragraph, by determining the maximum score associated with PEPTIDE and recording it as $Z_{i}$ (or $\tilde{Z}_{i}$ if it is a decoy).
-We define the labels, winning scores and side information in the same way to obtain our triplets $(L,W,\mathbf{x}_{i})$.
-Due to runtime considerations, we used 13 of the smallest narrow search files to assess Adaptive Knockoff’s power as described in Section 5.2.2 of the manuscript, and only one of the 10 target-decoy databases.
+We define the labels, winning scores and side information in the same way to obtain our triplets $(L,W,\mathbf{x})$.
+Due to runtime considerations, we used 13 of the spectrum files to assess Adaptive Knockoff’s power as described in Section 5.2.2 of the manuscript (each resulting in <33K<33K PSMs), and only one of the 10 target-decoy databases.
 To fairly compare with Adaptive Knockoffs, we used the same target-decoy databases for RESET ensemble.
-All 20 spectrum files and all 10 target-decoy databases for each spectrum file were used in the comparison of RESET ensemble and FDP-SD while controlling the FDP.
-S.15 Brief justification of the assumptions used with HEK293 and PRIDE-20 datasets
-A standard assumption of the mass spectrometry literature, which is also empirically validated, is that the label $L_{i}$, indicating whether an incorrect target or its decoy scored higher, is uniformly ±$\pm 1$ independently of all other target-decoy pairs and scores [21, 39].
-RESET additionally requires this statement to be true if we further condition on the side information.
-This is obviously true for some side information variables which are constant between the target and corresponding decoys — e.g., PepLen, Charge, lnNumSP, dm, absdM.
-The other side information variables are score-related — e.g., Xcorr, deltCn and deltLCn, and so conditioning on them assumes no more than the ‘standard assumption’.
-Hence, we believe Assumption 3 with $P(L_{i}=1)=c_{0}=1/2$ to be reasonably satisfied.
-S.16 Preparation of data with p-values
-Each of the 10 datasets from Zhang et al.
-were downloaded from https://github.com/patrickrchao/AdaPTGMM_Experiments/tree/main/AdaFDR_experiments/data_files, containing the p-value and side information pairs, $(p,\mathbf{x})$.
-All pairs were then used to report a list of discoveries by each p-value based method with the following exceptions.
-In the RNA-seq datasets, Airway, Bottomly and Pasilla, there exists ‘spikes’ of high p-values.
-These ‘spikes’ concentrate in certain regions of the side information, i.e., when the log-normalized gene counts are low.
-Hence we removed those p-values with log-normalized gene counts that were less than zero.
-Figure S1 show the histograms of the p-values before and after this removal.
-The gene-drug response data was obtained from the adaptMT package.
-Figure S1: Histograms of p-values in the RNA-seq data before and after p-value removal.
+All 20 spectrum files and all 10 target-decoy databases for each spectrum file were used in comparing RESET Ensemble’s FDX control with FDP-SD’s.
+D.3 Description of real data with p-values
+We provide the following description of each dataset used in Section 6.4 in the main text and in Section F.3.
+The following three RNA-seq datasets, the proteomics dataset, the two microbiome datasets, the two fMRI datasets, and the two eQTL datasets were downloaded from https://github.com/patrickrchao/AdaPTGMM_Experiments/tree/main/AdaFDR_experiments/data_files.
+The last gene-drug response data was obtained from the adaptMT package.
+  • Three RNA-seq datasets (Airway (Himes et al.
+(2014)), Bottomly (Bottomly et al.
+(2011)), Pasilla (Brooks et al.
+(2011))): The goal is to identify genes that are associated with varying gene expression levels in response to a change of conditions.
+In the Airway data, the differential expression analysis is conducted in response to a drug, dexamethasone; in the Bottomly data, the analysis is conducted between two mouse strains; and in the Pasilla data, the analysis is conducted between normal and so-called ‘Pasilla knockdown’ conditions (Huber and Reyes ).
+In all cases the side information consists of the log-normalized gene expression read counts.
+  • A proteomics dataset (Proteomics (Dephoure and Gygi (2012))): The goal is to identify proteins that have different protein abundances when treated with rapamycin versus DMSO in yeast cells.
+Each protein consists of side information equal to the natural log of the number of peptides belonging to the protein that were ‘quantified’ in all six samples.
+  • Two microbiome datasets (Microbiome Enigma Ph and Microbiome Enigma Al (Smith et al.
+(2015); Korthauer et al.
+(2019))): The goal is to identify microbiome organisms, recorded as OTUs (Operational Taxonomic Units), that are associated with certain conditions (pH and Al).
+Each hypothesis (OTU) is equipped with a two-dimensional side information — the ubiquity, which is defined as the percentage of samples detected with the OTU, and the mean nonzero abundance, which is defined as the average abundance of each OTU across the samples in which it was detected.
+  • Two fMRI datasets (Auditory and Imagination (Tabelow and Polzehl (2011))): The goal is to identify voxels that are activated in response to different stimuli.
+In the first dataset, a single individual received auditory stimulus and in the second dataset, a single individual was instructed to imagine playing tennis.
+Each hypothesis (voxel) is equipped with four dimensional side information: the spatial position of the voxel (in three dimensions) and a categorical variable with the Brodmann label (Brodmann (1909)), which is used to delineate different areas of the brain with different functions.
+  • Two eQTL datasets (Subcutaneous and Omentum (Consortium et al.
+(2015); Consortium (2020))): The goal is to identify single nucleotide polymorphisms (SNPs) that are associated with changes in gene expression levels.
+Such an association is referred to as an expression quantitative trait locus (eQTL).
+Each hypothesis (SNP) is equipped with the following four side information variables: (1) the distance between the SNP and the gene transcription’s start site (TSS), (2) the log gene expression level, (3) the alternative allele frequency (AAF) of the SNP and (4) the chromatin state of the SNP.
+  • Gene-drug response data (Estrogen (Davis and Meltzer (2007); Li and Barber (2017))): The goal is to identify genes that are differentially expressed in response to a low-dosage estrogen treatment applied to breast cancer cells.
+Each hypothesis (gene) is assigned a ‘rank’ that was generated using an analysis of the gene expression levels in response to (1) a high-dosage estrogen treatment and (2) a medium-dosage estrogen treatment.
+Here, the ranks are in terms of the strength of association between the expression level of each gene and the level of dosage in (1) and (2).
+Thus, we can analyse the data twice, once for each side information variable based on (1) and (2).
+D.4 Preparation of data with p-values
+All pairs of p-values and side information from the above datasets were used to report a list of discoveries using each p-value based method with the following exceptions.
+In the RNA-seq datasets, Airway, Bottomly and Pasilla, there exists ‘spikes’ of high p-values (Figure D1, top row).
+These spikes are associated with genes with low log-normalized read count.
+Hence, removing all genes with log-normalized read count less than zero — whose information content is questionable anyhow — also eliminated those undesirable spikes (Figure D1 bottom row).
+The number of genes removed is given in Table D5.
+Dataset # of genes removed % of genes removed
+Airway 10616 32%
+Bottomly 2203 16%
+Pasila 1575 13%
+Table D5: The number and percentage of genes removed from the RNA-seq datasets.
+Figure D1: Histograms of p-values in the RNA-seq data before and after removal of genes with negative log-normalized read count.
 Top row: the original histograms of the p-values.
-Bottom row: the corresponding histograms after removal of the p-values associated with low counts.
-S.17 Tables
-Table S6: The number of discoveries and computations times for 13 of the PRIDE-20 dataset at the 1% FDR level.
-We calculated the number of discoveries at the 1% FDR using each method of Adaptive Knockoffs and TDC.
-We also computed computation times for Adaptive Knockoffs (in minutes).
-We used both the actual times and the estimated times according to Section S.13.
-Project ID Method Estimated Time Actual Time Discoveries
-PXD002470 AdaKO EM 0.58 1.17 300
-AdaKO GAM 0.03 0.08 272
-AdaKO LR 0.07 0.14 483
-AdaKO RF 7.50 9.04 488
-TDC - - 455
-PXD006856 AdaKO EM 24.72 59.80 0
-AdaKO GAM 0.80 2.17 379
-AdaKO LR 3.67 4.63 428
-AdaKO RF 427.01 486.70 421
-TDC - - 425
-PXD008920 AdaKO EM 33.81 48.63 8485
-AdaKO GAM 0.95 2.39 8443
-AdaKO LR 2.98 6.11 9143
-AdaKO RF 250.66 270.67 9352
-TDC - - 9086
-PXD008996 AdaKO EM 119.73 188.54 7464
-AdaKO GAM 3.39 7.55 7958
-AdaKO LR 18.77 20.18 8481
-AdaKO RF 1430.16 1626.96 8625
-TDC - - 8413
-PXD010504 AdaKO EM 134.07 185.59 459
-AdaKO GAM 4.03 5.11 1358
-AdaKO LR 9.25 11.37 1585
-AdaKO RF 2134.38 2420.31 1626
-TDC - - 1551
-PXD012528 AdaKO EM 202.99 241.66 0
-AdaKO GAM 6.37 8.34 239
-AdaKO LR 13.57 16.34 314
-AdaKO RF 3769.54 4143.21 405
-TDC - - 342
-PXD012611 AdaKO EM 4.79 7.59 2850
-AdaKO GAM 0.15 0.26 3152
-AdaKO LR 0.62 0.98 3250
-AdaKO RF 43.53 47.01 3262
-TDC - - 3260
-PXD013274 AdaKO EM 34.65 49.52 14025
-AdaKO GAM 0.81 1.70 14271
-AdaKO LR 4.59 3.92 14791
-AdaKO RF 205.10 248.82 14765
-TDC - - 14769.0
-PXD016724 AdaKO EM 499.32 598.31 3617
-AdaKO GAM 13.76 24.01 3802
-AdaKO LR 36.84 58.90 4536
-AdaKO RF 8070.09 8882.65 4809
-TDC - - 4745
-PXD019186 AdaKO EM 247.15 389.51 0
-AdaKO GAM 7.04 13.49 1020
-AdaKO LR 34.77 30.35 942
-AdaKO RF 4143.83 4427.20 1217
-TDC - - 912
-PXD022257 AdaKO EM 151.82 244.22 1619
-AdaKO GAM 15.50 9.94 1918
-AdaKO LR 12.34 23.81 2500
-AdaKO RF 2475.32 2638.35 2501
-TDC - - 2509
-PXD023571 AdaKO EM 30.23 43.86 1030
-AdaKO GAM 3.04 4.40 1743
-AdaKO LR 2.92 10.34 0
-AdaKO RF 437.75 474.17 2161
-TDC - - 1505
-PXD026895 AdaKO EM 62.89 85.62 3255
-AdaKO GAM 8.94 4.94 3676
-AdaKO LR 10.02 12.04 4044
-AdaKO RF 847.23 835.02 4047
-TDC - - 3980.0
-S.18 Figures
-Figure S2: Estimated FDR of each method in numerical simulations.
-Each panel plots the estimated FDR for each method at FDR thresholds ranging from 5% to 30%.
-The first row corresponds to Simulation 1 with three values of $k\in\{50,150,300\}$, the second row corresponds to Simulation 2 with three values of $k\in\{150,300,450\}$, and the last row corresponds to Simulation 3 and 4.
+Bottom row: the corresponding histograms after the removal of genes with low read counts.
+Appendix E Supplementary results
+E.1 Further comparisons between RESET and AdaKO in simulations
+For each simulation and FDR threshold in Figure 1 of the main text, we determined the maximum power across all RESET and AdaKO methods and computed the relative power of each method to this maximum.
+The resulting relative powers are reported as boxplots in Figure E1A, giving an indication of the performance of each method across the simulations as a whole.
+Additional plots showing the performance of each method in each separate simulation are given in Figures E2 (power) and E3 (estimated FDR).
+While in Figure E2 we often find a version of AdaKO that is more powerful, or at least comparable to RESET Ensemble, this varies between the versions.
+On the other hand, RESET Ensemble’s performance is consistent across these simulations, and overall typically achieves the highest relative power when the results are aggregated.
+Interestingly, RESET NN seems to be marginally more powerful than RESET Ensemble.
+While this may be the case, we still recommend RESET Ensemble as the default approach to alleviate the burden of selecting any single classification algorithm, as for example, RESET Ensemble appears to perform marginally better in the p-value setting (see Figure 6 of the main text).
+Figure E1: Relative power and runtimes for each method.
+(A) Each boxplot describes the relative power for each method, computed across the combination of simulations and FDR thresholds in Figure 1.
+(B) The log10 runtimes of each method applied at the 5% FDR level for 5 runs of Simulation 4.
 For readability, the points are jittered in the horizontal direction.
-Figure S3: Estimated power of each method in numerical simulations.
-Each panel plots the power for each method at FDR thresholds ranging from 5% to 30%.
-The first row corresponds to Simulation 1 with three values of $k\in\{50,150,300\}$, the second row corresponds to Simulation 2 with three values of $k\in\{150,300,450\}$, and the last row corresponds to Simulation 3 and 4.
+One advantage of RESET is its computational speed.
+Indeed, our implementation of RESET, as described in Section 4.4 of the main text, has many steps that can be parallelized: the $r$ evaluations of each classification algorithm on the $K$ folds can all be done in parallel using multiple cores.
+In contrast, Adaptive Knockoffs must iteratively apply a classification algorithm for every nonzero scoring hypothesis in the candidate set $\mathcal{\hat{S}}_{t}$ until the estimated FDR is $\leq\alpha$, or all hypotheses are revealed.
+Figure E1B demonstrates this advantage showing the (log10) runtimes of RESET (using 20 CPUs) and of Adaptive Knockoffs at the 5% FDR level across 5 runs of Simulation 4.
+We find that each version of RESET is significantly faster than each version of Adaptive Knockoffs.
+RESET Ensemble is around $10^{3}$ to $10^{4}$ times faster than AdaKO EM, $10^{2}$ to $10^{3}$ times faster than AdaKO RF, and about 3-5 times faster than AdaKO LR and AdaKO GAM.
+Figure E2: Estimated power of each method in numerical simulations.
+Each panel shows the power for each method at FDR thresholds ranging from 5% to 30%.
+The first row corresponds to Simulation 1 with three values of $k\in\{50,150,300\}$, the second row corresponds to Simulation 2 with three values of $k\in\{150,300,450\}$, and the last row corresponds to Simulations 3 and 4.
 For readability, the points are jittered in the horizontal direction.
 A description of the AdaKO methods and KO can be found in Section 2.
+Figure E3: Estimated FDR of each method in numerical simulations.
+Each panel shows the estimated FDR for each method at FDR thresholds ranging from 5% to 30%.
+The estimated FDR is calculated as the proportion of irrelevant variables in the discovery list, averaged over the runs.
+The first row corresponds to Simulation 1 with three values of $k\in\{50,150,300\}$, the second row corresponds to Simulation 2 with three values of $k\in\{150,300,450\}$, and the last row corresponds to Simulations 3 and 4.
+For readability, the points are jittered in the horizontal direction.
+E.2 Comparison of AdaKO RF and RESET Ensemble using PRIDE data
+We looked at the performance of RESET Ensemble versus AdaKO RF because in this setup it was the most powerful implementation of Adaptive Knockoffs.
+Due to the computational constraints of AdaKO RF, we focused on the results using a narrow search.
+In Figure E4A, the left boxplot displays the average number of discoveries by RESET Ensemble, divided by the number of discoveries by AdaKO RF using thirteen of the PRIDE-20 spectrum files.
+In six of the thirteen PRIDE spectrum files, we found RESET Ensemble to be the superior method, while in the remaining seven spectrum files, AdaKO RF produced the same or more discoveries.
+In one dataset, AdaKO RF discovered 1,217 peptides — approximately an 11% increase over the average of 1,084 peptides discovered by RESET Ensemble.
+Figure E4: Comparison of RESET Ensemble and AdaKO RF.
+For each of the 13 PRIDE datasets, we computed the average number of discoveries using 10 applications of RESET Ensemble, varying RESET’s internal seed.
+The left panel shows a boxplot of those RESET Ensemble averages divided by the number of discoveries using AdaKO RF (left boxplot) and TDC (right boxplot) where each circle corresponds to one of the 13 smaller PRIDE-20 datasets.
+In the centre panel, we compare the estimated and actual runtimes for AdaKO RF.
+Lastly, in the right panel, we compare the actual runtimes of RESET Ensemble and AdaKO RF (in log10 scale) on the same 13 PRIDE-20 datasets.
+We also assessed the computational times of AdaKO RF in Figure E4B, using both the actual and estimated runtimes as outlined in Section C.14.
+Reassuringly, our estimated times were less than the actual times taken by AdaKO RF to complete.
+Although we find that, when analysing these smaller spectrum sets, AdaKO RF is the most powerful method of Adaptive Knockoffs, and it is competitive to RESET Ensemble power-wise, it is again clear that AdaKO RF is generally too slow with some runtimes exceeding multiple days even for these small datasets.
+In fact, on the dataset where AdaKO RF outperforms RESET Ensemble by about 11%, it takes 3.07 days for AdaKO RF to discover its peptides while RESET Ensemble takes about half a minute to complete.
+Figure E4C shows that in the worst cases, RESET takes roughly 1 minute while AdaKO RF takes over 6 days.
+The issue of runtimes is especially problematic considering the fact that multiple spectrum files are often analysed jointly, as in our HEK293 analysis in Section 5.2.1.
+E.3 FDX control in mass spectrometry data
+We evaluated RESET Ensemble’s FDX control using the HEK293 and all twenty PRIDE-20 spectrum files (see Sections D.1 and D.2 for data preparation and search settings).
+As in Section 5.2.1, the joint HEK293 spectrum files were searched against 5 distinct target-decoy databases, where each target-decoy database was generated by concatenating the target database to a decoy database consisting of shuffled target peptides.
+Similarly, each PRIDE-20 spectrum file was searched against 10 distinct target-decoy databases.
+We used the coinflip version of FDP-SD (Luo et al.
+(2023)), which provides a uniform increase in the number of discoveries (at the cost of some randomness in the discovery list, see Supplementary Algorithm A2).
+Our results using RESET Ensemble and FDP-SD are given in Table E1 and Figure E5.
+Table E1: The average number of (HEK293) discoveries while controlling the FDX.
+The average number of discoveries reported for RESET Ensemble and FDP-SD on the combined HEK293 spectrum files at different confidence values $1-\gamma=0.5,0.8,0.9$ and an FDP threshold of 1%.
+The averages are taken over 5 applications of each method, once for each of the 5 combined target-decoy databases.
+Confidence $1-\gamma$
+Method 0.5 0.8 0.9
+RESET Ensemble 82930 82677 82508
+FDP-SD 75977 75450 75255
+RESET Ensemble delivers on average approximately 9-10% more HEK293 discoveries than FDP-SD across all confidence levels (Table E1).
+The picture is somewhat murkier when considering the PRIDE-20 datasets: while RESET Ensemble appears to be generally more powerful (Figure E5, zoomed-in right panel), it fails to make any discoveries at high confidence levels in a few datasets (Figure E5, left panel).
+Upon a closer look, the numbers of discoveries made by FDP-SD in these datasets are relatively small, with the largest average number of discoveries among these datasets being 417.
+When the number of discoveries is expected to be small and the confidence parameter is high, RESET may yield less power, even if the side information is reasonably informative — a phenomenon we dwell on in Section F.4.
+RESET Ensemble advantage over FDP-SD is clearer when applying the two procedures to open as opposed to narrow searches.
+As discussed in Section D.2, an open search allows for the detection of peptides whose masses have been altered by post-translational modifications (PTMs).
+In this setting, side information variables such as dm, which records the mass difference between the peptide and the spectrum, can be highly informative given that some differences will correspond to masses of common PTMs.
+Thus, it is not surprising that RESET, which can take advantage of such side information, improves on FDP-SD even further across all confidence levels in this context.
+Figure E5: Comparison of RESET Ensemble and FDP-SD.
+The left panel shows boxplots of the average number of discoveries using RESET divided by the average number of discoveries using FDP-SD (the right panel is just a ‘zoomed-in’ version).
+We used all twenty PRIDE-20 datasets at an $\alpha=1\%$ FDP threshold and varying confidence parameters with both narrow and open search modes.
+The averages are taken over 10 applications, one for each of the 10 target-decoy database pairs we generated for each of the PRIDE-20 dataset.
+E.4 Result-related figures
+Figure E6: Comparison of RESET and other methods using a collection of publicly available datasets.
+Each method was evaluated at the FDR thresholds of 1%, 5%, 10%, and 20%.
+The number of discoveries reported by each RESET method is averaged over 10 applications at FDR threshold and dataset.
+Each colour and shape combination corresponds to a unique method.
+For readability, each point is jittered in the horizontal direction.
+Appendix F Supplementary analyses
+F.1 Analysis of AdaDetect and empirical p-values
+The main assumption of AdaDetect’s FDR control in the novelty detection problem, is that the null observations $\left(Y_{1},\dots,Y_{n}\right)$ and the null test set $\left(X_{i}:X_{i}\sim P_{0}\right)$ are exchangeable conditioned on the novelties $\left(X_{i}:X_{i}\not\sim P_{0}\right)$ (Marandon et al.
+(2024)).
+In the multiple testing setup, the closest analogue to the null observations is the collection of decoy information $\{(\mathbf{x}_{i},W_{i}):L_{i}=-1\}$, and the analogue to the null test set is the true null target information $\{(\mathbf{x}_{i},W_{i}):L_{i}=1,i\in N\}$.
+However the exchangeability assumption of AdaDetect in this setting fails for two reasons: (1) the collection of decoy information $\{(W_{i},\mathbf{x}_{i}):L_{i}=-1\}$ contains a mixture of true and false null hypotheses which cannot be discerned and (2) the true null hypotheses can be heterogeneous.
+Despite failing to meet the assumptions for guaranteed FDR control, we nonetheless verified this in practice through the following simulation inspired by Zhang et al.
+(2019).
+We generated $m=20K$ hypotheses with side information $\mathbf{x}_{i}\sim U[0,1]$, each drawn independently.
+Each hypothesis is declared a false null with probability $\pi_{i}$ and a true null otherwise, where
+$\pi_{i}=\pi(\mathbf{x}_{i}):=0.1\cdot\varphi^{c}_{[0.75,0.05]}\left(\mathbf{x}_{i}\right),$
+and $\varphi^{c}_{[\mu,\sigma]}$ denotes the 0-1 clamping (truncating between 0 and 1) of the normal density with mean $\mu$ and standard deviation $\sigma$.
+For each hypothesis we drew $z_{i}\sim\mathcal{N}(\mu_{i},1)$ where
+[…]
+where $B_{i}\in\{0,-0.5\}$ with equal probability and independent of everything else.
+One sided p-values are then constructed as $p_{i}=1-\Phi(z_{i})$, where $\Phi$ is the standard normal CDF.
+This construction allows us to create a heterogeneous mixture of true nulls that violate the exchangeability assumptions.
+We evaluated two versions of RESET on the simulated data, the first is the standard RESET Ensemble as described in Section 4.4 in the main text, and the second adopts AdaDetect’s empirical p-value approach, which we now describe.
+For each target ($L_{j}=1$), we defined the empirical p-value by comparing the target’s score with the scores of the estimating decoys:
+$p_{j}:=\frac{1+\#\{k:\widetilde{W}_{k}>\widetilde{W}_{j},\tilde{L}_{k}=+1,L_{k}=-1\}}{1+\ell},$ (8)
+where $\ell:=\#\{k:\tilde{L}_{k}=+1,L_{k}=-1\}$ is the number of estimating decoys.
+Next, we used Storey’s procedure which estimates the fraction of true null hypotheses, $\pi_{0}$, and applies the BH procedure at the adjusted level of $\alpha/\hat{\pi}_{0}^{Storey}$.
+Storey’s estimate is given by
+$\hat{\pi}_{0}^{Storey}:=\frac{1+H(\lambda)}{(1-\lambda)\cdot\#\{L_{k}=+1\}},$
+where $H(\lambda)$ denotes the number of empirical p-values $\geq\lambda$.
+We used $\lambda=\frac{\lfloor\ell/2\rfloor}{\ell+1}$.
+This empirical p-value version of RESET is essentially applying the AdaDetect methodology.
+The only difference is that we can conduct a fair comparison in which both versions of RESET offer the same semi-supervised learning approach.
+Figure F1 shows the empirical FDRs using the empirical p-value version of RESET Ensemble (denoted as RESET Ensemble-p) versus the usual RESET Ensemble, averaged over 100 realizations of the data.
+At 10% and 20% FDR thresholds, we find that empirical FDRs calculated for RESET Ensemble-p are at least 5 standard errors above their respective thresholds, empirically demonstrating failure to control the FDR.
+Emery et al.
+(2020) show a similar failure to control the FDR using BH or Storey’s procedure using pooled empirical p-values when the true null hypotheses are heterogenous (which they call “noncalibrated”).
+Figure F1: Estimated FDRs with and without empirical p-values.
+The confidence intervals correspond to $\pm 5$ times the estimated standard error.
+For readability, points are jittered in the horizontal section.
+The horizontal segments mark the FDR thresholds on the y-axis.
+F.2 Brief justification of the assumptions used with HEK293 and PRIDE-20 datasets
+TDC (equivalently SSS+) has become the de facto method of FDR analysis in the proteomics literature since its introduction by Elias and Gygi (2007).
+An assumption of TDC for valid FDR control is that the label $L_{i}$, indicating whether an incorrect target or its decoy scored higher, is uniformly $\pm 1$ independently of all the scores as well as of the labels corresponding to the correct targets (Assumption 1 of the main text).
+This assumption relies heavily on the decoy construction which we now outline.
+There are two main approaches to decoy construction: randomly shuffling or reversing each target peptide sequence.
+This random shuffling or reversing is typically applied to the non-terminal amino acids, i.e., the amino acids between the first and last.
+The rationale is that the terminal amino acids of target peptides are typically of a certain type, e.g., lysine or arginine, and therefore they can be used to distinguish target and decoy peptides if they are not preserved.
+By either shuffling or reversing in this way, the result is a set of decoy peptides that have the exact same length, amino acid profile, and terminal amino acids as their corresponding target peptides, thus creating realistic competition for incorrect target discoveries in an attempt to satisfy the above assumption.
+While both approaches are commonly used, in this paper we use randomly shuffled decoys so that our results can be averaged across different decoy generations.
+Importantly, the above assumption has been corroborated empirically through entrapment experiments (Granholm et al.
+(2011)).
+In these experiments, a sample of known peptides is used so that the actual FDP can be gauged.
+This strategy has allowed others (He et al.
+(2015); Lin et al.
+(2022)) to show that the FDR is controlled using TDC, suggesting that Assumption 1 are reasonable in practice.
+RESET additionally requires this statement to be true if we further condition on the side information $\mathbf{x}$.
+This is obviously true for side information variables that are constant between the target and corresponding decoys — e.g., PepLen, Charge, lnNumSP, dm, absdM.
+The other side information variables are score-related — e.g., Xcorr, deltCn and deltLCn, and so conditioning on them assumes no more than the ‘standard assumption’ that already conditions on such a score in the first place.
+Hence, we believe Assumption 3 with $P(L_{i}=1)=c_{0}=1/2$ to be reasonably satisfied.
+Moreover, in our recent publication, we empirically demonstrated through entrapment experiments that RESET (combined with Percolator, a semi-supervised algorithm designed for peptide detection) controls the FDR (Freestone et al.
+(2025)).
+F.3 Analysis of Adipose datasets
+Figure F2: Comparison of power in Adipose data.
+In (A) and (B), we compare the number of discoveries using the two eQTL datasets.
+The number of discoveries reported by each RESET method is averaged over 10 applications at a given FDR threshold and dataset.
+In both panels, we evaluated each method at FDR thresholds of: 1%, 5%, 10%, 20%.
+Notably, we have excluded from Figure 6 the analysis of the two eQTL studies for the following reason.
+Due to linkage disequilibrium these datasets exhibit a high degree of p-value as well as side-information dependency, even among the true null hypotheses.
+Hence these datasets violate Assumption 3 which guarantees our — as well as the other considered methods’ — type-1 error control.
+It is thus unclear whether our comparisons would remain fair, particularly if one method violates FDR control more than the others.
+Notably, RESET Ensemble flagged this data as exhibiting suspicious dependence using the feature outlined in Section C.2.4.
+Figure F2(A-B) shows that overall RESET Ensemble is able to report a similar number of discoveries to other methods, with the exception of $\alpha=20\%$ in the Adipose Subcutaneous dataset, where RESET Ensemble reports substantially less discoveries.
+Even if we overlook the dependency in the data, and include the two eQTL studies in our power analysis in Figure 6, RESET Ensemble still ranks the highest.
+Indeed, Figure F3 shows the analogous results with the eQTL studies included.
+RESET Ensemble has the highest median at 93.6%, followed by AdaFDR at 93.2% and RESET NN at 91.8%.
+Figure F3: Relative power of each method.
+The boxplots show the relative power of each method, calculated for each dataset described in Section D.3 (including the two Adipose datasets) and across FDR thresholds of 1%, 5%, 10% and 20%.
+F.4 Further analysis of RESET with FDX control
+We point out the following limitation on the topic of RESET’s FDX control.
+When the number of discoveries is small, e.g., when the confidence is too high, RESET may yield less power, even if the side information is reasonably informative.
+To illustrate this phenomenon, we computed the bounds, $\delta_{i}$ on the number of decoy wins in the top $i$ scores from Algorithm A2, which are used to determine the reported list of discoveries in FDP-SD (with $c=1/2$) and RESET (using the default $c=2/3$).
+We considered a confidence level of $1-\gamma=90\%$ and an FDP threshold of $\alpha=1\%$.
+Since we removed approximately half the decoys for training purposes in RESET, the bounds need to be doubled to fairly compare between the two approaches.
+Refer to caption Refer to caption
+Figure F4: RESET’s and FDP-SD’s bounds on the number of decoy wins.
+We recorded, on the left y-axis, the bounds used to determine the list of FDP-SD discoveries and twice the bounds used to determine RESET’s discoveries at a range of indices at $\alpha=1\%$ and $1-\gamma=90\%$.
+On the right y-axis, we plot the ratio of these bounds (in black).
+The left panel looks at indices between 501 to 2K, while the right panel looks at indices from 10K to 20K.
+An index corresponds to the number of top scoring hypotheses (regardless of their labels).
+Figure F4 shows the bounds for FDP-SD, and double the bounds for RESET, along with the ratio of the two at a range of different indices.
+We can see that for smaller indices, this ratio is much smaller, initially at 0% at an index of 501 and 50% at an index of 1K, while at 10K this ratio is much higher at 94%.
+If the order of the hypotheses is the same, then RESET’s smaller bounds imply it will report fewer discoveries because RESET employs FDP-SD to control the FDX by searching for the largest index $i$ s.t.
+$A_{i_{0}}\leq\delta_{i_{0}},\dots,A_{i}\leq\delta_{i}$, and reports the top $i$ scoring positively labeled hypotheses.
+In other words, if RESET is to report roughly the same number of discoveries as FDP-SD, then RESET needs to rearrange the targets and estimating decoys successfully enough so as to make up for a 50% difference in the bounds when considering the top 1K scores, while only 6% when considering the top 10K.
+If the side information is not rich enough, then it is possible RESET will report fewer discoveries than FDP-SD in these cases.
+We intend to investigate this as part of future work.
+F.5 Variability of RESET
+RESET’s discovery list is variable due to the random split of the decoys.
+This variability is demonstrated in the histogram of the number of discoveries in Figure F5A using 100 applications of RESET Ensemble to the Airway dataset at an FDR threshold of 10%.
+The resulting number of discoveries has a mean of 6018 with a standard deviation of about 75.
+While this is undesirable, this analysis does not take into account the randomness of the data, and therefore that the added variance from RESET may be marginal.
+Indeed, in Figure F5B, we find that in the case of Simulation 5 from Section 6.2.1, RESET is overall arguably less variable than AdaPT, even though AdaPT does not employ any internal randomization.
+In practice, users may get around RESET’s randomization by setting the random number generator to be a function of the input data and parameters.
+We plan on adding this feature to a future update.
+Figure F5: Variability in the number of discoveries.
+(A) Histogram of the number of RESET Ensemble discoveries applied 100 times to the Airway data with a 10% FDR threshold and varying its internal seed.
+(B) Boxplots of RESET’s and AdaPT’s power in 100 applications of the ‘circle in the middle’ setup of Simulation 5 using a 10% FDR threshold.
+There is one major outlier in RESET’s case (which occurred when the signal was considerably low).